### v6

# Data-Driven Logistic Regression Ensembles With Applications in Genomics

Abstract
Advances in data collecting technologies in genomics have significantly increased the need for tools designed to study the genetic basis of many diseases. Effective statistical methods should excel in both prediction accuracy and biomarker identification. We introduce a novel approach to high-dimensional binary classification that integrates regularization with ensembling techniques. Our method constructs compact ensembles of interpretable models derived by optimizing a global objective function. In medical genomics applications, our approach identifies critical biomarkers overlooked by competing methods. We develop a variable importance ranking system to help researchers prioritize promising genes. The method’s asymptotic properties are established, and an efficient computational algorithm is provided. Through extensive simulations across complex scenarios and analysis of genomics datasets for cancer, multiple sclerosis, and psoriasis, we demonstrate strong predictive performance. Based on our numerical experiments, we offer practical guidelines for determining optimal ensemble size.
Keywords: Consistent ensemble; Diversity penalty; High-dimensional classification; Logistic regression models; Split modeling.
1 Introduction
The surge of genomic data collection through ever-evolving technologies has necessitated the development of sophisticated statistical methods to analyze high-dimensional gene expression data, unlocking its potential for breakthroughs in healthcare and scientific discovery. In many medical genomics applications, the goal is to train a classifier that can accurately predict the presence of a disease or particular subtypes of a disease based on the genetic profile of a patient. Furthermore, in the medical sciences and other fields predictions often have important consequences in decision-making processes, so there is a high demand for interpretable learning methods (see e.g. Murdoch et al., 2019; Rudin, 2019; Rudin et al., 2022). It should thus be clear how a classifier arrives at its predictions so that decisions can be fully explained. In addition, medical researchers aim to unravel the relation between the genetic profile and occurrence of a disease, so an appropriate classification method should also be able to identify patterns between the expression of key biomarkers and the presence of a disease.
The vast richness and availability of medical genomics data is evident from large publicly accessible databases such as the Gene Expression Omnibus (GEO) database (Edgar et al., 2002; Barrett et al., 2012). In many of these datasets, a large number of gene expression measurements are collected for a relatively small number of (cell tissue) samples. The status of the sample is typically given as well, specifying e.g. whether the sample collected was affected by a disease or not. Classification methods can then be used to discover patterns between the expression level of certain genes and the presence of the disease. The two types of methods that are generally deployed in such applications are sparse and ensemble methods, respectively.
On the one hand, sparse methods yield interpretable models that only use a subset of the genes to make a decision  (see e.g. Hastie et al., 2015, for a modern treatment of sparse methods). However, because the number of predictor genes is much larger than the number of samples, there may be several models comprised of different subsets of genes that are equally accurate in predicting the presence of the disease (a phenomenon coined “the multiplicity of good models” by McCullagh and Nelder, 1989). Thus, several potentially important genes may be erroneously discarded from the decision-making process when using a single sparse model. On the other hand, ensemble methods combine multiple diverse models and generally achieve superior prediction performance if the members of the ensemble are sufficiently diverse. Diversity is often achieved using randomization (see e.g. Ho, 1998; Breiman, 2001; Song et al., 2013) or sequentially fitting the residuals of the previous fit (see e.g. Friedman, 2001; Bühlmann and Yu, 2003; Schapire and Freund, 2012; Yu et al., 2020). Ensemble methods have been particularly successful in high-dimensional prediction tasks with genomic data (see e.g. Dorani et al., 2018; Zahoor and Zafar, 2020). However, while ad hoc methods have been developed to assess variable importance in some ensemble methods, such as the variable importance measure of  Breiman (2001), interpretation of the resulting prediction rules and identification of important predictor genes (i.e. key biomarkers) is less straightforward.
We propose a new approach to learn a diverse ensemble of sparse classification models that is especially well suited for high-dimensional medical genomics applications. Specifically, we extend the regression ensemble method proposed by Christidis et al. (2020) to the binary classification setting. We demonstrate the advantage of our ensemble method in terms of both prediction and identification of key biomarkers using several datasets from the GEO database. These datasets involve different types of common diseases such as cancer, multiple sclerosis and psoriasis. In particular, we show that the prediction accuracy of the sparse models in our ensembles matches the prediction accuracy of standard single-model sparse methods. Since the models in the ensembles are learned simultaneously and directly form the data (free of randomization or other heuristics) by optimizing a global objective function, they each provide an alternative explanation for the relationship between the genes and disease. The objective function contains a penalty that promotes diversity in the obtained models so that individual models may be driven by potentially different biological mechanisms. With respect to the identification of important predictor genes, our examples demonstrate that the proposed ensemble method can identify key biomarkers that are discarded by state-of-the-art sparse methods and ensemble variable ranking methods. At the same time, we show that the proposed method generally includes key biomarkers selected by sparse methods or flagged as important by ensemble variable ranking methods.
Each of the models in our proposed ensemble method is a penalized logistic regression with a sparsity inducing penalty such as the Lasso (Donoho and Johnstone, 1994; Tibshirani, 1996) or the elastic net (Zou and Hastie, 2005). Rather than resorting to randomization or other indirect methods to generate different models, we jointly learn the models in the ensemble on the training data and incorporate a diversity penalty (Christidis et al., 2020) in the objective function with the aim to diversify the models. The degree to which these models are sparse and diverse is driven directly by the data. In this way, the method efficiently exploits the so-called accuracy-diversity trade-off between the models, and generates an ensemble with high predictive performance that often matches or even outperforms popular state-of-the-art ensemble methods. Moreover, by ensembling the models at the level of their linear predictors we retain interpretability for the logistic regression coefficients in the ensembled model. We use several measures of diversity (Kuncheva and Whitaker, 2003) to study the accuracy-diversity trade-off in our ensemble method. This trade-off provides insight in the effect of the number of ensembled models on the performance of the resulting ensemble. This insight makes it possible to make some recommendations for the choice of the number of ensembled models.
Following the methodological development framework of Heinze et al. (2024), this work encompasses phases I and II, introducing a novel method with initial validation through comprehensive simulations across various scenarios (sample sizes, sparsity, correlation structures, class imbalance, non-linearity, and interactions).While the results of our numerical experiments and medical genomics data applications demonstrate the promising performance of the proposed classification ensembles across a variety of data scenarios, we present this work as a method development study with strong proof-of-concept validation, rather than a claim of universal superiority. The success observed across our diverse testing scenarios suggests that phases III and IV—independent validation and practical implementation—may naturally follow as the method gains adoption in the research community.
The remainder of the paper is organized as follows. In Section 2 we briefly review the literature on sparse and ensemble methods, and introduce the proposed method to learn a data-driven diverse ensemble of sparse logistic regression models. We show that this new ensemble method is consistent under mild regularity assumptions. In Section 3, we analyze lung and thyroid cancer genomic datasets to highlight the potential benefits of our hybrid approach over standard sparse and ensemble methods. We show how the method can uncover multiple mechanisms linking predictors to the response in the thyroid cancer dataset and how it can rank genes by importance. In Section 4 an efficient block coordinate descent algorithm to solve the multi-convex optimization problem is provided. Section 5 presents the results of an extensive simulation study to compare the proposed method with a large number of state-of-the-art competitors for high-dimensional binary classification. In Section 6 we further show the excellent performance of the proposed method in a numerical study using a large collection of high-dimensional medical genomic datasets. To provide insight in how to choose the number of models in the proposed method, we study the accuracy-diversity trade-off and computational cost of our method in Section 7. Potential directions for further research are provided in Section 8.
2 Split Logistic Regression
Training classification models on high-dimensional data, such as gene expression profiles where the number of measurements ($p$) exceeds the number of tissue samples ($n$), presents significant statistical and computational challenges. This $p>n$ scenario is common in bioinformatics and other fields. To address these challenges, two primary families of methods have emerged as dominant strategies in the literature: sparse modeling, which aims for feature selection and interpretability, and ensemble methods, which typically focus on maximizing predictive accuracy and stability by combining multiple models.
In this section, we briefly review representative state-of-the-art techniques from both families, discussing their strengths and limitations within the high-dimensional setting. Building upon this context, we then introduce Split Logistic Regression, our proposed hybrid approach specifically designed to combine the variable selection benefits of sparse methods with the generally superior predictive performance of ensemble learning.
2.1 Sparse Methods
Let $\mathcal{D}=\left(\mathbf{y},\mathbf{X}\right)$ denote the training data, where $\mathbf{y}\in\mathbb{R}^{n}$ is the vector of class labels and $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the design matrix which consists of $n$ measurements $\mathbf{x}_{i}$ on $p$ predictors. We consider the binary classification problem where the classes are labeled by $y_{i}\in\{0,1\}$. The predictor variables have been standardized, i.e. $\sum\limits_{i=1}^{n}x_{ij}/n=0$ and $\sum\limits_{i=1}^{n}x_{ij}^{2}/n=1$, $1\leq j\leq p$. Logistic regression models the class-conditional probabilities through a non-linear function of the predictor variables,
$p_{i}=\mathbb{P}(y_{i}=1|\mathbf{x}_{i})=S(\beta_{0}+\mathbf{x}_{i}^{T}\boldsymbol{\beta}),\quad 1\leq i\leq n,$ (1)
where $\beta_{0}$ and $\boldsymbol{\beta}\in\mathbb{R}^{p}$ are the intercept and vector of regression coefficients, respectively. The function $S(t)=e^{t}/(1+e^{t})$ is the well-known logistic function. With $f(\mathbf{x}_{i})=\beta_{0}+\mathbf{x}_{i}^{T}\boldsymbol{\beta}$, the negative log-likelihood for a single observation $(y_{i},\mathbf{x}_{i})$, often referred to as the logistic loss or log loss, is given by
$\mathcal{L}\left(f(\mathbf{x}_{i}),y_{i}\right)=\mathcal{L}(\beta_{0},\boldsymbol{\beta}\mid y_{i},\mathbf{x}_{i})=-y_{i}f(\mathbf{x}_{i})+\log(1+e^{f(\mathbf{x}_{i})}).$ (2)
This loss function arises directly from the Bernoulli likelihood where $\mathbb{P}(y_{i}=1|\mathbf{x}_{i})=p_{i}=S(f(\mathbf{x}_{i}))$ and $\mathbb{P}(y_{i}=0|\mathbf{x}_{i})=1-p_{i}=1-S(f(\mathbf{x}_{i}))=S(-f(\mathbf{x}_{i}))$. Minimizing the corresponding empirical loss (average negative log-likelihood)
$\mathcal{V}_{n}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(f(\mathbf{x}_{i}),y_{i}\right)$ (3)
via standard maximum likelihood estimation requires $p<n$. When $p\geq n$, the estimation problem is typically ill-posed. Even when $p<n$, if $p$ is large relative to $n$, minimizing (3) often leads to overfitting of the training data and poor out-of-sample prediction accuracy.
By restricting model complexity, sparse methods aim to find a single (sparse) model that achieves good prediction accuracy using only a small subset of the predictors. They have proven to be highly successful approaches for high-dimensional classification problems in the genomic sciences (see e.g. Zuo et al., 2017; Rejchel and Bogdan, 2020). Sparse regularization methods typically solve an optimization problem of the form
$\min_{\beta_{0}\in\mathbb{R},\,\boldsymbol{\beta}\in\mathbb{R}^{p}}\mathcal{V}_{n}(f)+\lambda P_{s}\left(\boldsymbol{\beta}\right),$ (4)
where $P_{s}(\boldsymbol{\beta})$ is a penalty function that induces sparsity in the coefficient vector $\boldsymbol{\beta}$. The tuning parameter $\lambda>0$ is usually determined in a data-driven way, typically by cross-validation (CV). A common and effective choice for the penalty term $P_{s}$ is the elastic net penalty (Zou and Hastie, 2005). It combines $\ell_{1}$ and $\ell_{2}$ regularization components:
$P_{s}(\boldsymbol{\beta})=\frac{1-\alpha}{2}\|\boldsymbol{\beta}\|_{2}^{2}+\alpha\|\boldsymbol{\beta}\|_{1},\quad\alpha\in[0,1].$ (5)
The mixing parameter $\alpha$ balances the sparsity-inducing $\ell_{1}$ norm $\|\boldsymbol{\beta}\|_{1}$ and the squared $\ell_{2}$ norm $\|\boldsymbol{\beta}\|_{2}^{2}$. This penalty includes Lasso (Tibshirani, 1996) ($\alpha=1$) and Ridge (Hoerl and Kennard, 1970) ($\alpha=0$) as special cases.
Next to Lasso and elastic net, some of the more popular sparse regularization methods are the adaptive Lasso (Zou, 2006), the relaxed Lasso (Meinshausen, 2007), the smoothly clipped absolute deviation (SCAD) estimator (Fan and Li, 2001) and the minimum concave penalized (MCP) estimator (Zhang, 2010). Additional approaches include Sure Independence Screening (SIS) (Fan and Lv, 2008), which first reduces dimensionality before applying a penalty like SCAD, and RuleFit (Friedman and Popescu, 2008), which combines sparse linear terms with rule-based terms derived from decision trees. A vast amount of asymptotic theory has been developed for a large class of regularized estimators, see e.g. Bühlmann and van de Geer (2011) for an extensive treatment. In summary, sparse regularization methods yield a single interpretable model with good prediction accuracy and extensive asymptotic theory that describes their behavior.
2.2 Ensemble Methods
Ensemble methods have proven to be very successful in high-dimensional classification tasks, often yielding higher prediction accuracy than their sparse single-model competitors. To better understand the behavior of ensemble models, Ueda and Nakano (1996) first developed a decomposition of its generalization error for the regression case. Brown et al. (2005) provided an in-depth analysis of the bias-variance-covariance trade-off in regression ensembles. In particular, if the ensemble of a collection of estimators $\hat{f}_{1},\dots,\hat{f}_{G}$ is their average $\bar{f}=\sum_{g=1}^{G}\hat{f}_{g}/G$, then its mean squared prediction error (MSPE) can be decomposed as $\text{MSPE}[\bar{f}]=\text{Bias}[\bar{f}]^{2}+\text{Var}[\bar{f}]+\sigma^{2}$, where $\sigma^{2}$ is the irreducible variance of the errors. The bias and variance of the regression ensemble can be decomposed further as
$\text{Bias}\left[\bar{f}\right]$ $=\overline{\text{Bias}}_{G},$ (6)
$\text{Var}\left[\bar{f}\right]$ $=\frac{1}{G}\ \overline{\text{Var}}_{G}+\frac{G-1}{G}\ \overline{\text{Cov}}_{G},$ (7)
where $\overline{\text{Bias}}_{G}$, $\overline{\text{Var}}_{G}$ and $\overline{\text{Cov}}_{G}$ are the average of the biases, variances, and pairwise covariances of the $G$ estimators in the ensemble, respectively. From (7) it becomes clear that if the number of estimators increases, their correlations play a much more critical role than their average variability to obtain a good ensemble estimator. A similar principle was derived for classifier ensembles by Tumer and Ghosh (1996) and later refined by Fumera and Roli (2003).
The importance of diversity among the constituent models within an ensemble is a well-established principle, exemplified by popular methods like random forests (Breiman, 2001). In random forests, diversity is achieved by constructing individual decision trees using random subsets of candidate features at each split point. This strategy results in less correlated trees, which collectively enhance the ensemble’s predictive accuracy compared to using highly similar trees. Other techniques also leverage randomization to foster diversity. For instance, the random generalized linear models (RGLM) method of Song et al. (2013) combines bagging (Breiman, 1996a) (random sampling of data points) with the random predictor subspace method (Ho, 1998) (random sampling of features). This dual randomization creates varied training conditions for each base model, leading to a diverse set of intentionally simpler, decorrelated models. Sequential ensemble methods employ a different strategy. In methods like (extreme) gradient boosting (Chen and Guestrin, 2016), diverse trees are generated iteratively, with each new tree trained to correct the errors (residuals) made by the preceding ensemble. While individual trees might be relatively weak predictors on their own, this sequential error correction process builds a powerful and accurate final ensemble. A third approach to ensemble learning is stacking (Breiman, 1996b), where several different algorithms are trained on the same dataset, and their predictions are combined through a meta-learning algorithm that learns optimal weights for aggregating base model outputs. The specific relationship between various measures of diversity in classifier ensembles and their resulting prediction accuracy has been formally investigated by Kuncheva and Whitaker (2003).
In summary, many conventional ensemble methods face challenges regarding interpretability, as they often aggregate predictions from numerous relatively simple or “weak” base models, where the contribution of individual predictors can be obscured within the combined structure. While the overall ensemble might perform well, understanding the driving factors requires specific tools. Ad hoc measures have indeed been introduced to gauge predictor importance in some widely used ”black-box” ensemble methods, such as the Gini importance index commonly employed in random forests (Breiman, 2001). Finally, while ensemble methodologies often incorporate heuristic elements in their construction, theoretical understanding of their behavior has advanced. Significant research has been dedicated to analyzing their properties, leading to important results such as proofs of consistency for methods like Random Forests (Biau et al., 2008).
2.3 Split Logistic Regression Models
We now introduce split logistic regression which combines the stability and interpretability of sparse methods with the high accuracy achieved by ensemble methods. Let $\left(\mathbf{y},\mathbf{X}\right)$ denote the training data, as before. Using the logistic loss function (14) for $\mathcal{L}(\beta_{0}^{g},\boldsymbol{\beta}^{g}\mid y_{i},\mathbf{x}_{i})$ in each model with $f_{g}(\mathbf{x}_{i})=\beta_{0}^{g}+\mathbf{x}_{i}^{T}\boldsymbol{\beta}^{g}$, the split logistic regression objective function to simultaneously fit $G$ models is given by
$\mathcal{J}\left(\beta_{0}^{1},\boldsymbol{\beta}^{1},\dots,\beta_{0}^{G},\boldsymbol{\beta}^{G}\right)=\sum_{g=1}^{G}\left[\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(\beta_{0}^{g},\boldsymbol{\beta}^{g}|y_{i},\mathbf{x}_{i}\right)+\lambda_{s}P_{s}(\boldsymbol{\beta}^{g})\right]+\frac{\lambda_{d}}{2}\sum_{h\neq g}P_{d}(\boldsymbol{\beta}^{h},\boldsymbol{\beta}^{g}),$ (8)
which needs to be minimized with respect to all regression coefficients. In the regression ensemble setting considered in Christidis et al. (2020), the loss function $\mathcal{L}(\beta_{0}^{g},\boldsymbol{\beta}^{g}|y_{i},\mathbf{x}_{i})$ used for each model is the squared error loss. The sparsity penalty function $P_{s}$ regularizes each of the $G$ individual model such that they contain a subset of the available predictors that work well together to model and predict the outcome. On the other hand, the goal of the diversity penalty $P_{d}$ is to restrict the models by discouraging each variable from appearing in several models. Hence, the diversity penalty encourages the different models in the ensemble to complement each other well.
Note that the diversity penalty $P_{d}$ needs to have two desirable properties. First, it should encourage the selection of uncorrelated models. Secondly, it should be computationally tractable so that the objective function (18) can be minimized in a stable and timely manner. Finally, for moderate values of $G$ the ensemble tends to be sparse in the sense that the set of predictors that appear in at least one of the models will be much smaller than the complete set of candidate predictors.
For the sparsity penalty we use the elastic net penalty in (16) but other penalties such as SCAD or MCP could be used as well. For the diversity penalty which encourages the individual models to be sufficiently different, we use the proposal of Christidis et al. (2020),
$P_{d}(\boldsymbol{\beta}^{h},\boldsymbol{\beta}^{g})=\sum_{j=1}^{p}|\beta_{j}^{g}||\beta_{j}^{h}|.$ (9)
This penalty directly measures the overlap between coefficient vectors of different models. When two models both include the same variable $j$ with non-zero coefficients, the penalty adds a cost proportional to the product $|\beta_{j}^{g}||\beta_{j}^{h}|$. Larger coefficient magnitudes in shared variables incur greater penalties, creating a strong incentive for models to utilize different sets of predictors.
This formulation encourages diversity by forcing models to focus on different aspects of the data: when one model captures a particular signal using certain variables, other models are directed to find alternative representations using different variables. We selected this specific penalty form because it maintains the multi-convex structure of the objective function enabling efficient optimization, operates directly on model coefficients rather than predictions preserving interpretability, and allows for a continuous trade-off between diversity and accuracy through the parameter $\lambda_{d}$. In high-dimensional settings with correlated predictors, this leads to models that capture complementary information, approaching the prediction problem from different perspectives while sharing variables only when their predictive value sufficiently outweighs the diversity penalty cost.
The tuning constants $\lambda_{s},\lambda_{d}\geq 0$ control the amount of shrinkage and diversity between the models, respectively. Letting $\lambda_{d}\to\infty$, enforces that the diversity penalty $P_{d}\left(\boldsymbol{\beta}^{h},\boldsymbol{\beta}^{g}\right)\to 0$ for all $g\neq h$ so that the active variables in each of the individual models are distinct. In this case the models in the ensemble are fully diverse in the sense that they do not share any predictor variables. On the other hand, it can be seen that for $\lambda_{d}=0$, the solution for all $G$ individual models is the same. In this case all models are equal to the logistic elastic net solution with penalty parameter $\lambda_{s}$, which is then also the split logistic regression ensemble solution. Hence, split logistic regression is a generalization of regularized logistic regression and allows for the selection of $G>1$ accurate and potentially diverse models that can be combined well in an ensemble. Note that since both $\lambda_{s}$ and $\lambda_{d}$ are chosen by CV, the degree of sparsity of the models and diversity between the models is driven by the data.
Minimizing the split logistic regression objective function (18) yields solutions $\hat{f}_{g}(\mathbf{x})=\hat{\beta}_{0}^{g}+\mathbf{x}^{T}\boldsymbol{\hat{\beta}}^{g}$ for $g=1,\dots,G$ which are well-suited for creating an ensemble. We use the ensembling function
$\hat{f}(\mathbf{x})=S\left(\frac{1}{G}\sum_{g=1}^{G}\hat{f}_{g}(\mathbf{x})\right)=S\left(\frac{1}{G}\sum_{g=1}^{G}\hat{\beta}_{0}^{g}+\mathbf{x}^{T}\left(\frac{1}{G}\sum_{g=1}^{G}{\boldsymbol{\hat{\beta}}}^{g}\right)\right).$ (10)
The advantage of this ensembling function is that the ensemble also becomes a logistic transformation of a linear function. For any predictor $j$, we denote its ensemble coefficient as $\bar{\beta}_{j}=\frac{1}{G}\sum_{g=1}^{G}\hat{\beta}_{j}^{g}$, which represents the average effect across all models in the ensemble. Let $\mathcal{M}_{g}=\{j:\hat{\beta}_{j}^{g}\neq 0\}$ denote the set of variables selected for model $g$. To investigate the importance of variables, we can then consider the sets:
$\mathcal{A}_{k}=\left\{j:\sum_{g=1}^{G}\mathbb{I}\left(j\in\mathcal{M}_{g}\right)\geq k\right\},\quad 1\leq k\leq G,$ (11)
where $\mathcal{A}_{G}\subseteq\mathcal{A}_{G-1}\subseteq\cdots\subseteq\mathcal{A}_{1}$. These sets identify variables in order of their importance, as those appearing in multiple models must contribute substantially to the loss function reduction to overcome the diversity penalty.
While the resulting model maintains the structure of a logistic regression, it is important to note that the interpretation of the ensemble coefficients requires careful consideration. As for penalized regression coefficients in general, they exhibit bias due to shrinkage effects. However, they do provide valuable information about the direction and relative magnitude of predictor effects. The combined information—average effect ($\bar{\beta}_{j}$) and selection frequency through the sets $\mathcal{A}_{k}$—enhance interpretability compared to black-box ensemble approaches while maintaining prediction accuracy.
For the ensembling function (10), consistency of split logistic regression is established in Theorem 1 below. The proof of Theorem 1 is provided in the supplementary material where we also study its prediction error in the more general case of model misspecification.
Theorem 1
Assume the data $\left(y_{i},\mathbf{x}_{i}\right)$, $1\leq i\leq n$, follow a logistic regression model for some $\boldsymbol{\beta}^{*}\in\mathbb{R}^{p}$, with $\lVert\boldsymbol{\beta}^{*}\rVert_{1}$ and $\lVert\boldsymbol{\beta}^{*}\rVert_{2}$ of order smaller than $\sqrt{n/\log(p)}$ and $\log(p)/n\to 0$. Let $\hat{f}_{1},\dots,\hat{f}_{G}$ be the solution of (18). If $\lambda_{s}$ and $\lambda_{d}$ are of order $\sqrt{\log(p)/n}$, then the ensemble prediction $\hat{f}$ given in (10) is consistent.
The proposed split logistic regression ensemble approach offers several potential empirical benefits compared to existing sparse regression and ensemble methods. By jointly estimating multiple sparse models with an explicit diversity penalty, our method can potentially discover different underlying mechanisms while maintaining the interpretability of each component model. By ensembling at the linear predictor level rather than the prediction level, we preserve coefficient interpretability while potentially improving predictive performance through the accuracy-diversity trade-off. This approach may also identify a broader set of relevant predictors across the ensemble compared to single sparse models. In the next section, we illustrate these benefits using cancer genomics datasets, where we demonstrate that our approach achieves strong predictive performance from a small number of models, each individually accurate and interpretable, while potentially revealing multiple biological pathways associated with cancer outcomes—a feature particularly valuable in high-dimensional biomedical applications.
3 Empirical Evaluation in Cancer Genomics
We demonstrate the effectiveness of split logistic regression through applications to two cancer genomics datasets, highlighting its strong performance in both prediction accuracy and biomarker identification compared to state-of-the-art sparse and ensemble methods. The comparison reveals that the proposed method achieves competitive or superior predictive performance while maintaining interpretable model structures. Furthermore, we illustrate how the variable importance sets defined in (11) can be leveraged to rank genes according to their biological relevance, providing researchers with a principled approach to prioritize potential biomarkers for further investigation.
In a typical medical genomics data application the training data consist of thousands of gene expression level measurements on a limited number of samples, collected from some cancer tumor tissue or from some adjacent normal tissue for example. In such an application the main goal is to train a classification model that accurately identifies the tumor status of the tissue samples using the available gene expression data. At the same time, medical researchers wish to identify important subsets of differentially expressed genes and the way they operate. As motivating examples we consider the lung and thyroid cancer genomic datasets, GSE10245 and GSE5364 respectively, from the Gene Expression Omnibus (GEO) database. GSE10245 consists of $n=58$ non-small cell lung cancer (NSCLC) tissue samples of subtypes adenocarcinoma (AC) or squamous cell carcinoma (SCC), with 40 samples of AC and 18 samples of SCC. After matching the probes to their corresponding genes, the dataset includes $p=$ 13,515 gene expression measurements. GSE5364 consists of $n=51$ cell tissue samples comprising 16 thyroid cancer tumor samples and 35 adjacent normal tissue samples, also with $p=$ 13,515 gene expression measurements after matching probes to their corresponding genes.
The objective is to construct a classifier that can accurately specify the cancer tumor status of the samples, AC or SCC subtype for the lung cancer dataset and cancer tumor or adjacent tissue sample for the thyroid cancer dataset, using their gene expression levels. The classifier should also identify, for each dataset, all possible relevant biomarkers that play a role in the differentiation between the two types of cell tissue samples. We acknowledge that several cancer datasets in this article, including the thyroid cancer dataset GSE5364, contain paired samples where both cancerous and adjacent normal tissues come from the same patients. This pairing introduces potential correlation between observations, while split logistic regression, like many classification methods, is developed under the assumption of sample independence. Nevertheless, such paired designs are ubiquitous in cancer genomics research, reflecting realistic clinical data collection practices. To ensure fair evaluation, all compared methods are assessed using identical data splits, consistent with established benchmarking protocols in the field. To provide a more comprehensive assessment, we deliberately included both paired datasets (such as GSE5364) and unpaired datasets (such as GSE10245 for lung cancer), enabling evaluation of classifier performance across these complementary experimental designs.
The data are pre-processed in a standard way by using a four step procedure similar to Dudoit et al. (2002): (1) thresholding of gene expression levels with a floor of 20 and a ceiling of 20,000; (2) exclusion of genes for which the ratio between the maximum and minimum expression level is less than 2; (3) exclusion of genes for which the difference between the maximum and minimum expression level is less than 200; (4) base-2 logarithmic transformation. For most methods we then retain from the remaining genes the $p=500$ genes which are most important according to their $q$-value (Storey, 2002; Storey et al., 2023). The $q$-value is a popular alternative to the $p$-value for high-dimensional data that controls for the false discovery rate and alleviates the issue of overestimating significance that arises from using multiple pairwise $t$-tests.
3.1 Methods and Performance Measures
We use the thyroid and lung cancer genomic datasets to illustrate the performance of several existing state-of-the-art sparse and ensemble methods and to compare them to our new proposal. We use the methods listed in Table 1 as implemented in their respective CRAN (R Core Team, 2024) R packages.
Table 1: Methods and Software Implementations
ID Method Abbreviation CRAN Package CRAN Reference
1 Split-Lasso Split-Lasso SplitGLM Christidis et al. (2021)
2 Split-Elastic Net Split-EN SplitGLM Christidis et al. (2021)
3 Lasso Lasso glmnet Friedman et al. (2010)
4 Elastic Net EN glmnet Friedman et al. (2010)
5 Adaptive Lasso Adaptive gcdnet Yang and Zou (2017)
6 Relaxed Lasso Relaxed glmnet Friedman et al. (2010)
7 Minimum concave penalized MCP ncvreg Breheny and Huang (2011)
8 Sure Independence Screening SIS-SCAD SIS Saldana and Feng (2018)
9 RuleFit RuleFit xrf Holub (2022)
10 Random Lasso RE-Lasso -- --
11 Random Elastic Net RE-EN -- --
12 Random GLM RGLM RGLM Song and Langfelder (2013)
13 Random Forest RF ranger Wright and Ziegler (2017)
14 Extreme Gradient Boosting XGB xgboost Chen et al. (2020)
Methods 1-2 are the split logistic regression ensembles proposed in Section 2.3, which combine sparse modeling with ensemble learning. The proposed method has been implemented in the R/C++ package SplitGLM (Christidis et al., 2021) available on CRAN, and uses either the Lasso or the elastic net as sparsity penalty. Methods 3-9 in Table 1 are sparse methods, methods 10-14 are “blackbox” ensemble methods. In particular, methods 11-12 (RE-Lasso and RE-EN) represent simple randomized ensembles that apply Lasso or elastic net to different bootstrap samples with random feature subsets, providing a direct comparison to our proposed approach.
[…]
To evaluate prediction performance we report the prediction accuracy (ACC) of all the methods. For the ensemble methods (Split-Lasso, Split-EN, RGLM, RF and XGB) we also report the average prediction accuracy of the individual models ($\overline{\text{ACC}}$) which measures the prediction accuracy of the individual models used in an ensemble. We will also report the sensitivity (SNS) and specificity (SPC) where these measures are important as for both datasets, data is unbalanced (AC much rarer in lung cancer dataset, cancer tissue much rarer in thyroid cancer dataset). To estimate ACC, $\overline{\text{ACC}}$, SNS and SPC, we randomly split the lung and thyroid datasets $N=50$ times into training sets of size $n=29$ and $n=26$, respectively, to fit the models, and corresponding tests sets of size $m=29$ and $m=25$, respectively, to evaluate the prediction performance of the methods. To analyze the most important genes for random forest via the variable importance measure, we note that the standard Gini importance measure is known to be biased (Strobl et al., 2007; Debeer and Strobl, 2020), particularly when predictors vary in scale or are highly correlated, which is commonly observed in genomic data. Therefore, we implement the unbiased Gini importance measure as proposed by Nembrini et al. (2018), which addresses these biases by accounting for predictor correlation structures and providing more reliable feature importance rankings.
To evaluate the performance of the methods with respect to biomarker discovery, we examine each method’s most frequently selected genes, verify whether they have been previously identified as cancer biomarkers in the literature, and assess the overlap of these top selections across different methods.
3.2 Prediction Performance
In Table 2, we observe that Split-Lasso and Split-EN methods achieve excellent ensemble prediction performance (ACC) on both lung and thyroid cancer datasets, with values of 0.93 and 0.90 respectively, despite using only a small number of models ($G=10$). Remarkably, the individual models within these ensembles maintain high prediction accuracy $\overline{\text{ACC}}$, with values of 0.90-0.92 for lung cancer and 0.85-0.86 for thyroid cancer, nearly matching their corresponding single-model counterparts. This strong individual model performance is particularly noteworthy given that Split-Lasso and Split-EN incorporate diversity penalties that deliberately encourage important genes to be distributed across different models in the ensemble. Despite this forced diversification, individual models in Split-EN remain highly sparse, containing on average only 24 and 21 genes for the lung and thyroid cancer datasets respectively across the $N=50$ replications, compared to 38 and 33 genes on average for single-model EN. Furthermore, the Split methods exhibit strong performance in both SNS and SPC measures, achieving values of 0.81 for SNS and 0.98-0.99 for SPC in the lung cancer dataset, and 0.81-0.82 for SNS and 0.94 for SPC in the thyroid cancer dataset. This balanced performance across independent samples (lung cancer) and paired samples (thyroid cancer) suggests that the methods can effectively handle both data structures while maintaining high classification accuracy for both majority and minority classes in these unbalanced datasets.
Table 2: Average ACC, $\overline{\text{ACC}}$, SNS, and SPC of sparse and ensemble methods estimated using $N=50$ random splits of the lung and thyroid cancer genomic datasets into training sets with $n=29$ and $n=26$ samples respectively, and test sets with the remaining samples. The standard errors are in parenthesis. The three best results for each criterion are highlighted in bold.
[…]
=2pt GSE10245 (Lung) GSE5364 (Thyroid) Method ACC $\mathbf{\overline{ACC}}$ SNS SPC ACC $\mathbf{\overline{ACC}}$ SNS SPC Split-Lasso-10 0.93 (0.05) 0.90 (0.03) 0.81 (0.15) 0.99 (0.02) 0.90 (0.05) 0.85 (0.05) 0.81 (0.11) 0.94 (0.08) Split-EN-10 0.93 (0.04) 0.92 (0.04) 0.81 (0.14) 0.98 (0.02) 0.90 (0.05) 0.86 (0.05) 0.82 (0.10) 0.94 (0.08) Lasso 0.91 (0.06) -- 0.76 (0.17) 0.98 (0.03) 0.86 (0.08) -- 0.77 (0.18) 0.91 (0.10) EN 0.92 (0.05) -- 0.80 (0.14) 0.99 (0.02) 0.88 (0.07) -- 0.79 (0.16) 0.93 (0.09) Adaptive 0.82 (0.10) -- 0.52 (0.30) 0.96 (0.05) 0.87 (0.10) -- 0.71 (0.27) 0.94 (0.11) Relaxed 0.90 (0.06) -- 0.79 (0.15) 0.96 (0.05) 0.84 (0.09) -- 0.77 (0.17) 0.88 (0.12) MCP 0.86 (0.09) -- 0.66 (0.18) 0.95 (0.06) 0.82 (0.11) -- 0.66 (0.24) 0.90 (0.12) SIS-SCAD 0.85 (0.08) -- 0.62 (0.15) 0.96 (0.06) 0.82 (0.11) -- 0.67 (0.23) 0.89 (0.11) RuleFit 0.81 (0.09) -- 0.61 (0.21) 0.91 (0.09) 0.85 (0.09) -- 0.70 (0.22) 0.93 (0.07) RE-Lasso-10 0.91 (0.05) 0.89 (0.04) 0.76 (0.15) 0.99 (0.02) 0.89 (0.07) 0.83 (0.05) 0.81 (0.13) 0.93 (0.09) RE-Lasso-100 0.92 (0.05) 0.88 (0.04) 0.77 (0.16) 0.99 (0.02) 0.90 (0.08) 0.82 (0.05) 0.79 (0.19) 0.95 (0.07) RE-EN-10 0.92 (0.05) 0.91 (0.04) 0.79 (0.16) 0.99 (0.02) 0.89 (0.06) 0.86 (0.06) 0.83 (0.11) 0.93 (0.09) RE-EN-100 0.92 (0.05) 0.91 (0.04) 0.79 (0.15) 0.99 (0.02) 0.90 (0.06) 0.86 (0.05) 0.84 (0.10) 0.93 (0.08) RGLM-10 0.90 (0.07) 0.82 (0.04) 0.74 (0.18) 0.98 (0.03) 0.89 (0.07) 0.76 (0.05) 0.82 (0.12) 0.92 (0.09) RGLM-100 0.91 (0.05) 0.82 (0.03) 0.76 (0.16) 0.99 (0.02) 0.90 (0.06) 0.76 (0.03) 0.82 (0.11) 0.94 (0.09) RF-10 0.88 (0.08) 0.76 (0.04) 0.71 (0.19) 0.96 (0.05) 0.87 (0.07) 0.73 (0.05) 0.74 (0.17) 0.94 (0.07) RF-500 0.90 (0.07) 0.77 (0.03) 0.72 (0.19) 1.00 (0.01) 0.91 (0.07) 0.73 (0.03) 0.79 (0.16) 0.97 (0.06) XGB 0.81 (0.08) 0.76 (0.07) 0.64 (0.21) 0.90 (0.09) 0.85 (0.09) 0.79 (0.09) 0.72 (0.21) 0.91 (0.08)
While RE-Lasso and RE-EN demonstrate competitive overall prediction performance, it is important to highlight several key limitations of these randomized ensemble approaches. To match the predictive performance of Split-Lasso and Split-EN, RE-Lasso and RE-EN required a substantially larger ensemble size of $G=100$ models. Furthermore, the individual models within RE-Lasso and RE-EN ensembles maintain approximately the same size as their base counterparts (Lasso and EN), containing substantially more predictors than the individual Split-Lasso and Split-EN models while achieving similar or slightly lower $\overline{\text{ACC}}$. A critical limitation of these randomized approaches is their reliance on arbitrary feature subsampling to generate model diversity, where each individual model is fitted using bootstrap samples and only 50% of predictors selected at random. This parameter setting is crucial, as using a smaller percentage would significantly degrade variable selection performance and consequently reduce individual model accuracy. As will be demonstrated in Section 3.3 and more extensively in our simulation studies in Section 5, this randomization-based approach to achieving diversity comes at a considerable cost to variable selection performance, potentially limiting the biological insights that can be derived from the models. In contrast, Split-Lasso and Split-EN achieve diversity through a principled optimization approach that directly identifies complementary gene subsets with strong predictive power.
RGLM and RF demonstrate decent ensemble prediction accuracy, particularly with larger ensemble sizes (RGLM-100 and RF-500 achieving ACC of 0.91 and 0.90 on lung data, and 0.90 and 0.91 on thyroid data), though it is evident that a substantial number of models are needed to match the performance achieved by Split-Lasso and Split-EN with merely $G=10$ models. A striking observation is the very poor individual model accuracy $\overline{\text{ACC}}$ of these methods—RF models average only 0.76-0.77 and 0.73 accuracy on lung and thyroid datasets respectively, while RGLM models achieve only 0.82 and 0.76. This poor individual performance is not surprising, as these methods rely heavily on randomization through bootstrapping and random feature selection to generate model diversity, sacrificing interpretability and individual model quality to achieve ensemble performance. The heavy reliance on randomization also negatively impacts variable selection capacity, as will be demonstrated in Section 3.3, where RF’s ability to identify important target biomarkers is substantially limited even when using unbiased importance measures. Among the remaining methods, the single-model sparse approaches (Adaptive, Relaxed, MCP, SIS-SCAD, and RuleFit) were not particularly competitive compared to our proposed methods, with accuracy typically ranging from 0.81 to 0.87. Similarly, XGB underperformed relative to our split methods, particularly at the individual model level where its $\overline{\text{ACC}}$ was notably lower (0.76 and 0.79 on lung and thyroid datasets, respectively), an expected result given that XGB models are sequentially fit to residuals rather than directly to the data, prioritizing complementarity over individual model performance.
3.3 Identification of Target Biomarkers
We examine the variable selection capabilities of the methods by analyzing their gene selection frequencies across the $N=50$ random data splits. First, we investigate the genes most frequently selected by Split-EN-10 and compare their selection patterns across other methods. This analysis reveals whether our proposed method identifies important genes that might be overlooked by conventional approaches. Second, we examine the top genes identified by RF, using the unbiased Gini importance measure proposed by Nembrini et al. (2018) to overcome known biases in standard RF variable importance metrics. Specifically, for each random split, we identify genes in the top 25% quantile of importance scores, then calculate how frequently these genes appear in this top quantile across all replications. This approach is necessary because the distribution of RF unbiased variable importance scores is highly skewed, with most scores either zero or very close to zero. The top 25% quantile selection results in sets of 125 to 139 genes with detectable variable importance signals, providing a meaningful basis for comparison with the more parsimonious models produced by regression-based methods.
Table 3 presents the top 10 genes most frequently selected by Split-EN-10 across the $N=50$ random splits, along with their selection frequencies in other methods. Similarly, Table 4 shows the top 10 genes most frequently appearing in RF’s top importance quantile, with corresponding selection frequencies across methods. From Table 3, it can be seen that there are clear patterns with certain biomarkers consistently chosen by Split-EN-10 that are overlooked by other methods. For example, PTGFRN is selected in 94% of Split-EN models but in only 38% of EN models, despite being recognized as potentially associated with the metastatic status of lung cancer (Aguila et al., 2019; Marquez et al., 2024). Even more striking is MISP, selected in 90% of Split-EN models but in only 8% of EN models and 32% of RF models, which has shown to be upregulated in certain forms of lung cancer (Zhang et al., 2025). Similarly, C11orf54 (88% in Split-EN vs. 18% in EN and 48% in RF) has been shown to promote DNA repair and has been identified as a biomarker for various cancers (Tan et al., 2023), while LACTB2 (86% in Split-EN vs. 18% in EN and 36% in RF) plays an important role in the suppression of carcinogenesis in lung cancer (Xu et al., 2022).
RE-EN with 10 models also shows moderately low selection rates for some of these genes (e.g., 49% for LACTB2), while the reverse pattern is not observed—Split-EN never falls below 84% selection frequency for the top 5 genes of RE-EN with 10 models. Moreover, for SLC16A1, which ranks 10th among RE-EN-10’s most selected genes and has been identified as a novel prognostic biomarker in NSCLC (Liu et al., 2020), Split-EN maintains a competitive 66% selection rate compared to RE-EN-10’s 74%. The consistent identification of potentially important biomarkers suggests that our proposed approach captures meaningful biological signals that traditional ensemble modeling methods may overlook—unless an impractically large number of models are used, in which case nearly all genes tend to be included such as in the case of RE-EN-100. A particularly striking example is TMC5 which has been shown to be a critical biomarker for differentiating SCC from ACC subtypes (Xiao et al., 2017). This gene appears in 86% of Split-EN models but is overlooked by standard EN (22%) and slightly less represented in RE-EN with 10 models (68%) whereas RE-EN-100 matches the selection rate but required 100 models. Surprisingly, this gene is only included about half the time as a top marker by RF (52%).
Table 3: Top 10 genes most frequently selected by Split-EN-10 across $N=50$ random splits of thyroid cancer data, and their selection frequencies in other methods.
Gene Split-EN-10 EN RE-EN-10 RE-EN-100 RF-500
CGN 1.00 0.98 1.00 1.00 0.90
PTGFRN 0.94 0.38 0.78 1.00 0.76
MISP 0.90 0.08 0.52 0.96 0.32
C11orf54 0.88 0.18 0.54 0.90 0.48
SLC6A8 0.88 0.54 0.86 0.92 0.78
LACTB2 0.86 0.18 0.46 0.84 0.36
TMC5 0.86 0.22 0.68 0.88 0.52
CSRP2 0.86 0.72 0.84 0.90 0.66
LPCAT1 0.86 0.04 0.50 0.90 0.60
CCDC68 0.84 0.36 0.66 0.86 0.34
Table 4 reveals that Split-EN-10 maintains impressively high selection frequencies for genes identified as important by RF-500, despite using substantially fewer models. Notably, Split-EN-10 selects these genes with frequencies comparable to or exceeding those of RE-EN-10, demonstrating its ability to capture diverse signals in the data without relying on randomization. The gene ranked 10th in RF’s importance measure, PRSS8, is known to potentially inhibit tumor growth in NSCLC (Ma et al., 2017), with a moderate selection rate of 64% across replications. This gene was selected in only 30% of the models by RE-EN-10, whereas Split-EN-10 maintained a stronger selection rate of 42%.
Table 4: Top 10 genes most frequently selected by RF-500 across $N=50$ random splits of thyroid cancer data, and their selection frequencies in other methods.
Gene Split-EN-10 EN RE-EN-10 RE-EN-100 RF-500
CGN 1.00 0.98 1.00 1.00 0.90
CD55 0.62 0.30 0.70 0.92 0.84
SLC6A8 0.88 0.54 0.86 0.92 0.78
TMEM125 0.74 0.42 0.76 0.98 0.78
PTGFRN 0.94 0.38 0.78 1.00 0.76
AGR2 0.76 0.30 0.62 0.84 0.68
CSRP2 0.86 0.72 0.84 0.90 0.66
SLC16A1 0.66 0.32 0.74 0.82 0.66
PRSS8 0.42 0.02 0.30 0.74 0.64
3.4 Ranking Genes
In split logistic regression genes may appear in more than one individual model, albeit at a cost in the global objective function (see Section 2.3). Hence, genes appearing in multiple individual models are expected to be more important biomarkers than others in terms of predicting the response. This information can thus be exploited to rank the genes in order of importance. The information of the individual models can be exploited to identify gene sets in order of importance using the sets defined in (11).
As an example, consider again the thyroid cancer dataset (GSE5364). With cross-validated values for $\lambda_{s}$, $\lambda_{d}$ on the full dataset, the Split-EN-10 solution for these data consists of $G=10$ sparse and diverse models. For this example, we find that $\lvert\mathcal{A}_{5}\rvert=0,\lvert\mathcal{A}_{4}\rvert=4,\lvert\mathcal{A}_{3}\rvert=3,\lvert\mathcal{A}_{2}\rvert=87$ and $\lvert\mathcal{A}_{1}\rvert=278$. Genes shared by multiple models consistently had the same sign, re-enforcing the understanding of their relationship with the thyroid cancer. This consistency in sign across models provides additional evidence for these genes’ effects on the outcome variable.
To further demonstrate Split-EN’s ability to identify clinically relevant biomarkers potentially overlooked by other methods, we examined the set $\mathcal{A}_{4}$ of our Split-EN-10 solutions across the $N=50$ random splits. This set consistently contained four genes with known biological significance in thyroid cancer. This set prominently featured TRPC1, which has been shown to play a crucial role in migration and proliferation of thyroid cancer cells (Asghar et al., 2015); APOD, consistently underexpressed in thyroid carcinomas compared to normal tissue (Huang et al., 2001); F11R, whose downregulation contributes to increased cellular motility in thyroid cancer (Czubak-Prowizor et al., 2022); and SPON2, whose elevated expression serves as a biomarker across multiple cancer types (Tang et al., 2023). Remarkably, TRPC1, F11R, and SPON2 were selected in fewer than 10% of standard EN models across the $N=50$ random splits, while APOD appeared in 46% of the models. Similarly, RF identified APOD as important in 74% of replications but rarely selected the other three genes (with F11R and SPON2 appearing in fewer than 10% of top importance sets). Only RE-EN with $G=100$ models—ten times the ensemble size of Split-EN-10—achieved comparable selection frequencies for these biologically significant markers. Conversely, examining RF’s top 10 markers by variable importance, Split-EN-10 successfully identified 7 of these markers in at least one of its component models, demonstrating substantial overlap despite methodological differences. These findings, obtained from just one specific Split-EN configuration with a modest ensemble size ($G=10$), underscore the method’s effectiveness in variable selection; further tuning could potentially enhance biomarker identification reliability even further.
4 Algorithm
The difficulty of obtaining a global minimizer of the objective function (18) is primarily due to the non-convexity of the diversity penalty $P_{d}$. Note that a global minimum of the nonnegative objective function (18) exists for any $\lambda_{s}>0$ because $\mathcal{J}\left(\beta_{0}^{1},\boldsymbol{\beta}^{1},\dots,\beta_{0}^{G},\boldsymbol{\beta}^{G}\right)\to\infty$ if $\|\boldsymbol{\beta}^{g}\|\to\infty$ for any $1\leq g\leq G$.
To construct an efficient algorithm, we observe that the objective function is multi-convex. That is, the parameters of the objective function can be partitioned in such a way that the problem is convex on each set when the others are kept fixed. A modern rigorous treatment of multi-convex programming can be found in Shen et al. (2017). In our case, the optimization problem (18) for the parameters $(\beta_{0}^{g},\boldsymbol{\beta}^{g})$ of a particular model reduces to a penalized logistic regression problem with a weighted elastic net penalty. Indeed, ignoring constant terms, the objective function for $(\beta_{0}^{g},\boldsymbol{\beta}^{g})$ reduces to
$\mathcal{J}\left(\beta_{0}^{g},\boldsymbol{\beta}^{g}\right)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\beta_{0}^{g},\boldsymbol{\beta}^{g}|y_{i},\mathbf{x}_{i})+\lambda_{s}\frac{(1-\alpha)}{2}\|\boldsymbol{\beta}^{g}\|_{2}^{2}+\sum\limits_{j=1}^{p}|\beta_{j}^{g}|u_{j,g},$
where the weights $u_{j,g}$ in the $\ell_{1}$ penalty term are given by $u_{j,g}=\alpha\lambda_{s}+\lambda_{d}/2\sum_{h\neq g}|\beta_{j}^{h}|.$ For each model the problem thus reduces to a weighted elastic net optimization where the weights in the Lasso penalty depend on the value of the coefficients in the other models. We exploit the multi-convex structure of the objective function to develop a block coordinate descent algorithm (Xu and Yin, 2013).
Recent work in non-convex optimization using block coordinate descent algorithms for applications in statistics and machine learning has been very promising, see Yang et al. (2019) for examples. The key idea is to sequentially update the current estimate for each model using a quadratic approximation $\mathcal{L}_{Q}$ for the logistic loss function in the objective function. To update the coefficients $(\beta_{0}^{g},\boldsymbol{\beta}^{g})$ of a particular model $g$ we thus need to solve
$\min_{\beta_{0}^{g}\in\mathbb{R},\,\boldsymbol{\beta}^{g}\in\mathbb{R}^{p}}\left\{\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{Q}\left(\beta_{0}^{g},\boldsymbol{\beta}^{g}\mid y_{i},\mathbf{x}_{i}\right)+\lambda_{s}P_{s}(\boldsymbol{\beta}^{g})+\frac{\lambda_{d}}{2}\sum_{\begin{subarray}{c}h=1\\ h\neq g\end{subarray}}^{G}P_{d}\left(\boldsymbol{\beta}^{h},\boldsymbol{\beta}^{g}\right)\right\}.$ (12)
Using the quadratic approximation $\mathcal{L}_{Q}$ for the logistic loss function (14) derived in the supplementary material, the update for each model is given in the proposition below.
Proposition 1
Let $(\tilde{\beta}_{0}^{1},\tilde{\boldsymbol{\beta}}^{1}),\dots,(\tilde{\beta}_{0}^{G},\tilde{\boldsymbol{\beta}}^{G})$ denote the current estimates. The coordinate descent updates for $\tilde{\beta}_{0}^{g}$ and $\tilde{\boldsymbol{\beta}}^{g}=(\tilde{\beta}_{1}^{g},\dots,\tilde{\beta}_{p}^{g})^{T}$ are given by
$\hat{\beta}_{0}^{g}$ $=\tilde{\beta}_{0}^{g}+\frac{\langle\mathbf{z}-\mathbf{\tilde{p}}^{g},\boldsymbol{1}_{n}\rangle}{\langle\mathbf{\tilde{w}}^{g},\boldsymbol{1}_{n}\rangle},$
$\hat{\beta}_{j}^{g}$ $=\frac{\text{Soft}\left(\frac{1}{n}\left({\tilde{r}_{j}^{g}}+\tilde{\beta}_{j}^{g}\langle\mathbf{x}_{j}^{2},\mathbf{\tilde{w}}^{g}\rangle\right),\,\alpha\lambda_{s}+\frac{\lambda_{d}}{2}\sum_{h\neq g}|\tilde{\beta}^{h}_{j}|\right)}{\frac{1}{n}\langle\mathbf{x}_{j}^{2},{\mathbf{\tilde{w}}}^{g}\rangle+(1-\alpha)\lambda_{s}}\quad j=1,\dots,p,$
where $\text{Soft}(\mu,\gamma)=\text{sign}(\mu)\times\text{max}(\lvert\mu\rvert-\gamma)$, $\boldsymbol{1}_{n}=(1,\dots,1)^{T}\in\mathbb{R}^{n}$ and $\tilde{r}_{j}^{g}=\left\langle\mathbf{x}_{j},{\mathbf{z}}\right\rangle-\langle\mathbf{x}_{j},\mathbf{\tilde{p}}^{g}\rangle$. The elements of the $n$-dimensional vectors $\mathbf{z}$, $\mathbf{\tilde{p}}^{g}$ and $\mathbf{\tilde{w}}^{g}$ are given by $z_{i}=(y_{i}+1)/2$, $\tilde{p}_{i}^{g}=S(\tilde{\beta_{0}}^{g}+\mathbf{x}_{i}^{T}\boldsymbol{\tilde{\beta}}^{g})$ and $\tilde{w}_{i}^{g}=\tilde{p}_{i}^{g}(1-\tilde{p}_{i}^{g}),1\leq i\leq n$, respectively.
The algorithm cycles through the components of $\left(\beta_{0}^{1},\boldsymbol{\beta}^{1}\right)$ by applying a single coordinate descent update to each parameter, then through those of $\left(\beta_{0}^{2},\boldsymbol{\beta}^{2}\right)$, and so on until we reach $\left(\beta_{0}^{G},\boldsymbol{\beta}^{G}\right)$. Then, we check for convergence. Convergence is declared when successive estimates of the coefficients in the ensemble model show little difference, i.e. $\max\limits_{0\leq j\leq p}|\tilde{\beta}_{j}-\hat{\beta}_{j}|^{2}<\delta,$ for some small tolerance level $\delta>0$, with $\tilde{\beta}_{j}=\sum_{g=1}^{G}\tilde{\beta}_{j}^{g}/G$ and $\hat{\beta}_{j}=\sum_{g=1}^{G}\hat{\beta}_{j}^{g}/G$ the respective estimates for the ensemble model. The algorithm converges to a coordinatewise minimizer of (18) by Theorem 4.1 of Tseng (2001). More details of the algorithm are given in the supplementary material.
To select the tuning parameters we alternate between a grid search for the sparsity penalty and a grid search for the diversity penalty, such that the cross-validated loss of the ensemble classifier is minimized. The details are available in the supplementary material. By default, $10$-fold CV is used. Note that the value $\lambda_{d}=0$ is included in the grid search for the diversity penalty, such that the (single model) elastic net is a possible solution of split logistic regression. The warm-start and active-set cycling strategies proposed by Friedman et al. (2010) are well suited for our algorithm, and have been incorporated to speed up the algorithm. Our choice for the ensembling function (10) also allows the construction of coefficient solution paths for the ensembled model which is illustrated in the supplementary material.
5 Simulation Study
Following the ADEMP framework Morris et al. (2019), we structure our simulation study to systematically evaluate our proposed method against established alternatives.
5.1 Aims
The primary aims of this simulation study are to:
  • Assess the predictive performance of split logistic regression compared to established sparse and ensemble methods under varying conditions.
  • Evaluate variable selection accuracy of our proposed approach.
  • Investigate how different correlation structures among predictors affect method performance.
  • Examine the impact of sample size, event probability, sparsity level, non-linear effects, and interaction effects on performance outcomes.
5.2 Data-Generating Mechanisms
We investigate five simulation scenarios based on the logistic regression framework as given in Equation (1). Across scenarios, we systematically vary several aspects of the data-generating process: correlation structure among predictors, correlation levels, sparsity level, sample size, class imbalance, and functional form (linear, interaction effects, and non-linear effects).
Base Configuration for All Scenarios:
  • Active coefficient values are randomly generated as $(-1)^{z}u$ where $z\sim\text{Bernoulli}(0.3)$ and $u\sim\text{Uniform}(0,1/2)$
  • Predictors follow multivariate normal distribution with mean zero and unit variance
  • Dimension $p=$ 1,000 with sparsity levels $\zeta\in\{0.1,0.2,0.4\}$ (proportion of active variables)
  • Training sample sizes $n\in\{50,100\}$ with event probability $\mathbb{P}(Y=1)\in\{0.2,0.3,0.4\}$
  • Test sample size $m=$ 5,000 for performance evaluation
  • $N=50$ Monte Carlo replications per configuration
To systematically assess method performance, we consider a range of data-generating mechanisms that span common scenarios encountered in high-dimensional biomedical data. Scenarios 1–3 are designed to explore the effects of different predictor covariance structures—namely, exchangeable correlation (Scenario 1), distinct correlation levels between active and inactive predictors (Scenario 2), and a block-wise (modular) correlation structure often observed in gene expression or genomics data (Scenario 3). To further challenge the methods, Scenario 4 adopts this realistic block-wise structure and incorporates within-block interaction effects, while Scenario 5 similarly uses the block correlation structure but introduces nonlinear (quadratic) effects for a subset of predictors.
Scenario 1: All predictors have equal pairwise correlation. Data are generated according to:
$\log\left(\frac{p_{i}}{1-p_{i}}\right)=\beta_{0}+\mathbf{x}_{A,i}^{T}\,\boldsymbol{\beta}_{A},\quad 1\leq i\leq n,$
where $\mathbf{x}_{A,i}^{T}$ are the active predictors and $\boldsymbol{\beta}_{A}$ the corresponding regression coefficients. All predictors are correlated with each other, with correlation $\rho\in\{0.2,0.5,0.8\}$.
Scenario 2: Differential correlation between active and inactive predictors. Data follow the same logistic model with correlation $\rho_{1}$ between active and inactive predictors and $\rho_{2}$ for all other correlations, where $\rho_{1}\in\{0,0.2,0.5\}$ and $\rho_{2}\in\{0.2,0.5,0.8\}$ such that $\rho_{1}<\rho_{2}$.
Scenario 3: Block structure among active predictors. Data are generated according to:
$\log\left(\frac{p_{i}}{1-p_{i}}\right)=\beta_{0}+\sum_{b=1}^{B}\mathbf{x}_{b,i}^{T}\,\boldsymbol{\beta}_{b},\quad 1\leq i\leq n,$
where $\mathbf{x}_{b,i}$ are the predictor variables for block $b$ and $\boldsymbol{\beta}_{b}$ are the corresponding regression coefficients. Each block contains 25 predictors, with $B=\zeta p/25$ blocks. Correlation between predictors in different blocks is $\rho_{1}\in\{0.2,0.5\}$, while correlation between predictors in the same block is $\rho_{2}\in\{0.5,0.8\}$, such that $\rho_{1}<\rho_{2}$.
Scenario 4: Block structure with interaction effects. Data are generated from:
$\log\left(\frac{p_{i}}{1-p_{i}}\right)=\beta_{0}+\sum_{b=1}^{B}\mathbf{x}_{b,i}^{T}\,\boldsymbol{\beta}_{b}+\sum_{j,k\in\mathcal{I}}\gamma_{jk}x_{ij}x_{ik},\quad 1\leq i\leq n,$
where $\mathcal{I}$ represents a subset of pairs of active predictors selected for interactions. Specifically, we randomly select $\lfloor p\zeta/10\rfloor$ pairs of active predictors to form interactions. The interaction coefficients $\gamma_{jk}$ are generated using the same mechanism as the main effects: $\gamma_{jk}=(-1)^{z}u$ where $z\sim\text{Bernoulli}(0.3)$ and $u\sim\text{Uniform}(0,1/4)$. The correlation structure remains identical to Scenario 3, with blocks of 25 predictors, between-block correlation $\rho_{1}\in\{0.2,0.5\}$, and within-block correlation $\rho_{2}\in\{0.5,0.8\}$.
Scenario 5: Block structure with non-linear effects. Data are generated from:
$\log\left(\frac{p_{i}}{1-p_{i}}\right)=\beta_{0}+\sum_{b=1}^{B}\mathbf{x}_{b,i}^{T}\,\boldsymbol{\beta}_{b}+\sum_{j\in\mathcal{N}}\delta_{j}x_{ij}^{2},\quad 1\leq i\leq n,$
where $\mathcal{N}$ represents a subset of active predictors selected to have additional non-linear effects. We select $\lfloor p\zeta/5\rfloor$ active predictors to have non-linear components. The non-linear coefficients $\delta_{j}$ are generated as $\delta_{j}=(-1)^{z}u$ where $z\sim\text{Bernoulli}(0.3)$ and $u\sim\text{Uniform}(0,1/4)$. The correlation structure remains identical to Scenario 3, with the same blocking and correlation parameters.
5.3 Estimands
The primary estimands of interest are:
  • The true coefficient vector $\boldsymbol{\beta}$ (for variable selection performance)
  • The true conditional class probabilities $p_{i}=P(Y_{i}=1|X_{i}=x_{i})$ (for prediction performance)
5.4 Methods
We compare the performance of split logistic regression to nine established methods as detailed in Section 3 and Table 1. For split logistic regression, we use $G=10$ models. For RGLM and RF we use their R package default configurations with $G=100$ and $G=500$ models respectively. For random forest variable importance, we implement the unbiased Gini importance measure (Nembrini et al., 2018) to address known biases in the standard Gini importance when predictors are correlated.
5.5 Performance measures
We evaluate methods using the following performance metrics:
Prediction performance metrics:
  • Accuracy (ACC): Proportion of correctly classified observations
  • Area Under the ROC Curve (AUC): Measure of discriminative ability across thresholds
  • Sensitivity (SNS): Proportion of true positive cases correctly classified
  • Specificity (SPC): Proportion of true negative cases correctly classified
  • Test-sample loss (TSL): Average negative log-likelihood (14) on test data
For all classification methods, we use a probability threshold of 0.5 to determine class assignment: observations with predicted probabilities above 0.5 are classified as class 1, and those below as class 0. While this standard threshold enables fair comparison across methods, we note that in practical applications, especially in medical contexts, this threshold could be optimized, e.g. based on the relative costs of false positives versus false negatives. The AUC metric provides a threshold-independent assessment of discriminative performance.
Variable selection performance metrics:
  • Recall (RCL): Proportion of truly active variables identified by the model
    $\text{RCL}=\frac{\sum_{j=1}^{p}\mathbb{I}(\beta_{j}\neq 0,\hat{\beta}_{j}\neq 0)}{\sum_{j=1}^{p}\mathbb{I}(\beta_{j}\neq 0)}$
  • Precision (PRC): Proportion of selected variables that are truly active
    $\text{PRC}=\frac{\sum_{j=1}^{p}\mathbb{I}(\beta_{j}\neq 0,\hat{\beta}_{j}\neq 0)}{\sum_{j=1}^{p}\mathbb{I}(\hat{\beta}_{j}\neq 0)}$
Since RF with its default number of models tends to use all predictors, and XGB builds sequential models on residuals making feature selection interpretation challenging, we do not compute RCL and PRC for these methods.
5.6 Results
Table 5.6 presents the average ranks of the prediction performance metrics across all simulation scenarios. The split-ensemble approaches, particularly Split-EN-10 and Split-Lasso-10, demonstrate remarkable predictive performance across all scenarios. Despite using only 10 base models, these methods consistently outperform ensemble approaches that employ 100 models or more (RE-Lasso-100, RE-EN-100, RGLM-100, and RF-500). Split-EN-10 achieves the best overall performance, achieving the best average rank in ACC, SNS, AUC and TSL metrics across nearly all scenarios, making it particularly valuable for class-imbalanced datasets where both overall prediction accuracy and detection of the minority class cases are critical.
A particularly valuable finding is the consistency of the split-ensemble methods across different data structures. Whether dealing with simple main effects, complex interactions, or non-linear relationships, both Split-EN-10 and Split-Lasso-10 maintain their superior performance. This demonstrates that these methods adapt effectively to various data complexities without requiring scenario-specific adjustments. Importantly, this excellent performance is achieved with just 10 models and without relying on randomization techniques employed by RE-Lasso, RE-EN, RGLM, and RF methods. Instead, split-ensemble methods use a well-formulated objective function that balances model fit with variable selection and diversity in a mathematically sound framework. While RF achieved the highest specificity across all scenarios, this came at a significant cost in terms of sensitivity. In contrast, the split-ensemble methods maintained competitive specificity rankings while excelling in sensitivity measures, offering a more balanced classification approach.
Figure 1 illustrates SNS and SPC performance across 50 random training sets under challenging high-dimensional conditions with class imbalance. The split-ensemble methods demonstrate an impressive balance of both metrics, maintaining high sensitivity while preserving competitive specificity. Despite using only 10 base models, Split-Lasso-10 and Split-EN-10 consistently outperform more complex methods. Notably, the closest competitors, RE-Lasso-100 and RE-EN-100 with 100 base models each, show lower sensitivity distributions with their 75th percentiles roughly corresponding to the split methods’ median performance. While RF and Adaptive Lasso achieve high specificity, they sacrifice sensitivity considerably. The split-ensemble approaches offer the best overall classification balance in this challenging scenario.
Table 5: Average prediction performance ranks of the methods across all simulation configurations for main effects, interactions, and non-linear scenarios. The three best results for each criterion are highlighted in bold.
[…]
=2pt {tabu}l lllll lllll lllll Main Effects Interactions Non-Linear
Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 2.68 2.37 5.65 3.11 2.00 2.02 2.09 5.44 2.19 1.94 2.37 2.33 5.75 2.48 2.46
Split-EN-10 2.06 2.14 4.84 2.06 1.60 1.72 2.00 5.26 1.59 1.67 1.81 1.96 5.10 1.43 1.72
Lasso 7.68 8.04 9.07 8.14 6.94 7.39 7.74 9.23 7.96 6.91 7.28 7.55 9.35 8.06 7.06
EN 6.32 5.91 7.53 6.99 5.52 5.93 5.52 7.76 6.89 5.48 5.87 5.54 7.89 7.00 5.52
Adaptive 10.91 12.47 3.56 9.88 9.76 11.52 13.09 3.26 10.52 10.07 11.41 13.20 2.81 11.15 9.98
Relaxed 8.92 7.21 11.14 9.48 11.18 8.57 6.81 11.13 9.04 10.44 8.35 6.57 11.56 9.00 10.74
MCP 12.48 11.76 12.23 12.58 11.24 11.85 11.15 11.78 12.11 11.09 11.98 11.20 11.37 12.31 11.26
SIS-SCAD 13.28 12.81 11.21 12.51 12.06 13.37 12.35 11.76 12.76 12.46 13.65 12.44 11.22 12.54 12.50
RuleFit 11.54 10.74 12.09 12.90 13.48 11.72 10.43 12.13 12.78 13.69 11.85 9.94 12.48 12.81 13.85
RE-Lasso-100 4.28 4.40 4.66 4.77 3.81 4.16 4.30 4.67 4.67 3.87 3.72 3.83 5.22 4.50 3.52
RE-EN-100 3.35 3.46 4.94 3.81 3.41 3.38 3.20 5.63 3.91 3.56 2.89 2.91 6.26 3.44 3.02
RGLM-100 3.67 4.35 3.69 4.15 4.99 3.98 5.13 3.73 4.31 4.61 4.54 6.05 3.20 4.72 4.89
RF-500 7.09 9.67 2.67 3.15 9.04 8.35 11.54 1.26 4.70 8.98 8.50 11.63 1.19 4.43 8.44
XGB 10.76 9.66 11.72 11.47 9.96 11.04 9.65 11.96 11.57 10.22 10.78 9.83 11.59 11.13 10.04
[…]
Table 6: Average variable selection performance ranks of the methods across all simulation configurations for main effects, interactions, and non-linear scenarios. The three best results for each criterion are highlighted in bold.
[…]
=2pt {tabu}l ll ll ll Main Effects Interactions Non-Linear
Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 4.19 6.36 4.26 6.48 4.11 5.85
Split-EN-10 2.67 7.23 2.74 7.50 2.74 7.11
Lasso 8.11 6.64 8.11 7.26 8.11 7.72
EN 6.69 6.12 6.93 5.39 6.93 5.61
Adaptive 9.00 7.30 9.16 7.91 9.37 8.28
Relaxed 9.81 4.90 9.73 5.35 9.52 6.06
MCP 11.42 4.20 11.33 3.97 11.40 3.81
SIS-SCAD 11.58 3.12 11.67 1.53 11.60 1.57
RuleFit 6.39 3.30 6.07 3.61 6.07 2.98
RE-Lasso-100 2.48 10.50 2.39 10.72 2.35 10.81
RE-EN-100 1.01 11.23 1.00 11.09 1.00 11.11
RGLM-100 4.65 7.09 4.61 7.19 4.80 7.07
RF-500 -- -- -- -- -- --
XGB -- -- -- -- -- --
Table 5.6 presents the average ranks of the variable selection performance metrics across all simulation scenarios. The results reveal a clear trade-off between RCL and PRC among the methods. While RE-EN-100 and RE-Lasso-100 achieve the highest recall, they do so at a significant expense of precision, ranking near the bottom (10-11th place) for PRC across all scenarios. This suggests that these methods tend toward excessive variable inclusion when using 100 base models. In contrast, Split-EN-10 demonstrates an impressive balance, ranking among the top three methods for recall while maintaining mid-range precision performance, all with just 10 models. SIS-SCAD, MCP, and RuleFit excel in precision but perform poorly in recall, indicating overly conservative variable selection. Split-Lasso-10 shows competitive recall comparable to traditional sparse methods while offering better precision than randomization-based ensemble methods that use 10 times more base models. This indicates that Split-Lasso-10 is an effective approach for both predictive and interpretable modeling across various data complexity scenarios.
Figure 2 further illustrates the variable selection performance trade-offs under challenging high-dimensional conditions. The plot confirms that RE-Lasso-100 and RE-EN-100 achieve high recall but at a severe cost to precision, indicating they include numerous irrelevant variables. In contrast, Split-EN-10 and Split-Lasso-10 demonstrate a more balanced profile, maintaining competitive recall with substantially better precision than the randomization-based ensembles, despite using only 10 models. The split-ensemble methods’ favorable position in this recall-precision trade-off space highlights their ability to balance identifying true signals while limiting false discoveries, making them particularly valuable when both accurate prediction and meaningful variable selection are required.
[…]
6 Medical Genomics Data Applications
To demonstrate the competitive performance of split logistic regression in practice, we apply split logistic regression and its competitors in Section 3 to ten medical genomics datasets available from the Gene Expression Omnibus (GEO) database. Some characteristics of the datasets are given in Table 6.
Table 7: GEO identification (ID) codes, dataset descriptions, sample sizes ($n$), class distributions, pairing status, and the number of genes retained after matching probes to genes.
[…]
=2pt {tabu}llrrrr GEO ID Description $\boldsymbol{n}$ Class Distribution Paired Genes
GSE20347 Esophageal cancerous cell tissue 34 17 cancer, 17 normal Yes 13,515
GSE23400 Esophageal cancerous cell tissue 106 53 cancer, 53 normal Yes 13,515
GSE23400 Esophageal cancerous cell tissue 102 51 cancer, 51 normal Yes 11,271
GSE5364 Esophageal cancerous cell tissue 29 16 cancer, 13 normal Yes 13,515
GSE25869 Gastric cancerous cell tissue 64 32 cancer, 32 normal Yes 14,476
GSE5364 Lung cancerous cell tissue 30 16 cancer, 14 normal Yes 13,515
GSE10245 Lung cancerous cell tissue 58 40 adeno, 18 squamous No 13,515
GSE5364 Thyroid cancerous cell tissue 51 16 cancer, 35 normal Yes 13,515
GSE21942 Multiple sclerosis cell tissue 29 14 sclerosis, 15 normal No 23,520
GSE14905 Psoriasis cell tissue 54 34 psoriasis, 20 normal No 23,520
The first eight datasets involve the classification of different types of cancerous cell tissues, whereas the last two datasets involve the separation of psoriasis and multiple sclerosis cell tissues from adjacent normal cell tissue. Our benchmark includes both paired samples, where tissue from the same patient contributes to both disease and control groups, and unpaired samples from different individuals. While most classification methods are designed for independent samples, including paired samples in our benchmark is valuable since such designs are common in real-world biomedical applications. This mixture enables evaluation of how classifiers perform across varied experimental designs encountered in practice. All datasets are first pre-processed as described in Section 3 using the choices $p=100,250,500$, and $1,000$, respectively, for the number of retained genes according to their $q$-value.
Each dataset is randomly split $N=50$ times into a training set and test set. For the proportion of training data we considered both 35% and 50%, where proportion 35% is only used if the resulting training set contains at least 20 observations. Each method is applied to the training data and evaluated on the test set using both ACC and TSL. For the split logistic regression methods we use cross-validation (CV) to determine the optimal number of models, with candidates $G=5$, $10$, and $25$. For $p\in\{100,250,500,1{,}000\}$, Table 8 shows the average ranks across all datasets and training proportions. The results reveal that Split-EN-CV consistently ranks among the top three methods across all feature dimensions, particularly impressive when $p\geq 500$ where it either matches or surpasses RE-EN-100 in ACC while using only 5–25 interpretable sparse models compared to RE-EN’s 100 models. In terms of overall performance, Split-EN-CV achieves the second-best average rank (2.96) for ACC, behind only RE-EN-100 (2.75) despite using far fewer models. Detailed results are available in the supplementary material. We note that some cancer datasets may contain paired samples from the same patients, potentially introducing correlation not explicitly modeled in our analytical approach.
Table 8: Average ranks for MR and TL over the ten gene expression datasets in Table 6 for different number of genes retained after pre-processing. The three best results for each column are highlighted in bold.
[…]
For each dataset we computed the average ranks of the methods based on the results for each training proportion and $p\in\{100,250,500,1,000\}$. Table 9 summarizes the results by showing for each method the number of times it belonged to the top one, top three and lowest three average ranks among the ten datasets. Split-EN-CV demonstrates remarkable performance, achieving the top rank in ACC for half of all datasets and in TSL for 3 datasets, while Split-Lasso-CV consistently appears in the top three for 7 datasets in ACC and 5 in TSL. Notably, neither split method ever falls into the worst three performing methods for either metric, highlighting their exceptional stability across diverse genomic datasets. While RE-EN-100 shows decent performance by appearing in the top three for several datasets, it achieves the top ACC rank in none of them despite utilizing 100 models compared to only 10 for our Split methods. Similarly, RE-Lasso-100 requires ten times the number of models to occasionally match the performance of Split methods. Traditional sparse methods generally underperform, with MCP ranking among the bottom three for 6 datasets in ACC, while more complex ensemble approaches like RF-500 and XGB show inconsistent performance, frequently appearing among the worst performers.
Table 9: Number of top and lowest ranks for MR and TL over the ten gene expression datasets in Table 6.
[…]
=2pt Top 1 Top 3 Low 3 Method ACC TSL ACC TSL ACC TSL Split-Lasso-CV 0 0 7 5 0 0 Split-EN-CV 5 3 6 6 0 0 Lasso 0 0 0 0 0 1 EN 1 1 2 2 0 0 Adaptive 0 0 1 0 2 0 Relaxed 0 0 0 0 0 10 MCP 0 0 0 0 6 1 SIS-SCAD 0 0 0 0 4 0 RuleFit 0 0 0 0 6 10 RE-Lasso-100 2 1 4 4 0 0 RE-EN-100 0 4 8 9 0 0 RGLM-100 0 1 0 2 0 0 RF-500 2 0 2 2 5 4 XGB 0 0 0 0 7 4
Figure 3 presents a comparative assessment of predictive performance using scaled test sample loss (TSL), which is essentially the deviance on test data, for the GSE10245 lung cancer dataset (differentiating between adenocarcinoma and squamous cell carcinoma). This dataset presents a particularly challenging case due to its class imbalance (SCC comprising only 18 samples and being rarer), making it more complex than many of the balanced datasets evaluated in Table 9. For each replication, we computed the scaled TSL as:
Scaled $\text{Scaled TSL}=\frac{\min(\text{TSL})}{\text{TSL}}$
where $\min(\text{TSL})$ represents the minimum test sample loss achieved by any method on that specific random training-test split. This scaling results in values between 0 and 1, where higher values represent superior performance. The Split-Lasso and Split-EN methods demonstrate remarkable efficiency, consistently achieving optimal or near-optimal performance despite leveraging only $G=10$ models. Their distributions are tightly concentrated near 1, indicating these methods either achieved the lowest TSL or came remarkably close across replications. In contrast, the ensembles based on randomization exhibited greater variability and often produced less reliable probability estimates, as evidenced by their substantially lower scaled TSL values. Split-logistic regression’s ability to maintain high-quality probability calibration with minimal ensemble complexity represents a significant advantage for implementation in clinical contexts, particularly for imbalanced datasets like this one where accurate probability estimation is crucial.
Figure 3: Comparison of model performance using scaled TSL (minimum TSL divided by each method’s TSL). Higher values indicate better performance, with 1.0 representing best probability calibration.
7 The Number of Models
Constructing accurate and diverse models for an ensemble are opposite objectives (Krogh and Vedelsby, 1995). We perform an empirical study to explore this accuracy-diversity trade-off for split logistic regression, which will drive the choice for the number of models in real medical genomics data applications.
The analysis of genomics data via high-throughput technologies has generated the need for classification algorithms that can handle high-dimensional data containing correlated predictors (genes) within different pathways or networks, see Yousefi et al. (2011) and Zhang and Coombes (2012) for example. In light of this, to investigate the accuracy-diversity trade-off of split logistic regression, we use the high-dimensional block correlation setting of Scenario 3 in Section 5 with configuration parameters $(n,p)=(50,\,$1,000), $(\rho_{1},\rho_{2})=(0.2,0.5)$, $\zeta\in\{0.1,0.2,0.4\}$ and $\mathbb{P}(Y=1)=0.2$.
To quantify diversity for ensemble classifiers, we adopt the entropy diversity measure of Kuncheva and Whitaker (2003). Given an ensemble comprised of $G$ individual classifiers, the entropy measure (EM) for a given $\mathbf{x}$ is defined as
$\text{EM}(\mathbf{x})=\frac{1}{G-\lceil G/2\rceil}\,\min\left(\ell(\mathbf{x}),G-\ell(\mathbf{x})\right),$ (13)
where $\ell(\mathbf{x})$ denotes the numbers of individual classifiers in the ensemble that correctly classify $\mathbf{x}$. The entropy measure ranges between 0 and 1, with $\text{EM}(\mathbf{x})=0$ and $\text{EM}(\mathbf{x})=1$ corresponding to no diversity and to the highest possible diversity between the individual classifiers, respectively. The overall entropy $\mathbb{E}[\text{EM}(\mathbf{x})]$ can be estimated by averaging $\text{EM}(\mathbf{x})$ over a test set.
7.1 Results
Table 10 shows the evolution of the ensemble prediction accuracy (ACC), the average prediction accuracy of the individual models ($\overline{\text{ACC}}$) and the entropy diversity measure (EM) averaged over the test sets, as a function of the number of models for split logistic regression. It also contains the overlap (OVP) between the individual models in the ensemble, defined as
$\text{OVP}=\frac{\sum\limits_{j=1}^{p}o_{j}\mathbb{I}\{\ o_{j}\neq 0\}}{\sum\limits_{j=1}^{p}\mathbb{I}\{o_{j}\neq 0\}},\quad o_{j}=\frac{1}{G}\sum_{g=1}^{G}\mathbb{I}\{\hat{\beta}_{j}^{g}\neq 0\}.$
It can be seen that in all three settings $\overline{\text{ACC}}$ decreases and EM increases with the number of models, while the ensemble ACC increases. Hence, as the number of models increases, the accuracy of the individual models has less impact on the ensemble ACC compared to their level of diversity. Split logistic regression manages to achieve a proper balance for this trade-off, resulting in a high ACC for the ensemble.
[…]
=2pt $\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$ $\mathbf{G}$ ACC $\mathbf{\overline{ACC}}$ EM OVP ACC $\mathbf{\overline{ACC}}$ EM OVP ACC $\mathbf{\overline{ACC}}$ EM OVP 1 0.86 -- -- -- 0.87 -- -- -- 0.87 -- -- -- 2 0.87 0.86 0.04 0.81 0.89 0.88 0.05 0.76 0.88 0.88 0.05 0.73 5 0.87 0.86 0.10 0.53 0.89 0.88 0.12 0.44 0.89 0.88 0.11 0.42 7 0.87 0.86 0.10 0.47 0.90 0.88 0.14 0.32 0.89 0.88 0.11 0.39 10 0.87 0.85 0.13 0.29 0.90 0.87 0.16 0.17 0.90 0.87 0.14 0.19 15 0.87 0.84 0.18 0.20 0.90 0.86 0.20 0.12 0.90 0.87 0.18 0.07 20 0.87 0.84 0.19 0.12 0.90 0.86 0.21 0.08 0.90 0.86 0.19 0.05 25 0.87 0.83 0.20 0.13 0.90 0.85 0.24 0.04 0.90 0.86 0.21 0.04
When $G$ is small, the average ${\overline{\text{MR}}}$ of the individual models in Table 10 is close to the MR of the logistic elastic net (the case $G=1$ in Table 10). The choice of diversity tuning parameter $\lambda_{d}$ is driven by the data based on a CV criterion. For a small number of models, a smaller value of $\lambda_{d}$ is selected such that the OVP of the models is large and they share a lot of important predictors. This results in accurate individual models but a relatively low diversity as seen from the EM values. As the number of models increases, the OVP becomes smaller, resulting in individual models that have a higher average EM. Indeed, for a large number of models it becomes beneficial to increase diversity between the models to decrease the misclassification rate of the ensemble. In this case the diversity penalty in split logistic regression reduces the overlap between individual models, leading to high diversity which results in high classification accuracy. In summary, split logistic regression thus successfully achieves the proper balance between individual model accuracy and diversity regardless of the number of models. Alternative diversity measures are considered in the supplementary material and lead to the same conclusions.
7.2 Computational Cost
Table 10 indicates that a larger number of models results in an ensemble with higher prediction accuracy. However, ACC stabilizes quickly, so there is a diminishing returns type of behavior in terms of prediction accuracy versus computational cost. Indeed, we also ran split logistic regression using $G=50$ models, but in all cases there is hardly any improvement in ACC compared to the ensemble with $G=25$ models shown in Table 10. In fact, with $G=25$ models split logistic regression already achieves nearly full diversity (OV $\approx 1/G$), so little gain is expected by increasing the number of models further while computation time does grow.
Table 7.2 shows the average computation time (in CPU seconds) across all sparsity levels of scenario 3 as a function of the number of models. As $G$ increases there is a price to pay in average individual ${\overline{\text{ACC}}}$ which may not be compensated by a higher ensemble ACC. The computation time seems to depend linearly on the number of models and is approximately given by $0.27+0.15\times G$ for $G\geq 2$. In real data applications such as the gene expression applications in the previous section it is generally a good strategy to use a data-driven choice of the number of models in the ensemble by increasing $G$ until the CV performance has stabilized.
Table 11: Computation time of R function call for split logistic regression in CPU seconds for varying number of models, using multithreading (5 threads). CPU seconds are on a 2.1 GHz Intel Xeon Platinum 8468 processor in a machine running CentOS Linux 7.9 with 32 GB of RAM.
[…]
=2pt {tabu}lrrrrrrr $\mathbf{G}$ 2 5 7 10 15 20 25
Time 0.55 1.05 1.39 1.76 2.53 3.40 4.06
8 Discussion and Future Directions
We presented a new approach to learn a diverse ensemble of sparse logistic regression models that is well suited for high-dimensional medical genomics data. The individual models for the ensemble are learned simultaneously by optimizing an objective function which balances between individual model strength and diversity between the models. The sparsity penalty in the objective function controls the stability of the individual models while the diversity penalty favorably exploits the accuracy-diversity trade-off to achieve excellent performance for the resulting ensemble. In contrast to other popular ensemble methods, split logistic regression models remain logistic regression models and thus are highly interpretable. Moreover, the individual models in the ensemble may be of interest in their own right because they each provide a relationship between the predictor genes and disease status that can provide insight to very complex biological mechanisms.
On several medical genomics datasets from the GEO database split logistic regression achieved state-of-the-art prediction accuracy. In several examples split logistic regression also identified unique key biomarkers that were not picked up by other methods, while at the same time selecting key biomarkers that were consistently picked up by other methods. A variable ranking method for split logistic regression was also developed to help guide researchers in which genes to investigate in the hope of developing effective drug therapies.
While this manuscript focuses on microarray gene expression data, which has been a canonical example of the $p\gg n$ scenario in statistical genomics for decades, the proposed split logistic regression approach is equally applicable to other types of high-dimensional omics data. Modern technologies like RNA-seq, single-cell sequencing, proteomics, and metabolomics all generate data with similar dimensional characteristics. For instance, single-cell RNA sequencing data can be transformed into pseudobulk profiles representing cell populations, methylation data can be summarized at the gene or region level, and proteomics datasets frequently contain thousands of measured proteins across limited samples. The common thread across these data types is that they all present the fundamental statistical challenge of extracting meaningful signals from thousands of features measured on relatively few subjects—precisely the scenario where our proposed method demonstrates advantages over traditional approaches.
Due to the diversity penalty, split logistic regression makes use of different groups of variables in the individual models to build an ensemble. Allowing interactions among predictors can be beneficial to further improve the prediction performance of classifiers. Since split logistic regression can have much higher recall than single-model methods such as Lasso and the elastic net, the proposed methodology can also be useful to detect interaction effects that would be missed by such single-model methods. This is important for example for genomics data applications where it is known that gene interaction effects are common. The diversity penalty can also be combined more generally with different sparsity penalties such as the group Lasso (Meier et al., 2008) for categorical variables or the fused Lasso (Tibshirani et al., 2005) for data exhibiting spatial or temporal structures.
Block coordinate descent is an effective approach to solve the multi-convex optimization problem underlying split logistic regression. Multi-convex programming is an emerging field in optimization with many applications in statistics and machine learning, see e.g. Shen et al. (2017) and Pardalos et al. (2017). In future research we will investigate whether alternative approaches can further decrease the computational cost of the method.
In split logistic regression the models are ensembled at the level of the linear predictors. This guarantees high interpretability of the ensemble model, but is not necessarily optimal from a prediction point of view. In future research it will be examined whether alternative ensembling functions can improve on the prediction accuracy of the ensemble.
Ensemble methods are very popular to analyze small sample data with a large number of predictor variables, and the proposed method provides a framework to build an optimal classification ensemble model. Similarly to logistic regression, the general split modeling framework could be applied to multi-class classification problems to obtain a powerful ensemble classifier. The split modeling framework could also be extended to generalized linear models in general.
Acknowledgments
Part of this work was conducted while Anthony-Alexander Christidis was a UBC Doctoral Researcher at KU Leuven’s Department of Mathematics under a Mitacs Globalink Research Award.
Supplementary Material
The supplementary material contains the consistency proof, details of our algorithm, the analysis with the alternative diversity measures and the full results of our simulation and medical genomics data experiments. The R/C++ package SplitGLM along with its reference manual is publicly available on CRAN at https://CRAN.R-project.org/package=SplitGLM. The data and scripts to replicate the numerical experiments are available at https://doi.org/10.5281/zenodo.15588653.
Appendix A: Ensemble Asymptotics
In this section, we prove a general result for the asymptotic behavior of the prediction error of the ensemble split regression method and show that it implies consistency of the prediction under the assumptions stated in Theorem 1 of the article.
8.1 Preliminaries
Consider data $\{y_{i},\mathbf{x}_{i}\}_{i=1}^{n}$ where $\mathbf{x}_{i}\in\mathbb{R}^{p}$ and $y_{i}\in\{-1,1\}$ for $i=1,2,\dots,n$. Without loss of generality, we assume each column of the design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ has been scaled by its maximum value such that $\underset{1\leq j\leq p}{\max}\lVert\mathbf{x}_{\cdot j}\rVert_{\infty}\leq 1$ where $\mathbf{x}_{\cdot j}$ is the $j$-th column of $\mathbf{X}$. Let $\mathcal{H}$ be a (rich) parameter space that includes the of space linear functions, and for each $f\in\mathcal{H}$ we take the (convex) logistic loss function :×$\mathcal{L}:\mathcal{H}\times\mathbb{R}^{p}\times\{-1,1\}\mapsto\mathbb{R}$ as defined in the main article,
$\mathcal{L}(f(\mathbf{x}_{i}),y_{i})=\log(1+e^{-y_{i}f(\mathbf{x}_{i})}).$ (14)
We denote the empirical risk by
$\mathcal{V}_{n}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f(\mathbf{x}_{i}),y_{i}),$
such that $\mathcal{V}_{n}$ is the empirical measure that puts mass $1/n$ for each observation $(y_{i},\mathbf{x}_{i})$, and the expected risk by
$\mathcal{V}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[\mathcal{L}(f(\mathbf{x}_{i}),y_{i})\right]$
We also denote the target function
$f^{*}=\operatorname*{arg\,min}_{f\in\mathcal{H}}\mathcal{V}(f)$
as minimizer of the expected risk. For any $f\in\mathcal{H}$, the excess risk is given by
$\mathcal{E}(f)=\mathcal{V}(f)-\mathcal{V}(f^{*}),$
where by definition $\mathcal{E}(f)\geq 0$ for all $f\in\mathcal{H}$. In the case of model misspecification where the target function $f^{*}$ is not necessarily linear, we define the linear subspace $\mathcal{H}_{\boldsymbol{\beta}}=\{f_{\boldsymbol{\beta}}:\boldsymbol{\beta}\in\mathbb{R}^{p+1}\}\subset\mathcal{H}$, where the map $\boldsymbol{\beta}\mapsto f_{\boldsymbol{\beta}}$ is linear. As in Section 6.6 of Bühlmann and van de Geer (2011) we consider the notationally simpler case without intercept. We denote the best linear approximation of the target function $f^{*}$ by
$f_{\boldsymbol{\beta}^{*}}=\operatorname*{arg\,min}_{f\in\mathcal{H}_{\boldsymbol{\beta}}}\mathcal{V}(f),$
We define the empirical process for the linear subspace as
$\left\{\mathcal{P}_{n}(f_{\boldsymbol{\beta}})=\mathcal{V}_{n}(f_{\boldsymbol{\beta}})-\mathcal{V}(f_{\boldsymbol{\beta}}):f_{\boldsymbol{\beta}}\in\mathcal{H}_{\boldsymbol{\beta}}\right\}.$
For a fixed (and arbitrary) $f_{\boldsymbol{\tilde{\beta}}}$, we define
$Z_{M}=\sup_{\lVert{\boldsymbol{\beta}}-\boldsymbol{\tilde{\beta}}\rVert_{1}\leq M}\bigg{\lvert}\mathcal{P}_{n}({f}_{\boldsymbol{\beta}})-\mathcal{P}_{n}({f}_{\boldsymbol{\tilde{\beta}}})\bigg{\rvert}$ (15)
and the set
=$\mathcal{B}=\{Z_{M}\leq\lambda_{0}M\},$
where
$\lambda_{0}$ $=4{T}\left(n,p\right)+\frac{t}{3n}+\sqrt{\frac{2t}{n}}\sqrt{1+8T(n,p)},$
$T(n,p)$ $=\sqrt{\frac{2\log(2p)}{n}}+\frac{\log(2p)}{3n}.$
By Lemma 14.20 and Theorem 14.5 of Bühlmann and van de Geer (2011), we have the probability inequality
$\mathbb{P}(\mathcal{B})\geq 1-\exp(-t),$
i.e. for some $M$ sufficiently small
$\mathbb{P}(Z_{M}\leq\lambda_{0}M)\geq 1-\eta$
for some $\lambda_{0}$ that depends on the sample size $n$, the dimensionality of the data $p$, and the confidence level $1-\eta$.
We denote the total sparsity and diversity penalties of the split logistic regression parameters for any set of $G$ linear functions $f_{\boldsymbol{\beta}^{1}},\dots,f_{\boldsymbol{\beta}^{G}}$ by
$P(f_{\boldsymbol{\beta}^{1}},\dots,f_{\boldsymbol{\beta}^{G}})=\sum_{g=1}^{G}P_{s}\left(\boldsymbol{\beta}^{g}\right)=\sum_{g=1}^{G}\left[\frac{1-\alpha}{2}\lVert\boldsymbol{\beta}^{g}\rVert_{2}^{2}+\alpha\lVert\boldsymbol{\beta}^{g}\rVert_{1}\right]$ (16)
where $\alpha\in[0,1]$, and
$Q(f_{\boldsymbol{\beta}^{1}},\dots,f_{\boldsymbol{\beta}^{G}})=\sum_{h\neq g}P_{d}\left(\boldsymbol{\beta}^{g},\boldsymbol{\beta}^{h}\right)=\sum_{h\neq g}\sum_{j=1}^{p}\lvert\beta_{j}^{g}\rvert\lvert\beta_{j}^{h}\rvert,$ (17)
respectively.
8.2 Ensemble Consistency
Let the solution to the split logistic regression objective function be the collection of functions
$\left({f}_{\boldsymbol{\hat{\beta}}^{1}},\dots,{f}_{\boldsymbol{\hat{\beta}}^{G}}\right)=\operatorname*{arg\,min}_{{f}_{\boldsymbol{\beta}^{1}},\dots,{f}_{\boldsymbol{\beta}^{G}}\in\mathcal{H}_{\boldsymbol{\beta}}}\left\{\sum_{g=1}^{G}\left[\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(f_{\boldsymbol{\beta}^{g}}(\mathbf{x}_{i}),y_{i}\right)+\lambda_{s}P_{s}(\boldsymbol{\beta}^{g})\right]+\frac{\lambda_{d}}{2}\sum_{h\neq g}P_{d}(\boldsymbol{\beta}^{h},\boldsymbol{\beta}^{g})\right\},$ (18)
and let
$({f}_{\boldsymbol{\tilde{\beta}}^{1}},\dots,{f}_{\boldsymbol{\tilde{\beta}}^{G}})=\operatorname*{arg\,min}_{{f}_{\boldsymbol{\beta}^{1}},\dots,{f}_{\boldsymbol{\beta}^{G}}\in\mathcal{H}_{\boldsymbol{\beta}}}\left\{\sum_{g=1}^{G}\left[\mathcal{E}(f_{\boldsymbol{\beta}^{g}})+\lambda_{s}P_{s}(\boldsymbol{\beta}^{g})\right]+\frac{\lambda_{d}}{2}\sum_{h\neq g}P_{d}(\boldsymbol{\beta}^{h},\boldsymbol{\beta}^{g})\right\},$ (19)
For $Z_{M}$ taken as (15) using the solution from (19), define
$\widetilde{M}=\frac{1}{G\lambda_{0}}\left[\sum_{g=1}^{G}\mathcal{E}({f}_{\boldsymbol{\tilde{\beta}}^{g}})+2\lambda_{s}P({f}_{\boldsymbol{\tilde{\beta}}^{1}},\dots,{f}_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q({f}_{\boldsymbol{\tilde{\beta}}^{1}},\dots,{f}_{\boldsymbol{\tilde{\beta}}^{G}})\right].$ (20)
Let the set
[…]
$\mathcal{V}\left(\frac{1}{G}\sum_{g=1}^{G}{f}_{\boldsymbol{\hat{\beta}}^{g}}\right)-\mathcal{V}(f^{*})\leq 2\left[\mathcal{E}({f}_{\boldsymbol{\beta}^{*}})+2\alpha\lambda_{s}\lVert\boldsymbol{\beta}^{*}\rVert_{1}+\frac{1-\alpha}{2}\lambda_{s}\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2}+\frac{\lambda_{d}(G-1)}{2}\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2}\right].$ (22)
Hence, if the target is linear, i.e. $f^{*}=f_{\boldsymbol{\beta}^{*}}$, then it holds that
$\mathcal{V}\left(\frac{1}{G}\sum_{g=1}^{G}{f}_{\boldsymbol{\hat{\beta}}^{g}}\right)-\mathcal{V}(f^{*})\leq 4\alpha\lambda_{s}\lVert\boldsymbol{\beta}^{*}\rVert_{1}+{(1-\alpha)}\lambda_{s}\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2}+{\lambda_{d}(G-1)}\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2}.$ (23)
Therefore, if the data come from a logistic model it follows that if we take $\lambda_{s}$ and $\lambda_{d}$ to be or order $\sqrt{log(p)/n}$, and we assume that $\lVert\boldsymbol{\beta}^{*}\rVert_{1}$ and $\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2}$ are of order smaller than $\sqrt{n/log(p)}$ and $log(p)/n\to 0$, then the ensemble prediction $(1/G)\sum_{g=1}^{G}{f}_{\boldsymbol{\hat{\beta}}^{g}}$ is consistent. In the more general case of model misspecification ($f^{*}\neq f_{\boldsymbol{\beta}^{*}}$), the prediction error converges to $2\,\mathcal{E}(f_{\boldsymbol{\beta}^{*}})$.
8.3 Ensemble Consistency Proof
Let ${f}_{\boldsymbol{\hat{\beta}}^{1}},\dots,{f}_{\boldsymbol{\hat{\beta}}^{G}}$ be the solution to split logistic regression with $G$ groups for data $\{y_{i},\mathbf{x}_{i}\}_{i=1}^{n}$. Then, for any $f_{\boldsymbol{\beta}^{1}},\dots,f_{\boldsymbol{\beta}^{G}}\in\mathcal{H}_{\boldsymbol{\beta}}$ it holds that
$\sum_{g=1}^{G}\mathcal{V}_{n}({f}_{\boldsymbol{\hat{\beta}}^{g}})+\lambda_{s}P({f}_{\boldsymbol{\hat{\beta}}^{1}},\dots,{f}_{\boldsymbol{\hat{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q({f}_{\boldsymbol{\hat{\beta}}^{1}},\dots,{f}_{\boldsymbol{\hat{\beta}}^{G}})$
$\leq$ $\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\beta}^{g}})+\lambda_{s}P(f_{\boldsymbol{\beta}^{1}},\dots,f_{\boldsymbol{\beta}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\beta}^{1}},\dots,f_{\boldsymbol{\beta}^{G}}).$
Note that if $\boldsymbol{\bar{\beta}}^{g}=t\boldsymbol{\hat{\beta}}^{g}+(1-t)\boldsymbol{\tilde{\beta}}^{g}$ for any $t\in[0,1]$, by a convexity argument
$\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})$
$\leq$ $t\left[\sum_{g=1}^{G}\mathcal{V}_{n}({f}_{\boldsymbol{\hat{\beta}}^{g}})+\lambda_{s}P({f}_{\boldsymbol{\hat{\beta}}^{1}},\dots,{f}_{\boldsymbol{\hat{\beta}}^{G}})\right]$
$+(1-t)\left[\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})\right]$
$\leq$ $t\left[\sum_{g=1}^{G}\mathcal{V}_{n}({f}_{\boldsymbol{\hat{\beta}}^{g}})+\lambda_{s}P({f}_{\boldsymbol{\hat{\beta}}^{1}},\dots,{f}_{\boldsymbol{\hat{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q({f}_{\boldsymbol{\hat{\beta}}^{1}},\dots,{f}_{\boldsymbol{\hat{\beta}}^{G}})\right]$
$+(1-t)\left[\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})\right]$
$\leq$ $\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}}).$
We can write
$\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})$
$=$ $\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})$
$+\left[\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\bar{\beta}}^{g}})-\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\bar{\beta}}^{g}})\right]+\left[\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})-\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})\right]$
$+\left[\sum_{g=1}^{G}\mathcal{V}(f_{\boldsymbol{\tilde{\beta}}^{g}})-\sum_{g=1}^{G}\mathcal{V}(f_{\boldsymbol{\tilde{\beta}}^{g}})\right]+\left[{\lambda_{s}}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})-\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})\right]$
$+\left[\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})-\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})\right]$
$=$ $-\left[\sum_{g=1}^{G}\mathcal{P}_{n}(f_{\boldsymbol{\bar{\beta}}^{g}})-\sum_{g=1}^{G}\mathcal{P}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})\right]+\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\tilde{\beta}}^{g}})$
$+\left[\left(\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})\right)-\left(\sum_{g=1}^{G}\mathcal{V}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})\right)\right]$
$+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}}).$
Thus we get the basic inequality
$\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})$
$\leq$ $-\left[\sum_{g=1}^{G}\mathcal{P}_{n}(f_{\boldsymbol{\bar{\beta}}^{g}})-\sum_{g=1}^{G}\mathcal{P}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})\right]+\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\tilde{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})$
$=$ $-\sum_{g=1}^{G}\left[\mathcal{P}_{n}(f_{\boldsymbol{\bar{\beta}}^{g}})-\mathcal{P}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})\right]+\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\tilde{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}}).$
In other words, to bound the sum of the excess risk $\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})$ we need to control the sum of the increments of the empirical processes $\mathcal{P}_{n}(f_{\boldsymbol{\bar{\beta}}^{g}})-\mathcal{P}_{n}(f_{\boldsymbol{\tilde{\beta}}^{g}})$, $1\leq g\leq G$.
Let
$t^{g}=\frac{\widetilde{M}}{\widetilde{M}+\lVert{\boldsymbol{\hat{\beta}}^{g}}-\boldsymbol{\tilde{\beta}}^{g}\rVert_{1}}.$
[…]
$\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})$
$\leq$ $\sum_{g=1}^{G}Z_{\widetilde{M}}+\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\tilde{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})$
$=$ $\sum_{g=1}^{G}\lambda_{0}\widetilde{M}+\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\tilde{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}}).$
For the case $\alpha\in(0,1]$ we obtain
$\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})$
$\leq$ $G\lambda_{0}\widetilde{M}+\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\tilde{\beta}}^{g}})+2\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})$
$=$ $2G\lambda_{0}\widetilde{M}\leq\alpha\lambda_{s}\frac{\widetilde{M}}{2}$
since $\lambda_{s}\geq 4G\lambda_{0}/\alpha$. Notice that
$\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})$
$\geq$ $\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}\alpha\sum_{g=1}^{G}\|\boldsymbol{\bar{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\|_{1}+\lambda_{s}\frac{1-\alpha}{2}\sum_{g=1}^{G}\|\boldsymbol{\bar{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\|_{2}^{2}$
$\geq$ $\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}\alpha\sum_{g=1}^{G}\|\boldsymbol{\bar{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\|_{1}.$
This implies
$\lVert\boldsymbol{\bar{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\rVert_{1}\leq\frac{\widetilde{M}}{2}$
for all $1\leq g\leq G$, which in turn implies
$\lVert\boldsymbol{\hat{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\rVert_{1}\leq\widetilde{M}$
for all $1\leq g\leq G$.
In the case $\alpha=0$, we obtain
$\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})$
$\leq$ $G\lambda_{0}\widetilde{M}+\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\tilde{\beta}}^{g}})+2\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})$
$=$ $2G\lambda_{0}\widetilde{M}\leq\frac{\lambda_{s}}{G^{2}}\frac{\widetilde{M}^{2}}{4},$
since $\lambda_{s}\geq 8G^{3}\lambda_{0}/\widetilde{M}$, and
$\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\lambda_{s}P(f_{\boldsymbol{\bar{\beta}}^{1}},\dots,f_{\boldsymbol{\bar{\beta}}^{G}})+\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})$
$\geq$ $\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\frac{\lambda_{s}}{2}\sum_{g=1}^{G}\|\boldsymbol{\bar{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\|_{2}^{2}$
$\geq$ $\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\frac{\lambda_{s}}{2{G}}\sum_{g=1}^{G}\|\boldsymbol{\bar{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\|_{1}^{2}$
$\geq$ $\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\bar{\beta}}^{g}})+\frac{\lambda_{s}}{2{G}^{2}}\left(\sum_{g=1}^{G}\|\boldsymbol{\bar{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\|_{1}\right)^{2}.$
This again implies that $\lVert\boldsymbol{\bar{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\rVert_{1}\leq{\widetilde{M}}/{2}$ and $\lVert\boldsymbol{\hat{\beta}}^{g}-\boldsymbol{\tilde{\beta}}^{g}\rVert_{1}\leq\widetilde{M}$ for $1\leq g\leq G$.
[…]
$\sum_{g=1}^{G}\mathcal{E}({f}_{\boldsymbol{\hat{\beta}}^{g}})\leq 2\left[\sum_{g=1}^{G}\mathcal{E}(f_{\boldsymbol{\tilde{\beta}}^{g}})+2\lambda_{s}P(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})+\frac{\lambda_{d}}{2}Q(f_{\boldsymbol{\tilde{\beta}}^{1}},\dots,f_{\boldsymbol{\tilde{\beta}}^{G}})\right]$
Notice that for the best linear predictor $f_{\boldsymbol{\beta}^{*}}$ used for all $G$ functions, we can rewrite (16) and (17) as
$P(f_{\boldsymbol{\beta}^{*}},\dots,f_{\boldsymbol{\beta}^{*}})$ $=\alpha G\lVert\boldsymbol{\beta}^{*}\rVert_{1}+\frac{1-\alpha}{2}G\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2},\,\text{and}$
$Q(f_{\boldsymbol{\beta}^{*}},\dots,f_{\boldsymbol{\beta}^{*}})$ $=G(G-1)\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2},$
respectively. Thus
$\frac{1}{G}\sum_{g=1}^{G}\mathcal{E}({f}_{\boldsymbol{\hat{\beta}}^{g}})\leq 2\left[\mathcal{E}(f_{\boldsymbol{\beta}^{*}})+2\alpha\lambda_{s}\lVert\boldsymbol{\beta}^{*}\rVert_{1}+\frac{1-\alpha}{2}\lambda_{s}\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2}+\frac{\lambda_{d}(G-1)}{2}\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2}\right].$
By the convexity of (14),
$\frac{1}{G}\sum_{g=1}^{G}\mathcal{E}({f}_{\boldsymbol{\hat{\beta}}^{g}})=\frac{1}{G}\sum_{g=1}^{G}\mathcal{V}({f}_{\boldsymbol{\hat{\beta}}^{g}})-\mathcal{V}(f^{*})\geq\mathcal{V}\left(\frac{1}{G}\sum_{g=1}^{G}{f}_{\boldsymbol{\hat{\beta}}^{g}}\right)-\mathcal{V}(f^{*}),$
so we get the desired inequality (22),
$\mathcal{V}\left(\frac{1}{G}\sum_{g=1}^{G}{f}_{\boldsymbol{\hat{\beta}}^{g}}\right)-\mathcal{V}(f^{*})\leq 2\left[\mathcal{E}(f_{\boldsymbol{\beta}^{*}})+2\alpha\lambda_{s}\lVert\boldsymbol{\beta}^{*}\rVert_{1}+\frac{1-\alpha}{2}\lambda_{s}\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2}+\frac{\lambda_{d}(G-1)}{2}\lVert\boldsymbol{\beta}^{*}\rVert_{2}^{2}\right].$
Appendix B: Details of the Algorithm
In this section, we provide the derivation for the quadratic approximation of the logistic regression loss, the high-level steps of the block coordinate descent algorithm, and a detailed description of the alternating grid search for the tuning parameters.
8.4 Quadratic Approximation
For the binary classification problem with the classes labeled as $Y=\{-1,1\}$, let $\mathbf{y}\in\mathbb{R}^{n}$ be the vector of class labels and $\mathbf{X}\in\mathbb{R}^{n\times p}$ be the design matrix with sample size $n$ and number of features $p$. The logistic regression loss function is given by
$\mathcal{L}\left(f(\mathbf{x}_{i}),y_{i}\right)=\mathcal{L}(\beta_{0},\boldsymbol{\beta}\mid y_{i},\mathbf{x}_{i})=\log\left(1+e^{-y_{i}f(\mathbf{x}_{i})}\right),\quad 1\leq i\leq n,$ (24)
where $f(\mathbf{x}_{i})=\beta_{0}+\mathbf{x}_{i}^{T}\boldsymbol{\beta}$ is a linear function of the predictor variables, $\beta_{0}\in\mathbb{R}$ and $\boldsymbol{\beta}\in\mathbb{R}^{p}$ are the intercept and vector of regression coefficients.
We denote ${\mathbf{X}_{A}}\in\mathbb{R}^{n\times(p+1)}$ the augmented design matrix whose first column is a column of ones and $\boldsymbol{\beta}_{A}\in\mathbb{R}^{p+1}=(\beta_{0},\boldsymbol{\beta}^{T})^{T}$ the vector with all regression parameters. The quadratic approximation for the logistic regression loss in (24) at the current estimates $\boldsymbol{\tilde{\beta}}_{A}$ is given by
$\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{Q}(\beta_{0},\boldsymbol{\beta}\mid y_{i},\mathbf{x}_{i})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\tilde{\beta}_{0},\boldsymbol{\tilde{\beta}}\mid y_{i},\mathbf{x}_{i})+\mathcal{V}\left(\tilde{\beta}_{0},\boldsymbol{\tilde{\beta}}\mid\mathbf{y},\mathbf{X}_{A}\right)\left(\boldsymbol{\beta}_{A}-\boldsymbol{\tilde{{\beta}}}_{A}\right)$
$+\frac{1}{2}\left(\boldsymbol{\beta}_{A}-\boldsymbol{\tilde{{\beta}}}_{A}\right)^{T}\mathcal{H}\left(\tilde{\beta}_{0},\boldsymbol{\tilde{\beta}}\mid\mathbf{y},\mathbf{X}_{A}\right)\left(\boldsymbol{\beta}_{A}-\boldsymbol{\tilde{{\beta}}}_{A}\right),$
where the gradient vector and hessian matrix are given by
$\mathcal{V}\left(\tilde{\beta}_{0},\boldsymbol{\tilde{\beta}}\mid\mathbf{y},\mathbf{X}\right)$ $=\nabla\left(\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}({\beta}_{0},\boldsymbol{{\beta}}\mid y_{i},\mathbf{x}_{i})\right)\Bigg{|}_{\left({\beta}_{0},\,\boldsymbol{{\beta}}\right)=\left(\tilde{\beta}_{0},\,\boldsymbol{\tilde{\beta}}\right)}$
$=\frac{1}{n}\mathbf{X}_{A}^{T}(\mathbf{z}-\boldsymbol{\tilde{p}}),$
$\mathcal{H}\left(\tilde{\beta}_{0},\boldsymbol{\tilde{\beta}}\mid\mathbf{y},\mathbf{X}\right)$ $=\nabla\left(\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}({\beta}_{0},\boldsymbol{{\beta}}\mid y_{i},\mathbf{x}_{i})\right)\nabla^{T}\Bigg{|}_{\left({\beta}_{0},\,\boldsymbol{{\beta}}\right)=\left(\tilde{\beta}_{0},\,\boldsymbol{\tilde{\beta}}\right)}$
$=-\frac{1}{n}\mathbf{X}_{A}^{T}\mathbf{\tilde{W}}\mathbf{X}_{A}.$
The elements of the $n$-dimensional vectors $\mathbf{z}$, $\mathbf{\tilde{p}}$ and $\mathbf{\tilde{w}}$ are given by $z_{i}=(y_{i}+1)/2$, $\tilde{p}_{i}=S(\tilde{\beta}_{0}+\mathbf{x}_{i}^{T}\boldsymbol{\tilde{\beta}})$ and $\tilde{w}_{i}=p_{i}(1-p_{i})$, $1\leq i\leq n$ respectively. The $n\times n$ weight matrix at the current parameter estimates is given by $\mathbf{\tilde{W}}=\text{diag}(\mathbf{\tilde{w}})$. The quadratic approximation can subsequently be rewritten as a weighted least-squares problem
$\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{Q}(\beta_{0},\boldsymbol{\beta}\mid y_{i},\mathbf{x}_{i})$ $=\frac{1}{2n}\left(\boldsymbol{\tilde{y}}-\mathbf{X}_{A}\boldsymbol{\beta}_{A}\right)^{T}\boldsymbol{\tilde{W}}\left(\boldsymbol{\tilde{y}}-\mathbf{X}_{A}\boldsymbol{\beta}_{A}\right)+C\left(\tilde{\beta}_{0},\boldsymbol{\tilde{\beta}}\right)$
$=\frac{1}{2n}\sum_{i=1}^{n}\tilde{{w}}_{i}\left(\tilde{y}_{i}-f(\mathbf{x}_{i})\right)^{2}+C\left(\tilde{\beta}_{0},\boldsymbol{\tilde{\beta}}\right),$ (25)
where the elements of the $n$-dimensional vector $\mathbf{\tilde{y}}$ are given by $\tilde{y}_{i}=\tilde{\beta}_{0}+\mathbf{x}_{i}^{T}\boldsymbol{\tilde{\beta}}+(z_{i}-\tilde{{p}}_{i})/\tilde{{w}}_{i}$, $1\leq i\leq n$, and $C\left(\tilde{\beta}_{0},\boldsymbol{\tilde{\beta}}\right)$ is a constant term.
8.5 Block Coordinate Descent Algorithm
The objective function is multi-convex and can be written as a weighed elastic net problem for each individual model, where the $L_{1}$ penalty depends on the parameters in the other models. In particular, for a given model $g$, the objective function is given by
$\mathcal{J}\left(\beta_{0}^{g},\boldsymbol{\beta}^{g}\bigm{|}\mathbf{y},\mathbf{X}\right)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\beta_{0}^{g},\boldsymbol{\beta}^{g}\mid y_{i},\mathbf{x}_{i})+\lambda_{s}\frac{(1-\alpha)}{2}\|\boldsymbol{\beta}^{g}\|_{2}^{2}+\sum\limits_{j=1}^{p}|\beta_{j}^{g}|u_{j,g},\quad 1\leq g\leq G.$ (26)
with weights
$u_{j,g}=\alpha\lambda_{s}+\frac{\lambda_{d}}{2}\sum_{h\neq g}|\beta_{j}^{h}|.$
We apply a block coordinate descent algorithm by cycling through the parameters of one model at a time and we apply the coordinate descent updates in a deterministic, cyclic order. When updating the parameters of an individual model, a single coordinate descent update is applied for each parameter as follows. For notational convenience, we denote by $\mathbf{\tilde{{p}}}^{g}$, $\mathbf{\tilde{{w}}}^{g}$ and $\mathbf{\tilde{{y}}}^{g}$ the $n$-dimensional vectors with elements $\tilde{p}_{i}^{g}=S(\tilde{\beta_{0}}^{g}+\mathbf{x}_{i}^{T}\boldsymbol{\tilde{\beta}}^{g})$, $\tilde{w}_{i}^{g}=\tilde{p}_{i}^{g}(1-\tilde{p}_{i}^{g})$ and $\tilde{y}_{i}^{g}=\tilde{\beta}_{0}^{g}+\mathbf{x}_{i}^{T}\boldsymbol{\tilde{\beta}}^{g}+(z_{i}-\tilde{{p}}_{i}^{g})/\tilde{{w}}_{i}^{g}$, $1\leq i\leq n$, respectively. To obtain the coordinate descent updates we replace the logistic loss in the objective function (26) by its quadratic approximation (25) at the current parameter estimates for the ensemble. For parameter $j$ of a particular model $g$, $1\leq j\leq p$, the coordinate descent update is then given by
$\hat{\beta}_{j}^{g}$ $=\operatorname*{arg\,min}_{\beta_{j}^{g}\in\mathbb{R}}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{Q}(\beta_{0}^{g},\boldsymbol{\beta}^{g}\mid y_{i},\mathbf{x}_{i})+\lambda_{s}\frac{(1-\alpha)}{2}\|\boldsymbol{\beta}^{g}\|_{2}^{2}+\sum\limits_{j=1}^{p}|\beta_{j}^{g}|u_{j,g}$
$=\operatorname*{arg\,min}_{\beta_{j}^{g}\in\mathbb{R}}\frac{1}{2n}\sum_{i=1}^{n}\tilde{{w}}_{i}\left(\tilde{y}_{i}^{g}-\beta_{0}^{g}-\sum_{i=1}^{n}\sum_{k\neq j}^{p}x_{ik}\tilde{\beta}_{k}^{g}-{\beta}_{j}^{g}\,x_{ij}\right)^{2}+\lambda_{s}\frac{(1-\alpha)}{2}\left(\beta_{j}^{g}\right)^{2}+|\beta_{j}^{g}|u_{j,g}$
$=\frac{\text{Soft}\left(\frac{1}{n}\left({\tilde{r}_{j}^{g}}+\tilde{\beta}_{j}^{g}\langle\mathbf{x}_{j}^{2},\mathbf{\tilde{w}}^{g}\rangle\right),\,\alpha\lambda_{s}+\frac{\lambda_{d}}{2}\sum_{h\neq g}|\tilde{\beta}^{h}_{j}|\right)}{\frac{1}{n}\langle\mathbf{x}_{j}^{2},\mathbf{\tilde{w}}^{g}\rangle+(1-\alpha)\lambda_{s}},$
where $\tilde{r}_{j}^{g}=\left\langle\mathbf{x}_{j},{\mathbf{z}}\right\rangle-\langle\mathbf{x}_{j},\mathbf{\tilde{p}}^{g}\rangle$, and the last equality follows from the optimality condition for subgradients. A similar derivation can be made for the coordinate descent update of the intercept term
$\hat{\beta}_{0}^{g}$ $=\tilde{\beta}_{0}^{g}+\frac{\langle\mathbf{z}-\mathbf{\tilde{p}}^{g},\boldsymbol{1}_{n}\rangle}{\langle\mathbf{\tilde{w}}^{g},\boldsymbol{1}_{n}\rangle},$
which yields the results in Proposition 1 of the article. When all parameter estimates of model $g$ have been updated, also the vectors $\mathbf{\tilde{{p}}}^{g}$ and $\mathbf{\tilde{{w}}}^{g}$ are updated. The active set cycling strategy (Friedman et al., 2010) is also adopted and available in our software implementation. In Algorithm 1 we provide the steps to generate solutions for split logistic regression when $\lambda_{s}$ and $\lambda_{d}$ are fixed.
Algorithm 1 Split Logistic Regression for Fixed $\lambda_{s}$ and $\lambda_{d}$
1:Design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$, response vector $\mathbf{y}\in\mathbb{R}^{n}$, current solutions $\boldsymbol{\tilde{\beta}}_{1:G}$, $\ell_{1}$-$\ell_{2}$ mixing parameter $\alpha\in[0,1]$, sparsity and diversity tuning parameters $\lambda_{s},\lambda_{d}\geq 0$, and convergence tolerance parameter $\delta>0$.
2:
3:Compute the current probabilities $\tilde{p}_{i}^{g}=S(\tilde{\beta_{0}}^{g}+\mathbf{x}_{i}^{T}\boldsymbol{\tilde{\beta}}^{g})$, weights $\tilde{w}_{i}^{g}=\tilde{p}_{i}^{g}(1-\tilde{p}_{i}^{g})$ and residuals $\tilde{r}_{j}^{g}=\left\langle\mathbf{x}_{j},{\mathbf{z}}\right\rangle-\langle\mathbf{x}_{j},\mathbf{\tilde{p}}^{g}\rangle$, $1\leq i\leq n$, $1\leq j\leq p$, $1\leq g\leq G$.
4:
5:Repeat the following steps until convergence.
  1. 2.1:
    For each model $g$, $1\leq g\leq G$:
    1. 2.1.1:
      Perform a single (block) coordinate descent update for the intercept and each predictor $j$, $1\leq j\leq p$.
      1. 2.1.1.1:
        Compute the new intercept in model $g$,
        $\hat{\beta}_{0}^{g}=\tilde{\beta}_{0}^{g}+\frac{\langle\mathbf{z}-\mathbf{\tilde{p}}^{g},\boldsymbol{1}_{n}\rangle}{\langle\mathbf{\tilde{w}}^{g},\boldsymbol{1}_{n}\rangle}.$
      2. 2.1.1.2:
        If $\hat{\beta}_{0}^{g}\neq\tilde{\beta}_{0}^{g}$, then update the probabilities $\mathbf{\tilde{p}}^{g}$, weights $\mathbf{\tilde{w}}^{g}$ and residuals $\mathbf{\tilde{r}}^{g}$ for model $g$.
      3. 2.1.1.3:
        Update $j$-th coefficient in model $g$,
        $\hat{\beta}^{g}_{j}=\frac{\operatorname{Soft}\left(\frac{1}{n}\left({\tilde{r}_{j}^{g}}+\tilde{\beta}_{j}^{g}\langle\mathbf{x}_{j}^{2},\mathbf{\tilde{w}}^{g}\rangle\right),\,\alpha\lambda_{s}+\frac{\lambda_{d}}{2}\sum_{h\neq g}|\tilde{\beta}^{h}_{j}|\right)}{\frac{1}{n}\langle\mathbf{x}_{j}^{2},{\mathbf{\tilde{w}}}^{g}\rangle+(1-\alpha)\lambda_{s}}.$
      4. 2.1.1.4:
        If $\hat{\beta}_{j}^{g}\neq\tilde{\beta}_{j}^{g}$, then update the probabilities $\mathbf{\tilde{p}}^{g}$, weights $\mathbf{\tilde{w}}^{g}$ and residuals $\mathbf{\tilde{r}}^{g}$ for model $g$.
6:If successive estimates of the coefficients in the ensemble model show little difference, i.e.
$\max_{1\leq j\leq p}\left(\frac{1}{G}\sum_{g=1}^{G}\tilde{\beta}_{j}^{g}-\frac{1}{G}\sum_{g=1}^{G}\hat{\beta}_{j}^{g}\right)^{2}<\delta,$
then convergence is declared.
7:
8:Return the coefficients for each model $(\hat{\beta}_{0}^{g},\boldsymbol{\hat{\beta}}^{g})$, $1\leq g\leq G$.
8.6 Alternating Grid Search for Tuning Parameters
The selection of the sparsity and diversity tuning parameters, $\lambda_{s}$ and $\lambda_{d}$, is done by an alternating grid search. The first grid search is over $\lambda_{s}$ with the diversity tuning parameter fixed at $\lambda_{d}^{(0)}=0$, which yields a first value $\lambda_{s}^{\text{opt}}$ minimizing the cross-validated loss. Keeping the sparsity parameter fixed at value $\lambda_{s}^{\text{opt}}$, we now perform a grid search over $\lambda_{d}$ which yields $\lambda_{d}^{\text{opt}}$. This process is repeated until the cross-validated loss no longer decreases. The high-level steps of the alternating grid search are given in Algorithm 2.
To construct a grid for $\lambda_{s}$, we estimate a value $\lambda_{s}^{\max}$ that makes all models null. In the special case where $\lambda_{d}=0$ and $\alpha>0$, it can easily be shown that $\lambda_{s}^{\text{max}}=\frac{1}{2\alpha}\max_{1\leq j\leq p}|\bar{\mathbf{x}}_{j}|$. For $\lambda_{d}>0$, we estimate the smallest $\lambda_{s}^{\text{max}}$ that makes all models null by performing an internal grid search. Based on this maximal value $\lambda_{s}^{\text{max}}$ we then construct the grid for the sparsity penalty $\lambda_{s}$ similarly to the case of (single-model) penalized logistic regression. that is, we use (by default) 100 log-equispaced points between $\epsilon\lambda_{s}^{\text{max}}$ and $\lambda_{s}^{\text{max}}$, where $\epsilon=10^{-4}$ if $p<n$ and $10^{-2}$ otherwise.
The smallest diversity penalty $\lambda_{d}^{\text{max}}$ that makes the models fully disjoint for some fixed $\lambda_{s}\geq 0$ is similarly estimated via a grid search. We then analogously generate the diversity penalty grid using (by default) 100 log-equispaced points between $\epsilon\lambda_{d}^{\text{max}}$ and $\lambda_{d}^{\text{max}}$. For a grid search over one of the tuning parameters while keeping the other one fixed, we use warm-starts by computing solutions for a decreasing sequence of $\lambda_{s}$ or $\lambda_{d}$, leading to a more stable algorithm.
Algorithm 2 Alternating CV Procedure
1:Design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$, response vector $\mathbf{y}\in\mathbb{R}^{n}$, $\ell_{1}$-$\ell_{2}$ mixing parameter $\alpha\in[0,1]$ and convergence tolerance parameter $\delta>0$.
2:
3:Set $\lambda_{d}^{\text{opt}}=0$ and the next search is for the sparsity tuning parameter $\lambda_{s}^{\text{opt}}$.
4:
5:Alternate between a search for $\lambda_{s}^{\text{opt}}$ or $\lambda_{d}^{\text{opt}}$ until CV MSPE no longer decreases.
  1. 2.1:
    If the search is for $\lambda_{s}^{\text{opt}}$:
    1. 2.1.1:
      If $\lambda_{d}^{\text{opt}}=0$, set $\lambda_{s}^{\text{max}}=({1}/{2\alpha})\max_{1\leq j\leq p}|\bar{\mathbf{x}}_{j}|$. Otherwise perform a grid search to find the smallest $\lambda_{s}^{\text{max}}$ such that each model is null.
    2. 2.1.2:
      Generate the log-equispaced grid between $\epsilon\lambda_{s}$ and $\lambda_{s}^{\text{max}}$.
    3. 2.1.3:
      For each $\lambda_{s}$ in the log-equispaced grid compute $\boldsymbol{\hat{\beta}}_{1:G}(\lambda_{s})=(\boldsymbol{\hat{\beta}}^{1}(\lambda_{s}),\dots,\boldsymbol{\hat{\beta}}^{G}(\lambda_{s}))$ with Algorithm 1, using the previous solution in the grid as a warm-start.
    4. 2.1.4:
      Set $\lambda_{s}^{\text{opt}}$ using the value in the grid that minimized the CV MSPE.
    Otherwise if the search is for $\lambda_{d}^{\text{opt}}$:
    1. 2.1.1:
      Perform a grid search to find the smallest $\lambda_{d}^{\text{max}}$ such that makes models fully disjoint.
    2. 2.1.2:
      Generate the log-equispaced grid between $\epsilon\lambda_{d}$ and $\lambda_{d}^{\text{max}}$.
    3. 2.1.3:
      For each $\lambda_{d}$ in the log-equispaced grid compute $\boldsymbol{\hat{\beta}}_{1:G}(\lambda_{d})=(\boldsymbol{\hat{\beta}}^{1}(\lambda_{d}),\dots,\boldsymbol{\hat{\beta}}^{G}(\lambda_{d}))$ with Algorithm 1, using the previous solution in the grid as a warm-start.
    4. 2.1.4:
      Set $\lambda_{d}^{\text{opt}}$ using the value in the grid that minimized the CV MSPE.
6:For $\lambda_{d}^{\text{opt}}$ and the smallest $\lambda_{s}^{\text{max}}$ such that each model is null, generate the log-equispaced grid between $\epsilon\lambda_{s}$ and $\lambda_{s}^{\text{max}}$.
7:
8:For each $\lambda_{s}$ in the log-equispaced grid compute $\boldsymbol{\hat{\beta}}_{1:G}(\lambda_{s})=(\boldsymbol{\hat{\beta}}^{1}(\lambda_{s}),\dots,\boldsymbol{\hat{\beta}}^{G}(\lambda_{s}))$ with Algorithm 1, using the previous solution in the grid as a warm-start.
9:
10:Return the coefficients of the models $\boldsymbol{\hat{\beta}}_{1:G}(\lambda_{s})=(\boldsymbol{\hat{\beta}}^{1}(\lambda_{s}),\dots,\boldsymbol{\hat{\beta}}^{G}(\lambda_{s}))$ for each $\lambda_{s}$ in the grid.
Appendix C: Alternative Diversity Measures
In this section, we investigate the accuracy-diversity trade-off using several alternative diversity measures to complement and consolidate the results obtained in Section 6 of the main article based on the entropy diversity measure.
8.7 Disagreement Measure
The disagreement (DIS) diversity measure (Skalak et al., 1996; Ho, 1998) of an ensemble comprised of $G$ individual classifiers for a given input $\mathbf{x}$ is defined as
$\text{DIS}(\mathbf{x})=\frac{1}{G(G-1)}\sum_{g=1}^{G}\sum_{h\neq g}^{G}\text{DIS}_{g,h}(\mathbf{x})$
where the disagreement between between classifiers $g$ and $h$ is given by
$\text{DIS}_{g,h}(\mathbf{x})=\begin{cases}1,&\text{if classifiers$g$ and $h$ disagree on the class of }\mathbf{x},\\ 0,&\text{if classifiers $g$ and $h$ agree on the class of }\mathbf{x}.\end{cases}$
The disagreement measure is a pairwise diversity measure and ranges between 0 and 1, where $\text{DIS}(\mathbf{x})=0$ corresponds to no disagreement and increasing values of $\text{DIS}(\mathbf{x})$ correspond to more disagreement between the individual classifiers.
8.8 Double-Fault Measure
The double-fault (DF) diversity measure (Giacinto and Roli, 2001) of an ensemble comprised of $G$ individual classifiers for some a given $\mathbf{x}$ is defined as
$\text{DF}(\mathbf{x})=\frac{1}{G(G-1)}\sum_{g=1}^{G}\sum_{h\neq g}^{G}\text{DF}_{g,h}(\mathbf{x})$
where the double-fault between between classifiers $g$ and $h$ is given by
$\text{DF}_{g,h}(\mathbf{x})=\begin{cases}1,&\text{if classifiers $g$ and $h$ both misclassify }\mathbf{x},\\ 0,&\text{if at most one of classifiers $g$ and $h$ misclassify }\mathbf{x}.\end{cases}$
The double-fault measure is a pairwise diversity measure and ranges between 0 and 1, where $\text{DF}(\mathbf{x})=0$ corresponds to no double-faults and increasing values of $\text{DF}(\mathbf{x})$ correspond to more double-faults between the individual classifiers.
8.9 Kohavi-Wolpert Variance
The Kohavi-Wolpert variance (KW) diversity measure (Kohavi et al., 1996) of an ensemble comprised of $G$ individual classifiers for a given input $\mathbf{x}$ is defined as
$\text{KW}(\mathbf{x})=\frac{1}{G^{2}}\,\ell(\mathbf{x})(G-\ell(\mathbf{x}))$
where $\ell(\mathbf{x})$ denotes the numbers of individual classifiers that correctly classified input $\mathbf{x}$. The Kohavi-Wolpert variance is a non-pairwise diversity measure with $\text{KW}(\mathbf{x})=0$ corresponding to no diversity and increasing values of $\text{KW}(\mathbf{x})$ corresponding to more diversity between the individual classifiers. Kuncheva and Whitaker (2003) have shown that
$\text{KW}(\mathbf{x})=\frac{G-1}{2G}\,\text{DIS}(\mathbf{x}).$
8.10 Generalized Diversity
The generalized diversity (GD) measure (Partridge and Krzanowski, 1997) of an ensemble comprised of $G$ individual classifiers for a given input $\mathbf{x}$ is defined as
$\text{GD}(\mathbf{x})=1-\frac{\sum_{g=1}^{G}\frac{g(g-1)}{G(G-1)}\,\mathbb{P}(\ell(\mathbf{x})=g)}{\sum_{g=1}^{G}\frac{g}{G}\,\mathbb{P}(\ell(\mathbf{x})=g)}$
where $\ell(\mathbf{x})$ denotes the numbers of individual classifiers from the ensemble that correctly classify input $\mathbf{x}$. The generalized diversity is a non-pairwise diversity measure and ranges between 0 and 1, where $\text{GD}(\mathbf{x})=0$ corresponds to no diversity and $\text{GD}(\mathbf{x})=1$ corresponds to maximum diversity between the individual classifiers.
8.11 Results
In Table 12 we report the results of the alternative diversity measures for the same simulation settings as in Table 2 in the article. Similarly to the entropy diversity in the article, the DIS, DF, KW and GD measures are reported as a function of the number of models in split logistic regression, averaged over the test sets. It can be seen that the DIS, KW and GD diversity measures all increase with the number of models, while the DF diversity measure decreases. Hence, all measures confirm that the individual models become more diverse when the number of models increases.
Table 12: DIS, DF, KW and GD as a function of the number of models under Scenario 3.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{G}$ DIS DF KW GD DIS DF KW GD DIS DF KW GD
2 0.04 0.84 0.01 0.37 0.05 0.86 0.01 0.50 0.05 0.85 0.01 0.45
3 0.07 0.83 0.03 0.59 0.08 0.84 0.03 0.63 0.07 0.84 0.03 0.63
4 0.07 0.82 0.03 0.61 0.10 0.83 0.04 0.72 0.08 0.84 0.03 0.68
5 0.10 0.80 0.05 0.72 0.13 0.81 0.06 0.78 0.11 0.82 0.05 0.75
6 0.12 0.78 0.06 0.75 0.14 0.79 0.07 0.79 0.13 0.80 0.06 0.79
7 0.14 0.77 0.07 0.77 0.15 0.78 0.07 0.80 0.14 0.79 0.07 0.80
8 0.14 0.76 0.07 0.77 0.17 0.77 0.08 0.81 0.15 0.78 0.07 0.81
Appendix D: Full Results of Simulation Study
In this section, the full results of the simulation study in Section 5 of the article are reported. For the five scenarios described in the article and the considered values of the correlation parameters ($\rho$, $\rho_{1}$ and $\rho_{2}$), sample size ($n$), sparsity level ($\zeta$), and probability of positive events ($\pi_{1}$), the tables report the prediction accuracy (ACC), sensitivity (SNS), specificity (SPC), area under the ROC Curve (AUC), test-sample loss (TSL), recall (RCL) and precision (PRC) for the fourteen methods listed in Section 3 of the article.
  • Scenario 1:
    • Tables 2-7 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 8-13 contain the results for the RCL and PR.
  • Scenario 2:
    • Tables 14-19 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 20-25 contain the results for the RCL and PR.
  • Scenario 3:
    • Tables 26-31 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 32-37 contain the results for the RCL and PR.
  • Scenario 4:
    • Tables 38-43 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 44-49 contain the results for the RCL and PR.
  • Scenario 5:
    • Tables 50-55 contain the results for the ACC, SNS, SPC, AUC and TSL.
    • Tables 56-61 contain the results for the RCL and PR.
Appendix E: Full Results for Medical Genomics Data
Tables 62-75 contain the full results for the gene expression data benchmark study in Section 6 of the article. For each data set we report the performance of the fourteen methods listed in Section 3 of the article for the considered training set proportions and the different numbers of genes $p$ preserved after the pre-processing step, where we considered $p=$ 100, 250, 500 and 1,000. To simplify the comparison for each setting, we report relative performance of the methods in terms of ACC and TSL. That is, the performance reported for each method is relative to the best performing method, i.e. the ratio of its performance to the best performing method. Thus, the best performing method in each column of the tables corresponds the value 1.00.
Table 13: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.73 0.58 0.84 0.81 1.11 0.87 0.81 0.91 0.96 0.57 0.90 0.83 0.94 0.97 0.47
Split-EN-10 0.73 0.58 0.84 0.82 1.09 0.88 0.82 0.91 0.96 0.56 0.90 0.84 0.94 0.97 0.46
Lasso 0.69 0.54 0.81 0.76 1.23 0.81 0.72 0.87 0.90 0.85 0.82 0.72 0.89 0.91 0.77
Elastic Net 0.71 0.55 0.82 0.78 1.17 0.83 0.76 0.89 0.92 0.74 0.86 0.77 0.91 0.94 0.64
Adaptive 0.60 0.20 0.92 0.65 1.32 0.76 0.55 0.89 0.86 1.00 0.78 0.58 0.91 0.88 0.95
Relaxed 0.69 0.54 0.81 0.76 1.28 0.80 0.73 0.85 0.89 1.11 0.82 0.72 0.88 0.90 0.96
0.4 MCP 0.66 0.46 0.81 0.71 1.32 0.74 0.59 0.83 0.82 1.10 0.75 0.58 0.85 0.82 1.06
SIS-SCAD 0.65 0.51 0.75 0.70 1.56 0.71 0.58 0.79 0.78 1.27 0.72 0.59 0.80 0.78 1.31
RuleFit 0.67 0.55 0.76 0.72 1.84 0.73 0.61 0.82 0.80 1.40 0.75 0.60 0.85 0.83 1.20
RE-Lasso-100 0.73 0.60 0.84 0.82 1.10 0.86 0.78 0.92 0.95 0.61 0.88 0.79 0.95 0.96 0.54
RE-EN-100 0.74 0.60 0.84 0.82 1.11 0.87 0.80 0.92 0.96 0.58 0.89 0.81 0.95 0.97 0.50
RGLM-100 0.71 0.46 0.90 0.82 1.15 0.83 0.66 0.95 0.96 0.94 0.85 0.66 0.97 0.97 0.91
RF-500 0.73 0.61 0.83 0.82 1.06 0.85 0.75 0.92 0.95 0.79 0.87 0.75 0.95 0.96 0.76
XGB 0.65 0.53 0.74 0.69 1.29 0.73 0.61 0.81 0.81 1.07 0.74 0.60 0.83 0.81 1.04
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.76 0.44 0.92 0.82 1.01 0.88 0.72 0.95 0.96 0.53 0.90 0.78 0.95 0.97 0.44
Split-EN-10 0.76 0.46 0.91 0.82 1.00 0.89 0.73 0.95 0.96 0.51 0.91 0.79 0.96 0.97 0.42
Lasso 0.73 0.36 0.91 0.76 1.11 0.83 0.60 0.93 0.90 0.77 0.84 0.62 0.93 0.91 0.73
Elastic Net 0.74 0.38 0.91 0.78 1.08 0.85 0.66 0.94 0.93 0.66 0.87 0.70 0.94 0.94 0.60
Adaptive 0.68 0.06 0.98 0.60 1.25 0.77 0.32 0.96 0.83 0.96 0.78 0.34 0.96 0.84 0.93
Relaxed 0.72 0.34 0.90 0.75 1.17 0.82 0.63 0.91 0.90 0.94 0.83 0.63 0.91 0.90 1.27
0.3 MCP 0.70 0.26 0.91 0.69 1.23 0.76 0.44 0.90 0.80 1.05 0.77 0.46 0.90 0.81 1.01
SIS-SCAD 0.69 0.31 0.88 0.70 1.26 0.74 0.42 0.89 0.78 1.20 0.75 0.46 0.87 0.79 1.09
RuleFit 0.71 0.41 0.86 0.73 1.58 0.76 0.44 0.90 0.80 1.20 0.76 0.50 0.88 0.79 1.35
RE-Lasso-100 0.77 0.49 0.91 0.83 1.01 0.87 0.65 0.96 0.96 0.57 0.89 0.70 0.97 0.97 0.51
RE-EN-100 0.77 0.50 0.90 0.83 1.01 0.88 0.69 0.96 0.96 0.53 0.90 0.73 0.96 0.97 0.47
RGLM-100 0.74 0.27 0.96 0.82 1.06 0.81 0.40 0.99 0.96 0.85 0.83 0.45 0.99 0.97 0.83
RF-500 0.77 0.50 0.90 0.82 0.97 0.85 0.56 0.98 0.95 0.73 0.87 0.62 0.98 0.96 0.70
XGB 0.70 0.41 0.84 0.70 1.19 0.77 0.45 0.91 0.81 1.00 0.78 0.50 0.89 0.83 0.95
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.79 0.29 0.95 0.81 0.92 0.89 0.54 0.98 0.96 0.47 0.91 0.62 0.98 0.97 0.40
Split-EN-10 0.79 0.29 0.96 0.82 0.91 0.90 0.57 0.98 0.97 0.45 0.92 0.64 0.98 0.98 0.38
Lasso 0.78 0.24 0.95 0.75 1.04 0.85 0.40 0.97 0.90 0.68 0.86 0.43 0.97 0.91 0.64
Elastic Net 0.78 0.26 0.95 0.78 0.98 0.87 0.45 0.97 0.93 0.59 0.88 0.50 0.98 0.94 0.54
Adaptive 0.76 0.03 1.00 0.60 1.09 0.82 0.10 0.99 0.74 0.87 0.82 0.09 0.99 0.77 0.86
Relaxed 0.77 0.24 0.94 0.73 1.32 0.85 0.46 0.95 0.89 1.03 0.85 0.49 0.94 0.90 0.97
0.2 MCP 0.76 0.17 0.94 0.68 1.17 0.81 0.20 0.96 0.75 0.93 0.81 0.25 0.95 0.80 0.87
SIS-SCAD 0.75 0.16 0.94 0.70 1.17 0.81 0.17 0.96 0.77 0.93 0.81 0.23 0.95 0.78 0.88
RuleFit 0.76 0.27 0.91 0.69 1.50 0.80 0.30 0.93 0.74 1.28 0.80 0.35 0.92 0.75 1.35
RE-Lasso-100 0.80 0.35 0.94 0.83 0.91 0.89 0.51 0.98 0.96 0.48 0.90 0.54 0.99 0.97 0.43
RE-EN-100 0.80 0.35 0.94 0.83 0.91 0.90 0.56 0.98 0.96 0.44 0.91 0.57 0.99 0.97 0.40
RGLM-100 0.78 0.11 0.99 0.81 0.95 0.83 0.15 1.00 0.96 0.70 0.84 0.18 1.00 0.97 0.68
RF-500 0.80 0.31 0.95 0.82 0.87 0.86 0.32 0.99 0.95 0.61 0.87 0.37 1.00 0.96 0.59
XGB 0.76 0.26 0.91 0.71 1.07 0.82 0.29 0.95 0.82 0.83 0.83 0.32 0.95 0.83 0.80
Table 14: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.83 0.76 0.88 0.92 0.76 0.92 0.91 0.93 0.98 0.35 0.94 0.91 0.96 0.99 0.30
Split-EN-10 0.83 0.76 0.88 0.92 0.76 0.93 0.91 0.93 0.99 0.35 0.94 0.91 0.96 0.99 0.29
Lasso 0.81 0.73 0.87 0.90 0.85 0.88 0.86 0.90 0.96 0.52 0.90 0.85 0.93 0.97 0.47
Elastic Net 0.82 0.75 0.87 0.91 0.81 0.90 0.88 0.92 0.97 0.43 0.92 0.88 0.94 0.98 0.38
Adaptive 0.76 0.55 0.91 0.87 0.99 0.85 0.75 0.92 0.95 0.69 0.87 0.79 0.94 0.96 0.62
Relaxed 0.81 0.73 0.86 0.90 1.08 0.88 0.86 0.89 0.96 0.77 0.89 0.86 0.91 0.97 0.72
0.4 MCP 0.77 0.68 0.84 0.86 0.98 0.82 0.76 0.86 0.91 0.80 0.84 0.77 0.89 0.93 0.72
SIS-SCAD 0.76 0.64 0.85 0.85 1.05 0.79 0.70 0.85 0.89 0.89 0.81 0.71 0.88 0.90 0.82
RuleFit 0.77 0.68 0.84 0.84 1.33 0.80 0.72 0.86 0.88 1.17 0.81 0.71 0.88 0.90 1.04
RE-Lasso-100 0.83 0.77 0.88 0.92 0.76 0.92 0.89 0.93 0.98 0.39 0.93 0.89 0.96 0.99 0.34
RE-EN-100 0.83 0.77 0.88 0.92 0.76 0.92 0.90 0.93 0.98 0.37 0.94 0.90 0.96 0.99 0.32
RGLM-100 0.83 0.72 0.91 0.92 0.83 0.91 0.85 0.95 0.98 0.65 0.92 0.86 0.97 0.99 0.60
RF-500 0.83 0.76 0.89 0.92 0.74 0.91 0.87 0.94 0.98 0.56 0.92 0.87 0.96 0.99 0.53
XGB 0.78 0.70 0.84 0.87 0.95 0.83 0.77 0.87 0.92 0.77 0.84 0.75 0.90 0.93 0.73
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.84 0.69 0.92 0.92 0.72 0.93 0.83 0.97 0.99 0.33 0.94 0.88 0.97 0.99 0.28
Split-EN-10 0.85 0.69 0.92 0.92 0.71 0.93 0.84 0.97 0.99 0.32 0.94 0.89 0.97 0.99 0.27
Lasso 0.82 0.64 0.92 0.90 0.82 0.89 0.75 0.95 0.97 0.50 0.90 0.79 0.95 0.97 0.45
Elastic Net 0.83 0.66 0.92 0.91 0.77 0.91 0.79 0.96 0.98 0.42 0.92 0.84 0.96 0.98 0.35
Adaptive 0.77 0.37 0.96 0.89 0.93 0.84 0.54 0.98 0.95 0.68 0.86 0.63 0.97 0.96 0.61
Relaxed 0.82 0.65 0.91 0.90 0.99 0.88 0.77 0.93 0.96 0.96 0.89 0.79 0.93 0.96 0.95
0.3 MCP 0.78 0.52 0.91 0.85 0.96 0.82 0.57 0.93 0.89 0.84 0.83 0.61 0.92 0.90 0.82
SIS-SCAD 0.77 0.49 0.91 0.85 0.97 0.81 0.52 0.93 0.89 0.83 0.81 0.55 0.93 0.89 0.82
RuleFit 0.79 0.57 0.90 0.85 1.15 0.83 0.57 0.94 0.88 1.00 0.83 0.60 0.93 0.90 0.88
RE-Lasso-100 0.85 0.70 0.92 0.92 0.72 0.93 0.82 0.97 0.99 0.34 0.93 0.84 0.97 0.99 0.32
RE-EN-100 0.85 0.71 0.92 0.93 0.72 0.93 0.83 0.97 0.99 0.33 0.94 0.87 0.97 0.99 0.30
RGLM-100 0.84 0.61 0.95 0.92 0.77 0.90 0.69 0.99 0.99 0.58 0.91 0.73 0.99 0.99 0.56
RF-500 0.84 0.68 0.93 0.92 0.70 0.91 0.75 0.98 0.98 0.51 0.92 0.79 0.98 0.99 0.49
XGB 0.79 0.59 0.90 0.87 0.91 0.84 0.63 0.94 0.92 0.74 0.85 0.65 0.94 0.92 0.71
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.86 0.60 0.94 0.93 0.63 0.95 0.79 0.98 0.99 0.25 0.95 0.85 0.98 0.99 0.23
Split-EN-10 0.86 0.61 0.94 0.93 0.63 0.95 0.81 0.98 0.99 0.24 0.95 0.86 0.98 0.99 0.22
Lasso 0.84 0.53 0.94 0.91 0.72 0.91 0.66 0.97 0.97 0.42 0.92 0.74 0.96 0.97 0.38
Elastic Net 0.85 0.56 0.94 0.92 0.68 0.93 0.72 0.98 0.98 0.33 0.93 0.79 0.97 0.98 0.30
Adaptive 0.80 0.20 0.98 0.84 0.88 0.87 0.37 0.99 0.95 0.57 0.88 0.46 0.99 0.96 0.54
Relaxed 0.84 0.53 0.94 0.90 1.24 0.90 0.70 0.95 0.96 0.74 0.91 0.76 0.94 0.96 0.68
0.2 MCP 0.80 0.35 0.94 0.83 0.91 0.85 0.41 0.96 0.88 0.69 0.85 0.47 0.95 0.89 0.67
SIS-SCAD 0.79 0.31 0.94 0.84 0.95 0.84 0.32 0.97 0.89 0.72 0.85 0.41 0.95 0.89 0.73
RuleFit 0.81 0.47 0.91 0.80 1.34 0.86 0.47 0.95 0.85 0.99 0.86 0.52 0.95 0.87 0.97
RE-Lasso-100 0.87 0.66 0.93 0.93 0.62 0.94 0.78 0.98 0.99 0.27 0.95 0.83 0.98 0.99 0.26
RE-EN-100 0.87 0.66 0.93 0.93 0.62 0.95 0.81 0.98 0.99 0.25 0.95 0.84 0.98 0.99 0.25
RGLM-100 0.86 0.51 0.97 0.93 0.66 0.91 0.55 1.00 0.99 0.46 0.92 0.61 1.00 0.99 0.47
RF-500 0.87 0.61 0.95 0.93 0.61 0.92 0.63 0.99 0.99 0.41 0.93 0.71 0.99 0.99 0.41
XGB 0.82 0.49 0.92 0.86 0.84 0.87 0.51 0.95 0.91 0.66 0.87 0.56 0.94 0.92 0.66
Table 15: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.8, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.88 0.85 0.90 0.96 0.55 0.95 0.94 0.96 1.00 0.22 0.96 0.94 0.98 1.00 0.19
Split-EN-10 0.88 0.85 0.90 0.96 0.55 0.96 0.94 0.96 1.00 0.22 0.97 0.95 0.98 1.00 0.18
Lasso 0.87 0.84 0.89 0.95 0.61 0.93 0.90 0.95 0.99 0.32 0.94 0.91 0.96 0.99 0.28
Elastic Net 0.88 0.84 0.90 0.96 0.58 0.95 0.92 0.96 0.99 0.26 0.96 0.93 0.98 1.00 0.22
Adaptive 0.85 0.78 0.91 0.95 0.71 0.93 0.88 0.95 0.99 0.42 0.93 0.88 0.97 0.99 0.38
Relaxed 0.86 0.83 0.89 0.95 0.94 0.92 0.90 0.94 0.98 0.58 0.93 0.90 0.94 0.98 0.64
0.4 MCP 0.83 0.78 0.87 0.92 0.73 0.87 0.81 0.91 0.95 0.60 0.87 0.80 0.92 0.95 0.60
SIS-SCAD 0.84 0.77 0.88 0.93 0.74 0.88 0.82 0.92 0.96 0.59 0.88 0.80 0.94 0.96 0.59
RuleFit 0.84 0.80 0.88 0.91 1.01 0.88 0.83 0.91 0.91 1.01 0.88 0.83 0.91 0.90 1.17
RE-Lasso-100 0.88 0.85 0.90 0.96 0.56 0.95 0.93 0.96 0.99 0.25 0.96 0.94 0.98 1.00 0.22
RE-EN-100 0.88 0.86 0.90 0.96 0.56 0.95 0.94 0.96 1.00 0.23 0.96 0.94 0.98 1.00 0.21
RGLM-100 0.88 0.85 0.91 0.96 0.57 0.95 0.93 0.97 1.00 0.38 0.96 0.92 0.99 1.00 0.36
RF-500 0.88 0.85 0.90 0.96 0.54 0.95 0.93 0.96 0.99 0.35 0.96 0.93 0.98 1.00 0.32
XGB 0.84 0.80 0.87 0.91 0.80 0.87 0.83 0.90 0.90 0.75 0.87 0.83 0.89 0.89 0.80
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.89 0.81 0.93 0.96 0.51 0.95 0.92 0.97 0.99 0.21 0.96 0.94 0.97 1.00 0.19
Split-EN-10 0.89 0.81 0.93 0.96 0.51 0.96 0.93 0.97 1.00 0.21 0.97 0.95 0.97 1.00 0.18
Lasso 0.88 0.78 0.93 0.95 0.59 0.93 0.87 0.96 0.99 0.31 0.94 0.89 0.96 0.99 0.28
Elastic Net 0.89 0.80 0.93 0.96 0.54 0.95 0.90 0.97 0.99 0.25 0.95 0.91 0.97 0.99 0.22
Adaptive 0.85 0.62 0.96 0.95 0.69 0.92 0.80 0.97 0.98 0.41 0.93 0.82 0.97 0.99 0.41
Relaxed 0.88 0.78 0.92 0.95 0.86 0.92 0.88 0.94 0.98 0.53 0.93 0.89 0.94 0.98 0.77
0.3 MCP 0.84 0.69 0.91 0.92 0.71 0.87 0.72 0.93 0.94 0.60 0.87 0.74 0.93 0.95 0.59
SIS-SCAD 0.84 0.63 0.93 0.93 0.74 0.87 0.68 0.95 0.95 0.62 0.88 0.73 0.95 0.96 0.59
RuleFit 0.87 0.75 0.92 0.93 0.77 0.87 0.76 0.92 0.87 1.29 0.88 0.79 0.92 0.88 1.21
RE-Lasso-100 0.89 0.82 0.93 0.96 0.50 0.95 0.91 0.97 0.99 0.24 0.96 0.93 0.98 1.00 0.22
RE-EN-100 0.89 0.82 0.93 0.96 0.50 0.95 0.92 0.97 1.00 0.23 0.96 0.93 0.98 1.00 0.21
RGLM-100 0.89 0.80 0.94 0.96 0.52 0.95 0.88 0.98 1.00 0.35 0.96 0.90 0.99 1.00 0.35
RF-500 0.89 0.81 0.93 0.96 0.50 0.95 0.90 0.97 0.99 0.32 0.96 0.91 0.98 1.00 0.32
XGB 0.86 0.76 0.91 0.92 0.70 0.87 0.76 0.91 0.90 0.77 0.87 0.80 0.91 0.89 0.77
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.91 0.74 0.96 0.97 0.42 0.96 0.88 0.98 1.00 0.18 0.97 0.91 0.99 1.00 0.15
Split-EN-10 0.91 0.74 0.96 0.97 0.42 0.96 0.88 0.99 1.00 0.18 0.97 0.91 0.99 1.00 0.15
Lasso 0.90 0.69 0.96 0.96 0.49 0.94 0.80 0.98 0.99 0.28 0.95 0.82 0.98 0.99 0.24
Elastic Net 0.91 0.72 0.97 0.96 0.45 0.95 0.83 0.98 0.99 0.22 0.96 0.86 0.99 1.00 0.18
Adaptive 0.86 0.42 0.99 0.93 0.63 0.91 0.63 0.99 0.98 0.38 0.92 0.66 0.99 0.99 0.36
Relaxed 0.90 0.70 0.95 0.95 0.84 0.93 0.81 0.96 0.98 0.68 0.94 0.83 0.97 0.98 0.62
0.2 MCP 0.86 0.52 0.96 0.92 0.63 0.88 0.60 0.96 0.94 0.52 0.89 0.62 0.96 0.95 0.49
SIS-SCAD 0.85 0.44 0.97 0.93 0.65 0.88 0.54 0.97 0.95 0.54 0.89 0.55 0.98 0.96 0.51
RuleFit 0.87 0.66 0.94 0.85 1.19 0.88 0.71 0.93 0.82 1.30 0.90 0.71 0.94 0.83 1.16
RE-Lasso-100 0.91 0.76 0.96 0.97 0.41 0.96 0.87 0.98 1.00 0.20 0.97 0.91 0.99 1.00 0.17
RE-EN-100 0.91 0.77 0.96 0.97 0.41 0.96 0.88 0.98 1.00 0.19 0.98 0.92 0.99 1.00 0.15
RGLM-100 0.91 0.72 0.97 0.97 0.44 0.95 0.78 0.99 1.00 0.31 0.96 0.81 1.00 1.00 0.28
RF-500 0.91 0.74 0.97 0.97 0.43 0.95 0.81 0.99 0.99 0.29 0.96 0.83 0.99 1.00 0.26
XGB 0.87 0.68 0.93 0.91 0.69 0.88 0.71 0.93 0.93 0.65 0.90 0.71 0.94 0.94 0.58
Table 16: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.75 0.66 0.82 0.83 1.02 0.88 0.84 0.91 0.96 0.52 0.91 0.88 0.93 0.97 0.42
Split-EN-10 0.75 0.66 0.82 0.83 1.02 0.88 0.84 0.91 0.96 0.52 0.91 0.88 0.93 0.97 0.41
Lasso 0.72 0.61 0.80 0.80 1.10 0.84 0.77 0.89 0.93 0.71 0.86 0.81 0.90 0.94 0.64
Elastic Net 0.73 0.62 0.81 0.80 1.09 0.85 0.79 0.89 0.94 0.65 0.88 0.83 0.91 0.95 0.57
Adaptive 0.67 0.39 0.88 0.74 1.21 0.82 0.69 0.90 0.91 0.83 0.84 0.74 0.91 0.93 0.76
Relaxed 0.72 0.61 0.80 0.80 1.12 0.84 0.77 0.88 0.92 0.76 0.86 0.81 0.89 0.94 0.71
0.4 MCP 0.69 0.55 0.79 0.75 1.18 0.79 0.68 0.86 0.87 0.90 0.79 0.71 0.85 0.88 0.90
SIS-SCAD 0.68 0.56 0.77 0.74 1.21 0.75 0.65 0.82 0.82 1.06 0.75 0.66 0.82 0.83 1.05
RuleFit 0.71 0.61 0.78 0.77 1.64 0.79 0.68 0.87 0.88 1.01 0.80 0.69 0.87 0.88 0.97
RE-Lasso-100 0.75 0.67 0.81 0.83 1.04 0.88 0.82 0.92 0.96 0.54 0.90 0.85 0.94 0.97 0.45
RE-EN-100 0.76 0.67 0.82 0.83 1.05 0.88 0.82 0.92 0.96 0.52 0.91 0.86 0.94 0.97 0.44
RGLM-100 0.75 0.59 0.86 0.83 1.10 0.86 0.72 0.95 0.96 0.88 0.89 0.77 0.97 0.97 0.85
RF-500 0.76 0.69 0.80 0.83 1.02 0.87 0.80 0.92 0.95 0.68 0.90 0.83 0.94 0.97 0.65
XGB 0.68 0.59 0.75 0.73 1.21 0.79 0.68 0.86 0.88 0.88 0.80 0.70 0.87 0.89 0.84
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.77 0.52 0.89 0.83 0.96 0.90 0.79 0.94 0.96 0.46 0.91 0.79 0.96 0.98 0.39
Split-EN-10 0.77 0.51 0.89 0.83 0.96 0.90 0.79 0.94 0.97 0.46 0.92 0.80 0.96 0.98 0.38
Lasso 0.75 0.47 0.88 0.80 1.05 0.86 0.70 0.92 0.93 0.65 0.87 0.69 0.95 0.94 0.59
Elastic Net 0.75 0.49 0.88 0.81 1.03 0.87 0.73 0.93 0.94 0.58 0.88 0.73 0.95 0.96 0.53
Adaptive 0.70 0.17 0.96 0.72 1.15 0.83 0.53 0.96 0.92 0.75 0.84 0.53 0.97 0.93 0.71
Relaxed 0.74 0.47 0.88 0.79 1.06 0.86 0.72 0.92 0.93 0.73 0.87 0.70 0.94 0.94 0.67
0.3 MCP 0.72 0.40 0.88 0.76 1.11 0.80 0.55 0.91 0.87 0.85 0.81 0.53 0.93 0.87 0.87
SIS-SCAD 0.71 0.41 0.86 0.74 1.16 0.78 0.53 0.88 0.83 0.95 0.79 0.52 0.90 0.84 0.96
RuleFit 0.73 0.49 0.85 0.77 1.60 0.81 0.53 0.93 0.87 0.99 0.82 0.52 0.95 0.88 0.97
RE-Lasso-100 0.77 0.54 0.88 0.83 0.99 0.89 0.74 0.95 0.96 0.49 0.90 0.73 0.98 0.97 0.44
RE-EN-100 0.77 0.55 0.88 0.84 1.00 0.89 0.76 0.95 0.96 0.47 0.91 0.75 0.97 0.98 0.42
RGLM-100 0.75 0.37 0.94 0.83 1.03 0.84 0.50 0.99 0.96 0.79 0.84 0.46 1.00 0.97 0.77
RF-500 0.77 0.58 0.86 0.83 0.96 0.88 0.69 0.96 0.96 0.62 0.89 0.66 0.98 0.97 0.60
XGB 0.71 0.46 0.84 0.74 1.15 0.81 0.53 0.92 0.87 0.83 0.82 0.52 0.95 0.89 0.79
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.80 0.36 0.94 0.83 0.84 0.92 0.69 0.97 0.97 0.37 0.92 0.70 0.98 0.98 0.34
Split-EN-10 0.80 0.36 0.94 0.84 0.84 0.92 0.70 0.97 0.97 0.36 0.93 0.72 0.98 0.98 0.33
Lasso 0.79 0.30 0.94 0.80 0.92 0.89 0.57 0.96 0.93 0.52 0.89 0.58 0.97 0.94 0.52
Elastic Net 0.79 0.33 0.94 0.81 0.90 0.90 0.62 0.96 0.95 0.47 0.90 0.61 0.97 0.96 0.46
Adaptive 0.77 0.04 1.00 0.69 1.03 0.86 0.28 0.99 0.90 0.65 0.85 0.27 0.99 0.90 0.66
Relaxed 0.79 0.31 0.94 0.79 0.97 0.89 0.60 0.95 0.93 0.61 0.89 0.61 0.95 0.94 0.64
0.2 MCP 0.78 0.24 0.94 0.75 0.98 0.84 0.34 0.96 0.84 0.75 0.83 0.35 0.95 0.85 0.76
SIS-SCAD 0.77 0.24 0.93 0.74 1.02 0.83 0.36 0.94 0.83 0.83 0.83 0.34 0.95 0.83 0.84
RuleFit 0.78 0.33 0.92 0.76 1.48 0.86 0.37 0.97 0.86 0.84 0.85 0.34 0.97 0.87 0.95
RE-Lasso-100 0.81 0.38 0.94 0.84 0.88 0.91 0.59 0.98 0.97 0.41 0.91 0.57 0.99 0.97 0.40
RE-EN-100 0.81 0.39 0.94 0.84 0.88 0.91 0.62 0.98 0.97 0.39 0.91 0.61 0.99 0.98 0.37
RGLM-100 0.79 0.18 0.98 0.83 0.90 0.86 0.27 1.00 0.97 0.62 0.85 0.23 1.00 0.98 0.63
RF-500 0.81 0.45 0.92 0.84 0.86 0.90 0.50 0.99 0.96 0.50 0.89 0.46 0.99 0.97 0.51
XGB 0.77 0.31 0.92 0.75 0.99 0.86 0.36 0.97 0.88 0.67 0.85 0.34 0.97 0.88 0.70
Table 17: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.84 0.80 0.87 0.92 0.72 0.93 0.91 0.94 0.99 0.31 0.95 0.93 0.96 0.99 0.26
Split-EN-10 0.84 0.80 0.87 0.92 0.72 0.93 0.91 0.94 0.99 0.30 0.95 0.93 0.96 0.99 0.25
Lasso 0.82 0.77 0.86 0.91 0.78 0.91 0.88 0.92 0.97 0.43 0.92 0.88 0.94 0.98 0.39
Elastic Net 0.83 0.78 0.86 0.91 0.76 0.92 0.89 0.93 0.98 0.38 0.93 0.90 0.95 0.98 0.33
Adaptive 0.81 0.71 0.88 0.90 0.88 0.90 0.84 0.93 0.97 0.53 0.90 0.85 0.95 0.97 0.50
Relaxed 0.82 0.77 0.85 0.90 0.84 0.90 0.88 0.92 0.97 0.51 0.91 0.88 0.93 0.97 0.49
0.4 MCP 0.80 0.75 0.84 0.89 0.87 0.86 0.82 0.89 0.94 0.63 0.86 0.81 0.90 0.94 0.63
SIS-SCAD 0.79 0.72 0.83 0.87 0.97 0.84 0.78 0.88 0.93 0.70 0.84 0.78 0.89 0.93 0.71
RuleFit 0.81 0.75 0.85 0.89 1.11 0.86 0.80 0.91 0.95 0.67 0.87 0.79 0.92 0.95 0.66
RE-Lasso-100 0.84 0.80 0.87 0.92 0.74 0.93 0.90 0.95 0.99 0.33 0.94 0.91 0.96 0.99 0.29
RE-EN-100 0.84 0.80 0.87 0.92 0.75 0.93 0.91 0.95 0.99 0.32 0.94 0.92 0.96 0.99 0.28
RGLM-100 0.84 0.79 0.88 0.92 0.79 0.93 0.87 0.96 0.99 0.57 0.94 0.89 0.98 0.99 0.56
RF-500 0.84 0.80 0.86 0.92 0.74 0.93 0.89 0.95 0.98 0.45 0.94 0.90 0.96 0.99 0.45
XGB 0.81 0.75 0.85 0.89 0.86 0.87 0.82 0.91 0.95 0.57 0.88 0.82 0.92 0.96 0.54
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.73 0.91 0.92 0.69 0.93 0.89 0.96 0.99 0.29 0.96 0.90 0.98 0.99 0.22
Split-EN-10 0.85 0.73 0.91 0.92 0.68 0.94 0.89 0.96 0.99 0.29 0.96 0.91 0.98 0.99 0.21
Lasso 0.84 0.70 0.90 0.91 0.74 0.91 0.83 0.94 0.97 0.41 0.93 0.85 0.96 0.98 0.34
Elastic Net 0.84 0.71 0.91 0.91 0.72 0.92 0.85 0.95 0.98 0.36 0.94 0.87 0.97 0.99 0.29
Adaptive 0.81 0.53 0.95 0.90 0.84 0.89 0.74 0.96 0.97 0.52 0.91 0.77 0.97 0.98 0.45
Relaxed 0.83 0.70 0.90 0.91 0.81 0.90 0.83 0.93 0.97 0.53 0.92 0.86 0.95 0.98 0.46
0.3 MCP 0.81 0.65 0.89 0.88 0.83 0.86 0.71 0.92 0.93 0.65 0.87 0.72 0.94 0.94 0.60
SIS-SCAD 0.80 0.63 0.88 0.87 0.94 0.84 0.67 0.92 0.92 0.72 0.86 0.70 0.93 0.93 0.63
RuleFit 0.82 0.67 0.89 0.89 1.07 0.86 0.70 0.94 0.94 0.69 0.88 0.72 0.96 0.96 0.58
RE-Lasso-100 0.85 0.74 0.91 0.92 0.71 0.93 0.86 0.96 0.99 0.31 0.95 0.88 0.98 0.99 0.26
RE-EN-100 0.85 0.74 0.91 0.92 0.71 0.93 0.87 0.96 0.99 0.30 0.95 0.89 0.98 0.99 0.24
RGLM-100 0.85 0.70 0.93 0.92 0.73 0.92 0.78 0.98 0.99 0.53 0.93 0.80 0.99 0.99 0.50
RF-500 0.85 0.75 0.90 0.92 0.70 0.93 0.83 0.97 0.98 0.43 0.94 0.84 0.99 0.99 0.40
XGB 0.82 0.68 0.89 0.89 0.81 0.87 0.71 0.94 0.95 0.57 0.89 0.74 0.96 0.96 0.51
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.87 0.62 0.95 0.93 0.59 0.95 0.84 0.98 0.99 0.22 0.95 0.86 0.98 0.99 0.21
Split-EN-10 0.87 0.62 0.95 0.93 0.59 0.95 0.84 0.98 0.99 0.22 0.96 0.86 0.98 0.99 0.20
Lasso 0.86 0.57 0.95 0.91 0.65 0.93 0.74 0.97 0.98 0.34 0.93 0.77 0.97 0.98 0.33
Elastic Net 0.86 0.59 0.94 0.92 0.63 0.94 0.78 0.98 0.98 0.29 0.94 0.80 0.98 0.99 0.27
Adaptive 0.82 0.29 0.98 0.90 0.75 0.91 0.58 0.98 0.97 0.43 0.91 0.62 0.98 0.97 0.43
Relaxed 0.86 0.57 0.94 0.91 0.70 0.92 0.76 0.96 0.97 0.48 0.92 0.77 0.96 0.97 0.53
0.2 MCP 0.83 0.44 0.94 0.87 0.80 0.87 0.50 0.95 0.91 0.79 0.86 0.55 0.95 0.91 0.65
SIS-SCAD 0.82 0.38 0.95 0.87 0.85 0.86 0.40 0.97 0.92 0.63 0.86 0.47 0.97 0.93 0.64
RuleFit 0.84 0.52 0.94 0.88 0.98 0.88 0.54 0.96 0.89 0.74 0.88 0.55 0.97 0.93 0.63
RE-Lasso-100 0.87 0.63 0.94 0.93 0.62 0.94 0.77 0.98 0.99 0.26 0.95 0.80 0.99 0.99 0.24
RE-EN-100 0.87 0.63 0.94 0.93 0.62 0.95 0.79 0.98 0.99 0.24 0.95 0.82 0.99 0.99 0.23
RGLM-100 0.87 0.54 0.96 0.93 0.62 0.92 0.62 1.00 0.99 0.43 0.92 0.65 1.00 0.99 0.43
RF-500 0.87 0.63 0.94 0.93 0.62 0.94 0.70 0.99 0.99 0.36 0.94 0.73 0.99 0.99 0.36
XGB 0.85 0.53 0.94 0.89 0.72 0.90 0.57 0.97 0.95 0.49 0.90 0.61 0.97 0.96 0.47
Table 18: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.8, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.89 0.87 0.91 0.96 0.50 0.96 0.95 0.97 1.00 0.19 0.96 0.96 0.97 1.00 0.18
Split-EN-10 0.89 0.87 0.91 0.96 0.50 0.96 0.95 0.97 1.00 0.19 0.97 0.96 0.97 1.00 0.18
Lasso 0.88 0.85 0.91 0.96 0.54 0.94 0.92 0.96 0.99 0.27 0.94 0.93 0.95 0.99 0.25
Elastic Net 0.89 0.86 0.91 0.96 0.52 0.95 0.93 0.97 0.99 0.23 0.96 0.94 0.96 0.99 0.22
Adaptive 0.88 0.81 0.92 0.96 0.64 0.94 0.90 0.96 0.99 0.34 0.94 0.92 0.96 0.99 0.33
Relaxed 0.88 0.85 0.91 0.96 0.55 0.93 0.92 0.95 0.99 0.47 0.94 0.93 0.94 0.99 0.39
0.4 MCP 0.86 0.82 0.90 0.94 0.63 0.90 0.86 0.93 0.97 0.46 0.90 0.86 0.92 0.97 0.47
SIS-SCAD 0.86 0.81 0.90 0.95 0.64 0.90 0.86 0.93 0.97 0.45 0.90 0.87 0.93 0.97 0.46
RuleFit 0.88 0.84 0.90 0.95 0.74 0.90 0.86 0.92 0.94 0.74 0.89 0.86 0.92 0.95 0.70
RE-Lasso-100 0.89 0.87 0.91 0.96 0.52 0.96 0.94 0.97 0.99 0.21 0.96 0.95 0.97 1.00 0.21
RE-EN-100 0.89 0.87 0.91 0.96 0.52 0.96 0.94 0.97 1.00 0.21 0.96 0.95 0.97 1.00 0.20
RGLM-100 0.89 0.87 0.91 0.96 0.53 0.96 0.93 0.98 1.00 0.33 0.96 0.95 0.97 1.00 0.34
RF-500 0.89 0.87 0.91 0.96 0.54 0.96 0.94 0.97 0.99 0.29 0.96 0.95 0.97 1.00 0.30
XGB 0.88 0.85 0.90 0.95 0.59 0.92 0.88 0.94 0.98 0.39 0.92 0.90 0.94 0.98 0.36
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.90 0.84 0.93 0.97 0.46 0.96 0.94 0.97 1.00 0.17 0.97 0.95 0.98 1.00 0.17
Split-EN-10 0.90 0.84 0.93 0.97 0.46 0.97 0.94 0.97 1.00 0.17 0.97 0.95 0.98 1.00 0.17
Lasso 0.89 0.81 0.93 0.96 0.50 0.95 0.90 0.97 0.99 0.23 0.95 0.90 0.97 0.99 0.24
Elastic Net 0.90 0.82 0.93 0.96 0.48 0.96 0.91 0.97 0.99 0.20 0.96 0.92 0.98 0.99 0.20
Adaptive 0.88 0.71 0.96 0.96 0.60 0.94 0.86 0.98 0.99 0.32 0.94 0.87 0.98 0.99 0.31
Relaxed 0.89 0.81 0.93 0.96 0.59 0.94 0.90 0.96 0.99 0.34 0.94 0.91 0.96 0.99 0.37
0.3 MCP 0.86 0.73 0.92 0.93 0.64 0.89 0.77 0.94 0.96 0.49 0.88 0.76 0.93 0.95 0.54
SIS-SCAD 0.86 0.68 0.94 0.94 0.66 0.90 0.77 0.96 0.97 0.50 0.89 0.75 0.96 0.97 0.54
RuleFit 0.88 0.79 0.92 0.94 0.75 0.92 0.81 0.96 0.96 0.52 0.91 0.80 0.96 0.95 0.56
RE-Lasso-100 0.90 0.84 0.93 0.97 0.47 0.96 0.92 0.98 1.00 0.19 0.96 0.93 0.98 1.00 0.19
RE-EN-100 0.90 0.84 0.93 0.97 0.47 0.96 0.93 0.98 1.00 0.18 0.97 0.93 0.98 1.00 0.19
RGLM-100 0.90 0.83 0.93 0.96 0.48 0.96 0.90 0.98 1.00 0.30 0.96 0.91 0.99 1.00 0.31
RF-500 0.90 0.84 0.93 0.96 0.47 0.96 0.92 0.98 1.00 0.26 0.96 0.92 0.98 1.00 0.28
XGB 0.88 0.80 0.92 0.95 0.56 0.93 0.85 0.97 0.98 0.33 0.92 0.84 0.96 0.98 0.37
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.91 0.76 0.96 0.97 0.41 0.97 0.92 0.98 1.00 0.16 0.97 0.93 0.98 1.00 0.15
Split-EN-10 0.91 0.76 0.96 0.97 0.40 0.97 0.92 0.98 1.00 0.16 0.97 0.93 0.98 1.00 0.15
Lasso 0.91 0.73 0.96 0.96 0.46 0.95 0.86 0.98 0.99 0.22 0.95 0.86 0.98 0.99 0.21
Elastic Net 0.91 0.74 0.96 0.96 0.43 0.96 0.88 0.98 0.99 0.18 0.96 0.88 0.98 1.00 0.17
Adaptive 0.88 0.56 0.98 0.96 0.52 0.94 0.78 0.99 0.99 0.29 0.94 0.78 0.99 0.99 0.28
Relaxed 0.90 0.72 0.96 0.96 0.53 0.95 0.85 0.97 0.99 0.36 0.94 0.86 0.97 0.99 0.44
0.2 MCP 0.87 0.60 0.95 0.92 0.66 0.89 0.67 0.95 0.95 0.50 0.89 0.66 0.95 0.94 0.50
SIS-SCAD 0.86 0.44 0.98 0.94 0.63 0.89 0.57 0.97 0.96 0.51 0.88 0.52 0.98 0.96 0.53
RuleFit 0.89 0.70 0.95 0.93 0.67 0.92 0.71 0.98 0.96 0.49 0.92 0.73 0.97 0.95 0.50
RE-Lasso-100 0.91 0.76 0.96 0.97 0.42 0.97 0.89 0.99 1.00 0.17 0.97 0.89 0.99 1.00 0.16
RE-EN-100 0.91 0.77 0.96 0.97 0.42 0.97 0.90 0.99 1.00 0.16 0.97 0.90 0.99 1.00 0.16
RGLM-100 0.91 0.74 0.96 0.97 0.42 0.96 0.85 0.99 1.00 0.26 0.96 0.85 0.99 1.00 0.26
RF-500 0.91 0.76 0.96 0.97 0.44 0.96 0.88 0.99 1.00 0.23 0.96 0.87 0.99 1.00 0.23
XGB 0.90 0.71 0.95 0.95 0.50 0.93 0.76 0.98 0.98 0.35 0.93 0.78 0.97 0.98 0.34
Table 19: Mean recall (RCL) and precision (PRC) for Scenario 1, $\rho=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.18 0.14 0.29 0.25 0.28 0.45
Split-EN-10 0.22 0.14 0.35 0.24 0.36 0.44
Lasso 0.02 0.15 0.03 0.24 0.03 0.43
Elastic Net 0.04 0.15 0.05 0.24 0.05 0.43
Adaptive 0.02 0.16 0.03 0.25 0.03 0.42
Relaxed 0.02 0.16 0.03 0.23 0.02 0.42
0.4 MCP 0.01 0.14 0.01 0.27 0.01 0.44
SIS-SCAD 0.01 0.19 0.00 0.22 0.00 0.46
RuleFit 0.09 0.16 0.07 0.31 0.06 0.48
RE-Lasso-100 0.44 0.11 0.42 0.21 0.41 0.41
RE-EN-100 0.57 0.10 0.57 0.21 0.56 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.23 0.12 0.17 0.21 0.16 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.16 0.15 0.27 0.25 0.27 0.46
Split-EN-10 0.21 0.15 0.35 0.24 0.35 0.45
Lasso 0.02 0.13 0.03 0.25 0.03 0.43
Elastic Net 0.03 0.15 0.05 0.23 0.05 0.42
Adaptive 0.02 0.15 0.03 0.24 0.03 0.43
Relaxed 0.02 0.15 0.03 0.26 0.02 0.44
0.3 MCP 0.01 0.12 0.01 0.26 0.01 0.45
SIS-SCAD 0.00 0.17 0.00 0.27 0.00 0.50
RuleFit 0.09 0.18 0.07 0.34 0.06 0.51
RE-Lasso-100 0.42 0.11 0.40 0.21 0.39 0.41
RE-EN-100 0.55 0.11 0.55 0.21 0.52 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.21 0.11 0.15 0.21 0.15 0.40
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.14 0.15 0.26 0.27 0.24 0.47
Split-EN-10 0.18 0.14 0.33 0.25 0.32 0.46
Lasso 0.02 0.14 0.02 0.24 0.02 0.40
Elastic Net 0.03 0.14 0.04 0.24 0.04 0.40
Adaptive 0.02 0.14 0.02 0.25 0.02 0.41
Relaxed 0.02 0.12 0.02 0.22 0.02 0.42
0.2 MCP 0.01 0.18 0.00 0.26 0.00 0.41
SIS-SCAD 0.00 0.18 0.00 0.24 0.00 0.37
RuleFit 0.08 0.21 0.06 0.39 0.04 0.54
RE-Lasso-100 0.39 0.11 0.37 0.22 0.35 0.40
RE-EN-100 0.51 0.10 0.49 0.21 0.49 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.17 0.11 0.13 0.22 0.13 0.42
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 20: Mean recall (RCL) and precision (PRC) for Scenario 1, $\rho=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.17 0.14 0.23 0.25 0.22 0.46
Split-EN-10 0.20 0.13 0.31 0.23 0.30 0.45
Lasso 0.02 0.12 0.02 0.21 0.02 0.41
Elastic Net 0.04 0.13 0.05 0.21 0.05 0.41
Adaptive 0.02 0.13 0.02 0.20 0.02 0.40
Relaxed 0.02 0.12 0.02 0.19 0.02 0.41
0.4 MCP 0.01 0.12 0.00 0.15 0.01 0.49
SIS-SCAD 0.00 0.08 0.00 0.19 0.00 0.43
RuleFit 0.07 0.18 0.05 0.34 0.04 0.52
RE-Lasso-100 0.37 0.11 0.36 0.21 0.36 0.41
RE-EN-100 0.49 0.10 0.51 0.20 0.52 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.16 0.11 0.12 0.20 0.12 0.40
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.16 0.16 0.25 0.28 0.23 0.50
Split-EN-10 0.20 0.14 0.32 0.26 0.32 0.47
Lasso 0.02 0.12 0.02 0.22 0.02 0.41
Elastic Net 0.04 0.13 0.04 0.22 0.05 0.42
Adaptive 0.02 0.13 0.02 0.24 0.02 0.42
Relaxed 0.02 0.14 0.02 0.22 0.02 0.42
0.3 MCP 0.00 0.14 0.00 0.21 0.00 0.38
SIS-SCAD 0.00 0.13 0.00 0.22 0.00 0.38
RuleFit 0.08 0.23 0.05 0.36 0.04 0.55
RE-Lasso-100 0.34 0.11 0.35 0.21 0.34 0.41
RE-EN-100 0.49 0.11 0.50 0.21 0.50 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.16 0.12 0.13 0.21 0.12 0.40
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.17 0.16 0.25 0.29 0.24 0.52
Split-EN-10 0.21 0.15 0.34 0.27 0.34 0.49
Lasso 0.02 0.14 0.02 0.21 0.02 0.44
Elastic Net 0.03 0.13 0.04 0.21 0.04 0.43
Adaptive 0.02 0.13 0.02 0.22 0.02 0.43
Relaxed 0.02 0.14 0.01 0.21 0.02 0.43
0.2 MCP 0.00 0.12 0.00 0.26 0.00 0.40
SIS-SCAD 0.00 0.11 0.00 0.25 0.00 0.40
RuleFit 0.06 0.22 0.04 0.33 0.03 0.56
RE-Lasso-100 0.32 0.11 0.32 0.21 0.32 0.41
RE-EN-100 0.46 0.11 0.48 0.21 0.49 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.14 0.11 0.11 0.21 0.11 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 21: Mean recall (RCL) and precision (PRC) for Scenario 1, $\rho=$ 0.8, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.13 0.13 0.20 0.27 0.18 0.50
Split-EN-10 0.19 0.13 0.32 0.24 0.30 0.46
Lasso 0.01 0.11 0.01 0.18 0.02 0.39
Elastic Net 0.03 0.11 0.04 0.21 0.04 0.40
Adaptive 0.01 0.11 0.02 0.19 0.02 0.38
Relaxed 0.01 0.11 0.01 0.17 0.01 0.40
0.4 MCP 0.00 0.17 0.00 0.19 0.00 0.58
SIS-SCAD 0.00 0.16 0.00 0.17 0.00 0.41
RuleFit 0.04 0.24 0.02 0.48 0.01 0.72
RE-Lasso-100 0.27 0.11 0.28 0.21 0.27 0.40
RE-EN-100 0.45 0.10 0.47 0.20 0.47 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.12 0.11 0.10 0.21 0.10 0.40
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.14 0.14 0.23 0.31 0.21 0.53
Split-EN-10 0.16 0.14 0.33 0.27 0.33 0.50
Lasso 0.01 0.11 0.02 0.21 0.02 0.39
Elastic Net 0.02 0.10 0.04 0.22 0.04 0.40
Adaptive 0.01 0.11 0.01 0.20 0.02 0.41
Relaxed 0.01 0.11 0.01 0.22 0.01 0.40
0.3 MCP 0.00 0.15 0.00 0.24 0.00 0.44
SIS-SCAD 0.00 0.06 0.00 0.19 0.00 0.36
RuleFit 0.04 0.20 0.02 0.51 0.01 0.64
RE-Lasso-100 0.25 0.10 0.26 0.21 0.27 0.41
RE-EN-100 0.44 0.10 0.46 0.20 0.47 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.12 0.10 0.10 0.20 0.10 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.22 0.19 0.28 0.37 0.25 0.60
Split-EN-10 0.28 0.16 0.39 0.32 0.36 0.55
Lasso 0.01 0.14 0.01 0.21 0.01 0.41
Elastic Net 0.03 0.13 0.04 0.20 0.04 0.40
Adaptive 0.01 0.13 0.01 0.22 0.01 0.42
Relaxed 0.01 0.12 0.01 0.25 0.01 0.47
0.2 MCP 0.00 0.17 0.00 0.32 0.00 0.52
SIS-SCAD 0.00 0.08 0.00 0.25 0.00 0.53
RuleFit 0.04 0.23 0.03 0.47 0.02 0.64
RE-Lasso-100 0.23 0.11 0.24 0.21 0.23 0.41
RE-EN-100 0.43 0.10 0.45 0.21 0.44 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.11 0.11 0.08 0.20 0.08 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 22: Mean recall (RCL) and precision (PRC) for Scenario 1, $\rho=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.32 0.14 0.41 0.24 0.42 0.44
Split-EN-10 0.32 0.14 0.47 0.24 0.48 0.43
Lasso 0.04 0.17 0.05 0.27 0.05 0.43
Elastic Net 0.06 0.17 0.08 0.26 0.07 0.42
Adaptive 0.04 0.17 0.05 0.27 0.05 0.44
Relaxed 0.04 0.17 0.05 0.27 0.05 0.44
0.4 MCP 0.02 0.19 0.02 0.28 0.01 0.43
SIS-SCAD 0.01 0.23 0.01 0.26 0.01 0.44
RuleFit 0.16 0.15 0.13 0.28 0.11 0.46
RE-Lasso-100 0.65 0.11 0.56 0.21 0.54 0.41
RE-EN-100 0.75 0.11 0.66 0.21 0.65 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.43 0.11 0.28 0.22 0.26 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.26 0.15 0.40 0.24 0.40 0.45
Split-EN-10 0.28 0.14 0.47 0.24 0.48 0.44
Lasso 0.04 0.17 0.05 0.26 0.05 0.45
Elastic Net 0.06 0.17 0.07 0.25 0.07 0.45
Adaptive 0.04 0.17 0.05 0.25 0.04 0.44
Relaxed 0.04 0.18 0.05 0.26 0.04 0.45
0.3 MCP 0.02 0.19 0.02 0.27 0.01 0.44
SIS-SCAD 0.01 0.22 0.01 0.26 0.01 0.42
RuleFit 0.16 0.15 0.12 0.27 0.10 0.48
RE-Lasso-100 0.62 0.11 0.53 0.21 0.51 0.41
RE-EN-100 0.73 0.11 0.63 0.21 0.62 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.40 0.11 0.26 0.22 0.24 0.42
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.25 0.15 0.41 0.26 0.37 0.47
Split-EN-10 0.28 0.14 0.48 0.24 0.45 0.46
Lasso 0.04 0.18 0.04 0.25 0.04 0.43
Elastic Net 0.05 0.18 0.07 0.25 0.06 0.43
Adaptive 0.04 0.19 0.04 0.24 0.04 0.43
Relaxed 0.04 0.20 0.04 0.24 0.04 0.44
0.2 MCP 0.02 0.21 0.01 0.25 0.01 0.47
SIS-SCAD 0.01 0.20 0.01 0.26 0.01 0.46
RuleFit 0.14 0.15 0.10 0.29 0.08 0.48
RE-Lasso-100 0.57 0.11 0.47 0.21 0.46 0.40
RE-EN-100 0.68 0.11 0.58 0.21 0.56 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.34 0.11 0.21 0.22 0.20 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 23: Mean recall (RCL) and precision (PRC) for Scenario 1, $\rho=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.20 0.14 0.33 0.25 0.33 0.46
Split-EN-10 0.25 0.14 0.39 0.24 0.41 0.44
Lasso 0.04 0.15 0.04 0.23 0.04 0.41
Elastic Net 0.05 0.15 0.07 0.23 0.06 0.41
Adaptive 0.04 0.15 0.04 0.23 0.04 0.41
Relaxed 0.03 0.18 0.04 0.24 0.03 0.43
0.4 MCP 0.01 0.17 0.01 0.22 0.01 0.42
SIS-SCAD 0.01 0.20 0.01 0.21 0.00 0.40
RuleFit 0.14 0.17 0.10 0.28 0.08 0.46
RE-Lasso-100 0.51 0.11 0.46 0.21 0.46 0.41
RE-EN-100 0.65 0.11 0.58 0.21 0.58 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.32 0.11 0.20 0.22 0.18 0.40
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.22 0.15 0.36 0.27 0.34 0.48
Split-EN-10 0.25 0.15 0.42 0.25 0.42 0.46
Lasso 0.03 0.15 0.04 0.23 0.03 0.41
Elastic Net 0.05 0.14 0.06 0.24 0.06 0.41
Adaptive 0.03 0.15 0.04 0.23 0.03 0.42
Relaxed 0.03 0.16 0.03 0.24 0.03 0.41
0.3 MCP 0.01 0.16 0.01 0.22 0.01 0.40
SIS-SCAD 0.01 0.18 0.01 0.22 0.00 0.40
RuleFit 0.13 0.17 0.09 0.29 0.07 0.46
RE-Lasso-100 0.50 0.11 0.44 0.21 0.43 0.40
RE-EN-100 0.62 0.10 0.57 0.21 0.56 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.29 0.11 0.19 0.22 0.17 0.40
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.26 0.16 0.38 0.29 0.35 0.52
Split-EN-10 0.28 0.15 0.47 0.27 0.43 0.49
Lasso 0.03 0.15 0.03 0.23 0.03 0.43
Elastic Net 0.05 0.14 0.05 0.22 0.05 0.42
Adaptive 0.03 0.14 0.03 0.23 0.03 0.41
Relaxed 0.03 0.15 0.03 0.24 0.03 0.42
0.2 MCP 0.01 0.18 0.00 0.19 0.00 0.33
SIS-SCAD 0.01 0.20 0.00 0.20 0.00 0.36
RuleFit 0.10 0.17 0.08 0.32 0.07 0.50
RE-Lasso-100 0.43 0.11 0.39 0.20 0.40 0.40
RE-EN-100 0.57 0.10 0.52 0.20 0.53 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.25 0.11 0.15 0.21 0.15 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 24: Mean recall (RCL) and precision (PRC) for Scenario 1, $\rho=$ 0.8, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.17 0.14 0.27 0.26 0.26 0.48
Split-EN-10 0.22 0.13 0.37 0.24 0.38 0.45
Lasso 0.02 0.13 0.02 0.21 0.03 0.41
Elastic Net 0.04 0.13 0.05 0.21 0.05 0.41
Adaptive 0.02 0.14 0.02 0.21 0.03 0.40
Relaxed 0.02 0.13 0.02 0.20 0.02 0.41
0.4 MCP 0.01 0.13 0.00 0.16 0.00 0.42
SIS-SCAD 0.01 0.17 0.00 0.16 0.00 0.41
RuleFit 0.09 0.16 0.07 0.36 0.06 0.52
RE-Lasso-100 0.39 0.11 0.35 0.21 0.36 0.40
RE-EN-100 0.57 0.10 0.53 0.20 0.54 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.24 0.11 0.14 0.20 0.14 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.21 0.16 0.35 0.32 0.29 0.55
Split-EN-10 0.27 0.14 0.44 0.29 0.41 0.52
Lasso 0.02 0.13 0.03 0.22 0.02 0.41
Elastic Net 0.03 0.12 0.06 0.22 0.05 0.42
Adaptive 0.02 0.13 0.02 0.21 0.02 0.42
Relaxed 0.02 0.14 0.02 0.22 0.02 0.40
0.3 MCP 0.00 0.14 0.00 0.22 0.00 0.44
SIS-SCAD 0.01 0.15 0.00 0.23 0.00 0.35
RuleFit 0.08 0.17 0.07 0.36 0.05 0.55
RE-Lasso-100 0.38 0.11 0.34 0.20 0.34 0.40
RE-EN-100 0.54 0.10 0.51 0.21 0.51 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.22 0.11 0.14 0.21 0.13 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.29 0.17 0.41 0.37 0.34 0.59
Split-EN-10 0.35 0.16 0.52 0.33 0.45 0.56
Lasso 0.02 0.12 0.02 0.22 0.02 0.42
Elastic Net 0.03 0.12 0.05 0.22 0.05 0.41
Adaptive 0.02 0.13 0.02 0.21 0.02 0.41
Relaxed 0.02 0.11 0.02 0.21 0.02 0.41
0.2 MCP 0.00 0.10 0.00 0.27 0.00 0.57
SIS-SCAD 0.00 0.10 0.00 0.15 0.00 0.46
RuleFit 0.07 0.19 0.05 0.35 0.04 0.54
RE-Lasso-100 0.32 0.11 0.31 0.21 0.32 0.41
RE-EN-100 0.49 0.10 0.49 0.20 0.49 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.17 0.11 0.12 0.21 0.12 0.41
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 25: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 2b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.79 0.89 0.93 0.70 0.90 0.86 0.93 0.97 0.45 0.93 0.91 0.95 0.99 0.31
Split-EN-10 0.85 0.79 0.89 0.93 0.69 0.90 0.86 0.93 0.97 0.44 0.93 0.91 0.95 0.99 0.32
Lasso 0.82 0.75 0.87 0.90 0.84 0.87 0.81 0.90 0.95 0.60 0.89 0.85 0.92 0.97 0.51
Elastic Net 0.83 0.76 0.88 0.92 0.79 0.88 0.83 0.91 0.96 0.53 0.91 0.87 0.93 0.98 0.42
Adaptive 0.79 0.66 0.89 0.90 0.95 0.84 0.72 0.92 0.94 0.75 0.87 0.76 0.93 0.96 0.65
Relaxed 0.82 0.76 0.86 0.90 1.04 0.87 0.82 0.89 0.95 0.92 0.88 0.85 0.90 0.96 0.78
0.4 MCP 0.79 0.71 0.84 0.88 0.90 0.82 0.73 0.88 0.91 0.79 0.83 0.75 0.88 0.92 0.77
SIS-SCAD 0.77 0.69 0.83 0.86 1.01 0.80 0.70 0.86 0.89 0.86 0.81 0.71 0.87 0.90 0.83
RuleFit 0.78 0.69 0.84 0.86 1.13 0.81 0.73 0.86 0.88 0.99 0.82 0.73 0.88 0.88 1.14
RE-Lasso-100 0.84 0.78 0.89 0.93 0.69 0.90 0.84 0.93 0.97 0.47 0.93 0.89 0.95 0.99 0.37
RE-EN-100 0.85 0.78 0.89 0.93 0.68 0.90 0.85 0.93 0.97 0.46 0.93 0.90 0.95 0.99 0.35
RGLM-100 0.83 0.71 0.92 0.93 0.95 0.89 0.78 0.96 0.98 0.77 0.92 0.85 0.97 0.99 0.66
RF-500 0.85 0.79 0.89 0.94 0.73 0.90 0.85 0.94 0.97 0.57 0.93 0.89 0.96 0.99 0.52
XGB 0.79 0.71 0.85 0.88 0.91 0.83 0.75 0.89 0.92 0.74 0.85 0.76 0.90 0.94 0.69
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.66 0.94 0.93 0.66 0.90 0.75 0.97 0.97 0.44 0.94 0.86 0.97 0.99 0.30
Split-EN-10 0.85 0.66 0.94 0.93 0.65 0.90 0.75 0.97 0.98 0.43 0.94 0.87 0.97 0.99 0.29
Lasso 0.83 0.61 0.92 0.90 0.79 0.87 0.68 0.95 0.95 0.60 0.90 0.80 0.94 0.97 0.47
Elastic Net 0.84 0.64 0.93 0.91 0.73 0.89 0.71 0.96 0.96 0.53 0.91 0.83 0.95 0.98 0.39
Adaptive 0.78 0.36 0.97 0.87 0.92 0.82 0.44 0.98 0.92 0.78 0.87 0.64 0.97 0.96 0.63
Relaxed 0.82 0.62 0.91 0.89 1.14 0.87 0.71 0.93 0.94 1.03 0.89 0.82 0.92 0.96 0.84
0.3 MCP 0.79 0.55 0.90 0.85 0.95 0.82 0.55 0.93 0.89 0.81 0.84 0.68 0.91 0.91 0.73
SIS-SCAD 0.79 0.53 0.90 0.86 0.93 0.81 0.52 0.93 0.89 0.85 0.82 0.61 0.91 0.90 0.77
RuleFit 0.79 0.60 0.88 0.83 1.23 0.82 0.57 0.93 0.87 0.97 0.83 0.64 0.92 0.89 0.95
RE-Lasso-100 0.85 0.64 0.94 0.92 0.67 0.90 0.73 0.97 0.97 0.47 0.93 0.84 0.97 0.99 0.34
RE-EN-100 0.85 0.64 0.94 0.93 0.66 0.90 0.74 0.97 0.97 0.45 0.93 0.84 0.97 0.99 0.33
RGLM-100 0.81 0.45 0.98 0.92 0.87 0.86 0.55 0.99 0.98 0.71 0.91 0.73 0.99 0.99 0.61
RF-500 0.86 0.66 0.94 0.93 0.69 0.90 0.72 0.98 0.98 0.54 0.93 0.82 0.98 0.99 0.48
XGB 0.81 0.61 0.90 0.87 0.86 0.85 0.63 0.94 0.92 0.70 0.86 0.69 0.93 0.93 0.65
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.88 0.52 0.97 0.94 0.55 0.92 0.64 0.98 0.98 0.38 0.94 0.76 0.98 0.99 0.28
Split-EN-10 0.88 0.52 0.97 0.94 0.55 0.92 0.65 0.98 0.98 0.38 0.94 0.77 0.98 0.99 0.28
Lasso 0.86 0.47 0.96 0.91 0.65 0.89 0.57 0.97 0.95 0.52 0.90 0.65 0.97 0.96 0.46
Elastic Net 0.87 0.50 0.96 0.92 0.62 0.90 0.61 0.98 0.96 0.45 0.92 0.70 0.97 0.98 0.37
Adaptive 0.82 0.18 0.99 0.86 0.79 0.83 0.21 0.99 0.87 0.74 0.85 0.35 0.99 0.91 0.67
Relaxed 0.86 0.51 0.95 0.89 0.93 0.88 0.59 0.95 0.94 0.94 0.89 0.68 0.95 0.95 0.80
0.2 MCP 0.83 0.38 0.95 0.86 0.76 0.84 0.36 0.96 0.87 0.80 0.84 0.45 0.95 0.88 0.74
SIS-SCAD 0.82 0.30 0.96 0.86 0.78 0.83 0.25 0.98 0.89 0.74 0.83 0.37 0.96 0.89 0.74
RuleFit 0.83 0.46 0.93 0.81 1.08 0.86 0.42 0.96 0.87 0.84 0.85 0.48 0.96 0.88 0.86
RE-Lasso-100 0.88 0.52 0.97 0.93 0.55 0.91 0.63 0.98 0.97 0.39 0.93 0.76 0.98 0.99 0.31
RE-EN-100 0.87 0.49 0.97 0.93 0.57 0.91 0.64 0.98 0.97 0.38 0.94 0.78 0.98 0.99 0.29
RGLM-100 0.84 0.24 0.99 0.93 0.74 0.86 0.34 1.00 0.98 0.60 0.89 0.51 1.00 0.99 0.53
RF-500 0.88 0.52 0.97 0.94 0.58 0.90 0.57 0.99 0.98 0.47 0.92 0.66 0.99 0.99 0.44
XGB 0.85 0.51 0.94 0.88 0.71 0.87 0.48 0.97 0.93 0.61 0.87 0.55 0.96 0.93 0.63
Table 26: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 2b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.90 0.85 0.93 0.97 0.49 0.93 0.91 0.95 0.99 0.32 0.95 0.92 0.97 0.99 0.24
Split-EN-10 0.90 0.84 0.93 0.97 0.49 0.93 0.91 0.95 0.99 0.33 0.95 0.93 0.97 1.00 0.25
Lasso 0.87 0.80 0.91 0.95 0.63 0.90 0.86 0.92 0.97 0.48 0.92 0.87 0.95 0.98 0.38
Elastic Net 0.88 0.81 0.92 0.96 0.60 0.91 0.87 0.94 0.98 0.42 0.93 0.89 0.96 0.99 0.33
Adaptive 0.85 0.73 0.93 0.95 0.78 0.89 0.82 0.95 0.97 0.58 0.92 0.85 0.96 0.98 0.47
Relaxed 0.86 0.81 0.90 0.94 1.16 0.89 0.87 0.90 0.96 0.80 0.91 0.88 0.93 0.98 0.99
0.4 MCP 0.83 0.76 0.89 0.92 0.75 0.86 0.80 0.89 0.94 0.64 0.87 0.80 0.92 0.95 0.63
SIS-SCAD 0.85 0.77 0.91 0.94 0.70 0.87 0.81 0.92 0.96 0.61 0.88 0.81 0.93 0.96 0.59
RuleFit 0.86 0.81 0.89 0.93 0.80 0.88 0.84 0.91 0.93 0.80 0.89 0.84 0.92 0.92 0.94
RE-Lasso-100 0.89 0.83 0.94 0.97 0.51 0.92 0.89 0.95 0.99 0.37 0.95 0.91 0.97 0.99 0.29
RE-EN-100 0.89 0.83 0.94 0.97 0.51 0.92 0.89 0.95 0.99 0.37 0.95 0.91 0.97 0.99 0.29
RGLM-100 0.90 0.85 0.94 0.97 0.64 0.94 0.91 0.96 0.99 0.47 0.96 0.93 0.98 1.00 0.40
RF-500 0.90 0.86 0.93 0.97 0.50 0.94 0.92 0.95 0.99 0.37 0.96 0.94 0.97 1.00 0.32
XGB 0.87 0.83 0.90 0.95 0.63 0.88 0.85 0.90 0.93 0.66 0.88 0.84 0.90 0.92 0.73
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.90 0.77 0.95 0.97 0.46 0.93 0.85 0.96 0.99 0.31 0.95 0.91 0.97 1.00 0.24
Split-EN-10 0.90 0.77 0.96 0.97 0.46 0.93 0.85 0.97 0.99 0.31 0.95 0.91 0.97 1.00 0.24
Lasso 0.87 0.72 0.94 0.95 0.58 0.90 0.79 0.95 0.97 0.44 0.93 0.85 0.96 0.98 0.35
Elastic Net 0.88 0.72 0.95 0.96 0.55 0.91 0.81 0.96 0.98 0.39 0.94 0.87 0.97 0.99 0.30
Adaptive 0.85 0.58 0.97 0.95 0.71 0.88 0.64 0.97 0.96 0.59 0.91 0.77 0.97 0.98 0.47
Relaxed 0.87 0.74 0.93 0.95 1.16 0.89 0.80 0.93 0.97 1.25 0.92 0.86 0.94 0.98 0.65
0.3 MCP 0.85 0.68 0.92 0.92 0.69 0.86 0.71 0.93 0.94 0.61 0.87 0.73 0.93 0.95 0.58
SIS-SCAD 0.85 0.64 0.95 0.94 0.69 0.87 0.68 0.95 0.95 0.62 0.88 0.70 0.95 0.96 0.60
RuleFit 0.87 0.74 0.93 0.92 0.81 0.88 0.77 0.93 0.92 0.90 0.88 0.78 0.93 0.88 1.12
RE-Lasso-100 0.89 0.74 0.96 0.97 0.49 0.93 0.83 0.97 0.99 0.34 0.95 0.88 0.97 0.99 0.27
RE-EN-100 0.89 0.73 0.96 0.97 0.49 0.93 0.83 0.96 0.98 0.34 0.95 0.88 0.97 0.99 0.27
RGLM-100 0.90 0.76 0.97 0.98 0.59 0.94 0.85 0.98 0.99 0.42 0.96 0.90 0.98 1.00 0.36
RF-500 0.91 0.81 0.95 0.97 0.47 0.94 0.87 0.97 0.99 0.35 0.96 0.91 0.97 1.00 0.31
XGB 0.87 0.77 0.92 0.94 0.62 0.88 0.79 0.92 0.93 0.66 0.88 0.80 0.92 0.92 0.72
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.92 0.65 0.98 0.97 0.38 0.94 0.77 0.98 0.99 0.27 0.95 0.81 0.99 1.00 0.23
Split-EN-10 0.92 0.64 0.98 0.97 0.38 0.94 0.77 0.99 0.99 0.28 0.95 0.81 0.99 1.00 0.23
Lasso 0.90 0.59 0.97 0.95 0.49 0.92 0.69 0.98 0.98 0.40 0.92 0.72 0.98 0.98 0.35
Elastic Net 0.90 0.59 0.98 0.96 0.46 0.93 0.71 0.98 0.98 0.35 0.93 0.74 0.98 0.99 0.30
Adaptive 0.85 0.24 0.99 0.94 0.65 0.87 0.42 0.99 0.97 0.55 0.88 0.46 0.99 0.96 0.52
Relaxed 0.89 0.62 0.96 0.95 0.86 0.92 0.75 0.96 0.97 0.84 0.92 0.76 0.96 0.97 0.70
0.2 MCP 0.88 0.55 0.96 0.93 0.57 0.89 0.60 0.96 0.95 0.53 0.89 0.61 0.96 0.95 0.51
SIS-SCAD 0.87 0.40 0.98 0.94 0.59 0.88 0.50 0.98 0.95 0.55 0.88 0.54 0.97 0.95 0.52
RuleFit 0.90 0.65 0.96 0.90 0.74 0.90 0.69 0.95 0.88 0.96 0.89 0.70 0.94 0.85 1.10
RE-Lasso-100 0.91 0.61 0.98 0.97 0.40 0.94 0.75 0.99 0.99 0.30 0.95 0.81 0.99 0.99 0.25
RE-EN-100 0.91 0.62 0.99 0.97 0.39 0.94 0.74 0.99 0.99 0.30 0.95 0.83 0.99 0.99 0.25
RGLM-100 0.91 0.60 0.99 0.98 0.48 0.94 0.76 0.99 0.99 0.36 0.95 0.79 0.99 1.00 0.31
RF-500 0.92 0.70 0.98 0.97 0.38 0.95 0.81 0.98 0.99 0.30 0.95 0.81 0.99 1.00 0.28
XGB 0.90 0.69 0.95 0.94 0.53 0.90 0.73 0.95 0.96 0.54 0.90 0.72 0.94 0.94 0.58
Table 27: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 2b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.89 0.85 0.93 0.97 0.50 0.93 0.91 0.94 0.98 0.34 0.95 0.93 0.97 0.99 0.24
Split-EN-10 0.89 0.84 0.93 0.97 0.52 0.93 0.91 0.94 0.98 0.35 0.95 0.93 0.97 0.99 0.25
Lasso 0.87 0.82 0.91 0.95 0.62 0.90 0.87 0.92 0.97 0.46 0.92 0.88 0.95 0.98 0.36
Elastic Net 0.88 0.83 0.92 0.96 0.57 0.91 0.88 0.93 0.98 0.40 0.93 0.90 0.96 0.99 0.31
Adaptive 0.86 0.75 0.93 0.95 0.75 0.89 0.82 0.94 0.97 0.57 0.92 0.85 0.96 0.98 0.46
Relaxed 0.87 0.81 0.90 0.95 1.26 0.89 0.87 0.90 0.97 0.85 0.91 0.88 0.93 0.98 0.96
0.4 MCP 0.84 0.77 0.89 0.93 0.72 0.86 0.81 0.90 0.94 0.63 0.87 0.81 0.92 0.95 0.58
SIS-SCAD 0.85 0.77 0.90 0.94 0.71 0.87 0.81 0.91 0.96 0.61 0.88 0.81 0.93 0.96 0.59
RuleFit 0.86 0.81 0.90 0.93 0.79 0.88 0.83 0.91 0.93 0.80 0.89 0.84 0.92 0.92 0.94
RE-Lasso-100 0.89 0.84 0.92 0.96 0.52 0.92 0.89 0.93 0.98 0.38 0.94 0.91 0.96 0.99 0.30
RE-EN-100 0.89 0.83 0.92 0.96 0.52 0.91 0.89 0.93 0.98 0.39 0.94 0.91 0.96 0.99 0.29
RGLM-100 0.90 0.84 0.94 0.97 0.63 0.94 0.91 0.96 0.99 0.47 0.96 0.93 0.98 1.00 0.40
RF-500 0.90 0.86 0.93 0.97 0.49 0.94 0.92 0.95 0.99 0.38 0.96 0.93 0.97 1.00 0.33
XGB 0.87 0.83 0.90 0.95 0.63 0.88 0.85 0.90 0.93 0.66 0.88 0.84 0.90 0.92 0.73
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.90 0.79 0.94 0.97 0.48 0.93 0.86 0.96 0.98 0.32 0.95 0.91 0.97 0.99 0.23
Split-EN-10 0.90 0.79 0.95 0.97 0.48 0.93 0.86 0.96 0.98 0.32 0.95 0.92 0.97 1.00 0.23
Lasso 0.88 0.75 0.94 0.95 0.57 0.91 0.80 0.95 0.97 0.44 0.93 0.86 0.95 0.98 0.34
Elastic Net 0.88 0.75 0.94 0.96 0.54 0.91 0.83 0.95 0.98 0.38 0.94 0.88 0.96 0.99 0.28
Adaptive 0.85 0.59 0.96 0.95 0.70 0.88 0.66 0.97 0.97 0.57 0.91 0.78 0.97 0.98 0.46
Relaxed 0.87 0.75 0.93 0.95 1.11 0.90 0.81 0.94 0.97 1.25 0.92 0.87 0.94 0.98 0.63
0.3 MCP 0.85 0.69 0.92 0.93 0.67 0.86 0.71 0.93 0.94 0.62 0.87 0.74 0.93 0.95 0.58
SIS-SCAD 0.85 0.64 0.95 0.94 0.68 0.87 0.67 0.95 0.95 0.63 0.88 0.70 0.95 0.96 0.60
RuleFit 0.87 0.74 0.93 0.93 0.80 0.88 0.78 0.93 0.92 0.90 0.88 0.78 0.93 0.88 1.12
RE-Lasso-100 0.89 0.77 0.94 0.96 0.51 0.92 0.84 0.95 0.98 0.36 0.94 0.89 0.97 0.99 0.27
RE-EN-100 0.89 0.77 0.94 0.96 0.52 0.92 0.84 0.95 0.98 0.37 0.95 0.89 0.97 0.99 0.26
RGLM-100 0.89 0.75 0.96 0.97 0.59 0.94 0.84 0.98 0.99 0.43 0.96 0.90 0.98 1.00 0.37
RF-500 0.90 0.81 0.95 0.97 0.46 0.94 0.87 0.97 0.99 0.36 0.96 0.91 0.98 1.00 0.31
XGB 0.87 0.77 0.92 0.94 0.63 0.88 0.79 0.92 0.94 0.66 0.88 0.80 0.92 0.92 0.72
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.92 0.70 0.98 0.97 0.38 0.95 0.82 0.98 0.99 0.26 0.95 0.84 0.98 0.99 0.22
Split-EN-10 0.92 0.69 0.97 0.97 0.38 0.95 0.81 0.98 0.99 0.26 0.96 0.85 0.98 0.99 0.22
Lasso 0.90 0.64 0.97 0.95 0.47 0.92 0.73 0.97 0.98 0.37 0.93 0.76 0.97 0.98 0.33
Elastic Net 0.91 0.66 0.97 0.96 0.43 0.93 0.77 0.98 0.98 0.32 0.94 0.80 0.98 0.99 0.27
Adaptive 0.86 0.29 0.99 0.95 0.62 0.89 0.49 0.99 0.97 0.51 0.89 0.49 0.99 0.97 0.48
Relaxed 0.90 0.66 0.96 0.95 0.69 0.92 0.77 0.96 0.97 0.77 0.92 0.77 0.96 0.98 0.61
0.2 MCP 0.88 0.54 0.96 0.92 0.59 0.89 0.60 0.96 0.94 0.54 0.89 0.60 0.96 0.94 0.51
SIS-SCAD 0.87 0.42 0.98 0.93 0.60 0.88 0.52 0.98 0.95 0.54 0.88 0.52 0.97 0.95 0.53
RuleFit 0.90 0.64 0.96 0.90 0.73 0.90 0.70 0.95 0.88 0.96 0.90 0.72 0.94 0.85 1.11
RE-Lasso-100 0.92 0.70 0.97 0.97 0.38 0.94 0.80 0.98 0.99 0.28 0.95 0.84 0.98 0.99 0.25
RE-EN-100 0.91 0.69 0.97 0.96 0.39 0.94 0.81 0.97 0.98 0.29 0.95 0.86 0.97 0.99 0.24
RGLM-100 0.91 0.61 0.99 0.97 0.47 0.94 0.76 0.99 0.99 0.36 0.95 0.79 0.99 1.00 0.32
RF-500 0.92 0.71 0.97 0.97 0.38 0.95 0.81 0.98 0.99 0.30 0.95 0.80 0.99 1.00 0.29
XGB 0.90 0.68 0.95 0.94 0.53 0.90 0.73 0.95 0.96 0.55 0.90 0.72 0.94 0.95 0.58
Table 28: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 2b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.80 0.89 0.94 0.65 0.91 0.88 0.94 0.98 0.39 0.95 0.92 0.96 0.99 0.26
Split-EN-10 0.85 0.80 0.89 0.94 0.65 0.91 0.88 0.94 0.98 0.39 0.95 0.92 0.96 0.99 0.26
Lasso 0.84 0.78 0.88 0.92 0.74 0.89 0.84 0.92 0.96 0.51 0.91 0.87 0.94 0.98 0.40
Elastic Net 0.84 0.78 0.88 0.93 0.72 0.90 0.85 0.92 0.97 0.47 0.92 0.89 0.95 0.98 0.35
Adaptive 0.83 0.73 0.90 0.92 0.82 0.89 0.82 0.93 0.96 0.59 0.90 0.84 0.95 0.97 0.50
Relaxed 0.83 0.76 0.87 0.92 0.82 0.88 0.84 0.91 0.96 0.60 0.90 0.87 0.93 0.97 0.58
0.4 MCP 0.81 0.75 0.86 0.90 0.89 0.86 0.80 0.90 0.94 0.64 0.86 0.81 0.90 0.95 0.62
SIS-SCAD 0.80 0.74 0.85 0.89 0.87 0.85 0.78 0.89 0.93 0.68 0.85 0.78 0.89 0.93 0.70
RuleFit 0.82 0.75 0.87 0.90 0.98 0.87 0.80 0.91 0.94 0.67 0.87 0.80 0.92 0.95 0.62
RE-Lasso-100 0.85 0.79 0.89 0.94 0.65 0.91 0.86 0.94 0.98 0.41 0.94 0.90 0.97 0.99 0.30
RE-EN-100 0.85 0.80 0.89 0.94 0.64 0.91 0.87 0.94 0.98 0.40 0.94 0.91 0.97 0.99 0.29
RGLM-100 0.84 0.74 0.91 0.94 0.88 0.91 0.83 0.96 0.98 0.68 0.94 0.87 0.98 0.99 0.60
RF-500 0.85 0.81 0.89 0.94 0.67 0.92 0.88 0.94 0.98 0.47 0.94 0.90 0.97 0.99 0.43
XGB 0.82 0.75 0.86 0.90 0.81 0.87 0.82 0.91 0.95 0.58 0.88 0.81 0.93 0.96 0.54
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.87 0.71 0.94 0.94 0.61 0.92 0.82 0.97 0.98 0.34 0.95 0.89 0.98 0.99 0.24
Split-EN-10 0.87 0.71 0.94 0.94 0.60 0.92 0.82 0.97 0.98 0.34 0.95 0.90 0.98 0.99 0.24
Lasso 0.85 0.68 0.93 0.92 0.68 0.90 0.77 0.95 0.97 0.46 0.92 0.83 0.96 0.98 0.37
Elastic Net 0.85 0.69 0.93 0.93 0.66 0.91 0.78 0.96 0.97 0.42 0.93 0.85 0.97 0.98 0.31
Adaptive 0.83 0.54 0.96 0.92 0.78 0.89 0.67 0.97 0.96 0.54 0.91 0.77 0.97 0.98 0.46
Relaxed 0.85 0.69 0.92 0.92 0.85 0.90 0.78 0.95 0.96 0.63 0.91 0.84 0.94 0.97 0.54
0.3 MCP 0.83 0.62 0.92 0.90 0.79 0.86 0.68 0.93 0.93 0.65 0.87 0.73 0.93 0.94 0.59
SIS-SCAD 0.83 0.63 0.92 0.90 0.80 0.86 0.65 0.94 0.93 0.65 0.86 0.70 0.93 0.93 0.63
RuleFit 0.83 0.66 0.91 0.90 0.93 0.87 0.68 0.95 0.94 0.69 0.88 0.71 0.96 0.96 0.58
RE-Lasso-100 0.86 0.70 0.94 0.94 0.60 0.92 0.79 0.97 0.98 0.37 0.94 0.86 0.98 0.99 0.28
RE-EN-100 0.86 0.70 0.94 0.94 0.61 0.92 0.79 0.97 0.98 0.36 0.95 0.87 0.98 0.99 0.27
RGLM-100 0.84 0.55 0.97 0.94 0.81 0.90 0.68 0.99 0.98 0.60 0.93 0.78 1.00 0.99 0.54
RF-500 0.87 0.72 0.93 0.94 0.62 0.92 0.80 0.97 0.98 0.43 0.94 0.86 0.98 0.99 0.39
XGB 0.84 0.66 0.92 0.91 0.74 0.88 0.70 0.95 0.95 0.55 0.89 0.73 0.96 0.96 0.51
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.89 0.58 0.97 0.95 0.51 0.93 0.76 0.97 0.98 0.30 0.95 0.83 0.98 0.99 0.23
Split-EN-10 0.89 0.57 0.97 0.95 0.51 0.93 0.76 0.98 0.98 0.30 0.95 0.84 0.98 0.99 0.22
Lasso 0.87 0.54 0.96 0.92 0.60 0.91 0.69 0.97 0.96 0.40 0.93 0.76 0.97 0.98 0.34
Elastic Net 0.88 0.54 0.97 0.93 0.57 0.92 0.72 0.97 0.97 0.36 0.94 0.79 0.97 0.98 0.29
Adaptive 0.84 0.28 0.99 0.91 0.70 0.89 0.48 0.99 0.96 0.51 0.91 0.60 0.98 0.97 0.44
Relaxed 0.87 0.53 0.95 0.91 0.91 0.91 0.71 0.96 0.96 0.51 0.92 0.76 0.96 0.97 0.49
0.2 MCP 0.84 0.42 0.95 0.88 0.78 0.87 0.52 0.95 0.91 0.64 0.87 0.57 0.95 0.91 0.62
SIS-SCAD 0.84 0.39 0.96 0.89 0.80 0.87 0.48 0.96 0.92 0.63 0.87 0.51 0.96 0.93 0.63
RuleFit 0.86 0.50 0.95 0.89 0.89 0.89 0.57 0.97 0.92 0.65 0.89 0.60 0.97 0.94 0.60
RE-Lasso-100 0.88 0.54 0.97 0.94 0.53 0.93 0.70 0.98 0.98 0.33 0.94 0.78 0.99 0.99 0.26
RE-EN-100 0.88 0.54 0.97 0.94 0.53 0.93 0.71 0.98 0.98 0.32 0.95 0.79 0.99 0.99 0.25
RGLM-100 0.85 0.30 0.99 0.94 0.67 0.90 0.50 1.00 0.98 0.51 0.92 0.62 1.00 0.99 0.47
RF-500 0.89 0.60 0.97 0.95 0.52 0.93 0.71 0.98 0.98 0.37 0.94 0.76 0.99 0.99 0.35
XGB 0.86 0.51 0.96 0.91 0.64 0.90 0.59 0.97 0.95 0.48 0.90 0.63 0.97 0.96 0.46
Table 29: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 2b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.90 0.87 0.92 0.97 0.46 0.94 0.92 0.96 0.99 0.26 0.96 0.96 0.96 1.00 0.20
Split-EN-10 0.90 0.86 0.92 0.97 0.46 0.94 0.92 0.96 0.99 0.26 0.96 0.96 0.97 1.00 0.21
Lasso 0.88 0.84 0.91 0.96 0.56 0.92 0.89 0.94 0.98 0.36 0.94 0.92 0.94 0.99 0.30
Elastic Net 0.89 0.84 0.91 0.96 0.54 0.93 0.90 0.95 0.99 0.33 0.95 0.93 0.95 0.99 0.26
Adaptive 0.88 0.82 0.93 0.96 0.64 0.92 0.88 0.95 0.98 0.43 0.94 0.91 0.95 0.99 0.36
Relaxed 0.88 0.83 0.91 0.95 0.64 0.92 0.89 0.93 0.98 0.54 0.93 0.92 0.93 0.98 0.45
0.4 MCP 0.86 0.81 0.89 0.94 0.64 0.89 0.84 0.92 0.96 0.53 0.90 0.87 0.92 0.97 0.47
SIS-SCAD 0.87 0.83 0.90 0.95 0.60 0.90 0.86 0.93 0.97 0.48 0.91 0.88 0.92 0.97 0.43
RuleFit 0.88 0.84 0.91 0.95 0.72 0.91 0.87 0.93 0.96 0.57 0.91 0.88 0.93 0.96 0.54
RE-Lasso-100 0.89 0.85 0.92 0.97 0.47 0.94 0.91 0.96 0.99 0.29 0.96 0.94 0.96 1.00 0.24
RE-EN-100 0.89 0.85 0.92 0.97 0.47 0.94 0.91 0.96 0.99 0.29 0.96 0.95 0.96 1.00 0.24
RGLM-100 0.90 0.87 0.93 0.97 0.59 0.95 0.92 0.97 0.99 0.41 0.97 0.95 0.97 1.00 0.35
RF-500 0.90 0.87 0.92 0.97 0.47 0.95 0.92 0.96 0.99 0.30 0.96 0.95 0.97 1.00 0.28
XGB 0.88 0.85 0.91 0.96 0.56 0.92 0.88 0.94 0.98 0.40 0.93 0.90 0.94 0.98 0.35
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.91 0.82 0.95 0.97 0.42 0.95 0.88 0.97 0.99 0.25 0.97 0.93 0.98 1.00 0.19
Split-EN-10 0.91 0.81 0.95 0.97 0.42 0.95 0.88 0.97 0.99 0.26 0.97 0.93 0.98 1.00 0.19
Lasso 0.89 0.77 0.95 0.96 0.50 0.93 0.82 0.97 0.98 0.36 0.94 0.88 0.97 0.99 0.28
Elastic Net 0.90 0.78 0.95 0.97 0.49 0.93 0.84 0.97 0.98 0.32 0.95 0.89 0.97 0.99 0.24
Adaptive 0.88 0.70 0.97 0.96 0.59 0.92 0.76 0.98 0.98 0.44 0.94 0.85 0.98 0.99 0.35
Relaxed 0.89 0.77 0.94 0.96 0.57 0.92 0.84 0.96 0.98 0.53 0.93 0.89 0.95 0.98 0.46
0.3 MCP 0.86 0.73 0.92 0.94 0.61 0.88 0.72 0.94 0.95 0.56 0.89 0.78 0.94 0.95 0.53
SIS-SCAD 0.87 0.72 0.94 0.95 0.60 0.89 0.68 0.97 0.97 0.56 0.90 0.77 0.95 0.97 0.51
RuleFit 0.89 0.79 0.94 0.95 0.62 0.91 0.79 0.95 0.95 0.58 0.91 0.82 0.96 0.96 0.50
RE-Lasso-100 0.90 0.79 0.96 0.97 0.43 0.94 0.84 0.98 0.99 0.29 0.96 0.90 0.99 1.00 0.22
RE-EN-100 0.90 0.79 0.96 0.97 0.43 0.94 0.84 0.98 0.99 0.29 0.96 0.91 0.99 1.00 0.22
RGLM-100 0.91 0.79 0.96 0.98 0.54 0.95 0.86 0.98 0.99 0.38 0.97 0.91 0.99 1.00 0.33
RF-500 0.91 0.83 0.95 0.97 0.42 0.95 0.88 0.98 0.99 0.28 0.97 0.93 0.99 1.00 0.25
XGB 0.89 0.80 0.94 0.96 0.51 0.92 0.82 0.96 0.98 0.37 0.93 0.85 0.97 0.98 0.33
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.92 0.71 0.98 0.98 0.35 0.96 0.84 0.99 0.99 0.22 0.97 0.89 0.99 1.00 0.17
Split-EN-10 0.92 0.70 0.98 0.98 0.36 0.96 0.83 0.99 0.99 0.22 0.97 0.89 0.99 1.00 0.17
Lasso 0.91 0.65 0.97 0.96 0.43 0.94 0.77 0.98 0.98 0.30 0.95 0.83 0.98 0.99 0.24
Elastic Net 0.91 0.66 0.97 0.97 0.41 0.94 0.78 0.98 0.99 0.27 0.96 0.84 0.99 0.99 0.21
Adaptive 0.89 0.51 0.99 0.97 0.50 0.92 0.65 0.99 0.98 0.38 0.94 0.72 0.99 0.99 0.34
Relaxed 0.90 0.67 0.97 0.96 0.58 0.93 0.78 0.97 0.98 0.45 0.94 0.82 0.97 0.98 0.47
0.2 MCP 0.88 0.59 0.95 0.93 0.57 0.89 0.64 0.96 0.94 0.51 0.89 0.66 0.96 0.95 0.52
SIS-SCAD 0.87 0.48 0.97 0.95 0.61 0.89 0.54 0.98 0.96 0.51 0.89 0.59 0.97 0.96 0.50
RuleFit 0.91 0.68 0.96 0.94 0.56 0.92 0.72 0.97 0.95 0.46 0.92 0.73 0.97 0.96 0.46
RE-Lasso-100 0.91 0.65 0.98 0.97 0.38 0.95 0.79 0.99 0.99 0.24 0.96 0.85 0.99 1.00 0.20
RE-EN-100 0.91 0.64 0.98 0.97 0.38 0.95 0.78 0.99 0.99 0.24 0.96 0.85 0.99 1.00 0.19
RGLM-100 0.92 0.66 0.99 0.98 0.44 0.96 0.81 0.99 0.99 0.31 0.97 0.86 1.00 1.00 0.27
RF-500 0.93 0.74 0.97 0.98 0.36 0.96 0.85 0.99 0.99 0.24 0.97 0.89 0.99 1.00 0.21
XGB 0.91 0.72 0.96 0.96 0.43 0.94 0.77 0.98 0.98 0.31 0.94 0.80 0.98 0.98 0.30
Table 30: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 2b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.90 0.87 0.92 0.97 0.48 0.94 0.92 0.96 0.99 0.26 0.96 0.96 0.97 1.00 0.19
Split-EN-10 0.90 0.86 0.92 0.97 0.48 0.94 0.92 0.96 0.99 0.26 0.96 0.96 0.97 1.00 0.19
Lasso 0.88 0.84 0.91 0.96 0.55 0.93 0.89 0.94 0.98 0.35 0.94 0.92 0.95 0.99 0.28
Elastic Net 0.89 0.85 0.91 0.96 0.53 0.93 0.90 0.95 0.99 0.32 0.95 0.94 0.96 0.99 0.24
Adaptive 0.88 0.81 0.93 0.96 0.63 0.92 0.88 0.95 0.98 0.43 0.94 0.92 0.96 0.99 0.35
Relaxed 0.88 0.84 0.91 0.96 0.61 0.92 0.90 0.94 0.98 0.52 0.93 0.92 0.94 0.99 0.53
0.4 MCP 0.86 0.81 0.90 0.94 0.63 0.90 0.85 0.92 0.97 0.48 0.90 0.87 0.93 0.97 0.45
SIS-SCAD 0.87 0.81 0.91 0.95 0.65 0.90 0.86 0.93 0.97 0.46 0.91 0.88 0.93 0.98 0.42
RuleFit 0.88 0.84 0.90 0.95 0.72 0.91 0.87 0.94 0.96 0.51 0.91 0.87 0.93 0.96 0.59
RE-Lasso-100 0.89 0.85 0.92 0.97 0.49 0.94 0.91 0.96 0.99 0.29 0.96 0.94 0.97 0.99 0.23
RE-EN-100 0.89 0.85 0.92 0.97 0.49 0.94 0.91 0.96 0.99 0.29 0.96 0.95 0.97 0.99 0.22
RGLM-100 0.53 0.52 0.54 0.97 1.98 0.63 0.62 0.63 0.99 1.91 0.52 0.51 0.52 1.00 2.79
RF-500 0.90 0.87 0.92 0.97 0.49 0.95 0.92 0.97 0.99 0.30 0.97 0.96 0.98 1.00 0.27
XGB 0.88 0.84 0.91 0.95 0.57 0.92 0.88 0.95 0.98 0.38 0.93 0.90 0.95 0.99 0.34
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.91 0.83 0.94 0.97 0.43 0.94 0.89 0.97 0.99 0.26 0.97 0.94 0.98 1.00 0.18
Split-EN-10 0.91 0.83 0.94 0.97 0.44 0.94 0.89 0.97 0.99 0.27 0.97 0.94 0.98 1.00 0.18
Lasso 0.89 0.80 0.94 0.96 0.50 0.93 0.84 0.96 0.98 0.34 0.94 0.89 0.97 0.99 0.26
Elastic Net 0.90 0.81 0.94 0.96 0.48 0.93 0.85 0.96 0.98 0.31 0.95 0.91 0.97 0.99 0.23
Adaptive 0.89 0.74 0.96 0.96 0.58 0.92 0.79 0.97 0.98 0.41 0.93 0.83 0.98 0.99 0.36
Relaxed 0.89 0.79 0.93 0.96 0.58 0.92 0.85 0.95 0.98 0.56 0.93 0.89 0.96 0.99 0.49
0.3 MCP 0.86 0.75 0.92 0.94 0.61 0.88 0.75 0.94 0.95 0.55 0.89 0.77 0.94 0.96 0.53
SIS-SCAD 0.87 0.74 0.93 0.95 0.62 0.88 0.70 0.96 0.96 0.55 0.90 0.76 0.96 0.97 0.50
RuleFit 0.89 0.79 0.93 0.95 0.65 0.90 0.79 0.95 0.95 0.60 0.91 0.82 0.95 0.96 0.51
RE-Lasso-100 0.90 0.82 0.94 0.97 0.45 0.94 0.86 0.97 0.99 0.29 0.96 0.91 0.98 0.99 0.21
RE-EN-100 0.90 0.82 0.94 0.97 0.45 0.94 0.86 0.97 0.99 0.29 0.96 0.92 0.98 0.99 0.21
RGLM-100 0.61 0.57 0.63 0.97 1.86 0.62 0.59 0.64 0.99 2.15 0.69 0.67 0.70 1.00 2.10
RF-500 0.91 0.84 0.94 0.97 0.44 0.95 0.89 0.97 0.99 0.29 0.97 0.93 0.98 1.00 0.25
XGB 0.89 0.81 0.93 0.96 0.51 0.92 0.82 0.96 0.98 0.38 0.93 0.85 0.96 0.98 0.33
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.92 0.76 0.96 0.97 0.39 0.96 0.87 0.98 0.99 0.21 0.97 0.91 0.98 1.00 0.16
Split-EN-10 0.92 0.75 0.96 0.97 0.39 0.96 0.87 0.98 0.99 0.21 0.97 0.91 0.99 1.00 0.16
Lasso 0.90 0.71 0.96 0.96 0.46 0.94 0.81 0.97 0.98 0.28 0.95 0.85 0.98 0.99 0.23
Elastic Net 0.91 0.72 0.96 0.96 0.43 0.95 0.82 0.97 0.99 0.25 0.96 0.87 0.98 0.99 0.19
Adaptive 0.89 0.57 0.98 0.96 0.52 0.93 0.70 0.98 0.98 0.34 0.94 0.76 0.99 0.99 0.31
Relaxed 0.90 0.71 0.95 0.96 0.77 0.93 0.81 0.96 0.98 0.56 0.94 0.86 0.97 0.99 0.40
0.2 MCP 0.87 0.63 0.94 0.93 0.61 0.90 0.66 0.95 0.95 0.49 0.89 0.68 0.95 0.95 0.50
SIS-SCAD 0.86 0.50 0.97 0.95 0.62 0.90 0.56 0.98 0.96 0.50 0.89 0.56 0.98 0.96 0.52
RuleFit 0.90 0.71 0.95 0.94 0.64 0.92 0.73 0.96 0.95 0.49 0.92 0.74 0.97 0.95 0.50
RE-Lasso-100 0.91 0.73 0.96 0.96 0.42 0.95 0.83 0.98 0.99 0.24 0.96 0.87 0.99 0.99 0.19
RE-EN-100 0.91 0.73 0.96 0.96 0.42 0.95 0.83 0.98 0.99 0.24 0.96 0.88 0.99 0.99 0.19
RGLM-100 0.67 0.58 0.69 0.97 1.81 0.74 0.68 0.75 0.99 2.01 0.72 0.66 0.74 1.00 2.74
RF-500 0.92 0.77 0.96 0.97 0.40 0.96 0.86 0.98 0.99 0.24 0.97 0.89 0.99 1.00 0.22
XGB 0.90 0.74 0.95 0.96 0.47 0.93 0.79 0.97 0.98 0.33 0.94 0.79 0.97 0.98 0.32
Table 31: Mean recall (RCL) and precision (PRC) for Scenario 2b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.51 0.35 0.45 0.48 0.34 0.67
Split-EN-10 0.54 0.35 0.50 0.46 0.43 0.66
Lasso 0.11 0.59 0.07 0.73 0.05 0.82
Elastic Net 0.16 0.55 0.13 0.68 0.09 0.80
Adaptive 0.10 0.58 0.07 0.69 0.04 0.80
Relaxed 0.10 0.64 0.07 0.81 0.04 0.90
0.4 MCP 0.04 0.77 0.02 0.82 0.01 0.89
SIS-SCAD 0.02 0.83 0.01 0.91 0.01 0.96
RuleFit 0.18 0.44 0.12 0.64 0.07 0.79
RE-Lasso-100 0.73 0.20 0.64 0.38 0.52 0.62
RE-EN-100 0.84 0.17 0.78 0.33 0.67 0.57
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.56 0.40 0.41 0.72 0.26 0.92
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.48 0.36 0.41 0.46 0.31 0.64
Split-EN-10 0.57 0.33 0.53 0.42 0.43 0.62
Lasso 0.09 0.58 0.07 0.70 0.04 0.80
Elastic Net 0.16 0.54 0.11 0.66 0.08 0.79
Adaptive 0.09 0.56 0.06 0.67 0.04 0.79
Relaxed 0.08 0.66 0.06 0.79 0.03 0.90
0.3 MCP 0.03 0.65 0.01 0.84 0.01 0.91
SIS-SCAD 0.02 0.89 0.01 0.95 0.01 0.98
RuleFit 0.17 0.46 0.11 0.69 0.06 0.81
RE-Lasso-100 0.71 0.21 0.60 0.38 0.49 0.60
RE-EN-100 0.83 0.17 0.76 0.33 0.65 0.55
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.53 0.40 0.37 0.68 0.25 0.92
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.51 0.35 0.39 0.48 0.29 0.66
Split-EN-10 0.53 0.38 0.53 0.43 0.41 0.63
Lasso 0.09 0.60 0.06 0.72 0.03 0.83
Elastic Net 0.15 0.54 0.11 0.68 0.07 0.80
Adaptive 0.08 0.60 0.05 0.67 0.03 0.82
Relaxed 0.08 0.63 0.05 0.80 0.03 0.91
0.2 MCP 0.02 0.79 0.01 0.85 0.00 0.90
SIS-SCAD 0.02 0.88 0.01 0.96 0.01 1.00
RuleFit 0.15 0.55 0.08 0.76 0.05 0.89
RE-Lasso-100 0.69 0.22 0.56 0.39 0.45 0.61
RE-EN-100 0.80 0.18 0.72 0.33 0.62 0.57
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.49 0.43 0.33 0.67 0.22 0.89
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 32: Mean recall (RCL) and precision (PRC) for Scenario 2b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.53 0.28 0.39 0.41 0.26 0.57
Split-EN-10 0.73 0.31 0.62 0.44 0.45 0.63
Lasso 0.08 0.51 0.05 0.55 0.03 0.66
Elastic Net 0.18 0.55 0.13 0.65 0.08 0.73
Adaptive 0.07 0.50 0.05 0.52 0.03 0.62
Relaxed 0.07 0.69 0.04 0.82 0.02 0.91
0.4 MCP 0.02 0.61 0.01 0.69 0.00 0.86
SIS-SCAD 0.02 0.91 0.01 0.96 0.01 0.99
RuleFit 0.14 0.51 0.06 0.85 0.02 0.96
RE-Lasso-100 0.69 0.19 0.55 0.33 0.41 0.50
RE-EN-100 0.89 0.17 0.79 0.32 0.66 0.51
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.52 0.50 0.35 0.84 0.21 0.99
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.51 0.30 0.37 0.42 0.26 0.58
Split-EN-10 0.73 0.30 0.60 0.44 0.45 0.62
Lasso 0.07 0.48 0.05 0.56 0.03 0.70
Elastic Net 0.18 0.54 0.13 0.63 0.08 0.75
Adaptive 0.07 0.44 0.05 0.55 0.03 0.67
Relaxed 0.06 0.64 0.04 0.77 0.02 0.89
0.3 MCP 0.01 0.71 0.01 0.81 0.00 0.89
SIS-SCAD 0.02 0.98 0.01 0.99 0.00 1.00
RuleFit 0.11 0.63 0.05 0.91 0.02 0.97
RE-Lasso-100 0.66 0.20 0.52 0.32 0.40 0.50
RE-EN-100 0.86 0.18 0.77 0.32 0.65 0.52
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.50 0.51 0.34 0.82 0.21 1.00
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.49 0.31 0.36 0.43 0.25 0.58
Split-EN-10 0.69 0.30 0.57 0.44 0.40 0.60
Lasso 0.07 0.50 0.04 0.63 0.02 0.65
Elastic Net 0.16 0.56 0.11 0.65 0.07 0.71
Adaptive 0.06 0.50 0.04 0.57 0.02 0.62
Relaxed 0.06 0.67 0.03 0.84 0.02 0.86
0.2 MCP 0.01 0.85 0.01 0.93 0.00 0.93
SIS-SCAD 0.02 0.99 0.01 1.00 0.00 1.00
RuleFit 0.11 0.79 0.07 0.94 0.04 0.98
RE-Lasso-100 0.60 0.20 0.48 0.33 0.35 0.47
RE-EN-100 0.84 0.19 0.75 0.33 0.62 0.51
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.45 0.51 0.30 0.81 0.18 0.96
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 33: Mean recall (RCL) and precision (PRC) for Scenario 2b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.45 0.35 0.35 0.47 0.26 0.63
Split-EN-10 0.61 0.36 0.53 0.48 0.42 0.66
Lasso 0.08 0.52 0.05 0.57 0.03 0.70
Elastic Net 0.15 0.55 0.11 0.62 0.08 0.74
Adaptive 0.07 0.49 0.04 0.56 0.03 0.68
Relaxed 0.07 0.62 0.04 0.78 0.02 0.84
0.4 MCP 0.02 0.58 0.01 0.64 0.00 0.78
SIS-SCAD 0.02 0.88 0.01 0.94 0.01 0.99
RuleFit 0.14 0.52 0.06 0.85 0.02 0.96
RE-Lasso-100 0.64 0.21 0.51 0.35 0.39 0.53
RE-EN-100 0.83 0.18 0.74 0.33 0.63 0.54
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.53 0.50 0.33 0.76 0.20 0.94
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.48 0.34 0.36 0.46 0.26 0.63
Split-EN-10 0.63 0.34 0.55 0.46 0.43 0.64
Lasso 0.07 0.48 0.04 0.58 0.03 0.72
Elastic Net 0.15 0.51 0.11 0.62 0.08 0.77
Adaptive 0.06 0.45 0.04 0.55 0.03 0.71
Relaxed 0.06 0.62 0.04 0.72 0.02 0.85
0.3 MCP 0.01 0.63 0.01 0.79 0.00 0.82
SIS-SCAD 0.02 0.94 0.01 0.97 0.00 0.99
RuleFit 0.11 0.61 0.05 0.89 0.02 0.97
RE-Lasso-100 0.61 0.21 0.49 0.34 0.38 0.54
RE-EN-100 0.80 0.18 0.73 0.33 0.62 0.54
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.48 0.50 0.32 0.75 0.21 0.94
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.49 0.32 0.37 0.45 0.25 0.61
Split-EN-10 0.61 0.33 0.53 0.46 0.38 0.62
Lasso 0.06 0.48 0.04 0.65 0.02 0.63
Elastic Net 0.14 0.53 0.10 0.64 0.06 0.70
Adaptive 0.06 0.49 0.04 0.58 0.02 0.62
Relaxed 0.06 0.55 0.03 0.74 0.02 0.74
0.2 MCP 0.01 0.86 0.01 0.90 0.00 0.95
SIS-SCAD 0.01 0.94 0.01 0.99 0.00 0.98
RuleFit 0.10 0.72 0.06 0.85 0.03 0.89
RE-Lasso-100 0.56 0.22 0.46 0.35 0.32 0.49
RE-EN-100 0.79 0.20 0.70 0.34 0.58 0.52
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.43 0.49 0.29 0.74 0.17 0.85
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 34: Mean recall (RCL) and precision (PRC) for Scenario 2b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.68 0.25 0.59 0.38 0.46 0.63
Split-EN-10 0.72 0.25 0.67 0.36 0.56 0.60
Lasso 0.15 0.60 0.11 0.67 0.07 0.83
Elastic Net 0.21 0.55 0.16 0.64 0.11 0.81
Adaptive 0.15 0.56 0.10 0.66 0.07 0.82
Relaxed 0.14 0.66 0.10 0.74 0.06 0.89
0.4 MCP 0.06 0.70 0.04 0.81 0.02 0.90
SIS-SCAD 0.04 0.80 0.02 0.92 0.01 0.92
RuleFit 0.32 0.43 0.22 0.71 0.14 0.87
RE-Lasso-100 0.81 0.15 0.75 0.32 0.63 0.59
RE-EN-100 0.89 0.13 0.84 0.28 0.75 0.55
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.83 0.34 0.62 0.71 0.39 0.97
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.60 0.31 0.58 0.40 0.45 0.61
Split-EN-10 0.70 0.27 0.68 0.36 0.56 0.59
Lasso 0.15 0.58 0.10 0.70 0.07 0.84
Elastic Net 0.20 0.53 0.16 0.66 0.11 0.80
Adaptive 0.14 0.55 0.10 0.69 0.07 0.82
Relaxed 0.13 0.69 0.10 0.76 0.06 0.89
0.3 MCP 0.05 0.74 0.03 0.81 0.02 0.92
SIS-SCAD 0.04 0.85 0.02 0.93 0.01 0.94
RuleFit 0.29 0.43 0.21 0.71 0.13 0.86
RE-Lasso-100 0.80 0.16 0.72 0.34 0.61 0.59
RE-EN-100 0.87 0.14 0.83 0.30 0.73 0.54
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.79 0.34 0.57 0.72 0.37 0.97
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.57 0.31 0.57 0.39 0.42 0.60
Split-EN-10 0.66 0.28 0.68 0.36 0.52 0.57
Lasso 0.12 0.57 0.09 0.72 0.06 0.80
Elastic Net 0.18 0.55 0.15 0.66 0.10 0.79
Adaptive 0.12 0.59 0.09 0.70 0.06 0.80
Relaxed 0.10 0.70 0.09 0.80 0.05 0.88
0.2 MCP 0.03 0.76 0.02 0.87 0.01 0.92
SIS-SCAD 0.03 0.83 0.02 0.94 0.01 0.98
RuleFit 0.26 0.45 0.17 0.71 0.11 0.84
RE-Lasso-100 0.77 0.17 0.69 0.36 0.57 0.59
RE-EN-100 0.86 0.15 0.81 0.31 0.69 0.55
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.71 0.37 0.51 0.74 0.33 0.96
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 35: Mean recall (RCL) and precision (PRC) for Scenario 2b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.65 0.23 0.51 0.36 0.37 0.55
Split-EN-10 0.82 0.20 0.72 0.37 0.55 0.57
Lasso 0.12 0.48 0.08 0.58 0.05 0.65
Elastic Net 0.22 0.50 0.16 0.61 0.10 0.71
Adaptive 0.12 0.50 0.08 0.55 0.05 0.64
Relaxed 0.09 0.69 0.06 0.84 0.04 0.87
0.4 MCP 0.03 0.52 0.02 0.69 0.01 0.77
SIS-SCAD 0.04 0.86 0.02 0.94 0.01 0.96
RuleFit 0.25 0.49 0.17 0.73 0.10 0.84
RE-Lasso-100 0.77 0.15 0.66 0.29 0.52 0.47
RE-EN-100 0.91 0.14 0.85 0.28 0.71 0.48
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.72 0.42 0.51 0.86 0.30 1.00
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.63 0.22 0.50 0.36 0.35 0.52
Split-EN-10 0.79 0.23 0.68 0.36 0.52 0.56
Lasso 0.11 0.52 0.08 0.60 0.04 0.67
Elastic Net 0.22 0.56 0.16 0.61 0.10 0.72
Adaptive 0.11 0.48 0.08 0.58 0.04 0.67
Relaxed 0.09 0.71 0.07 0.77 0.04 0.86
0.3 MCP 0.02 0.62 0.01 0.77 0.00 0.77
SIS-SCAD 0.03 0.88 0.02 0.96 0.01 0.99
RuleFit 0.23 0.50 0.15 0.73 0.08 0.82
RE-Lasso-100 0.75 0.16 0.63 0.29 0.49 0.47
RE-EN-100 0.90 0.15 0.82 0.29 0.70 0.49
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.70 0.44 0.47 0.85 0.29 1.00
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.62 0.22 0.50 0.37 0.35 0.53
Split-EN-10 0.78 0.23 0.66 0.37 0.51 0.55
Lasso 0.10 0.49 0.06 0.58 0.04 0.69
Elastic Net 0.20 0.53 0.14 0.63 0.09 0.72
Adaptive 0.10 0.46 0.06 0.56 0.04 0.69
Relaxed 0.09 0.61 0.06 0.77 0.03 0.86
0.2 MCP 0.01 0.73 0.01 0.87 0.00 0.91
SIS-SCAD 0.02 0.91 0.01 0.99 0.01 1.00
RuleFit 0.20 0.51 0.12 0.72 0.07 0.82
RE-Lasso-100 0.72 0.17 0.59 0.30 0.45 0.49
RE-EN-100 0.88 0.16 0.80 0.30 0.67 0.50
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.62 0.47 0.44 0.85 0.26 1.00
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 36: Mean recall (RCL) and precision (PRC) for Scenario 2b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.59 0.28 0.50 0.42 0.36 0.62
Split-EN-10 0.70 0.29 0.67 0.40 0.53 0.61
Lasso 0.11 0.54 0.07 0.59 0.05 0.73
Elastic Net 0.19 0.53 0.15 0.61 0.10 0.76
Adaptive 0.11 0.51 0.07 0.60 0.05 0.72
Relaxed 0.09 0.67 0.07 0.73 0.04 0.83
0.4 MCP 0.03 0.55 0.02 0.70 0.01 0.71
SIS-SCAD 0.04 0.85 0.02 0.93 0.01 0.96
RuleFit 0.23 0.46 0.16 0.70 0.09 0.80
RE-Lasso-100 0.74 0.17 0.64 0.33 0.51 0.53
RE-EN-100 0.88 0.15 0.81 0.31 0.70 0.53
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.75 0.41 0.50 0.80 0.30 0.99
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.56 0.29 0.50 0.39 0.36 0.59
Split-EN-10 0.66 0.29 0.65 0.39 0.52 0.59
Lasso 0.10 0.49 0.07 0.60 0.04 0.74
Elastic Net 0.17 0.52 0.14 0.62 0.10 0.77
Adaptive 0.10 0.49 0.07 0.60 0.04 0.73
Relaxed 0.08 0.70 0.06 0.71 0.04 0.85
0.3 MCP 0.02 0.58 0.01 0.69 0.00 0.79
SIS-SCAD 0.03 0.84 0.01 0.95 0.01 0.97
RuleFit 0.21 0.43 0.14 0.68 0.08 0.82
RE-Lasso-100 0.71 0.17 0.60 0.33 0.49 0.53
RE-EN-100 0.86 0.15 0.80 0.31 0.68 0.53
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.72 0.42 0.46 0.80 0.29 0.99
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.59 0.25 0.49 0.38 0.36 0.56
Split-EN-10 0.76 0.21 0.64 0.38 0.49 0.57
Lasso 0.09 0.52 0.06 0.60 0.04 0.75
Elastic Net 0.18 0.52 0.14 0.65 0.09 0.76
Adaptive 0.09 0.52 0.06 0.60 0.04 0.74
Relaxed 0.08 0.66 0.05 0.77 0.03 0.87
0.2 MCP 0.01 0.80 0.01 0.90 0.00 0.92
SIS-SCAD 0.02 0.92 0.01 0.98 0.01 1.00
RuleFit 0.19 0.48 0.11 0.71 0.07 0.84
RE-Lasso-100 0.68 0.18 0.57 0.34 0.44 0.53
RE-EN-100 0.85 0.16 0.77 0.32 0.65 0.53
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.65 0.45 0.41 0.80 0.26 0.97
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 37: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.78 0.68 0.84 0.86 0.95 0.85 0.76 0.90 0.93 0.69 0.89 0.84 0.93 0.97 0.49
Split-EN-10 0.78 0.68 0.84 0.87 0.94 0.85 0.76 0.90 0.93 0.69 0.90 0.85 0.93 0.97 0.47
Lasso 0.74 0.62 0.82 0.82 1.09 0.80 0.69 0.87 0.88 0.92 0.82 0.74 0.88 0.91 0.78
Elastic Net 0.75 0.64 0.82 0.84 1.03 0.82 0.72 0.88 0.90 0.83 0.85 0.78 0.90 0.94 0.66
Adaptive 0.67 0.40 0.86 0.76 1.20 0.74 0.48 0.91 0.82 1.06 0.78 0.60 0.90 0.87 0.95
Relaxed 0.72 0.62 0.79 0.80 1.32 0.79 0.69 0.86 0.88 1.10 0.81 0.75 0.86 0.90 1.04
0.4 MCP 0.70 0.58 0.78 0.77 1.20 0.74 0.58 0.84 0.81 1.09 0.75 0.64 0.83 0.83 1.05
SIS-SCAD 0.69 0.58 0.77 0.76 1.28 0.71 0.58 0.80 0.78 1.27 0.71 0.59 0.79 0.79 1.27
RuleFit 0.71 0.60 0.79 0.78 1.55 0.73 0.59 0.83 0.81 1.38 0.75 0.64 0.82 0.83 1.17
RE-Lasso-100 0.77 0.67 0.84 0.86 0.95 0.84 0.73 0.91 0.93 0.71 0.88 0.81 0.93 0.96 0.55
RE-EN-100 0.77 0.68 0.84 0.86 0.95 0.84 0.75 0.91 0.93 0.69 0.89 0.83 0.93 0.97 0.51
RGLM-100 0.74 0.55 0.87 0.85 1.11 0.81 0.60 0.95 0.93 0.98 0.85 0.70 0.96 0.97 0.91
RF-500 0.77 0.68 0.84 0.86 0.95 0.83 0.71 0.92 0.92 0.81 0.87 0.77 0.94 0.96 0.76
XGB 0.69 0.58 0.77 0.74 1.20 0.73 0.59 0.82 0.80 1.08 0.74 0.65 0.80 0.81 1.06
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.80 0.58 0.90 0.87 0.87 0.86 0.68 0.94 0.94 0.61 0.89 0.74 0.96 0.97 0.48
Split-EN-10 0.80 0.58 0.90 0.87 0.87 0.86 0.69 0.94 0.94 0.60 0.90 0.75 0.96 0.97 0.46
Lasso 0.77 0.51 0.89 0.83 1.00 0.82 0.60 0.91 0.89 0.81 0.84 0.62 0.93 0.91 0.74
Elastic Net 0.78 0.54 0.89 0.84 0.96 0.84 0.63 0.92 0.91 0.73 0.86 0.67 0.94 0.94 0.64
Adaptive 0.72 0.22 0.96 0.74 1.11 0.76 0.30 0.96 0.80 1.00 0.77 0.33 0.96 0.85 0.93
Relaxed 0.76 0.53 0.87 0.82 1.12 0.82 0.61 0.90 0.88 1.06 0.82 0.64 0.90 0.89 1.32
0.3 MCP 0.73 0.43 0.88 0.77 1.12 0.76 0.46 0.89 0.80 1.05 0.77 0.44 0.91 0.80 1.03
SIS-SCAD 0.72 0.44 0.85 0.76 1.25 0.75 0.45 0.87 0.79 1.11 0.74 0.42 0.88 0.78 1.17
RuleFit 0.74 0.49 0.85 0.77 1.38 0.76 0.51 0.87 0.79 1.35 0.76 0.48 0.87 0.77 1.42
RE-Lasso-100 0.80 0.57 0.91 0.87 0.87 0.86 0.65 0.94 0.94 0.63 0.88 0.66 0.97 0.96 0.54
RE-EN-100 0.80 0.58 0.90 0.87 0.88 0.86 0.68 0.94 0.94 0.60 0.88 0.69 0.97 0.97 0.50
RGLM-100 0.77 0.37 0.96 0.86 1.01 0.81 0.41 0.98 0.94 0.88 0.81 0.39 0.99 0.97 0.84
RF-500 0.80 0.58 0.91 0.87 0.88 0.85 0.60 0.95 0.93 0.73 0.86 0.57 0.97 0.96 0.71
XGB 0.72 0.50 0.82 0.75 1.12 0.76 0.51 0.87 0.81 1.00 0.77 0.48 0.89 0.82 0.98
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.83 0.35 0.97 0.87 0.78 0.88 0.51 0.97 0.94 0.58 0.90 0.56 0.98 0.97 0.45
Split-EN-10 0.83 0.35 0.97 0.88 0.74 0.88 0.52 0.97 0.95 0.55 0.90 0.58 0.98 0.97 0.43
Lasso 0.81 0.30 0.96 0.82 0.88 0.85 0.43 0.96 0.89 0.71 0.86 0.43 0.97 0.91 0.67
Elastic Net 0.82 0.31 0.96 0.83 0.83 0.87 0.48 0.96 0.92 0.63 0.87 0.48 0.97 0.94 0.57
Adaptive 0.79 0.07 0.99 0.70 0.97 0.81 0.09 0.99 0.74 0.89 0.81 0.11 0.99 0.76 0.87
Relaxed 0.81 0.32 0.95 0.81 1.05 0.85 0.47 0.94 0.88 1.20 0.85 0.48 0.95 0.89 0.99
0.2 MCP 0.79 0.21 0.95 0.73 0.99 0.82 0.26 0.95 0.79 0.93 0.80 0.24 0.95 0.76 0.99
SIS-SCAD 0.79 0.23 0.95 0.77 0.96 0.81 0.26 0.94 0.78 1.01 0.80 0.20 0.95 0.77 0.94
RuleFit 0.79 0.31 0.92 0.73 1.27 0.81 0.35 0.93 0.77 1.16 0.81 0.30 0.93 0.76 1.25
RE-Lasso-100 0.83 0.36 0.97 0.88 0.77 0.88 0.51 0.97 0.94 0.52 0.89 0.51 0.99 0.97 0.47
RE-EN-100 0.83 0.38 0.96 0.88 0.76 0.89 0.55 0.97 0.95 0.50 0.90 0.60 0.98 0.97 0.42
RGLM-100 0.80 0.12 0.99 0.86 0.86 0.84 0.18 1.00 0.94 0.74 0.83 0.15 1.00 0.97 0.70
RF-500 0.83 0.33 0.97 0.87 0.76 0.87 0.39 0.98 0.94 0.65 0.86 0.33 1.00 0.96 0.61
XGB 0.79 0.32 0.93 0.77 0.93 0.82 0.34 0.94 0.81 0.86 0.82 0.29 0.95 0.82 0.83
Table 38: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.81 0.75 0.86 0.91 0.81 0.85 0.79 0.90 0.94 0.65 0.88 0.82 0.92 0.96 0.55
Split-EN-10 0.81 0.75 0.86 0.91 0.82 0.86 0.79 0.91 0.94 0.64 0.88 0.83 0.92 0.96 0.53
Lasso 0.78 0.71 0.84 0.87 0.95 0.80 0.70 0.87 0.88 0.88 0.82 0.73 0.88 0.91 0.80
Elastic Net 0.79 0.71 0.85 0.88 0.91 0.82 0.72 0.88 0.90 0.81 0.84 0.77 0.89 0.93 0.71
Adaptive 0.73 0.53 0.87 0.84 1.09 0.75 0.56 0.89 0.84 1.02 0.77 0.59 0.89 0.88 0.96
Relaxed 0.77 0.70 0.83 0.86 1.14 0.79 0.70 0.86 0.88 1.02 0.82 0.75 0.86 0.90 1.05
0.4 MCP 0.76 0.68 0.82 0.85 0.99 0.76 0.64 0.84 0.84 1.03 0.74 0.61 0.83 0.82 1.08
SIS-SCAD 0.76 0.69 0.80 0.84 1.05 0.73 0.61 0.82 0.81 1.15 0.71 0.61 0.79 0.79 1.28
RuleFit 0.76 0.70 0.79 0.83 1.31 0.75 0.63 0.84 0.83 1.23 0.74 0.61 0.83 0.81 1.36
RE-Lasso-100 0.81 0.73 0.86 0.90 0.83 0.84 0.75 0.90 0.93 0.70 0.86 0.78 0.92 0.95 0.60
RE-EN-100 0.81 0.73 0.86 0.90 0.83 0.84 0.76 0.90 0.93 0.69 0.87 0.79 0.92 0.96 0.57
RGLM-100 0.79 0.64 0.89 0.90 1.04 0.81 0.64 0.93 0.93 0.96 0.83 0.67 0.95 0.96 0.92
RF-500 0.82 0.77 0.86 0.91 0.81 0.84 0.74 0.91 0.93 0.77 0.85 0.75 0.92 0.94 0.77
XGB 0.75 0.70 0.79 0.83 1.04 0.74 0.63 0.82 0.82 1.07 0.74 0.62 0.82 0.82 1.07
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.82 0.60 0.92 0.90 0.81 0.86 0.68 0.94 0.94 0.61 0.88 0.74 0.95 0.96 0.51
Split-EN-10 0.81 0.59 0.92 0.90 0.81 0.87 0.69 0.94 0.94 0.60 0.89 0.76 0.95 0.96 0.49
Lasso 0.78 0.53 0.91 0.86 0.97 0.82 0.59 0.92 0.90 0.82 0.83 0.61 0.92 0.90 0.77
Elastic Net 0.79 0.56 0.91 0.87 0.91 0.83 0.62 0.93 0.91 0.74 0.85 0.67 0.93 0.93 0.66
Adaptive 0.74 0.29 0.96 0.82 1.05 0.76 0.28 0.97 0.83 0.98 0.77 0.33 0.96 0.83 0.94
Relaxed 0.78 0.54 0.89 0.84 1.09 0.81 0.60 0.90 0.88 1.39 0.82 0.63 0.91 0.90 0.97
0.3 MCP 0.76 0.51 0.88 0.83 1.00 0.78 0.47 0.92 0.84 0.93 0.76 0.42 0.91 0.80 1.04
SIS-SCAD 0.75 0.50 0.88 0.83 1.18 0.75 0.47 0.88 0.80 1.17 0.75 0.45 0.87 0.78 1.15
RuleFit 0.76 0.54 0.87 0.80 1.27 0.77 0.47 0.90 0.80 1.24 0.77 0.48 0.89 0.79 1.30
RE-Lasso-100 0.81 0.57 0.92 0.89 0.84 0.85 0.64 0.95 0.93 0.65 0.87 0.66 0.96 0.95 0.57
RE-EN-100 0.81 0.57 0.92 0.89 0.84 0.85 0.65 0.94 0.93 0.64 0.88 0.69 0.96 0.96 0.53
RGLM-100 0.77 0.39 0.96 0.89 0.98 0.80 0.39 0.98 0.93 0.89 0.82 0.43 0.99 0.96 0.84
RF-500 0.82 0.61 0.92 0.90 0.80 0.85 0.61 0.95 0.93 0.72 0.85 0.60 0.96 0.94 0.71
XGB 0.76 0.55 0.87 0.81 1.02 0.77 0.50 0.89 0.82 0.98 0.77 0.50 0.88 0.82 0.99
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.86 0.46 0.97 0.91 0.65 0.89 0.58 0.97 0.95 0.53 0.90 0.61 0.97 0.96 0.44
Split-EN-10 0.86 0.45 0.97 0.91 0.65 0.89 0.58 0.97 0.95 0.53 0.91 0.64 0.97 0.97 0.41
Lasso 0.83 0.39 0.95 0.86 0.78 0.85 0.47 0.95 0.90 0.69 0.86 0.47 0.96 0.91 0.63
Elastic Net 0.84 0.39 0.96 0.88 0.74 0.86 0.50 0.96 0.92 0.63 0.88 0.53 0.96 0.93 0.55
Adaptive 0.80 0.09 0.99 0.76 0.91 0.81 0.10 0.99 0.74 0.91 0.82 0.10 0.99 0.76 0.86
Relaxed 0.82 0.43 0.93 0.85 1.14 0.85 0.48 0.94 0.89 0.93 0.85 0.50 0.94 0.89 1.11
0.2 MCP 0.82 0.34 0.95 0.83 0.84 0.82 0.31 0.95 0.82 0.85 0.82 0.29 0.95 0.81 0.83
SIS-SCAD 0.81 0.30 0.95 0.83 0.89 0.81 0.29 0.95 0.81 0.85 0.81 0.22 0.95 0.78 0.89
RuleFit 0.81 0.41 0.92 0.80 1.02 0.82 0.42 0.92 0.79 1.12 0.81 0.34 0.92 0.74 1.32
RE-Lasso-100 0.85 0.42 0.97 0.90 0.68 0.89 0.57 0.97 0.94 0.51 0.90 0.56 0.98 0.96 0.45
RE-EN-100 0.85 0.43 0.97 0.90 0.67 0.89 0.59 0.97 0.95 0.50 0.91 0.62 0.98 0.97 0.41
RGLM-100 0.82 0.17 0.99 0.90 0.81 0.84 0.23 1.00 0.94 0.73 0.85 0.22 1.00 0.97 0.67
RF-500 0.86 0.46 0.97 0.91 0.66 0.88 0.48 0.98 0.94 0.62 0.88 0.41 0.99 0.95 0.59
XGB 0.82 0.45 0.92 0.83 0.83 0.82 0.39 0.93 0.83 0.83 0.83 0.34 0.95 0.83 0.79
Table 39: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.86 0.82 0.89 0.94 0.65 0.91 0.88 0.92 0.98 0.43 0.94 0.91 0.95 0.99 0.31
Split-EN-10 0.86 0.82 0.89 0.94 0.64 0.91 0.89 0.92 0.98 0.42 0.94 0.92 0.96 0.99 0.29
Lasso 0.84 0.78 0.87 0.92 0.77 0.88 0.85 0.89 0.96 0.58 0.90 0.86 0.93 0.97 0.47
Elastic Net 0.85 0.80 0.88 0.93 0.70 0.89 0.86 0.91 0.97 0.50 0.92 0.88 0.94 0.98 0.39
Adaptive 0.81 0.70 0.89 0.91 0.88 0.85 0.75 0.91 0.94 0.72 0.87 0.76 0.94 0.96 0.64
Relaxed 0.83 0.78 0.86 0.91 1.04 0.87 0.84 0.88 0.95 0.85 0.88 0.85 0.91 0.96 0.86
0.4 MCP 0.80 0.73 0.85 0.89 0.89 0.83 0.78 0.87 0.92 0.73 0.84 0.77 0.88 0.92 0.73
SIS-SCAD 0.78 0.71 0.84 0.87 0.96 0.81 0.74 0.85 0.90 0.83 0.81 0.72 0.86 0.90 0.82
RuleFit 0.79 0.73 0.83 0.86 1.26 0.81 0.75 0.85 0.88 1.09 0.81 0.72 0.88 0.88 1.08
RE-Lasso-100 0.86 0.81 0.89 0.94 0.66 0.90 0.87 0.92 0.97 0.44 0.93 0.90 0.95 0.99 0.35
RE-EN-100 0.86 0.81 0.89 0.94 0.66 0.90 0.88 0.92 0.97 0.43 0.93 0.91 0.95 0.99 0.32
RGLM-100 0.85 0.78 0.90 0.94 0.77 0.90 0.84 0.93 0.97 0.65 0.92 0.86 0.97 0.99 0.60
RF-500 0.86 0.81 0.89 0.94 0.67 0.90 0.86 0.92 0.97 0.55 0.92 0.87 0.95 0.99 0.52
XGB 0.80 0.74 0.84 0.89 0.88 0.83 0.77 0.87 0.92 0.75 0.83 0.74 0.89 0.92 0.75
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.87 0.74 0.92 0.94 0.60 0.91 0.82 0.95 0.98 0.40 0.94 0.87 0.96 0.99 0.30
Split-EN-10 0.87 0.74 0.92 0.94 0.59 0.91 0.82 0.95 0.98 0.40 0.94 0.88 0.96 0.99 0.28
Lasso 0.85 0.71 0.91 0.92 0.69 0.88 0.75 0.94 0.96 0.56 0.90 0.80 0.94 0.97 0.46
Elastic Net 0.86 0.72 0.92 0.93 0.64 0.90 0.78 0.95 0.97 0.48 0.92 0.83 0.95 0.98 0.37
Adaptive 0.81 0.47 0.96 0.91 0.83 0.83 0.51 0.97 0.94 0.73 0.86 0.59 0.97 0.96 0.64
Relaxed 0.84 0.71 0.90 0.91 1.06 0.88 0.77 0.93 0.95 0.94 0.89 0.80 0.93 0.96 0.88
0.3 MCP 0.82 0.63 0.90 0.89 0.87 0.82 0.60 0.92 0.90 0.83 0.83 0.62 0.92 0.90 0.77
SIS-SCAD 0.80 0.58 0.90 0.87 0.85 0.81 0.52 0.93 0.89 0.83 0.81 0.56 0.92 0.89 0.80
RuleFit 0.81 0.62 0.89 0.85 1.12 0.83 0.61 0.92 0.89 0.97 0.83 0.63 0.92 0.88 1.02
RE-Lasso-100 0.87 0.74 0.92 0.94 0.60 0.91 0.80 0.95 0.97 0.42 0.93 0.86 0.96 0.99 0.32
RE-EN-100 0.87 0.74 0.92 0.94 0.60 0.91 0.81 0.95 0.97 0.41 0.94 0.87 0.97 0.99 0.30
RGLM-100 0.86 0.66 0.95 0.94 0.71 0.89 0.69 0.98 0.98 0.60 0.91 0.73 0.99 0.99 0.56
RF-500 0.87 0.73 0.93 0.94 0.60 0.90 0.75 0.96 0.97 0.51 0.92 0.79 0.97 0.98 0.49
XGB 0.82 0.65 0.90 0.89 0.81 0.84 0.63 0.93 0.91 0.73 0.84 0.65 0.93 0.92 0.73
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.89 0.66 0.96 0.95 0.50 0.93 0.74 0.97 0.98 0.34 0.94 0.82 0.98 0.99 0.27
Split-EN-10 0.89 0.66 0.96 0.95 0.51 0.93 0.75 0.97 0.98 0.33 0.94 0.83 0.98 0.99 0.26
Lasso 0.87 0.59 0.95 0.93 0.60 0.90 0.65 0.96 0.96 0.46 0.92 0.73 0.96 0.97 0.39
Elastic Net 0.88 0.61 0.95 0.94 0.57 0.91 0.70 0.97 0.97 0.40 0.93 0.78 0.97 0.98 0.32
Adaptive 0.83 0.28 0.98 0.88 0.74 0.85 0.31 0.99 0.91 0.66 0.87 0.46 0.98 0.94 0.57
Relaxed 0.87 0.60 0.94 0.92 0.95 0.90 0.66 0.96 0.95 0.62 0.90 0.75 0.94 0.96 0.80
0.2 MCP 0.83 0.42 0.94 0.86 0.79 0.84 0.39 0.96 0.87 0.74 0.84 0.47 0.94 0.88 0.74
SIS-SCAD 0.83 0.36 0.95 0.87 0.80 0.83 0.28 0.97 0.89 0.76 0.84 0.40 0.96 0.89 0.73
RuleFit 0.84 0.48 0.94 0.85 0.93 0.86 0.48 0.96 0.88 0.90 0.86 0.51 0.96 0.89 0.85
RE-Lasso-100 0.89 0.66 0.95 0.95 0.50 0.93 0.75 0.97 0.98 0.33 0.94 0.82 0.97 0.99 0.28
RE-EN-100 0.89 0.66 0.95 0.95 0.50 0.93 0.77 0.97 0.98 0.32 0.95 0.84 0.97 0.99 0.26
RGLM-100 0.88 0.52 0.98 0.95 0.58 0.90 0.53 0.99 0.98 0.50 0.91 0.61 1.00 0.99 0.48
RF-500 0.89 0.62 0.96 0.95 0.51 0.91 0.63 0.98 0.98 0.44 0.92 0.70 0.98 0.98 0.43
XGB 0.85 0.54 0.94 0.89 0.70 0.87 0.51 0.96 0.92 0.64 0.87 0.55 0.95 0.91 0.66
Table 40: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.79 0.72 0.85 0.88 0.88 0.89 0.84 0.92 0.96 0.51 0.90 0.87 0.93 0.97 0.44
Split-EN-10 0.79 0.72 0.85 0.88 0.88 0.88 0.83 0.92 0.96 0.52 0.91 0.87 0.93 0.97 0.43
Lasso 0.78 0.71 0.83 0.87 0.93 0.85 0.78 0.89 0.93 0.67 0.86 0.80 0.90 0.94 0.65
Elastic Net 0.78 0.71 0.84 0.87 0.92 0.86 0.79 0.90 0.94 0.62 0.87 0.82 0.91 0.95 0.57
Adaptive 0.77 0.63 0.87 0.87 1.01 0.83 0.69 0.91 0.92 0.80 0.84 0.73 0.91 0.93 0.75
Relaxed 0.77 0.70 0.82 0.86 0.97 0.84 0.77 0.89 0.92 0.77 0.86 0.81 0.89 0.94 0.73
0.4 MCP 0.76 0.69 0.82 0.84 0.99 0.80 0.69 0.86 0.88 0.88 0.80 0.72 0.86 0.89 0.86
SIS-SCAD 0.76 0.69 0.82 0.85 1.00 0.77 0.66 0.84 0.85 0.97 0.76 0.67 0.83 0.84 1.00
RuleFit 0.76 0.69 0.82 0.84 1.31 0.80 0.68 0.88 0.89 0.96 0.80 0.70 0.87 0.89 0.96
RE-Lasso-100 0.80 0.72 0.85 0.88 0.87 0.88 0.81 0.92 0.96 0.54 0.90 0.84 0.93 0.97 0.47
RE-EN-100 0.79 0.72 0.85 0.88 0.88 0.88 0.81 0.92 0.96 0.53 0.90 0.85 0.93 0.97 0.46
RGLM-100 0.76 0.61 0.87 0.86 1.08 0.85 0.69 0.95 0.95 0.88 0.88 0.75 0.96 0.97 0.84
RF-500 0.79 0.74 0.83 0.88 0.89 0.88 0.79 0.93 0.96 0.66 0.89 0.82 0.94 0.97 0.64
XGB 0.75 0.68 0.81 0.83 1.03 0.80 0.67 0.88 0.88 0.86 0.80 0.70 0.87 0.89 0.84
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.80 0.59 0.91 0.88 0.84 0.90 0.79 0.94 0.96 0.47 0.91 0.82 0.95 0.97 0.41
Split-EN-10 0.80 0.58 0.91 0.88 0.84 0.90 0.79 0.94 0.96 0.47 0.91 0.82 0.95 0.97 0.40
Lasso 0.79 0.57 0.90 0.86 0.90 0.87 0.72 0.93 0.94 0.62 0.87 0.73 0.93 0.94 0.59
Elastic Net 0.79 0.58 0.90 0.87 0.89 0.88 0.74 0.93 0.94 0.57 0.88 0.76 0.94 0.95 0.53
Adaptive 0.76 0.37 0.96 0.85 0.98 0.83 0.53 0.96 0.92 0.74 0.84 0.61 0.95 0.93 0.70
Relaxed 0.79 0.57 0.89 0.85 0.96 0.86 0.73 0.92 0.93 0.72 0.87 0.74 0.92 0.94 0.66
0.3 MCP 0.77 0.53 0.89 0.83 0.98 0.82 0.59 0.91 0.88 0.80 0.81 0.60 0.90 0.88 0.82
SIS-SCAD 0.78 0.54 0.89 0.84 0.97 0.80 0.56 0.89 0.85 0.90 0.78 0.56 0.88 0.84 0.96
RuleFit 0.77 0.56 0.88 0.83 1.34 0.82 0.56 0.93 0.88 0.96 0.82 0.56 0.93 0.88 0.97
RE-Lasso-100 0.80 0.58 0.91 0.88 0.86 0.89 0.74 0.95 0.96 0.50 0.90 0.76 0.96 0.97 0.45
RE-EN-100 0.80 0.58 0.91 0.88 0.86 0.89 0.75 0.95 0.96 0.49 0.90 0.78 0.96 0.97 0.43
RGLM-100 0.76 0.38 0.95 0.85 1.01 0.84 0.49 0.99 0.95 0.80 0.85 0.53 0.99 0.97 0.78
RF-500 0.80 0.63 0.89 0.88 0.84 0.88 0.71 0.96 0.96 0.61 0.89 0.72 0.97 0.97 0.60
XGB 0.76 0.56 0.86 0.82 0.99 0.82 0.56 0.93 0.88 0.79 0.81 0.56 0.92 0.88 0.81
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.83 0.41 0.96 0.88 0.73 0.91 0.68 0.97 0.97 0.39 0.93 0.74 0.98 0.98 0.33
Split-EN-10 0.83 0.40 0.96 0.88 0.74 0.91 0.67 0.97 0.97 0.39 0.93 0.75 0.98 0.98 0.32
Lasso 0.82 0.40 0.95 0.86 0.78 0.88 0.55 0.96 0.93 0.54 0.89 0.61 0.96 0.94 0.51
Elastic Net 0.83 0.40 0.96 0.87 0.77 0.89 0.58 0.97 0.94 0.50 0.90 0.65 0.97 0.96 0.45
Adaptive 0.79 0.12 0.99 0.84 0.90 0.85 0.26 0.99 0.90 0.67 0.85 0.34 0.99 0.92 0.65
Relaxed 0.82 0.41 0.95 0.86 0.80 0.88 0.57 0.96 0.93 0.58 0.89 0.62 0.96 0.94 0.58
0.2 MCP 0.81 0.33 0.95 0.82 0.89 0.84 0.37 0.96 0.86 0.74 0.84 0.40 0.95 0.86 0.75
SIS-SCAD 0.81 0.36 0.95 0.84 0.87 0.83 0.36 0.95 0.85 0.79 0.83 0.40 0.94 0.84 0.86
RuleFit 0.80 0.40 0.93 0.81 1.26 0.85 0.39 0.96 0.87 0.89 0.84 0.40 0.96 0.87 0.96
RE-Lasso-100 0.83 0.40 0.96 0.88 0.77 0.90 0.57 0.98 0.96 0.44 0.91 0.62 0.99 0.97 0.40
RE-EN-100 0.83 0.40 0.96 0.88 0.77 0.90 0.60 0.98 0.96 0.43 0.92 0.66 0.98 0.98 0.37
RGLM-100 0.80 0.18 0.99 0.85 0.87 0.85 0.24 1.00 0.96 0.66 0.85 0.28 1.00 0.98 0.65
RF-500 0.83 0.48 0.94 0.87 0.74 0.89 0.51 0.99 0.96 0.51 0.90 0.54 0.99 0.97 0.51
XGB 0.81 0.41 0.93 0.82 0.86 0.85 0.38 0.97 0.88 0.68 0.84 0.38 0.96 0.88 0.71
Table 41: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.84 0.76 0.89 0.92 0.71 0.89 0.85 0.91 0.96 0.51 0.90 0.88 0.92 0.97 0.44
Split-EN-10 0.83 0.76 0.88 0.92 0.72 0.89 0.84 0.91 0.96 0.52 0.91 0.88 0.92 0.97 0.43
Lasso 0.82 0.73 0.87 0.90 0.80 0.85 0.79 0.90 0.94 0.66 0.86 0.81 0.89 0.94 0.64
Elastic Net 0.82 0.73 0.88 0.91 0.79 0.86 0.80 0.90 0.94 0.63 0.87 0.83 0.90 0.95 0.58
Adaptive 0.80 0.65 0.91 0.90 0.90 0.84 0.72 0.91 0.93 0.78 0.84 0.75 0.90 0.93 0.75
Relaxed 0.81 0.73 0.87 0.90 0.84 0.85 0.79 0.89 0.93 0.72 0.86 0.82 0.89 0.94 0.73
0.4 MCP 0.80 0.71 0.86 0.89 0.85 0.83 0.75 0.88 0.92 0.75 0.81 0.73 0.86 0.89 0.83
SIS-SCAD 0.81 0.73 0.86 0.89 0.89 0.81 0.72 0.87 0.89 0.87 0.76 0.67 0.81 0.83 1.07
RuleFit 0.80 0.72 0.85 0.88 1.15 0.81 0.72 0.88 0.90 0.94 0.80 0.70 0.86 0.89 0.98
RE-Lasso-100 0.83 0.75 0.89 0.92 0.74 0.87 0.81 0.91 0.95 0.57 0.89 0.84 0.92 0.96 0.50
RE-EN-100 0.83 0.75 0.88 0.92 0.74 0.87 0.81 0.91 0.95 0.57 0.89 0.85 0.92 0.97 0.48
RGLM-100 0.47 0.45 0.48 0.92 1.65 0.61 0.57 0.64 0.95 1.38 0.59 0.57 0.60 0.96 1.46
RF-500 0.84 0.78 0.88 0.92 0.71 0.88 0.83 0.92 0.96 0.58 0.88 0.82 0.92 0.96 0.64
XGB 0.80 0.72 0.85 0.87 0.90 0.81 0.72 0.87 0.89 0.83 0.80 0.70 0.87 0.89 0.85
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.70 0.92 0.92 0.68 0.89 0.76 0.95 0.96 0.48 0.92 0.82 0.96 0.97 0.39
Split-EN-10 0.85 0.69 0.92 0.92 0.68 0.89 0.75 0.95 0.96 0.49 0.92 0.82 0.96 0.98 0.37
Lasso 0.82 0.65 0.91 0.90 0.77 0.86 0.69 0.93 0.93 0.63 0.87 0.72 0.94 0.94 0.58
Elastic Net 0.83 0.65 0.91 0.91 0.75 0.87 0.70 0.94 0.94 0.60 0.89 0.74 0.95 0.95 0.53
Adaptive 0.81 0.52 0.95 0.90 0.83 0.84 0.54 0.96 0.92 0.72 0.84 0.56 0.96 0.93 0.71
Relaxed 0.82 0.66 0.90 0.90 0.77 0.86 0.69 0.93 0.93 0.71 0.87 0.73 0.93 0.94 0.65
0.3 MCP 0.81 0.62 0.90 0.88 0.81 0.84 0.64 0.92 0.91 0.73 0.82 0.60 0.92 0.89 0.78
SIS-SCAD 0.81 0.62 0.90 0.89 0.86 0.82 0.59 0.92 0.89 0.81 0.79 0.55 0.90 0.85 0.99
RuleFit 0.81 0.64 0.90 0.88 1.09 0.83 0.57 0.93 0.89 0.95 0.82 0.54 0.94 0.88 0.95
RE-Lasso-100 0.84 0.67 0.92 0.92 0.71 0.87 0.69 0.95 0.95 0.55 0.90 0.75 0.97 0.97 0.44
RE-EN-100 0.84 0.66 0.92 0.92 0.72 0.87 0.70 0.95 0.95 0.55 0.91 0.77 0.97 0.97 0.42
RGLM-100 0.56 0.49 0.59 0.92 1.56 0.66 0.48 0.73 0.94 1.33 0.67 0.51 0.74 0.97 1.30
RF-500 0.85 0.72 0.91 0.92 0.66 0.88 0.72 0.95 0.96 0.55 0.89 0.71 0.97 0.96 0.57
XGB 0.82 0.65 0.90 0.88 0.83 0.83 0.57 0.93 0.89 0.76 0.82 0.54 0.94 0.89 0.78
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.87 0.54 0.96 0.93 0.59 0.91 0.62 0.98 0.96 0.43 0.92 0.74 0.97 0.97 0.36
Split-EN-10 0.87 0.53 0.96 0.93 0.59 0.91 0.61 0.98 0.96 0.43 0.92 0.75 0.97 0.98 0.35
Lasso 0.85 0.48 0.96 0.90 0.67 0.88 0.52 0.97 0.93 0.56 0.89 0.63 0.96 0.94 0.54
Elastic Net 0.85 0.48 0.96 0.91 0.65 0.89 0.53 0.97 0.93 0.53 0.90 0.65 0.96 0.95 0.48
Adaptive 0.81 0.20 0.99 0.89 0.78 0.85 0.22 0.99 0.90 0.67 0.85 0.37 0.98 0.92 0.65
Relaxed 0.85 0.51 0.94 0.89 0.75 0.88 0.53 0.96 0.92 0.65 0.89 0.65 0.95 0.94 0.63
0.2 MCP 0.84 0.49 0.94 0.88 0.71 0.86 0.41 0.96 0.88 0.67 0.84 0.45 0.95 0.87 0.73
SIS-SCAD 0.83 0.42 0.95 0.89 0.75 0.85 0.37 0.96 0.87 0.75 0.82 0.41 0.94 0.84 0.88
RuleFit 0.84 0.49 0.94 0.87 1.01 0.85 0.39 0.96 0.87 0.92 0.84 0.42 0.96 0.87 0.93
RE-Lasso-100 0.86 0.50 0.96 0.92 0.63 0.89 0.51 0.98 0.95 0.50 0.90 0.63 0.98 0.97 0.43
RE-EN-100 0.86 0.49 0.96 0.92 0.64 0.89 0.52 0.98 0.95 0.50 0.91 0.66 0.98 0.97 0.41
RGLM-100 0.68 0.34 0.78 0.92 1.28 0.71 0.31 0.80 0.94 1.27 0.70 0.36 0.80 0.97 1.19
RF-500 0.87 0.59 0.96 0.93 0.58 0.90 0.53 0.98 0.95 0.49 0.89 0.57 0.98 0.96 0.52
XGB 0.85 0.52 0.94 0.88 0.72 0.85 0.39 0.96 0.88 0.69 0.84 0.42 0.96 0.87 0.72
Table 42: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.87 0.82 0.90 0.95 0.60 0.92 0.89 0.93 0.98 0.39 0.94 0.92 0.95 0.99 0.27
Split-EN-10 0.86 0.82 0.90 0.95 0.60 0.92 0.89 0.94 0.98 0.38 0.94 0.92 0.95 0.99 0.27
Lasso 0.85 0.80 0.89 0.94 0.66 0.89 0.85 0.92 0.97 0.49 0.91 0.88 0.93 0.98 0.40
Elastic Net 0.86 0.80 0.90 0.94 0.65 0.90 0.86 0.93 0.97 0.46 0.92 0.89 0.94 0.98 0.35
Adaptive 0.84 0.75 0.91 0.93 0.75 0.88 0.81 0.93 0.96 0.59 0.90 0.84 0.94 0.97 0.50
Relaxed 0.85 0.79 0.89 0.93 0.72 0.89 0.85 0.91 0.96 0.58 0.91 0.88 0.92 0.97 0.54
0.4 MCP 0.83 0.78 0.88 0.92 0.75 0.86 0.80 0.90 0.94 0.64 0.87 0.82 0.91 0.95 0.59
SIS-SCAD 0.83 0.77 0.87 0.92 0.75 0.84 0.78 0.89 0.93 0.71 0.85 0.80 0.88 0.93 0.69
RuleFit 0.84 0.77 0.89 0.92 0.90 0.86 0.79 0.91 0.94 0.68 0.87 0.81 0.91 0.95 0.63
RE-Lasso-100 0.86 0.81 0.90 0.95 0.63 0.91 0.87 0.93 0.97 0.41 0.93 0.90 0.95 0.99 0.31
RE-EN-100 0.86 0.81 0.90 0.95 0.63 0.91 0.87 0.93 0.97 0.41 0.94 0.90 0.96 0.99 0.30
RGLM-100 0.86 0.79 0.90 0.94 0.71 0.90 0.83 0.95 0.97 0.61 0.93 0.88 0.97 0.99 0.55
RF-500 0.87 0.82 0.90 0.95 0.59 0.91 0.86 0.94 0.98 0.47 0.93 0.89 0.96 0.99 0.44
XGB 0.84 0.78 0.88 0.92 0.72 0.87 0.80 0.91 0.95 0.59 0.88 0.81 0.92 0.96 0.55
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.89 0.78 0.93 0.96 0.51 0.93 0.88 0.95 0.98 0.31 0.94 0.91 0.96 0.99 0.26
Split-EN-10 0.89 0.78 0.93 0.96 0.52 0.93 0.88 0.95 0.98 0.31 0.95 0.91 0.96 0.99 0.25
Lasso 0.87 0.75 0.93 0.94 0.59 0.91 0.82 0.94 0.97 0.42 0.92 0.85 0.95 0.98 0.38
Elastic Net 0.88 0.76 0.93 0.95 0.56 0.92 0.84 0.95 0.97 0.38 0.93 0.87 0.95 0.98 0.33
Adaptive 0.85 0.61 0.96 0.94 0.69 0.89 0.72 0.96 0.96 0.54 0.90 0.76 0.96 0.97 0.49
Relaxed 0.87 0.74 0.92 0.94 0.64 0.90 0.82 0.94 0.97 0.54 0.91 0.85 0.94 0.97 0.49
0.3 MCP 0.85 0.69 0.92 0.92 0.71 0.87 0.72 0.93 0.94 0.60 0.87 0.73 0.92 0.94 0.61
SIS-SCAD 0.85 0.68 0.92 0.92 0.76 0.86 0.69 0.92 0.93 0.66 0.85 0.70 0.92 0.93 0.66
RuleFit 0.86 0.71 0.92 0.93 0.82 0.87 0.72 0.94 0.94 0.66 0.87 0.72 0.94 0.94 0.65
RE-Lasso-100 0.88 0.77 0.93 0.95 0.54 0.92 0.85 0.95 0.98 0.34 0.94 0.88 0.96 0.99 0.29
RE-EN-100 0.88 0.77 0.93 0.95 0.54 0.92 0.85 0.95 0.98 0.34 0.94 0.89 0.96 0.99 0.28
RGLM-100 0.87 0.69 0.95 0.95 0.65 0.91 0.77 0.97 0.98 0.52 0.93 0.82 0.98 0.99 0.51
RF-500 0.89 0.78 0.94 0.96 0.52 0.92 0.83 0.96 0.98 0.41 0.93 0.85 0.97 0.99 0.41
XGB 0.86 0.71 0.93 0.93 0.65 0.88 0.73 0.94 0.95 0.54 0.88 0.75 0.94 0.95 0.54
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.90 0.69 0.96 0.96 0.44 0.94 0.81 0.97 0.98 0.28 0.95 0.85 0.98 0.99 0.23
Split-EN-10 0.90 0.69 0.96 0.96 0.45 0.94 0.81 0.97 0.98 0.28 0.95 0.86 0.98 0.99 0.22
Lasso 0.89 0.63 0.96 0.94 0.51 0.92 0.73 0.97 0.97 0.38 0.93 0.76 0.97 0.98 0.34
Elastic Net 0.89 0.65 0.96 0.95 0.49 0.93 0.75 0.97 0.97 0.34 0.93 0.78 0.97 0.98 0.29
Adaptive 0.86 0.38 0.99 0.94 0.61 0.90 0.52 0.99 0.96 0.47 0.90 0.58 0.99 0.97 0.45
Relaxed 0.89 0.63 0.96 0.94 0.60 0.92 0.74 0.96 0.97 0.47 0.92 0.77 0.96 0.97 0.50
0.2 MCP 0.86 0.47 0.96 0.90 0.65 0.87 0.52 0.96 0.91 0.63 0.86 0.53 0.95 0.91 0.65
SIS-SCAD 0.86 0.45 0.97 0.91 0.66 0.87 0.45 0.97 0.92 0.64 0.86 0.49 0.96 0.92 0.68
RuleFit 0.87 0.55 0.96 0.91 0.79 0.88 0.56 0.96 0.90 0.74 0.88 0.56 0.97 0.91 0.70
RE-Lasso-100 0.90 0.67 0.96 0.95 0.47 0.93 0.75 0.97 0.98 0.31 0.94 0.79 0.98 0.99 0.26
RE-EN-100 0.90 0.68 0.96 0.95 0.47 0.93 0.77 0.97 0.98 0.30 0.94 0.81 0.98 0.99 0.25
RGLM-100 0.89 0.55 0.98 0.95 0.54 0.92 0.61 0.99 0.98 0.44 0.92 0.64 1.00 0.99 0.43
RF-500 0.90 0.67 0.96 0.96 0.46 0.93 0.71 0.98 0.98 0.36 0.93 0.72 0.99 0.99 0.36
XGB 0.88 0.57 0.96 0.93 0.58 0.90 0.59 0.97 0.95 0.48 0.89 0.60 0.97 0.95 0.49
Table 43: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.30 0.22 0.30 0.31 0.29 0.49
Split-EN-10 0.33 0.22 0.34 0.31 0.39 0.48
Lasso 0.05 0.32 0.04 0.36 0.03 0.49
Elastic Net 0.08 0.28 0.07 0.36 0.06 0.49
Adaptive 0.05 0.27 0.04 0.36 0.03 0.48
Relaxed 0.04 0.34 0.04 0.37 0.03 0.50
0.4 MCP 0.02 0.36 0.01 0.35 0.01 0.51
SIS-SCAD 0.01 0.49 0.01 0.41 0.00 0.57
RuleFit 0.12 0.24 0.08 0.36 0.06 0.51
RE-Lasso-100 0.54 0.13 0.47 0.24 0.44 0.43
RE-EN-100 0.65 0.12 0.62 0.23 0.58 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.34 0.19 0.23 0.29 0.19 0.48
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.29 0.23 0.30 0.33 0.28 0.50
Split-EN-10 0.32 0.23 0.35 0.31 0.36 0.48
Lasso 0.05 0.29 0.04 0.36 0.03 0.47
Elastic Net 0.07 0.27 0.07 0.35 0.05 0.48
Adaptive 0.05 0.29 0.03 0.35 0.02 0.46
Relaxed 0.05 0.30 0.04 0.37 0.02 0.47
0.3 MCP 0.02 0.36 0.01 0.43 0.01 0.50
SIS-SCAD 0.01 0.49 0.01 0.50 0.00 0.51
RuleFit 0.12 0.26 0.09 0.41 0.06 0.55
RE-Lasso-100 0.53 0.14 0.46 0.25 0.41 0.44
RE-EN-100 0.64 0.12 0.59 0.23 0.55 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.32 0.19 0.23 0.31 0.17 0.48
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.25 0.27 0.26 0.35 0.25 0.52
Split-EN-10 0.35 0.21 0.32 0.32 0.34 0.50
Lasso 0.04 0.29 0.03 0.34 0.02 0.47
Elastic Net 0.07 0.26 0.06 0.34 0.05 0.48
Adaptive 0.04 0.29 0.03 0.35 0.02 0.47
Relaxed 0.04 0.31 0.03 0.40 0.02 0.50
0.2 MCP 0.01 0.35 0.01 0.39 0.00 0.49
SIS-SCAD 0.01 0.50 0.01 0.42 0.00 0.52
RuleFit 0.12 0.30 0.08 0.45 0.05 0.60
RE-Lasso-100 0.47 0.14 0.42 0.25 0.36 0.43
RE-EN-100 0.60 0.13 0.56 0.24 0.51 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.28 0.19 0.20 0.31 0.14 0.47
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 44: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.40 0.24 0.37 0.38 0.29 0.50
Split-EN-10 0.45 0.23 0.49 0.35 0.39 0.50
Lasso 0.05 0.31 0.04 0.37 0.02 0.40
Elastic Net 0.10 0.34 0.08 0.42 0.06 0.47
Adaptive 0.05 0.32 0.04 0.38 0.02 0.39
Relaxed 0.05 0.45 0.03 0.41 0.02 0.42
0.4 MCP 0.02 0.37 0.01 0.41 0.01 0.47
SIS-SCAD 0.02 0.70 0.01 0.65 0.00 0.56
RuleFit 0.14 0.29 0.10 0.41 0.06 0.53
RE-Lasso-100 0.53 0.13 0.48 0.25 0.40 0.41
RE-EN-100 0.67 0.12 0.64 0.24 0.56 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.40 0.25 0.30 0.39 0.20 0.52
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.37 0.24 0.33 0.37 0.27 0.49
Split-EN-10 0.43 0.23 0.44 0.35 0.38 0.49
Lasso 0.05 0.31 0.04 0.37 0.02 0.36
Elastic Net 0.10 0.34 0.07 0.40 0.05 0.43
Adaptive 0.05 0.28 0.03 0.34 0.02 0.35
Relaxed 0.04 0.44 0.03 0.40 0.02 0.39
0.3 MCP 0.02 0.38 0.01 0.45 0.00 0.41
SIS-SCAD 0.02 0.64 0.01 0.57 0.00 0.53
RuleFit 0.13 0.28 0.09 0.42 0.06 0.52
RE-Lasso-100 0.50 0.13 0.45 0.25 0.38 0.40
RE-EN-100 0.63 0.12 0.62 0.24 0.55 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.36 0.24 0.28 0.38 0.18 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.34 0.26 0.31 0.36 0.26 0.51
Split-EN-10 0.43 0.24 0.40 0.35 0.37 0.51
Lasso 0.04 0.30 0.03 0.33 0.02 0.37
Elastic Net 0.08 0.34 0.06 0.38 0.04 0.42
Adaptive 0.04 0.28 0.03 0.33 0.02 0.38
Relaxed 0.03 0.41 0.02 0.35 0.02 0.38
0.2 MCP 0.01 0.45 0.01 0.43 0.00 0.47
SIS-SCAD 0.02 0.68 0.01 0.55 0.00 0.49
RuleFit 0.12 0.34 0.08 0.42 0.05 0.55
RE-Lasso-100 0.47 0.14 0.41 0.24 0.35 0.41
RE-EN-100 0.64 0.13 0.58 0.24 0.53 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.34 0.25 0.24 0.36 0.15 0.48
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 45: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.28 0.22 0.30 0.34 0.24 0.50
Split-EN-10 0.34 0.22 0.41 0.33 0.36 0.49
Lasso 0.04 0.24 0.03 0.31 0.02 0.40
Elastic Net 0.08 0.25 0.06 0.34 0.05 0.43
Adaptive 0.04 0.24 0.03 0.31 0.02 0.39
Relaxed 0.04 0.29 0.03 0.31 0.02 0.39
0.4 MCP 0.01 0.30 0.01 0.35 0.00 0.32
SIS-SCAD 0.01 0.42 0.01 0.49 0.00 0.40
RuleFit 0.11 0.30 0.08 0.44 0.04 0.54
RE-Lasso-100 0.44 0.14 0.41 0.24 0.36 0.41
RE-EN-100 0.60 0.13 0.58 0.24 0.54 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.27 0.20 0.20 0.33 0.14 0.46
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.30 0.23 0.29 0.35 0.23 0.49
Split-EN-10 0.36 0.24 0.38 0.33 0.35 0.49
Lasso 0.04 0.27 0.03 0.32 0.02 0.38
Elastic Net 0.08 0.26 0.06 0.34 0.04 0.40
Adaptive 0.04 0.24 0.02 0.29 0.02 0.35
Relaxed 0.04 0.31 0.03 0.35 0.02 0.37
0.3 MCP 0.01 0.33 0.01 0.36 0.00 0.40
SIS-SCAD 0.01 0.44 0.01 0.44 0.00 0.46
RuleFit 0.10 0.31 0.07 0.44 0.04 0.54
RE-Lasso-100 0.43 0.14 0.38 0.25 0.33 0.40
RE-EN-100 0.59 0.13 0.55 0.24 0.51 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.27 0.21 0.19 0.32 0.13 0.44
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.30 0.23 0.28 0.36 0.24 0.51
Split-EN-10 0.36 0.24 0.37 0.34 0.32 0.51
Lasso 0.03 0.21 0.02 0.29 0.02 0.36
Elastic Net 0.06 0.23 0.05 0.32 0.04 0.38
Adaptive 0.03 0.22 0.02 0.29 0.01 0.36
Relaxed 0.03 0.23 0.02 0.29 0.01 0.38
0.2 MCP 0.01 0.29 0.00 0.40 0.00 0.43
SIS-SCAD 0.01 0.35 0.00 0.43 0.00 0.43
RuleFit 0.08 0.30 0.05 0.42 0.03 0.53
RE-Lasso-100 0.40 0.14 0.35 0.24 0.30 0.38
RE-EN-100 0.56 0.13 0.52 0.23 0.48 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.23 0.20 0.15 0.29 0.11 0.43
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 46: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.40 0.22 0.49 0.30 0.42 0.48
Split-EN-10 0.39 0.21 0.55 0.28 0.51 0.46
Lasso 0.09 0.37 0.08 0.41 0.05 0.49
Elastic Net 0.12 0.35 0.12 0.40 0.08 0.50
Adaptive 0.08 0.37 0.08 0.40 0.05 0.48
Relaxed 0.06 0.63 0.07 0.45 0.05 0.50
0.4 MCP 0.04 0.47 0.02 0.44 0.01 0.50
SIS-SCAD 0.03 0.64 0.02 0.66 0.01 0.62
RuleFit 0.22 0.24 0.18 0.39 0.12 0.52
RE-Lasso-100 0.67 0.11 0.63 0.25 0.56 0.43
RE-EN-100 0.75 0.10 0.72 0.23 0.67 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.46 0.13 0.39 0.32 0.30 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.32 0.25 0.49 0.31 0.44 0.49
Split-EN-10 0.34 0.23 0.55 0.29 0.51 0.47
Lasso 0.09 0.38 0.08 0.40 0.05 0.51
Elastic Net 0.11 0.35 0.11 0.40 0.08 0.51
Adaptive 0.09 0.37 0.07 0.40 0.05 0.51
Relaxed 0.07 0.58 0.07 0.41 0.05 0.52
0.3 MCP 0.03 0.48 0.02 0.43 0.01 0.54
SIS-SCAD 0.03 0.65 0.01 0.59 0.01 0.66
RuleFit 0.20 0.22 0.16 0.38 0.12 0.54
RE-Lasso-100 0.63 0.11 0.61 0.25 0.55 0.44
RE-EN-100 0.72 0.10 0.70 0.24 0.65 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.44 0.14 0.36 0.32 0.29 0.50
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.29 0.24 0.45 0.30 0.41 0.49
Split-EN-10 0.28 0.27 0.53 0.28 0.47 0.48
Lasso 0.08 0.40 0.06 0.36 0.04 0.48
Elastic Net 0.10 0.36 0.08 0.35 0.07 0.48
Adaptive 0.08 0.37 0.06 0.37 0.04 0.47
Relaxed 0.06 0.58 0.06 0.38 0.04 0.49
0.2 MCP 0.02 0.55 0.01 0.42 0.01 0.49
SIS-SCAD 0.03 0.68 0.01 0.52 0.01 0.53
RuleFit 0.18 0.24 0.13 0.38 0.09 0.54
RE-Lasso-100 0.58 0.11 0.54 0.24 0.49 0.43
RE-EN-100 0.67 0.11 0.65 0.23 0.60 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.39 0.14 0.31 0.31 0.24 0.48
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 47: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.49 0.22 0.52 0.32 0.45 0.50
Split-EN-10 0.59 0.19 0.60 0.31 0.54 0.49
Lasso 0.08 0.34 0.07 0.43 0.05 0.44
Elastic Net 0.13 0.36 0.12 0.46 0.08 0.50
Adaptive 0.08 0.35 0.07 0.41 0.04 0.44
Relaxed 0.07 0.51 0.07 0.47 0.04 0.44
0.4 MCP 0.03 0.36 0.02 0.45 0.01 0.46
SIS-SCAD 0.03 0.75 0.02 0.77 0.01 0.68
RuleFit 0.24 0.28 0.19 0.43 0.12 0.52
RE-Lasso-100 0.65 0.12 0.62 0.24 0.54 0.41
RE-EN-100 0.77 0.11 0.73 0.22 0.67 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.58 0.21 0.47 0.42 0.32 0.54
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.51 0.20 0.52 0.32 0.43 0.49
Split-EN-10 0.56 0.21 0.60 0.30 0.53 0.49
Lasso 0.08 0.31 0.07 0.41 0.04 0.43
Elastic Net 0.13 0.35 0.11 0.43 0.07 0.48
Adaptive 0.08 0.31 0.06 0.39 0.04 0.43
Relaxed 0.07 0.42 0.06 0.46 0.04 0.44
0.3 MCP 0.02 0.36 0.02 0.53 0.01 0.46
SIS-SCAD 0.03 0.70 0.02 0.77 0.01 0.58
RuleFit 0.22 0.28 0.18 0.45 0.11 0.51
RE-Lasso-100 0.65 0.12 0.58 0.24 0.52 0.41
RE-EN-100 0.75 0.11 0.70 0.23 0.65 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.54 0.22 0.44 0.43 0.31 0.53
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.44 0.24 0.47 0.33 0.40 0.51
Split-EN-10 0.54 0.21 0.56 0.32 0.51 0.50
Lasso 0.07 0.33 0.05 0.38 0.04 0.43
Elastic Net 0.12 0.35 0.09 0.41 0.07 0.47
Adaptive 0.07 0.34 0.05 0.38 0.03 0.43
Relaxed 0.06 0.50 0.05 0.42 0.04 0.43
0.2 MCP 0.02 0.43 0.01 0.60 0.01 0.49
SIS-SCAD 0.03 0.68 0.02 0.73 0.01 0.70
RuleFit 0.19 0.28 0.13 0.43 0.10 0.53
RE-Lasso-100 0.59 0.13 0.52 0.24 0.48 0.41
RE-EN-100 0.71 0.12 0.65 0.23 0.62 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.49 0.24 0.37 0.41 0.27 0.54
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 48: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.37 0.24 0.40 0.32 0.35 0.50
Split-EN-10 0.43 0.23 0.52 0.30 0.45 0.50
Lasso 0.06 0.26 0.05 0.33 0.03 0.41
Elastic Net 0.10 0.26 0.08 0.35 0.07 0.46
Adaptive 0.06 0.26 0.05 0.32 0.03 0.40
Relaxed 0.06 0.30 0.05 0.36 0.03 0.42
0.4 MCP 0.02 0.26 0.01 0.33 0.01 0.45
SIS-SCAD 0.03 0.55 0.01 0.48 0.01 0.48
RuleFit 0.19 0.27 0.14 0.38 0.09 0.50
RE-Lasso-100 0.59 0.13 0.52 0.24 0.46 0.42
RE-EN-100 0.71 0.12 0.65 0.23 0.61 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.46 0.20 0.31 0.33 0.22 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.42 0.23 0.42 0.32 0.35 0.51
Split-EN-10 0.47 0.22 0.51 0.30 0.45 0.50
Lasso 0.07 0.30 0.05 0.35 0.03 0.41
Elastic Net 0.11 0.31 0.08 0.37 0.06 0.44
Adaptive 0.06 0.29 0.05 0.33 0.03 0.39
Relaxed 0.07 0.35 0.04 0.35 0.03 0.41
0.3 MCP 0.02 0.34 0.01 0.44 0.01 0.43
SIS-SCAD 0.03 0.57 0.01 0.46 0.01 0.53
RuleFit 0.18 0.29 0.12 0.37 0.08 0.54
RE-Lasso-100 0.56 0.13 0.48 0.24 0.44 0.41
RE-EN-100 0.68 0.12 0.62 0.23 0.59 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.43 0.21 0.29 0.32 0.21 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.42 0.21 0.42 0.34 0.34 0.53
Split-EN-10 0.48 0.21 0.51 0.32 0.44 0.51
Lasso 0.05 0.25 0.04 0.30 0.03 0.38
Elastic Net 0.09 0.26 0.07 0.33 0.05 0.41
Adaptive 0.05 0.25 0.03 0.30 0.02 0.36
Relaxed 0.05 0.28 0.04 0.31 0.02 0.38
0.2 MCP 0.01 0.37 0.01 0.42 0.00 0.37
SIS-SCAD 0.02 0.46 0.01 0.47 0.00 0.44
RuleFit 0.14 0.25 0.10 0.40 0.07 0.52
RE-Lasso-100 0.50 0.13 0.44 0.24 0.39 0.40
RE-EN-100 0.63 0.12 0.59 0.23 0.54 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.35 0.20 0.25 0.32 0.18 0.47
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 49: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.78 0.68 0.85 0.87 0.93 0.84 0.77 0.89 0.93 0.69 0.89 0.84 0.92 0.97 0.49
Split-EN-10 0.78 0.68 0.85 0.87 0.93 0.85 0.78 0.89 0.93 0.68 0.89 0.85 0.92 0.97 0.47
Lasso 0.75 0.64 0.82 0.83 1.06 0.79 0.68 0.87 0.88 0.87 0.83 0.74 0.88 0.91 0.78
Elastic Net 0.76 0.66 0.83 0.84 1.04 0.81 0.72 0.88 0.90 0.80 0.85 0.78 0.90 0.94 0.65
Adaptive 0.68 0.41 0.86 0.76 1.18 0.72 0.47 0.89 0.81 1.10 0.78 0.57 0.91 0.88 0.96
Relaxed 0.74 0.63 0.81 0.81 1.18 0.78 0.67 0.86 0.86 1.12 0.82 0.75 0.86 0.91 1.00
0.4 MCP 0.71 0.57 0.80 0.78 1.17 0.73 0.58 0.83 0.81 1.07 0.75 0.59 0.85 0.82 1.04
SIS-SCAD 0.69 0.59 0.76 0.76 1.29 0.71 0.62 0.77 0.78 1.26 0.72 0.60 0.80 0.79 1.25
RuleFit 0.71 0.61 0.78 0.78 1.50 0.73 0.62 0.81 0.81 1.34 0.74 0.61 0.83 0.82 1.24
RE-Lasso-100 0.78 0.67 0.85 0.87 0.92 0.84 0.75 0.89 0.93 0.70 0.88 0.80 0.93 0.96 0.55
RE-EN-100 0.78 0.68 0.85 0.87 0.93 0.84 0.76 0.89 0.93 0.69 0.89 0.82 0.93 0.97 0.52
RGLM-100 0.76 0.57 0.88 0.86 1.07 0.81 0.64 0.92 0.93 0.98 0.85 0.68 0.96 0.97 0.90
RF-500 0.78 0.67 0.85 0.87 0.93 0.83 0.73 0.90 0.92 0.82 0.87 0.77 0.93 0.96 0.76
XGB 0.71 0.62 0.76 0.77 1.15 0.72 0.61 0.79 0.79 1.11 0.75 0.62 0.83 0.83 1.03
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.81 0.57 0.92 0.88 0.81 0.86 0.69 0.93 0.93 0.63 0.90 0.73 0.96 0.97 0.46
Split-EN-10 0.82 0.57 0.92 0.88 0.81 0.86 0.70 0.93 0.94 0.63 0.90 0.74 0.96 0.97 0.44
Lasso 0.79 0.51 0.90 0.84 0.95 0.81 0.58 0.91 0.88 0.83 0.84 0.58 0.94 0.90 0.73
Elastic Net 0.80 0.53 0.91 0.85 0.90 0.83 0.62 0.92 0.90 0.76 0.86 0.64 0.95 0.93 0.62
Adaptive 0.74 0.21 0.97 0.75 1.06 0.76 0.26 0.96 0.79 1.01 0.78 0.30 0.97 0.83 0.93
Relaxed 0.78 0.51 0.89 0.82 1.64 0.81 0.59 0.90 0.88 0.94 0.83 0.61 0.92 0.89 1.36
0.3 MCP 0.75 0.42 0.89 0.77 1.09 0.76 0.43 0.90 0.80 1.03 0.77 0.40 0.92 0.80 1.48
SIS-SCAD 0.74 0.43 0.87 0.77 1.11 0.74 0.46 0.86 0.78 1.13 0.75 0.40 0.90 0.78 1.06
RuleFit 0.75 0.46 0.88 0.77 1.37 0.76 0.47 0.89 0.79 1.31 0.77 0.45 0.90 0.79 1.23
RE-Lasso-100 0.81 0.56 0.92 0.88 0.82 0.85 0.63 0.94 0.93 0.66 0.88 0.66 0.97 0.97 0.52
RE-EN-100 0.82 0.57 0.92 0.88 0.82 0.85 0.66 0.94 0.93 0.64 0.89 0.68 0.97 0.97 0.49
RGLM-100 0.78 0.33 0.97 0.87 0.96 0.81 0.40 0.98 0.93 0.89 0.82 0.38 1.00 0.97 0.81
RF-500 0.81 0.55 0.92 0.88 0.82 0.84 0.59 0.95 0.92 0.75 0.86 0.56 0.98 0.96 0.69
XGB 0.75 0.46 0.87 0.77 1.05 0.76 0.49 0.87 0.79 1.02 0.77 0.45 0.90 0.81 0.97
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.40 0.96 0.87 0.71 0.88 0.50 0.97 0.94 0.55 0.90 0.57 0.98 0.97 0.42
Split-EN-10 0.85 0.40 0.96 0.87 0.71 0.88 0.51 0.97 0.94 0.55 0.91 0.59 0.98 0.97 0.39
Lasso 0.82 0.33 0.95 0.82 0.83 0.85 0.40 0.96 0.88 0.71 0.86 0.44 0.96 0.91 0.65
Elastic Net 0.83 0.36 0.95 0.84 0.79 0.86 0.44 0.96 0.91 0.65 0.88 0.47 0.97 0.94 0.55
Adaptive 0.81 0.06 0.99 0.68 0.92 0.82 0.12 0.99 0.75 0.86 0.83 0.15 0.99 0.80 0.80
Relaxed 0.81 0.36 0.93 0.81 1.62 0.85 0.45 0.95 0.88 0.86 0.86 0.48 0.95 0.90 0.95
0.2 MCP 0.80 0.23 0.94 0.73 0.99 0.81 0.24 0.95 0.76 0.90 0.82 0.26 0.95 0.77 0.99
SIS-SCAD 0.80 0.24 0.93 0.75 1.03 0.81 0.22 0.95 0.77 0.92 0.81 0.18 0.96 0.77 1.05
RuleFit 0.80 0.32 0.92 0.71 1.36 0.81 0.31 0.93 0.74 1.22 0.82 0.28 0.94 0.75 1.13
RE-Lasso-100 0.85 0.39 0.96 0.87 0.71 0.88 0.48 0.97 0.93 0.54 0.90 0.54 0.99 0.97 0.44
RE-EN-100 0.85 0.41 0.96 0.87 0.71 0.88 0.50 0.97 0.93 0.54 0.91 0.58 0.99 0.97 0.40
RGLM-100 0.82 0.14 0.99 0.87 0.80 0.83 0.15 1.00 0.93 0.73 0.84 0.18 1.00 0.97 0.66
RF-500 0.85 0.35 0.97 0.87 0.71 0.86 0.35 0.99 0.92 0.63 0.87 0.34 0.99 0.96 0.58
XGB 0.80 0.31 0.93 0.76 0.89 0.82 0.31 0.95 0.81 0.81 0.82 0.31 0.94 0.82 0.81
Table 50: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.81 0.72 0.88 0.90 0.81 0.85 0.80 0.89 0.94 0.66 0.88 0.80 0.93 0.96 0.53
Split-EN-10 0.81 0.72 0.88 0.90 0.81 0.86 0.80 0.89 0.94 0.65 0.89 0.82 0.94 0.97 0.49
Lasso 0.78 0.66 0.85 0.86 0.97 0.80 0.72 0.85 0.88 0.90 0.83 0.71 0.90 0.91 0.80
Elastic Net 0.79 0.67 0.86 0.88 0.92 0.81 0.74 0.86 0.90 0.83 0.85 0.74 0.91 0.93 0.69
Adaptive 0.74 0.52 0.90 0.84 1.05 0.75 0.57 0.87 0.84 1.02 0.77 0.55 0.91 0.87 0.96
Relaxed 0.77 0.67 0.84 0.85 1.12 0.79 0.72 0.84 0.87 1.14 0.81 0.71 0.88 0.90 1.15
0.4 MCP 0.76 0.64 0.84 0.84 1.02 0.75 0.64 0.82 0.83 1.03 0.76 0.60 0.86 0.84 1.03
SIS-SCAD 0.76 0.64 0.84 0.84 1.03 0.73 0.63 0.79 0.80 1.15 0.72 0.58 0.81 0.79 1.20
RuleFit 0.75 0.65 0.83 0.83 1.30 0.75 0.65 0.81 0.82 1.31 0.75 0.59 0.85 0.82 1.32
RE-Lasso-100 0.80 0.69 0.87 0.89 0.83 0.83 0.76 0.88 0.92 0.72 0.87 0.77 0.94 0.96 0.57
RE-EN-100 0.80 0.69 0.87 0.89 0.84 0.84 0.77 0.88 0.93 0.71 0.88 0.78 0.94 0.96 0.54
RGLM-100 0.79 0.60 0.91 0.89 1.01 0.82 0.67 0.92 0.93 0.97 0.84 0.64 0.97 0.96 0.89
RF-500 0.82 0.73 0.88 0.91 0.80 0.84 0.76 0.89 0.93 0.78 0.86 0.72 0.94 0.95 0.75
XGB 0.75 0.66 0.81 0.82 1.04 0.74 0.65 0.80 0.81 1.05 0.75 0.59 0.84 0.82 1.04
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.83 0.63 0.92 0.91 0.74 0.86 0.69 0.93 0.94 0.63 0.89 0.73 0.95 0.96 0.50
Split-EN-10 0.83 0.63 0.92 0.91 0.73 0.86 0.70 0.93 0.94 0.61 0.89 0.75 0.95 0.97 0.48
Lasso 0.80 0.58 0.90 0.86 0.89 0.82 0.60 0.91 0.89 0.82 0.83 0.59 0.93 0.91 0.76
Elastic Net 0.81 0.59 0.90 0.88 0.84 0.83 0.63 0.92 0.90 0.76 0.85 0.66 0.94 0.93 0.65
Adaptive 0.77 0.34 0.95 0.81 1.00 0.76 0.29 0.96 0.80 1.00 0.76 0.28 0.97 0.83 0.96
Relaxed 0.79 0.59 0.88 0.86 1.07 0.82 0.62 0.90 0.89 0.99 0.83 0.62 0.92 0.91 0.91
0.3 MCP 0.77 0.52 0.88 0.83 0.98 0.77 0.49 0.89 0.82 0.97 0.77 0.44 0.91 0.82 1.02
SIS-SCAD 0.77 0.51 0.89 0.83 0.99 0.75 0.50 0.86 0.79 1.33 0.75 0.39 0.90 0.79 1.08
RuleFit 0.77 0.55 0.87 0.81 1.29 0.77 0.50 0.88 0.80 1.23 0.76 0.49 0.88 0.79 1.37
RE-Lasso-100 0.82 0.61 0.92 0.90 0.77 0.84 0.63 0.94 0.92 0.67 0.87 0.65 0.96 0.96 0.56
RE-EN-100 0.82 0.61 0.92 0.90 0.77 0.85 0.65 0.93 0.93 0.66 0.88 0.68 0.96 0.96 0.52
RGLM-100 0.80 0.42 0.97 0.90 0.93 0.81 0.43 0.98 0.93 0.88 0.82 0.40 0.99 0.96 0.82
RF-500 0.83 0.64 0.92 0.91 0.75 0.84 0.62 0.94 0.93 0.73 0.85 0.57 0.97 0.95 0.70
XGB 0.77 0.56 0.86 0.82 0.98 0.77 0.51 0.88 0.81 1.00 0.77 0.48 0.90 0.82 0.96
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.86 0.41 0.97 0.91 0.63 0.88 0.55 0.96 0.94 0.52 0.90 0.57 0.98 0.96 0.43
Split-EN-10 0.86 0.41 0.97 0.91 0.64 0.89 0.56 0.96 0.94 0.51 0.91 0.58 0.98 0.97 0.41
Lasso 0.84 0.33 0.96 0.84 0.76 0.85 0.40 0.96 0.88 0.68 0.87 0.41 0.97 0.91 0.63
Elastic Net 0.85 0.36 0.96 0.87 0.72 0.86 0.45 0.96 0.90 0.64 0.88 0.46 0.97 0.93 0.56
Adaptive 0.82 0.07 0.99 0.70 0.90 0.82 0.08 0.99 0.74 0.88 0.83 0.13 0.99 0.79 0.80
Relaxed 0.83 0.36 0.95 0.83 0.97 0.85 0.45 0.94 0.86 0.91 0.86 0.45 0.96 0.90 0.98
0.2 MCP 0.82 0.28 0.96 0.79 0.83 0.82 0.26 0.95 0.78 0.84 0.82 0.22 0.96 0.79 0.85
SIS-SCAD 0.82 0.23 0.96 0.81 0.86 0.82 0.25 0.95 0.79 0.87 0.82 0.18 0.96 0.78 0.85
RuleFit 0.82 0.39 0.92 0.76 1.12 0.81 0.35 0.92 0.77 1.20 0.82 0.31 0.94 0.77 1.17
RE-Lasso-100 0.86 0.44 0.96 0.89 0.64 0.88 0.49 0.97 0.93 0.55 0.90 0.51 0.98 0.96 0.45
RE-EN-100 0.86 0.44 0.96 0.89 0.65 0.88 0.54 0.96 0.93 0.53 0.90 0.56 0.98 0.96 0.41
RGLM-100 0.83 0.15 0.99 0.89 0.76 0.84 0.20 1.00 0.93 0.72 0.85 0.18 1.00 0.96 0.66
RF-500 0.86 0.41 0.97 0.90 0.64 0.87 0.42 0.98 0.93 0.61 0.87 0.36 0.99 0.95 0.57
XGB 0.83 0.40 0.93 0.82 0.79 0.82 0.36 0.93 0.81 0.82 0.83 0.31 0.95 0.83 0.78
Table 51: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.78 0.90 0.93 0.68 0.91 0.89 0.93 0.98 0.40 0.93 0.90 0.95 0.99 0.32
Split-EN-10 0.85 0.78 0.90 0.93 0.69 0.92 0.90 0.93 0.98 0.39 0.93 0.91 0.95 0.99 0.31
Lasso 0.83 0.76 0.89 0.92 0.76 0.88 0.84 0.90 0.96 0.56 0.89 0.85 0.92 0.97 0.49
Elastic Net 0.84 0.76 0.89 0.92 0.73 0.89 0.87 0.91 0.97 0.48 0.91 0.87 0.93 0.98 0.41
Adaptive 0.81 0.66 0.92 0.92 0.86 0.86 0.78 0.91 0.95 0.69 0.87 0.77 0.93 0.96 0.66
Relaxed 0.82 0.75 0.88 0.91 1.29 0.87 0.84 0.88 0.95 1.00 0.88 0.85 0.90 0.96 0.91
0.4 MCP 0.81 0.72 0.87 0.90 0.83 0.83 0.77 0.87 0.91 0.76 0.83 0.76 0.88 0.92 0.78
SIS-SCAD 0.82 0.73 0.88 0.90 0.82 0.80 0.73 0.85 0.89 0.86 0.81 0.72 0.86 0.90 0.83
RuleFit 0.81 0.73 0.87 0.88 1.11 0.81 0.73 0.86 0.88 1.06 0.81 0.72 0.86 0.88 1.16
RE-Lasso-100 0.84 0.78 0.89 0.93 0.71 0.91 0.88 0.93 0.97 0.43 0.92 0.89 0.95 0.99 0.36
RE-EN-100 0.84 0.78 0.89 0.93 0.72 0.91 0.88 0.93 0.98 0.42 0.93 0.90 0.95 0.99 0.34
RGLM-100 0.83 0.73 0.90 0.92 0.82 0.91 0.86 0.94 0.98 0.64 0.92 0.85 0.96 0.99 0.60
RF-500 0.85 0.79 0.90 0.94 0.66 0.90 0.87 0.93 0.97 0.55 0.92 0.87 0.95 0.98 0.52
XGB 0.82 0.76 0.87 0.90 0.82 0.83 0.76 0.87 0.92 0.76 0.84 0.77 0.89 0.93 0.70
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.86 0.70 0.92 0.94 0.64 0.92 0.85 0.96 0.98 0.35 0.94 0.88 0.96 0.99 0.29
Split-EN-10 0.85 0.69 0.92 0.93 0.65 0.92 0.85 0.95 0.98 0.35 0.94 0.88 0.97 0.99 0.27
Lasso 0.84 0.67 0.92 0.92 0.71 0.89 0.79 0.94 0.96 0.51 0.90 0.80 0.95 0.97 0.44
Elastic Net 0.85 0.67 0.92 0.93 0.69 0.90 0.81 0.95 0.97 0.44 0.92 0.83 0.96 0.98 0.36
Adaptive 0.80 0.41 0.96 0.90 0.85 0.85 0.59 0.96 0.95 0.68 0.86 0.61 0.97 0.96 0.62
Relaxed 0.84 0.67 0.91 0.92 0.87 0.89 0.80 0.92 0.96 0.81 0.89 0.80 0.93 0.96 0.72
0.3 MCP 0.83 0.63 0.91 0.90 0.77 0.83 0.62 0.92 0.90 0.78 0.84 0.64 0.92 0.91 0.72
SIS-SCAD 0.83 0.58 0.94 0.91 0.77 0.81 0.58 0.92 0.89 0.81 0.81 0.56 0.92 0.89 0.83
RuleFit 0.82 0.65 0.90 0.87 1.03 0.83 0.63 0.92 0.90 0.86 0.84 0.60 0.93 0.90 0.85
RE-Lasso-100 0.85 0.70 0.91 0.93 0.67 0.91 0.82 0.95 0.98 0.39 0.93 0.86 0.97 0.99 0.32
RE-EN-100 0.85 0.71 0.91 0.92 0.68 0.92 0.84 0.95 0.98 0.37 0.94 0.87 0.97 0.99 0.30
RGLM-100 0.83 0.60 0.93 0.92 0.77 0.90 0.73 0.98 0.98 0.59 0.92 0.75 0.99 0.99 0.55
RF-500 0.86 0.73 0.92 0.93 0.62 0.91 0.79 0.97 0.98 0.50 0.92 0.80 0.97 0.98 0.48
XGB 0.84 0.68 0.90 0.90 0.77 0.85 0.67 0.93 0.93 0.68 0.85 0.66 0.93 0.92 0.69
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.88 0.55 0.96 0.94 0.54 0.94 0.77 0.98 0.98 0.29 0.94 0.81 0.98 0.99 0.26
Split-EN-10 0.88 0.55 0.96 0.94 0.54 0.94 0.78 0.98 0.99 0.28 0.95 0.82 0.98 0.99 0.25
Lasso 0.87 0.51 0.96 0.92 0.61 0.91 0.66 0.97 0.96 0.44 0.91 0.71 0.96 0.97 0.40
Elastic Net 0.88 0.53 0.96 0.93 0.58 0.92 0.71 0.97 0.97 0.37 0.93 0.75 0.97 0.98 0.33
Adaptive 0.83 0.19 0.99 0.84 0.76 0.86 0.34 0.99 0.92 0.62 0.86 0.36 0.99 0.95 0.60
Relaxed 0.87 0.54 0.95 0.92 1.16 0.90 0.68 0.95 0.95 0.90 0.91 0.73 0.95 0.96 0.67
0.2 MCP 0.85 0.40 0.96 0.88 0.75 0.85 0.46 0.95 0.88 0.73 0.85 0.44 0.95 0.88 0.68
SIS-SCAD 0.84 0.29 0.97 0.89 0.79 0.84 0.34 0.97 0.89 0.71 0.84 0.34 0.97 0.89 0.75
RuleFit 0.86 0.47 0.95 0.85 0.95 0.87 0.49 0.96 0.87 0.87 0.86 0.50 0.95 0.86 1.00
RE-Lasso-100 0.88 0.63 0.95 0.93 0.53 0.94 0.78 0.97 0.98 0.30 0.94 0.80 0.97 0.99 0.28
RE-EN-100 0.88 0.64 0.94 0.93 0.54 0.94 0.80 0.97 0.98 0.28 0.94 0.84 0.97 0.99 0.26
RGLM-100 0.87 0.43 0.98 0.93 0.62 0.90 0.54 0.99 0.98 0.48 0.91 0.58 1.00 0.99 0.47
RF-500 0.88 0.57 0.96 0.94 0.53 0.92 0.64 0.99 0.98 0.42 0.92 0.67 0.99 0.98 0.41
XGB 0.86 0.53 0.95 0.90 0.68 0.87 0.53 0.96 0.91 0.62 0.87 0.55 0.95 0.92 0.66
Table 52: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.80 0.73 0.85 0.88 0.86 0.87 0.80 0.91 0.95 0.61 0.90 0.85 0.94 0.97 0.43
Split-EN-10 0.80 0.73 0.85 0.88 0.87 0.87 0.80 0.91 0.94 0.61 0.90 0.85 0.94 0.97 0.43
Lasso 0.78 0.69 0.83 0.86 0.94 0.83 0.74 0.89 0.92 0.75 0.86 0.77 0.91 0.94 0.65
Elastic Net 0.78 0.71 0.84 0.87 0.92 0.84 0.76 0.90 0.93 0.71 0.87 0.80 0.92 0.95 0.58
Adaptive 0.76 0.58 0.87 0.84 1.03 0.80 0.64 0.91 0.90 0.88 0.83 0.70 0.92 0.92 0.76
Relaxed 0.77 0.69 0.82 0.85 0.99 0.83 0.74 0.88 0.91 0.81 0.86 0.78 0.90 0.94 0.72
0.4 MCP 0.75 0.65 0.82 0.82 1.06 0.78 0.66 0.86 0.87 0.91 0.80 0.68 0.87 0.88 0.88
SIS-SCAD 0.74 0.64 0.81 0.82 1.06 0.76 0.64 0.84 0.84 1.00 0.76 0.64 0.83 0.83 1.05
RuleFit 0.75 0.66 0.81 0.82 1.37 0.78 0.65 0.88 0.87 1.09 0.80 0.65 0.89 0.88 0.99
RE-Lasso-100 0.80 0.72 0.85 0.88 0.89 0.86 0.78 0.92 0.94 0.63 0.89 0.81 0.94 0.97 0.48
RE-EN-100 0.80 0.72 0.85 0.88 0.90 0.86 0.78 0.91 0.94 0.63 0.90 0.82 0.94 0.97 0.46
RGLM-100 0.78 0.62 0.88 0.87 1.03 0.83 0.64 0.95 0.94 0.93 0.87 0.69 0.97 0.97 0.84
RF-500 0.80 0.74 0.84 0.88 0.86 0.86 0.77 0.92 0.94 0.70 0.89 0.78 0.95 0.97 0.65
XGB 0.74 0.65 0.80 0.81 1.06 0.78 0.64 0.87 0.87 0.92 0.80 0.66 0.89 0.88 0.85
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.82 0.63 0.91 0.89 0.78 0.87 0.71 0.94 0.95 0.56 0.91 0.80 0.96 0.97 0.39
Split-EN-10 0.82 0.63 0.91 0.89 0.78 0.87 0.72 0.94 0.95 0.57 0.92 0.81 0.96 0.98 0.38
Lasso 0.80 0.58 0.90 0.87 0.85 0.84 0.64 0.93 0.91 0.71 0.87 0.70 0.94 0.94 0.59
Elastic Net 0.81 0.58 0.91 0.87 0.83 0.85 0.66 0.93 0.92 0.67 0.88 0.73 0.94 0.95 0.52
Adaptive 0.76 0.30 0.96 0.84 0.98 0.80 0.38 0.97 0.89 0.85 0.83 0.49 0.97 0.92 0.74
Relaxed 0.80 0.56 0.90 0.86 0.89 0.84 0.65 0.92 0.91 0.75 0.87 0.70 0.93 0.94 0.63
0.3 MCP 0.78 0.51 0.90 0.83 0.94 0.80 0.52 0.91 0.86 0.87 0.82 0.54 0.92 0.88 0.81
SIS-SCAD 0.77 0.51 0.89 0.83 0.98 0.78 0.48 0.91 0.83 1.00 0.79 0.52 0.90 0.84 0.97
RuleFit 0.78 0.53 0.89 0.83 1.28 0.81 0.52 0.92 0.86 1.08 0.82 0.52 0.94 0.88 0.95
RE-Lasso-100 0.82 0.61 0.91 0.89 0.79 0.86 0.67 0.95 0.94 0.60 0.90 0.73 0.97 0.97 0.44
RE-EN-100 0.82 0.62 0.91 0.89 0.80 0.87 0.68 0.94 0.94 0.59 0.91 0.75 0.97 0.97 0.42
RGLM-100 0.79 0.40 0.96 0.87 0.94 0.82 0.41 0.99 0.94 0.83 0.85 0.48 0.99 0.97 0.75
RF-500 0.82 0.64 0.90 0.89 0.78 0.86 0.64 0.95 0.94 0.65 0.89 0.68 0.97 0.97 0.59
XGB 0.77 0.52 0.88 0.82 0.97 0.80 0.51 0.92 0.86 0.87 0.82 0.53 0.94 0.89 0.77
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.44 0.96 0.89 0.67 0.90 0.57 0.97 0.95 0.47 0.92 0.67 0.98 0.98 0.34
Split-EN-10 0.85 0.45 0.96 0.89 0.66 0.90 0.57 0.97 0.95 0.47 0.93 0.69 0.98 0.98 0.33
Lasso 0.83 0.40 0.95 0.86 0.76 0.87 0.49 0.96 0.91 0.61 0.89 0.55 0.97 0.94 0.53
Elastic Net 0.84 0.41 0.95 0.87 0.73 0.88 0.51 0.97 0.93 0.57 0.90 0.59 0.97 0.96 0.46
Adaptive 0.81 0.12 0.99 0.81 0.84 0.84 0.20 0.99 0.88 0.71 0.85 0.27 0.99 0.91 0.64
Relaxed 0.83 0.40 0.94 0.86 0.81 0.87 0.49 0.96 0.91 0.69 0.88 0.57 0.96 0.94 0.64
0.2 MCP 0.82 0.32 0.95 0.81 0.87 0.84 0.29 0.97 0.84 0.75 0.84 0.31 0.96 0.85 0.77
SIS-SCAD 0.82 0.33 0.95 0.82 0.87 0.83 0.29 0.96 0.83 0.82 0.83 0.33 0.95 0.84 0.84
RuleFit 0.82 0.37 0.94 0.81 1.16 0.84 0.33 0.97 0.85 1.01 0.85 0.32 0.97 0.86 0.96
RE-Lasso-100 0.85 0.43 0.96 0.89 0.69 0.89 0.49 0.98 0.94 0.51 0.90 0.54 0.99 0.97 0.41
RE-EN-100 0.85 0.44 0.96 0.89 0.70 0.89 0.51 0.98 0.95 0.50 0.91 0.58 0.99 0.98 0.39
RGLM-100 0.82 0.19 0.99 0.88 0.78 0.84 0.18 1.00 0.94 0.67 0.84 0.18 1.00 0.98 0.62
RF-500 0.85 0.48 0.95 0.89 0.67 0.88 0.44 0.98 0.94 0.54 0.89 0.44 0.99 0.97 0.50
XGB 0.82 0.37 0.93 0.81 0.83 0.84 0.32 0.97 0.86 0.73 0.85 0.31 0.98 0.88 0.69
Table 53: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.84 0.79 0.87 0.92 0.72 0.89 0.84 0.92 0.96 0.52 0.90 0.85 0.94 0.97 0.44
Split-EN-10 0.84 0.78 0.87 0.92 0.72 0.89 0.83 0.92 0.96 0.52 0.91 0.85 0.94 0.97 0.42
Lasso 0.81 0.75 0.86 0.90 0.82 0.85 0.78 0.90 0.93 0.67 0.86 0.78 0.90 0.94 0.65
Elastic Net 0.82 0.75 0.86 0.91 0.80 0.86 0.79 0.90 0.94 0.64 0.87 0.80 0.91 0.95 0.59
Adaptive 0.80 0.67 0.89 0.90 0.91 0.83 0.69 0.92 0.92 0.81 0.83 0.69 0.92 0.92 0.76
Relaxed 0.80 0.74 0.85 0.89 0.85 0.85 0.78 0.90 0.93 0.74 0.85 0.78 0.90 0.94 0.74
0.4 MCP 0.79 0.73 0.84 0.88 0.88 0.83 0.75 0.88 0.91 0.76 0.80 0.69 0.87 0.89 0.85
SIS-SCAD 0.79 0.71 0.85 0.88 0.90 0.80 0.70 0.87 0.88 0.95 0.76 0.64 0.85 0.84 1.02
RuleFit 0.79 0.73 0.84 0.87 1.18 0.80 0.70 0.87 0.89 0.99 0.80 0.67 0.89 0.89 0.99
RE-Lasso-100 0.83 0.77 0.87 0.92 0.74 0.87 0.80 0.91 0.95 0.58 0.89 0.81 0.94 0.96 0.50
RE-EN-100 0.83 0.77 0.87 0.92 0.75 0.87 0.80 0.91 0.95 0.58 0.89 0.82 0.94 0.97 0.48
RGLM-100 0.82 0.69 0.91 0.91 0.94 0.84 0.69 0.94 0.94 0.88 0.86 0.69 0.97 0.97 0.83
RF-500 0.84 0.80 0.87 0.92 0.71 0.88 0.81 0.92 0.96 0.61 0.88 0.79 0.94 0.96 0.63
XGB 0.79 0.72 0.83 0.87 0.93 0.80 0.69 0.87 0.89 0.85 0.80 0.66 0.88 0.89 0.86
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.86 0.71 0.93 0.93 0.64 0.89 0.75 0.95 0.96 0.49 0.91 0.82 0.95 0.97 0.39
Split-EN-10 0.86 0.70 0.93 0.93 0.64 0.89 0.75 0.95 0.96 0.49 0.92 0.82 0.96 0.97 0.38
Lasso 0.84 0.67 0.92 0.91 0.72 0.85 0.67 0.93 0.93 0.66 0.87 0.72 0.93 0.94 0.59
Elastic Net 0.84 0.67 0.92 0.91 0.71 0.86 0.68 0.94 0.94 0.62 0.88 0.75 0.94 0.95 0.53
Adaptive 0.81 0.47 0.96 0.91 0.84 0.82 0.49 0.96 0.91 0.77 0.84 0.54 0.96 0.92 0.70
Relaxed 0.83 0.66 0.91 0.90 0.80 0.85 0.67 0.93 0.92 0.82 0.87 0.73 0.92 0.94 0.66
0.3 MCP 0.82 0.62 0.91 0.89 0.79 0.83 0.60 0.92 0.90 0.76 0.82 0.58 0.92 0.88 0.79
SIS-SCAD 0.82 0.60 0.92 0.89 0.86 0.81 0.56 0.92 0.88 0.86 0.79 0.53 0.90 0.84 0.95
RuleFit 0.82 0.64 0.90 0.88 1.07 0.82 0.55 0.94 0.88 1.02 0.82 0.53 0.94 0.88 0.98
RE-Lasso-100 0.85 0.68 0.93 0.92 0.66 0.87 0.68 0.95 0.95 0.57 0.90 0.74 0.96 0.97 0.46
RE-EN-100 0.85 0.68 0.93 0.92 0.66 0.87 0.69 0.95 0.95 0.56 0.90 0.76 0.96 0.97 0.44
RGLM-100 0.83 0.52 0.97 0.92 0.84 0.82 0.45 0.99 0.94 0.81 0.85 0.49 0.99 0.97 0.75
RF-500 0.86 0.73 0.92 0.93 0.64 0.88 0.70 0.95 0.95 0.58 0.89 0.69 0.97 0.96 0.58
XGB 0.82 0.64 0.90 0.88 0.81 0.81 0.54 0.93 0.88 0.81 0.82 0.53 0.93 0.88 0.78
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.88 0.56 0.96 0.93 0.54 0.91 0.63 0.97 0.96 0.43 0.92 0.68 0.98 0.97 0.35
Split-EN-10 0.88 0.55 0.96 0.93 0.54 0.90 0.62 0.98 0.96 0.43 0.93 0.69 0.98 0.97 0.33
Lasso 0.86 0.49 0.96 0.91 0.63 0.88 0.51 0.97 0.93 0.56 0.89 0.53 0.97 0.94 0.50
Elastic Net 0.87 0.51 0.96 0.91 0.61 0.88 0.53 0.97 0.93 0.53 0.90 0.58 0.97 0.95 0.46
Adaptive 0.83 0.22 0.99 0.89 0.74 0.84 0.21 0.99 0.88 0.71 0.85 0.20 0.99 0.89 0.66
Relaxed 0.86 0.50 0.96 0.90 0.72 0.88 0.52 0.96 0.92 0.65 0.89 0.56 0.96 0.94 0.55
0.2 MCP 0.85 0.45 0.95 0.88 0.71 0.85 0.39 0.96 0.87 0.73 0.85 0.32 0.96 0.86 0.70
SIS-SCAD 0.84 0.42 0.95 0.88 0.77 0.84 0.37 0.96 0.86 0.75 0.84 0.28 0.96 0.83 0.79
RuleFit 0.85 0.50 0.94 0.87 0.99 0.84 0.38 0.96 0.85 1.02 0.86 0.34 0.97 0.87 0.87
RE-Lasso-100 0.87 0.51 0.97 0.93 0.57 0.89 0.52 0.98 0.95 0.51 0.91 0.55 0.98 0.96 0.41
RE-EN-100 0.87 0.51 0.97 0.93 0.58 0.89 0.54 0.98 0.95 0.50 0.91 0.59 0.98 0.97 0.39
RGLM-100 0.84 0.28 0.99 0.92 0.70 0.84 0.22 1.00 0.94 0.66 0.86 0.22 1.00 0.97 0.59
RF-500 0.88 0.60 0.96 0.94 0.54 0.89 0.53 0.98 0.95 0.50 0.90 0.48 0.99 0.96 0.48
XGB 0.85 0.51 0.94 0.88 0.70 0.85 0.39 0.96 0.86 0.72 0.86 0.34 0.97 0.88 0.65
Table 54: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.86 0.80 0.89 0.94 0.65 0.92 0.90 0.93 0.98 0.35 0.94 0.92 0.96 0.99 0.27
Split-EN-10 0.85 0.80 0.89 0.93 0.65 0.92 0.90 0.93 0.98 0.35 0.94 0.92 0.96 0.99 0.26
Lasso 0.84 0.78 0.88 0.93 0.70 0.90 0.87 0.92 0.97 0.47 0.91 0.87 0.94 0.98 0.40
Elastic Net 0.85 0.78 0.89 0.93 0.68 0.91 0.88 0.93 0.97 0.43 0.92 0.89 0.95 0.98 0.34
Adaptive 0.84 0.74 0.91 0.93 0.78 0.89 0.83 0.93 0.96 0.56 0.90 0.83 0.95 0.97 0.48
Relaxed 0.84 0.79 0.88 0.92 0.72 0.90 0.87 0.91 0.97 0.58 0.91 0.87 0.93 0.97 0.49
0.4 MCP 0.83 0.76 0.87 0.91 0.78 0.86 0.81 0.90 0.94 0.63 0.87 0.80 0.91 0.95 0.58
SIS-SCAD 0.84 0.77 0.88 0.92 0.74 0.85 0.79 0.89 0.93 0.70 0.85 0.78 0.90 0.93 0.67
RuleFit 0.84 0.79 0.87 0.92 0.96 0.87 0.81 0.91 0.95 0.66 0.87 0.79 0.92 0.95 0.64
RE-Lasso-100 0.85 0.79 0.89 0.93 0.67 0.91 0.88 0.93 0.98 0.39 0.94 0.90 0.96 0.99 0.30
RE-EN-100 0.85 0.79 0.89 0.93 0.68 0.91 0.89 0.93 0.98 0.38 0.94 0.90 0.96 0.99 0.29
RGLM-100 0.83 0.75 0.89 0.92 0.80 0.91 0.86 0.94 0.98 0.59 0.93 0.87 0.98 0.99 0.54
RF-500 0.86 0.81 0.89 0.94 0.65 0.92 0.88 0.94 0.98 0.46 0.93 0.88 0.96 0.99 0.44
XGB 0.84 0.80 0.87 0.92 0.73 0.87 0.81 0.91 0.95 0.60 0.88 0.81 0.93 0.96 0.53
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.87 0.71 0.93 0.94 0.58 0.94 0.88 0.96 0.99 0.30 0.95 0.90 0.97 0.99 0.24
Split-EN-10 0.87 0.71 0.93 0.94 0.59 0.94 0.88 0.96 0.99 0.30 0.95 0.91 0.97 0.99 0.23
Lasso 0.86 0.69 0.93 0.93 0.63 0.91 0.82 0.95 0.97 0.42 0.92 0.83 0.96 0.98 0.37
Elastic Net 0.86 0.69 0.93 0.93 0.62 0.92 0.84 0.96 0.98 0.37 0.93 0.86 0.96 0.98 0.31
Adaptive 0.84 0.55 0.96 0.93 0.71 0.90 0.73 0.97 0.97 0.51 0.90 0.75 0.97 0.97 0.47
Relaxed 0.86 0.69 0.92 0.93 0.68 0.91 0.82 0.94 0.97 0.53 0.92 0.85 0.95 0.98 0.45
0.3 MCP 0.84 0.64 0.92 0.91 0.72 0.87 0.74 0.93 0.94 0.60 0.87 0.71 0.94 0.94 0.59
SIS-SCAD 0.84 0.62 0.93 0.92 0.78 0.86 0.69 0.93 0.93 0.66 0.85 0.66 0.94 0.93 0.66
RuleFit 0.85 0.69 0.92 0.92 0.89 0.88 0.71 0.95 0.95 0.62 0.88 0.70 0.96 0.95 0.62
RE-Lasso-100 0.87 0.70 0.93 0.94 0.62 0.93 0.85 0.96 0.98 0.33 0.94 0.87 0.97 0.99 0.28
RE-EN-100 0.86 0.70 0.93 0.93 0.63 0.93 0.85 0.96 0.98 0.33 0.94 0.88 0.97 0.99 0.26
RGLM-100 0.85 0.62 0.94 0.92 0.71 0.92 0.77 0.98 0.98 0.53 0.93 0.79 0.99 0.99 0.49
RF-500 0.87 0.74 0.93 0.94 0.58 0.93 0.83 0.97 0.98 0.41 0.93 0.83 0.98 0.99 0.40
XGB 0.86 0.71 0.92 0.92 0.67 0.88 0.73 0.95 0.96 0.53 0.88 0.72 0.95 0.96 0.53
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.89 0.58 0.96 0.94 0.49 0.94 0.82 0.97 0.98 0.27 0.96 0.87 0.98 0.99 0.20
Split-EN-10 0.89 0.57 0.96 0.94 0.50 0.94 0.82 0.97 0.98 0.26 0.96 0.88 0.98 0.99 0.20
Lasso 0.88 0.54 0.96 0.93 0.54 0.92 0.74 0.97 0.97 0.37 0.93 0.77 0.97 0.98 0.32
Elastic Net 0.89 0.54 0.96 0.93 0.52 0.93 0.77 0.97 0.98 0.33 0.94 0.80 0.98 0.98 0.27
Adaptive 0.86 0.28 0.99 0.92 0.62 0.90 0.54 0.98 0.96 0.47 0.90 0.58 0.99 0.97 0.44
Relaxed 0.88 0.55 0.95 0.92 0.62 0.92 0.74 0.96 0.97 0.44 0.93 0.79 0.96 0.98 0.40
0.2 MCP 0.87 0.47 0.96 0.90 0.64 0.87 0.53 0.95 0.91 0.64 0.87 0.55 0.95 0.92 0.62
SIS-SCAD 0.86 0.41 0.96 0.91 0.67 0.87 0.46 0.97 0.92 0.62 0.86 0.46 0.96 0.92 0.66
RuleFit 0.88 0.54 0.95 0.89 0.82 0.89 0.58 0.97 0.92 0.63 0.89 0.58 0.96 0.92 0.69
RE-Lasso-100 0.89 0.58 0.96 0.93 0.53 0.93 0.76 0.98 0.98 0.30 0.95 0.81 0.98 0.99 0.24
RE-EN-100 0.89 0.57 0.96 0.93 0.53 0.94 0.78 0.97 0.98 0.29 0.95 0.83 0.98 0.99 0.23
RGLM-100 0.88 0.45 0.97 0.92 0.58 0.92 0.62 0.99 0.98 0.44 0.93 0.67 1.00 0.99 0.42
RF-500 0.89 0.61 0.96 0.94 0.49 0.93 0.72 0.98 0.98 0.36 0.94 0.75 0.99 0.99 0.34
XGB 0.88 0.58 0.95 0.92 0.57 0.90 0.61 0.97 0.95 0.48 0.90 0.62 0.97 0.96 0.46
Table 55: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.28 0.22 0.30 0.32 0.30 0.50
Split-EN-10 0.35 0.20 0.38 0.30 0.40 0.48
Lasso 0.05 0.27 0.04 0.36 0.03 0.47
Elastic Net 0.08 0.27 0.07 0.36 0.06 0.47
Adaptive 0.04 0.26 0.04 0.35 0.03 0.45
Relaxed 0.04 0.31 0.04 0.36 0.03 0.48
0.4 MCP 0.02 0.34 0.01 0.40 0.01 0.47
SIS-SCAD 0.01 0.38 0.01 0.45 0.00 0.55
RuleFit 0.12 0.24 0.09 0.35 0.06 0.54
RE-Lasso-100 0.51 0.13 0.48 0.25 0.44 0.44
RE-EN-100 0.65 0.12 0.61 0.23 0.58 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.32 0.18 0.24 0.30 0.18 0.48
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.28 0.20 0.32 0.31 0.27 0.49
Split-EN-10 0.32 0.19 0.39 0.29 0.36 0.48
Lasso 0.04 0.25 0.04 0.33 0.03 0.45
Elastic Net 0.07 0.24 0.06 0.33 0.05 0.46
Adaptive 0.04 0.25 0.04 0.33 0.03 0.46
Relaxed 0.04 0.30 0.03 0.35 0.02 0.49
0.3 MCP 0.02 0.31 0.01 0.33 0.01 0.48
SIS-SCAD 0.01 0.44 0.01 0.39 0.00 0.50
RuleFit 0.11 0.26 0.08 0.36 0.06 0.56
RE-Lasso-100 0.48 0.13 0.45 0.24 0.40 0.43
RE-EN-100 0.60 0.12 0.59 0.23 0.55 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.28 0.17 0.23 0.30 0.16 0.46
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.22 0.20 0.28 0.33 0.25 0.51
Split-EN-10 0.27 0.19 0.35 0.31 0.33 0.49
Lasso 0.04 0.22 0.03 0.31 0.02 0.47
Elastic Net 0.06 0.22 0.05 0.32 0.04 0.48
Adaptive 0.03 0.24 0.03 0.32 0.02 0.47
Relaxed 0.03 0.24 0.03 0.33 0.02 0.50
0.2 MCP 0.01 0.29 0.01 0.35 0.00 0.53
SIS-SCAD 0.01 0.32 0.00 0.39 0.00 0.52
RuleFit 0.10 0.30 0.07 0.41 0.04 0.59
RE-Lasso-100 0.43 0.13 0.42 0.25 0.36 0.42
RE-EN-100 0.57 0.12 0.54 0.23 0.50 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.24 0.17 0.18 0.29 0.14 0.47
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 56: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.38 0.24 0.34 0.34 0.28 0.50
Split-EN-10 0.46 0.24 0.43 0.33 0.41 0.50
Lasso 0.06 0.31 0.04 0.31 0.02 0.41
Elastic Net 0.10 0.34 0.07 0.36 0.05 0.46
Adaptive 0.05 0.30 0.03 0.29 0.02 0.39
Relaxed 0.05 0.47 0.03 0.35 0.02 0.43
0.4 MCP 0.02 0.33 0.01 0.39 0.01 0.43
SIS-SCAD 0.02 0.67 0.01 0.57 0.00 0.54
RuleFit 0.13 0.29 0.09 0.37 0.06 0.51
RE-Lasso-100 0.54 0.14 0.46 0.23 0.40 0.40
RE-EN-100 0.68 0.13 0.62 0.23 0.57 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.39 0.25 0.26 0.35 0.20 0.51
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.35 0.24 0.34 0.35 0.28 0.49
Split-EN-10 0.45 0.22 0.45 0.32 0.39 0.50
Lasso 0.05 0.27 0.03 0.32 0.02 0.38
Elastic Net 0.09 0.31 0.07 0.35 0.05 0.42
Adaptive 0.04 0.26 0.03 0.30 0.02 0.38
Relaxed 0.04 0.34 0.03 0.33 0.02 0.37
0.3 MCP 0.01 0.35 0.01 0.37 0.01 0.45
SIS-SCAD 0.02 0.61 0.01 0.53 0.00 0.50
RuleFit 0.12 0.28 0.09 0.38 0.06 0.53
RE-Lasso-100 0.50 0.13 0.44 0.24 0.37 0.40
RE-EN-100 0.64 0.12 0.60 0.23 0.55 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.35 0.24 0.26 0.35 0.18 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.34 0.25 0.30 0.33 0.26 0.49
Split-EN-10 0.39 0.25 0.40 0.32 0.34 0.50
Lasso 0.04 0.26 0.02 0.27 0.02 0.34
Elastic Net 0.07 0.30 0.05 0.32 0.04 0.40
Adaptive 0.04 0.30 0.02 0.26 0.02 0.38
Relaxed 0.03 0.34 0.02 0.30 0.01 0.36
0.2 MCP 0.01 0.43 0.01 0.34 0.00 0.39
SIS-SCAD 0.01 0.54 0.01 0.51 0.00 0.51
RuleFit 0.10 0.34 0.07 0.41 0.04 0.55
RE-Lasso-100 0.46 0.13 0.38 0.23 0.33 0.39
RE-EN-100 0.61 0.13 0.56 0.23 0.50 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.31 0.24 0.21 0.33 0.15 0.48
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 57: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.22 0.24 0.27 0.31 0.23 0.48
Split-EN-10 0.26 0.25 0.37 0.31 0.34 0.48
Lasso 0.04 0.29 0.03 0.28 0.02 0.37
Elastic Net 0.08 0.30 0.06 0.31 0.05 0.41
Adaptive 0.04 0.29 0.03 0.27 0.02 0.34
Relaxed 0.03 0.50 0.02 0.30 0.02 0.35
0.4 MCP 0.01 0.45 0.01 0.31 0.00 0.37
SIS-SCAD 0.02 0.66 0.01 0.37 0.00 0.41
RuleFit 0.09 0.31 0.07 0.38 0.05 0.54
RE-Lasso-100 0.38 0.12 0.38 0.23 0.35 0.39
RE-EN-100 0.52 0.11 0.55 0.23 0.52 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.27 0.20 0.17 0.29 0.14 0.45
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.23 0.24 0.29 0.33 0.24 0.50
Split-EN-10 0.25 0.27 0.39 0.31 0.34 0.49
Lasso 0.04 0.31 0.03 0.29 0.02 0.36
Elastic Net 0.07 0.32 0.06 0.31 0.04 0.40
Adaptive 0.04 0.31 0.02 0.27 0.02 0.35
Relaxed 0.03 0.45 0.02 0.32 0.02 0.36
0.3 MCP 0.01 0.41 0.01 0.33 0.00 0.44
SIS-SCAD 0.02 0.73 0.01 0.44 0.00 0.49
RuleFit 0.09 0.28 0.07 0.41 0.04 0.55
RE-Lasso-100 0.38 0.12 0.37 0.23 0.32 0.39
RE-EN-100 0.52 0.11 0.54 0.23 0.50 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.26 0.20 0.18 0.30 0.13 0.44
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.24 0.24 0.28 0.34 0.23 0.52
Split-EN-10 0.28 0.24 0.40 0.32 0.33 0.50
Lasso 0.03 0.28 0.02 0.28 0.01 0.34
Elastic Net 0.06 0.29 0.05 0.29 0.04 0.39
Adaptive 0.03 0.29 0.02 0.26 0.01 0.33
Relaxed 0.02 0.36 0.02 0.31 0.01 0.35
0.2 MCP 0.01 0.55 0.00 0.39 0.00 0.41
SIS-SCAD 0.01 0.61 0.00 0.41 0.00 0.35
RuleFit 0.08 0.42 0.05 0.42 0.03 0.52
RE-Lasso-100 0.34 0.12 0.33 0.23 0.30 0.39
RE-EN-100 0.49 0.12 0.51 0.22 0.48 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.20 0.19 0.15 0.29 0.11 0.44
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 58: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.47 0.20 0.50 0.30 0.44 0.49
Split-EN-10 0.47 0.21 0.55 0.29 0.50 0.48
Lasso 0.09 0.33 0.08 0.42 0.05 0.51
Elastic Net 0.13 0.31 0.11 0.41 0.08 0.52
Adaptive 0.09 0.34 0.08 0.41 0.05 0.50
Relaxed 0.08 0.42 0.07 0.43 0.05 0.51
0.4 MCP 0.04 0.39 0.03 0.46 0.02 0.50
SIS-SCAD 0.03 0.63 0.02 0.67 0.01 0.59
RuleFit 0.23 0.24 0.18 0.38 0.13 0.53
RE-Lasso-100 0.68 0.12 0.64 0.25 0.57 0.44
RE-EN-100 0.77 0.11 0.72 0.23 0.67 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.54 0.18 0.43 0.33 0.30 0.50
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.45 0.21 0.48 0.31 0.44 0.49
Split-EN-10 0.48 0.21 0.55 0.29 0.52 0.47
Lasso 0.09 0.34 0.07 0.39 0.05 0.51
Elastic Net 0.12 0.33 0.10 0.39 0.08 0.51
Adaptive 0.09 0.34 0.06 0.39 0.05 0.51
Relaxed 0.08 0.42 0.07 0.39 0.05 0.51
0.3 MCP 0.04 0.43 0.02 0.42 0.01 0.54
SIS-SCAD 0.03 0.63 0.01 0.59 0.01 0.62
RuleFit 0.21 0.24 0.16 0.39 0.11 0.53
RE-Lasso-100 0.67 0.13 0.60 0.25 0.54 0.44
RE-EN-100 0.75 0.12 0.70 0.23 0.65 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.52 0.18 0.39 0.32 0.28 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.37 0.24 0.42 0.31 0.39 0.50
Split-EN-10 0.40 0.23 0.49 0.29 0.48 0.48
Lasso 0.08 0.32 0.05 0.36 0.04 0.48
Elastic Net 0.11 0.31 0.08 0.36 0.07 0.49
Adaptive 0.08 0.30 0.05 0.36 0.04 0.48
Relaxed 0.08 0.36 0.05 0.36 0.04 0.49
0.2 MCP 0.02 0.36 0.01 0.45 0.01 0.48
SIS-SCAD 0.03 0.55 0.01 0.51 0.01 0.58
RuleFit 0.20 0.26 0.12 0.38 0.09 0.54
RE-Lasso-100 0.62 0.13 0.52 0.24 0.47 0.43
RE-EN-100 0.71 0.12 0.63 0.23 0.59 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.46 0.19 0.31 0.31 0.22 0.47
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 59: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.54 0.20 0.54 0.31 0.46 0.51
Split-EN-10 0.58 0.20 0.62 0.29 0.55 0.50
Lasso 0.09 0.33 0.07 0.40 0.04 0.43
Elastic Net 0.13 0.35 0.12 0.44 0.08 0.48
Adaptive 0.09 0.32 0.06 0.40 0.04 0.44
Relaxed 0.07 0.51 0.06 0.44 0.04 0.44
0.4 MCP 0.02 0.32 0.02 0.47 0.01 0.47
SIS-SCAD 0.03 0.72 0.02 0.80 0.01 0.65
RuleFit 0.24 0.28 0.20 0.43 0.12 0.53
RE-Lasso-100 0.66 0.12 0.62 0.24 0.55 0.42
RE-EN-100 0.75 0.11 0.74 0.23 0.68 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.58 0.22 0.48 0.42 0.33 0.53
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.54 0.20 0.51 0.31 0.44 0.49
Split-EN-10 0.59 0.20 0.61 0.30 0.54 0.49
Lasso 0.09 0.36 0.06 0.38 0.04 0.44
Elastic Net 0.13 0.37 0.11 0.42 0.08 0.48
Adaptive 0.08 0.34 0.06 0.37 0.04 0.43
Relaxed 0.07 0.52 0.06 0.41 0.04 0.45
0.3 MCP 0.02 0.42 0.02 0.46 0.01 0.49
SIS-SCAD 0.04 0.74 0.02 0.73 0.01 0.60
RuleFit 0.23 0.30 0.18 0.42 0.11 0.53
RE-Lasso-100 0.65 0.12 0.60 0.24 0.51 0.41
RE-EN-100 0.76 0.12 0.71 0.23 0.65 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.56 0.23 0.45 0.41 0.31 0.54
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.52 0.21 0.47 0.32 0.40 0.51
Split-EN-10 0.58 0.21 0.54 0.32 0.50 0.51
Lasso 0.07 0.32 0.05 0.36 0.03 0.41
Elastic Net 0.12 0.34 0.09 0.40 0.06 0.46
Adaptive 0.07 0.32 0.05 0.36 0.03 0.40
Relaxed 0.06 0.46 0.05 0.37 0.03 0.41
0.2 MCP 0.02 0.45 0.01 0.49 0.01 0.48
SIS-SCAD 0.03 0.63 0.01 0.64 0.01 0.54
RuleFit 0.19 0.29 0.13 0.40 0.09 0.52
RE-Lasso-100 0.59 0.13 0.52 0.24 0.46 0.41
RE-EN-100 0.72 0.12 0.65 0.23 0.61 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.50 0.25 0.37 0.39 0.26 0.53
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 60: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.39 0.19 0.42 0.33 0.34 0.49
Split-EN-10 0.40 0.20 0.51 0.31 0.45 0.48
Lasso 0.06 0.32 0.06 0.36 0.03 0.39
Elastic Net 0.11 0.34 0.10 0.38 0.06 0.43
Adaptive 0.07 0.33 0.05 0.34 0.03 0.39
Relaxed 0.05 0.62 0.05 0.37 0.03 0.40
0.4 MCP 0.02 0.31 0.01 0.37 0.01 0.40
SIS-SCAD 0.03 0.66 0.01 0.52 0.01 0.56
RuleFit 0.18 0.26 0.12 0.38 0.09 0.51
RE-Lasso-100 0.53 0.11 0.52 0.24 0.44 0.40
RE-EN-100 0.67 0.11 0.65 0.24 0.59 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.40 0.15 0.31 0.35 0.21 0.47
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.33 0.21 0.42 0.32 0.35 0.51
Split-EN-10 0.36 0.23 0.52 0.30 0.44 0.50
Lasso 0.06 0.34 0.05 0.33 0.03 0.40
Elastic Net 0.10 0.34 0.09 0.35 0.06 0.44
Adaptive 0.06 0.32 0.04 0.31 0.03 0.39
Relaxed 0.05 0.58 0.04 0.34 0.03 0.40
0.3 MCP 0.01 0.35 0.01 0.37 0.01 0.46
SIS-SCAD 0.02 0.64 0.01 0.48 0.01 0.54
RuleFit 0.15 0.25 0.12 0.38 0.08 0.51
RE-Lasso-100 0.50 0.11 0.48 0.24 0.43 0.40
RE-EN-100 0.63 0.11 0.62 0.23 0.57 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.36 0.16 0.28 0.33 0.20 0.47
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.37 0.19 0.44 0.34 0.34 0.52
Split-EN-10 0.39 0.19 0.54 0.32 0.44 0.51
Lasso 0.05 0.31 0.04 0.33 0.03 0.39
Elastic Net 0.09 0.32 0.08 0.36 0.05 0.41
Adaptive 0.05 0.30 0.04 0.32 0.03 0.37
Relaxed 0.04 0.50 0.04 0.33 0.02 0.38
0.2 MCP 0.01 0.58 0.01 0.35 0.00 0.43
SIS-SCAD 0.02 0.63 0.01 0.53 0.00 0.43
RuleFit 0.13 0.28 0.10 0.41 0.07 0.51
RE-Lasso-100 0.45 0.12 0.46 0.25 0.38 0.39
RE-EN-100 0.57 0.11 0.61 0.24 0.53 0.40
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.32 0.18 0.26 0.34 0.18 0.45
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 61: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.80 0.48 0.93 0.86 0.87 0.86 0.70 0.93 0.93 0.64 0.89 0.81 0.94 0.97 0.48
Split-EN-10 0.80 0.49 0.93 0.87 0.85 0.86 0.71 0.92 0.94 0.63 0.90 0.82 0.94 0.97 0.46
Lasso 0.77 0.41 0.92 0.82 0.98 0.81 0.60 0.91 0.88 0.86 0.83 0.70 0.90 0.91 0.78
Elastic Net 0.78 0.44 0.92 0.83 0.95 0.83 0.64 0.91 0.91 0.76 0.85 0.74 0.92 0.94 0.65
Adaptive 0.73 0.15 0.98 0.72 1.10 0.75 0.29 0.96 0.82 1.01 0.75 0.48 0.92 0.85 0.98
Relaxed 0.77 0.44 0.90 0.81 1.45 0.81 0.62 0.89 0.88 1.43 0.81 0.69 0.88 0.89 1.02
0.4 MCP 0.74 0.32 0.91 0.73 1.13 0.76 0.48 0.89 0.81 1.04 0.75 0.55 0.87 0.83 1.05
SIS-SCAD 0.73 0.36 0.88 0.74 1.30 0.74 0.49 0.85 0.78 1.22 0.73 0.53 0.84 0.79 1.16
RuleFit 0.75 0.42 0.88 0.75 1.51 0.75 0.51 0.87 0.79 1.37 0.75 0.57 0.86 0.81 1.30
RE-Lasso-100 0.81 0.51 0.93 0.87 0.85 0.85 0.67 0.93 0.93 0.66 0.88 0.76 0.95 0.96 0.54
RE-EN-100 0.81 0.51 0.93 0.87 0.86 0.85 0.68 0.93 0.93 0.64 0.89 0.78 0.95 0.97 0.51
RGLM-100 0.77 0.25 0.98 0.86 0.98 0.81 0.45 0.97 0.93 0.90 0.83 0.57 0.98 0.97 0.89
RF-500 0.80 0.48 0.94 0.87 0.85 0.84 0.62 0.94 0.93 0.76 0.86 0.70 0.95 0.96 0.74
XGB 0.73 0.43 0.86 0.74 1.10 0.75 0.50 0.87 0.80 1.03 0.74 0.57 0.84 0.81 1.04
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.84 0.34 0.97 0.88 0.73 0.87 0.55 0.96 0.94 0.59 0.90 0.73 0.97 0.97 0.44
Split-EN-10 0.84 0.35 0.97 0.88 0.73 0.87 0.57 0.96 0.94 0.57 0.91 0.75 0.97 0.97 0.42
Lasso 0.82 0.29 0.96 0.82 0.85 0.83 0.44 0.95 0.88 0.77 0.84 0.58 0.94 0.91 0.72
Elastic Net 0.83 0.30 0.96 0.84 0.81 0.85 0.48 0.96 0.91 0.70 0.86 0.64 0.95 0.94 0.61
Adaptive 0.80 0.08 0.99 0.65 0.96 0.79 0.11 0.99 0.75 0.95 0.79 0.29 0.97 0.82 0.92
Relaxed 0.81 0.30 0.95 0.80 1.06 0.83 0.45 0.94 0.87 0.88 0.83 0.58 0.93 0.91 0.94
0.3 MCP 0.80 0.20 0.95 0.70 0.98 0.79 0.27 0.95 0.77 0.98 0.78 0.40 0.92 0.81 0.96
SIS-SCAD 0.79 0.18 0.96 0.75 0.98 0.78 0.26 0.94 0.78 1.00 0.76 0.40 0.90 0.79 1.06
RuleFit 0.79 0.31 0.91 0.71 1.53 0.79 0.34 0.93 0.76 1.20 0.77 0.46 0.89 0.77 1.48
RE-Lasso-100 0.85 0.39 0.97 0.89 0.70 0.87 0.52 0.97 0.93 0.59 0.88 0.64 0.97 0.97 0.51
RE-EN-100 0.85 0.42 0.96 0.89 0.71 0.87 0.55 0.97 0.94 0.58 0.89 0.68 0.97 0.97 0.47
RGLM-100 0.81 0.11 0.99 0.87 0.82 0.81 0.18 1.00 0.93 0.79 0.82 0.37 1.00 0.97 0.79
RF-500 0.84 0.31 0.98 0.88 0.72 0.85 0.40 0.98 0.93 0.67 0.86 0.55 0.98 0.96 0.67
XGB 0.80 0.31 0.93 0.77 0.91 0.80 0.35 0.94 0.80 0.90 0.79 0.44 0.91 0.82 0.93
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.88 0.14 0.99 0.88 0.63 0.90 0.35 0.98 0.93 0.49 0.91 0.42 0.99 0.97 0.40
Split-EN-10 0.88 0.15 0.99 0.88 0.62 0.90 0.35 0.99 0.94 0.48 0.91 0.44 0.99 0.98 0.39
Lasso 0.87 0.14 0.99 0.82 0.65 0.88 0.25 0.98 0.87 0.62 0.88 0.33 0.98 0.90 0.57
Elastic Net 0.88 0.15 0.99 0.84 0.63 0.88 0.28 0.98 0.90 0.57 0.89 0.36 0.99 0.93 0.50
Adaptive 0.87 0.01 1.00 0.60 0.79 0.86 0.02 1.00 0.66 0.77 0.85 0.06 1.00 0.71 0.76
Relaxed 0.87 0.16 0.98 0.80 0.96 0.87 0.26 0.97 0.84 1.83 0.88 0.37 0.97 0.90 0.69
0.2 MCP 0.86 0.06 0.99 0.69 0.83 0.86 0.09 0.98 0.70 0.89 0.85 0.14 0.98 0.75 0.77
SIS-SCAD 0.86 0.06 0.98 0.73 0.84 0.86 0.10 0.98 0.76 0.79 0.85 0.10 0.98 0.77 0.80
RuleFit 0.85 0.18 0.96 0.72 1.05 0.85 0.22 0.95 0.73 1.04 0.85 0.22 0.96 0.75 1.02
RE-Lasso-100 0.89 0.29 0.98 0.90 0.53 0.91 0.45 0.98 0.94 0.43 0.91 0.46 0.99 0.97 0.38
RE-EN-100 0.89 0.33 0.98 0.90 0.52 0.90 0.45 0.98 0.94 0.44 0.92 0.49 0.99 0.97 0.36
RGLM-100 0.87 0.03 1.00 0.88 0.65 0.87 0.06 1.00 0.93 0.60 0.86 0.08 1.00 0.97 0.60
RF-500 0.88 0.12 0.99 0.87 0.62 0.88 0.20 1.00 0.92 0.54 0.88 0.21 1.00 0.96 0.52
XGB 0.87 0.14 0.98 0.77 0.71 0.86 0.20 0.97 0.80 0.71 0.86 0.20 0.97 0.83 0.70
Table 62: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.82 0.57 0.93 0.90 0.78 0.86 0.71 0.93 0.94 0.62 0.89 0.80 0.93 0.96 0.52
Split-EN-10 0.82 0.56 0.93 0.90 0.78 0.86 0.71 0.93 0.94 0.61 0.89 0.82 0.94 0.97 0.49
Lasso 0.79 0.51 0.91 0.85 0.94 0.81 0.59 0.91 0.89 0.84 0.82 0.70 0.90 0.91 0.78
Elastic Net 0.80 0.53 0.92 0.87 0.87 0.82 0.62 0.92 0.90 0.78 0.85 0.74 0.91 0.94 0.67
Adaptive 0.74 0.21 0.97 0.76 1.06 0.74 0.27 0.96 0.81 1.03 0.77 0.50 0.92 0.87 0.95
Relaxed 0.79 0.53 0.90 0.85 1.16 0.80 0.61 0.90 0.88 1.06 0.82 0.71 0.89 0.91 1.07
0.4 MCP 0.77 0.46 0.90 0.82 1.00 0.76 0.45 0.90 0.80 0.99 0.75 0.57 0.87 0.82 1.02
SIS-SCAD 0.76 0.45 0.90 0.82 1.04 0.75 0.46 0.88 0.79 1.10 0.73 0.58 0.82 0.79 1.29
RuleFit 0.76 0.50 0.88 0.79 1.38 0.76 0.50 0.88 0.80 1.32 0.75 0.59 0.85 0.82 1.22
RE-Lasso-100 0.82 0.56 0.93 0.89 0.79 0.84 0.64 0.93 0.93 0.68 0.87 0.76 0.94 0.96 0.57
RE-EN-100 0.82 0.56 0.93 0.89 0.80 0.84 0.66 0.93 0.93 0.67 0.88 0.78 0.94 0.96 0.53
RGLM-100 0.78 0.32 0.97 0.89 0.94 0.80 0.42 0.98 0.93 0.90 0.84 0.62 0.97 0.96 0.88
RF-500 0.82 0.57 0.93 0.90 0.77 0.84 0.63 0.94 0.93 0.74 0.86 0.72 0.94 0.95 0.74
XGB 0.77 0.52 0.87 0.81 1.00 0.76 0.50 0.88 0.80 1.01 0.74 0.59 0.83 0.81 1.05
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.38 0.97 0.90 0.68 0.87 0.55 0.97 0.94 0.57 0.90 0.73 0.96 0.96 0.47
Split-EN-10 0.85 0.38 0.97 0.90 0.67 0.87 0.55 0.97 0.94 0.56 0.90 0.75 0.96 0.97 0.44
Lasso 0.83 0.31 0.97 0.86 0.79 0.84 0.43 0.95 0.88 0.75 0.85 0.59 0.94 0.91 0.70
Elastic Net 0.84 0.34 0.97 0.87 0.75 0.85 0.47 0.96 0.91 0.69 0.87 0.65 0.94 0.94 0.61
Adaptive 0.80 0.06 1.00 0.72 0.94 0.79 0.10 0.99 0.72 0.96 0.79 0.28 0.97 0.83 0.90
Relaxed 0.83 0.36 0.95 0.85 0.98 0.83 0.46 0.95 0.88 1.10 0.84 0.63 0.91 0.90 0.90
0.3 MCP 0.82 0.26 0.96 0.81 0.86 0.80 0.31 0.94 0.78 1.12 0.79 0.39 0.92 0.80 0.95
SIS-SCAD 0.81 0.21 0.96 0.81 0.95 0.79 0.27 0.94 0.79 0.97 0.77 0.37 0.91 0.79 1.04
RuleFit 0.81 0.34 0.93 0.75 1.20 0.79 0.39 0.91 0.76 1.40 0.78 0.45 0.89 0.77 1.32
RE-Lasso-100 0.85 0.41 0.97 0.90 0.67 0.87 0.52 0.97 0.93 0.59 0.89 0.65 0.97 0.96 0.51
RE-EN-100 0.85 0.43 0.96 0.90 0.67 0.87 0.55 0.97 0.94 0.57 0.89 0.70 0.96 0.96 0.47
RGLM-100 0.82 0.14 0.99 0.89 0.79 0.82 0.20 1.00 0.93 0.78 0.84 0.39 0.99 0.96 0.77
RF-500 0.85 0.38 0.97 0.90 0.67 0.85 0.45 0.97 0.92 0.66 0.86 0.57 0.97 0.95 0.66
XGB 0.82 0.35 0.94 0.81 0.86 0.80 0.38 0.92 0.82 0.89 0.79 0.47 0.90 0.82 0.92
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.90 0.21 0.99 0.91 0.53 0.90 0.37 0.99 0.94 0.49 0.90 0.48 0.98 0.95 0.48
Split-EN-10 0.90 0.21 0.99 0.91 0.52 0.90 0.37 0.99 0.95 0.48 0.90 0.50 0.98 0.96 0.43
Lasso 0.89 0.18 0.99 0.84 0.59 0.88 0.27 0.98 0.88 0.60 0.87 0.36 0.97 0.90 0.63
Elastic Net 0.89 0.20 0.99 0.87 0.55 0.88 0.27 0.99 0.90 0.56 0.88 0.39 0.98 0.93 0.56
Adaptive 0.88 0.01 1.00 0.62 0.73 0.86 0.03 1.00 0.65 0.78 0.84 0.09 0.99 0.77 0.78
Relaxed 0.89 0.20 0.98 0.84 0.71 0.88 0.28 0.98 0.86 0.88 0.86 0.38 0.96 0.88 1.02
0.2 MCP 0.88 0.12 0.98 0.73 0.71 0.86 0.12 0.98 0.75 0.81 0.84 0.18 0.97 0.76 0.87
SIS-SCAD 0.88 0.07 0.99 0.77 0.67 0.86 0.12 0.98 0.78 0.79 0.83 0.12 0.98 0.77 0.85
RuleFit 0.88 0.24 0.97 0.77 0.86 0.85 0.29 0.94 0.74 1.09 0.83 0.27 0.94 0.74 1.23
RE-Lasso-100 0.90 0.35 0.98 0.91 0.46 0.91 0.46 0.98 0.94 0.43 0.90 0.50 0.98 0.96 0.42
RE-EN-100 0.90 0.34 0.98 0.90 0.48 0.91 0.45 0.98 0.94 0.43 0.91 0.55 0.98 0.96 0.40
RGLM-100 0.88 0.03 1.00 0.89 0.59 0.87 0.07 1.00 0.94 0.61 0.85 0.13 1.00 0.96 0.62
RF-500 0.89 0.17 0.99 0.90 0.53 0.89 0.25 0.99 0.93 0.53 0.87 0.28 0.99 0.94 0.56
XGB 0.88 0.19 0.98 0.83 0.62 0.87 0.22 0.97 0.83 0.68 0.84 0.25 0.96 0.82 0.76
Table 63: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.87 0.78 0.92 0.95 0.61 0.90 0.86 0.93 0.97 0.44 0.93 0.91 0.95 0.99 0.31
Split-EN-10 0.87 0.78 0.92 0.95 0.61 0.91 0.87 0.93 0.97 0.44 0.94 0.92 0.95 0.99 0.30
Lasso 0.85 0.74 0.90 0.93 0.72 0.88 0.82 0.91 0.95 0.58 0.89 0.84 0.92 0.97 0.49
Elastic Net 0.86 0.76 0.91 0.94 0.66 0.89 0.83 0.92 0.96 0.52 0.91 0.87 0.94 0.98 0.40
Adaptive 0.81 0.58 0.94 0.91 0.83 0.84 0.70 0.93 0.94 0.74 0.87 0.76 0.93 0.96 0.64
Relaxed 0.84 0.75 0.89 0.92 1.01 0.87 0.82 0.90 0.95 0.87 0.88 0.84 0.90 0.95 1.03
0.4 MCP 0.81 0.66 0.89 0.89 0.86 0.82 0.71 0.89 0.91 0.78 0.83 0.76 0.88 0.92 0.74
SIS-SCAD 0.79 0.61 0.89 0.88 0.88 0.80 0.68 0.87 0.89 0.86 0.81 0.72 0.87 0.89 0.86
RuleFit 0.80 0.66 0.88 0.87 1.06 0.82 0.72 0.87 0.89 1.05 0.81 0.72 0.87 0.88 1.09
RE-Lasso-100 0.87 0.77 0.92 0.94 0.62 0.90 0.86 0.92 0.97 0.45 0.93 0.89 0.95 0.99 0.34
RE-EN-100 0.87 0.78 0.91 0.94 0.62 0.90 0.85 0.93 0.97 0.45 0.93 0.90 0.95 0.99 0.32
RGLM-100 0.86 0.71 0.94 0.94 0.72 0.89 0.80 0.95 0.97 0.65 0.93 0.85 0.97 0.99 0.60
RF-500 0.87 0.76 0.92 0.94 0.62 0.90 0.83 0.93 0.97 0.56 0.92 0.87 0.95 0.98 0.52
XGB 0.81 0.68 0.88 0.89 0.84 0.83 0.74 0.89 0.92 0.75 0.84 0.76 0.89 0.92 0.72
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.88 0.70 0.94 0.95 0.55 0.91 0.76 0.97 0.97 0.41 0.94 0.86 0.97 0.99 0.29
Split-EN-10 0.88 0.70 0.94 0.95 0.55 0.91 0.76 0.97 0.98 0.40 0.94 0.87 0.97 0.99 0.28
Lasso 0.86 0.63 0.93 0.93 0.66 0.88 0.68 0.95 0.95 0.56 0.90 0.78 0.95 0.97 0.46
Elastic Net 0.87 0.66 0.94 0.94 0.61 0.89 0.72 0.96 0.96 0.50 0.92 0.82 0.96 0.98 0.37
Adaptive 0.82 0.36 0.97 0.87 0.79 0.82 0.41 0.98 0.94 0.72 0.87 0.59 0.97 0.95 0.62
Relaxed 0.86 0.66 0.92 0.92 0.92 0.88 0.70 0.94 0.95 0.73 0.89 0.78 0.93 0.96 0.88
0.3 MCP 0.82 0.46 0.93 0.87 0.81 0.82 0.50 0.94 0.89 0.81 0.84 0.60 0.93 0.91 0.72
SIS-SCAD 0.81 0.39 0.95 0.88 0.82 0.81 0.42 0.95 0.88 0.81 0.81 0.50 0.94 0.89 0.79
RuleFit 0.83 0.56 0.92 0.86 1.04 0.83 0.51 0.95 0.89 0.92 0.84 0.59 0.94 0.88 1.00
RE-Lasso-100 0.88 0.72 0.93 0.95 0.54 0.90 0.74 0.97 0.97 0.43 0.93 0.84 0.97 0.99 0.32
RE-EN-100 0.88 0.73 0.93 0.95 0.54 0.91 0.76 0.96 0.97 0.41 0.94 0.86 0.97 0.99 0.30
RGLM-100 0.86 0.56 0.96 0.95 0.64 0.88 0.59 0.99 0.97 0.59 0.92 0.72 0.99 0.99 0.54
RF-500 0.88 0.67 0.95 0.95 0.56 0.89 0.68 0.97 0.97 0.52 0.92 0.78 0.98 0.98 0.47
XGB 0.84 0.59 0.92 0.89 0.77 0.84 0.57 0.94 0.91 0.73 0.85 0.64 0.93 0.92 0.72
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.91 0.48 0.98 0.96 0.44 0.93 0.66 0.98 0.98 0.31 0.95 0.81 0.98 0.99 0.25
Split-EN-10 0.91 0.49 0.98 0.96 0.43 0.93 0.67 0.98 0.98 0.30 0.95 0.83 0.98 0.99 0.23
Lasso 0.90 0.44 0.98 0.93 0.50 0.91 0.55 0.97 0.95 0.43 0.92 0.70 0.97 0.97 0.38
Elastic Net 0.90 0.47 0.98 0.94 0.46 0.92 0.59 0.98 0.97 0.37 0.93 0.74 0.97 0.98 0.31
Adaptive 0.86 0.07 1.00 0.79 0.72 0.88 0.19 0.99 0.91 0.59 0.87 0.32 0.99 0.94 0.59
Relaxed 0.89 0.45 0.97 0.92 0.70 0.91 0.60 0.96 0.94 0.96 0.91 0.72 0.95 0.96 0.74
0.2 MCP 0.87 0.24 0.98 0.85 0.69 0.87 0.30 0.97 0.87 0.65 0.86 0.41 0.95 0.88 0.70
SIS-SCAD 0.86 0.14 0.99 0.86 0.68 0.87 0.22 0.98 0.88 0.65 0.85 0.29 0.97 0.89 0.67
RuleFit 0.87 0.36 0.96 0.82 0.96 0.87 0.41 0.95 0.81 1.07 0.87 0.47 0.95 0.85 1.00
RE-Lasso-100 0.91 0.65 0.96 0.96 0.39 0.93 0.75 0.97 0.98 0.30 0.95 0.80 0.98 0.99 0.26
RE-EN-100 0.91 0.66 0.96 0.96 0.39 0.94 0.76 0.97 0.98 0.30 0.95 0.83 0.98 0.99 0.25
RGLM-100 0.89 0.33 0.99 0.95 0.48 0.91 0.42 0.99 0.98 0.42 0.92 0.55 1.00 0.99 0.44
RF-500 0.90 0.44 0.99 0.95 0.45 0.92 0.51 0.99 0.97 0.39 0.93 0.65 0.99 0.98 0.39
XGB 0.88 0.39 0.97 0.91 0.58 0.88 0.44 0.95 0.92 0.58 0.87 0.54 0.95 0.91 0.63
Table 64: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.83 0.60 0.92 0.89 0.76 0.87 0.74 0.93 0.95 0.56 0.91 0.85 0.94 0.97 0.41
Split-EN-10 0.82 0.59 0.92 0.89 0.76 0.87 0.74 0.93 0.95 0.57 0.91 0.85 0.94 0.97 0.40
Lasso 0.81 0.55 0.91 0.87 0.85 0.84 0.68 0.91 0.92 0.70 0.86 0.76 0.91 0.94 0.63
Elastic Net 0.81 0.55 0.91 0.87 0.83 0.85 0.70 0.92 0.92 0.67 0.88 0.79 0.92 0.95 0.56
Adaptive 0.77 0.28 0.97 0.85 0.95 0.81 0.50 0.95 0.89 0.81 0.84 0.67 0.93 0.92 0.74
Relaxed 0.80 0.54 0.91 0.86 0.87 0.84 0.69 0.91 0.92 0.72 0.86 0.77 0.91 0.94 0.71
0.4 MCP 0.78 0.49 0.90 0.83 0.93 0.80 0.56 0.91 0.86 0.85 0.80 0.66 0.88 0.88 0.85
SIS-SCAD 0.78 0.48 0.91 0.83 0.95 0.78 0.54 0.89 0.84 0.96 0.77 0.61 0.86 0.84 1.02
RuleFit 0.78 0.51 0.90 0.83 1.26 0.81 0.54 0.92 0.86 1.05 0.81 0.62 0.91 0.88 0.97
RE-Lasso-100 0.82 0.58 0.92 0.89 0.78 0.87 0.70 0.94 0.94 0.59 0.90 0.81 0.95 0.97 0.46
RE-EN-100 0.82 0.59 0.92 0.89 0.78 0.87 0.72 0.93 0.94 0.58 0.90 0.82 0.95 0.97 0.44
RGLM-100 0.79 0.36 0.97 0.88 0.91 0.83 0.47 0.98 0.94 0.84 0.87 0.65 0.98 0.97 0.82
RF-500 0.83 0.62 0.91 0.89 0.76 0.87 0.69 0.94 0.94 0.65 0.89 0.77 0.96 0.97 0.63
XGB 0.78 0.50 0.90 0.83 0.93 0.80 0.54 0.91 0.86 0.87 0.80 0.62 0.90 0.88 0.85
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.45 0.96 0.89 0.66 0.89 0.65 0.96 0.95 0.51 0.91 0.77 0.97 0.98 0.39
Split-EN-10 0.85 0.44 0.96 0.89 0.67 0.89 0.64 0.96 0.95 0.50 0.92 0.78 0.97 0.98 0.38
Lasso 0.84 0.41 0.95 0.86 0.73 0.86 0.55 0.95 0.92 0.64 0.87 0.67 0.95 0.94 0.59
Elastic Net 0.84 0.42 0.95 0.87 0.72 0.87 0.58 0.95 0.93 0.61 0.89 0.70 0.95 0.95 0.53
Adaptive 0.81 0.10 0.99 0.82 0.84 0.82 0.29 0.98 0.89 0.75 0.84 0.48 0.97 0.92 0.70
Relaxed 0.84 0.40 0.95 0.86 0.76 0.86 0.55 0.95 0.91 0.68 0.87 0.69 0.94 0.94 0.68
0.3 MCP 0.82 0.31 0.95 0.81 0.83 0.82 0.41 0.94 0.85 0.83 0.82 0.50 0.94 0.88 0.80
SIS-SCAD 0.81 0.31 0.94 0.82 0.91 0.81 0.39 0.94 0.84 0.86 0.79 0.46 0.92 0.84 0.91
RuleFit 0.82 0.35 0.94 0.81 1.19 0.83 0.41 0.95 0.86 1.03 0.82 0.47 0.95 0.88 1.03
RE-Lasso-100 0.85 0.43 0.96 0.89 0.71 0.88 0.59 0.96 0.94 0.54 0.90 0.69 0.98 0.97 0.45
RE-EN-100 0.85 0.44 0.96 0.89 0.71 0.88 0.60 0.96 0.94 0.53 0.90 0.72 0.98 0.97 0.42
RGLM-100 0.82 0.17 0.99 0.87 0.78 0.83 0.27 0.99 0.94 0.74 0.83 0.41 1.00 0.97 0.74
RF-500 0.85 0.46 0.95 0.89 0.68 0.88 0.56 0.97 0.94 0.58 0.88 0.62 0.98 0.97 0.58
XGB 0.82 0.35 0.94 0.81 0.82 0.83 0.41 0.95 0.87 0.77 0.82 0.47 0.95 0.88 0.79
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.89 0.29 0.98 0.90 0.52 0.91 0.47 0.98 0.95 0.40 0.93 0.63 0.99 0.98 0.31
Split-EN-10 0.89 0.28 0.98 0.90 0.52 0.91 0.49 0.98 0.96 0.40 0.93 0.64 0.99 0.98 0.30
Lasso 0.88 0.26 0.98 0.87 0.60 0.90 0.39 0.98 0.92 0.51 0.90 0.48 0.98 0.94 0.49
Elastic Net 0.89 0.25 0.98 0.88 0.57 0.90 0.41 0.98 0.93 0.48 0.91 0.52 0.98 0.96 0.43
Adaptive 0.87 0.03 1.00 0.79 0.67 0.87 0.07 1.00 0.83 0.67 0.86 0.18 1.00 0.90 0.61
Relaxed 0.88 0.26 0.98 0.86 0.63 0.89 0.41 0.97 0.92 0.56 0.90 0.52 0.97 0.94 0.57
0.2 MCP 0.87 0.13 0.98 0.78 0.67 0.87 0.20 0.98 0.81 0.69 0.85 0.25 0.97 0.84 0.71
SIS-SCAD 0.87 0.12 0.98 0.80 0.72 0.87 0.18 0.98 0.83 0.71 0.85 0.23 0.97 0.84 0.72
RuleFit 0.87 0.21 0.97 0.78 0.93 0.88 0.27 0.97 0.83 0.85 0.87 0.30 0.98 0.85 0.88
RE-Lasso-100 0.89 0.27 0.98 0.90 0.56 0.91 0.42 0.99 0.95 0.42 0.91 0.49 0.99 0.97 0.38
RE-EN-100 0.89 0.27 0.98 0.90 0.56 0.91 0.43 0.99 0.95 0.42 0.92 0.53 0.99 0.98 0.36
RGLM-100 0.88 0.06 1.00 0.89 0.59 0.87 0.09 1.00 0.95 0.56 0.86 0.15 1.00 0.98 0.56
RF-500 0.89 0.26 0.98 0.90 0.52 0.90 0.33 0.99 0.95 0.46 0.90 0.37 1.00 0.97 0.46
XGB 0.87 0.20 0.97 0.81 0.64 0.88 0.26 0.98 0.86 0.61 0.87 0.28 0.98 0.87 0.64
Table 65: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.85 0.68 0.92 0.92 0.66 0.89 0.77 0.94 0.96 0.52 0.91 0.85 0.94 0.97 0.42
Split-EN-10 0.85 0.67 0.92 0.92 0.67 0.89 0.77 0.94 0.96 0.52 0.91 0.86 0.94 0.97 0.41
Lasso 0.83 0.63 0.92 0.90 0.76 0.85 0.70 0.92 0.93 0.68 0.87 0.78 0.91 0.94 0.61
Elastic Net 0.83 0.63 0.92 0.90 0.75 0.86 0.71 0.93 0.93 0.64 0.88 0.80 0.92 0.95 0.56
Adaptive 0.81 0.47 0.95 0.89 0.84 0.82 0.55 0.95 0.91 0.79 0.84 0.67 0.93 0.93 0.74
Relaxed 0.82 0.63 0.90 0.89 0.82 0.85 0.70 0.92 0.93 0.71 0.86 0.78 0.91 0.94 0.66
0.4 MCP 0.81 0.61 0.90 0.88 0.81 0.82 0.64 0.91 0.90 0.81 0.81 0.67 0.89 0.89 0.81
SIS-SCAD 0.81 0.58 0.91 0.88 0.87 0.80 0.59 0.90 0.87 0.90 0.77 0.60 0.87 0.85 1.00
RuleFit 0.81 0.60 0.90 0.87 1.11 0.81 0.59 0.92 0.88 1.02 0.81 0.64 0.90 0.89 0.98
RE-Lasso-100 0.84 0.64 0.93 0.92 0.70 0.87 0.72 0.94 0.94 0.59 0.89 0.81 0.94 0.97 0.48
RE-EN-100 0.84 0.64 0.93 0.92 0.71 0.87 0.73 0.94 0.95 0.59 0.90 0.82 0.94 0.97 0.46
RGLM-100 0.81 0.47 0.97 0.91 0.86 0.83 0.53 0.98 0.94 0.83 0.86 0.66 0.98 0.97 0.81
RF-500 0.85 0.70 0.92 0.93 0.66 0.87 0.73 0.94 0.95 0.60 0.88 0.78 0.94 0.96 0.61
XGB 0.81 0.61 0.90 0.87 0.84 0.81 0.60 0.92 0.89 0.81 0.80 0.63 0.90 0.88 0.84
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.87 0.54 0.96 0.92 0.57 0.91 0.71 0.96 0.96 0.44 0.91 0.78 0.96 0.97 0.39
Split-EN-10 0.87 0.53 0.96 0.92 0.57 0.90 0.70 0.96 0.96 0.44 0.92 0.79 0.96 0.97 0.37
Lasso 0.86 0.50 0.95 0.90 0.66 0.87 0.61 0.95 0.93 0.57 0.87 0.67 0.94 0.94 0.58
Elastic Net 0.86 0.51 0.95 0.90 0.63 0.88 0.63 0.95 0.94 0.54 0.88 0.70 0.95 0.95 0.52
Adaptive 0.84 0.26 0.98 0.89 0.72 0.85 0.38 0.98 0.91 0.69 0.84 0.46 0.97 0.92 0.70
Relaxed 0.85 0.50 0.94 0.89 0.78 0.87 0.62 0.94 0.93 0.64 0.87 0.68 0.94 0.94 0.63
0.3 MCP 0.84 0.45 0.94 0.87 0.72 0.85 0.53 0.94 0.89 0.68 0.82 0.50 0.94 0.88 0.81
SIS-SCAD 0.84 0.43 0.94 0.87 0.82 0.83 0.48 0.93 0.87 0.81 0.80 0.46 0.92 0.83 0.99
RuleFit 0.84 0.47 0.94 0.86 1.00 0.84 0.47 0.95 0.87 0.94 0.83 0.49 0.94 0.88 0.96
RE-Lasso-100 0.87 0.50 0.96 0.92 0.61 0.89 0.63 0.97 0.95 0.50 0.90 0.69 0.97 0.96 0.45
RE-EN-100 0.87 0.50 0.96 0.92 0.61 0.89 0.64 0.96 0.95 0.50 0.90 0.72 0.97 0.97 0.43
RGLM-100 0.84 0.27 0.99 0.91 0.71 0.85 0.34 0.99 0.95 0.70 0.85 0.41 1.00 0.97 0.72
RF-500 0.88 0.58 0.95 0.93 0.56 0.89 0.64 0.97 0.95 0.52 0.89 0.64 0.97 0.96 0.56
XGB 0.85 0.48 0.94 0.87 0.71 0.84 0.47 0.95 0.88 0.70 0.82 0.47 0.94 0.87 0.78
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.91 0.36 0.99 0.93 0.44 0.92 0.54 0.98 0.96 0.37 0.92 0.62 0.98 0.97 0.35
Split-EN-10 0.91 0.35 0.99 0.93 0.45 0.92 0.54 0.98 0.96 0.37 0.93 0.64 0.98 0.97 0.34
Lasso 0.89 0.29 0.98 0.90 0.52 0.90 0.40 0.98 0.92 0.49 0.89 0.47 0.97 0.94 0.50
Elastic Net 0.90 0.30 0.99 0.91 0.50 0.90 0.42 0.98 0.93 0.47 0.90 0.51 0.98 0.95 0.45
Adaptive 0.88 0.06 1.00 0.83 0.63 0.87 0.08 1.00 0.85 0.63 0.85 0.14 0.99 0.89 0.66
Relaxed 0.89 0.31 0.98 0.89 0.60 0.90 0.41 0.97 0.92 0.59 0.89 0.46 0.97 0.92 0.64
0.2 MCP 0.88 0.25 0.98 0.85 0.65 0.87 0.24 0.98 0.84 0.66 0.85 0.25 0.97 0.84 0.71
SIS-SCAD 0.88 0.22 0.98 0.86 0.66 0.87 0.20 0.98 0.85 0.67 0.85 0.21 0.97 0.83 0.74
RuleFit 0.89 0.29 0.97 0.83 0.79 0.88 0.29 0.98 0.84 0.79 0.86 0.29 0.97 0.85 0.94
RE-Lasso-100 0.90 0.32 0.99 0.92 0.49 0.91 0.41 0.99 0.95 0.43 0.91 0.48 0.99 0.97 0.41
RE-EN-100 0.90 0.31 0.99 0.92 0.51 0.91 0.45 0.99 0.95 0.42 0.91 0.52 0.99 0.97 0.39
RGLM-100 0.88 0.10 1.00 0.91 0.55 0.88 0.11 1.00 0.94 0.54 0.86 0.15 1.00 0.97 0.57
RF-500 0.91 0.38 0.99 0.93 0.45 0.91 0.40 0.99 0.95 0.43 0.90 0.40 0.99 0.96 0.47
XGB 0.89 0.31 0.98 0.87 0.57 0.88 0.27 0.98 0.87 0.59 0.86 0.28 0.98 0.87 0.65
Table 66: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.88 0.80 0.92 0.95 0.55 0.92 0.89 0.94 0.98 0.37 0.94 0.92 0.96 0.99 0.26
Split-EN-10 0.88 0.80 0.92 0.95 0.55 0.92 0.89 0.94 0.98 0.37 0.95 0.93 0.96 0.99 0.26
Lasso 0.86 0.77 0.91 0.94 0.62 0.90 0.85 0.92 0.97 0.48 0.91 0.87 0.94 0.98 0.40
Elastic Net 0.87 0.78 0.91 0.94 0.60 0.90 0.86 0.93 0.97 0.44 0.92 0.89 0.95 0.98 0.35
Adaptive 0.85 0.69 0.94 0.94 0.70 0.89 0.81 0.93 0.96 0.57 0.90 0.82 0.95 0.97 0.51
Relaxed 0.86 0.76 0.91 0.93 0.74 0.90 0.86 0.92 0.96 0.58 0.91 0.87 0.93 0.97 0.54
0.4 MCP 0.84 0.73 0.90 0.92 0.71 0.87 0.80 0.90 0.94 0.63 0.87 0.80 0.91 0.95 0.59
SIS-SCAD 0.84 0.72 0.90 0.91 0.76 0.84 0.76 0.89 0.92 0.74 0.85 0.77 0.90 0.93 0.69
RuleFit 0.85 0.73 0.91 0.92 0.86 0.86 0.77 0.91 0.94 0.75 0.87 0.78 0.92 0.95 0.64
RE-Lasso-100 0.87 0.79 0.92 0.95 0.59 0.91 0.87 0.94 0.98 0.39 0.94 0.90 0.96 0.99 0.30
RE-EN-100 0.87 0.79 0.92 0.95 0.59 0.91 0.87 0.94 0.98 0.39 0.94 0.91 0.96 0.99 0.29
RGLM-100 0.87 0.74 0.93 0.94 0.67 0.91 0.83 0.95 0.98 0.59 0.94 0.87 0.97 0.99 0.55
RF-500 0.88 0.79 0.92 0.95 0.55 0.91 0.86 0.94 0.98 0.46 0.93 0.89 0.96 0.99 0.44
XGB 0.85 0.74 0.91 0.93 0.69 0.87 0.79 0.92 0.95 0.58 0.88 0.80 0.93 0.96 0.54
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.90 0.75 0.95 0.96 0.47 0.93 0.85 0.95 0.98 0.33 0.95 0.89 0.97 0.99 0.24
Split-EN-10 0.90 0.74 0.95 0.96 0.48 0.93 0.85 0.95 0.98 0.33 0.95 0.90 0.97 0.99 0.24
Lasso 0.88 0.69 0.94 0.94 0.55 0.91 0.78 0.95 0.97 0.43 0.92 0.83 0.96 0.98 0.36
Elastic Net 0.89 0.70 0.95 0.95 0.52 0.91 0.80 0.95 0.97 0.40 0.93 0.85 0.96 0.98 0.31
Adaptive 0.86 0.50 0.98 0.94 0.64 0.89 0.65 0.97 0.96 0.53 0.91 0.74 0.97 0.97 0.45
Relaxed 0.88 0.69 0.94 0.94 0.68 0.90 0.79 0.94 0.96 0.55 0.91 0.84 0.94 0.97 0.46
0.3 MCP 0.86 0.62 0.94 0.92 0.65 0.86 0.65 0.94 0.92 0.62 0.87 0.70 0.93 0.94 0.60
SIS-SCAD 0.86 0.58 0.95 0.92 0.65 0.86 0.64 0.93 0.92 0.66 0.86 0.66 0.94 0.93 0.65
RuleFit 0.87 0.63 0.94 0.92 0.80 0.87 0.66 0.95 0.94 0.65 0.87 0.67 0.95 0.94 0.67
RE-Lasso-100 0.89 0.73 0.95 0.95 0.50 0.92 0.81 0.95 0.98 0.36 0.94 0.86 0.97 0.99 0.28
RE-EN-100 0.89 0.73 0.94 0.95 0.51 0.92 0.82 0.95 0.98 0.36 0.94 0.87 0.97 0.99 0.27
RGLM-100 0.88 0.64 0.96 0.95 0.58 0.91 0.72 0.98 0.98 0.52 0.93 0.78 0.99 0.99 0.49
RF-500 0.90 0.74 0.95 0.95 0.49 0.92 0.79 0.96 0.98 0.42 0.93 0.82 0.98 0.99 0.41
XGB 0.87 0.65 0.94 0.93 0.60 0.88 0.69 0.95 0.95 0.54 0.88 0.70 0.96 0.95 0.52
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.93 0.64 0.97 0.96 0.35 0.94 0.76 0.98 0.98 0.26 0.95 0.82 0.98 0.99 0.21
Split-EN-10 0.93 0.64 0.97 0.96 0.35 0.94 0.76 0.98 0.98 0.26 0.95 0.83 0.98 0.99 0.20
Lasso 0.91 0.57 0.97 0.95 0.42 0.92 0.67 0.97 0.97 0.36 0.93 0.72 0.98 0.98 0.33
Elastic Net 0.92 0.60 0.97 0.95 0.39 0.93 0.70 0.98 0.97 0.33 0.94 0.75 0.98 0.98 0.28
Adaptive 0.89 0.24 0.99 0.94 0.52 0.89 0.40 0.99 0.96 0.46 0.91 0.55 0.99 0.97 0.41
Relaxed 0.91 0.58 0.96 0.94 0.50 0.92 0.67 0.97 0.96 0.50 0.93 0.74 0.97 0.97 0.44
0.2 MCP 0.88 0.35 0.97 0.89 0.61 0.87 0.40 0.96 0.88 0.72 0.87 0.51 0.95 0.90 0.63
SIS-SCAD 0.88 0.33 0.97 0.91 0.66 0.87 0.28 0.98 0.91 0.64 0.86 0.41 0.97 0.91 0.66
RuleFit 0.90 0.49 0.96 0.87 0.73 0.89 0.49 0.96 0.87 0.76 0.88 0.53 0.96 0.90 0.73
RE-Lasso-100 0.92 0.62 0.97 0.96 0.38 0.94 0.70 0.98 0.98 0.29 0.94 0.74 0.99 0.99 0.25
RE-EN-100 0.92 0.63 0.97 0.96 0.38 0.94 0.73 0.98 0.98 0.28 0.95 0.77 0.99 0.99 0.24
RGLM-100 0.91 0.46 0.99 0.96 0.42 0.92 0.50 1.00 0.98 0.41 0.92 0.59 1.00 0.99 0.40
RF-500 0.92 0.60 0.98 0.96 0.36 0.93 0.62 0.99 0.98 0.35 0.93 0.68 0.99 0.99 0.34
XGB 0.91 0.51 0.97 0.93 0.47 0.90 0.51 0.98 0.94 0.47 0.90 0.56 0.98 0.95 0.47
Table 67: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.26 0.22 0.30 0.33 0.29 0.49
Split-EN-10 0.29 0.22 0.36 0.31 0.36 0.47
Lasso 0.04 0.26 0.04 0.35 0.03 0.45
Elastic Net 0.07 0.26 0.06 0.34 0.05 0.46
Adaptive 0.04 0.24 0.04 0.35 0.03 0.46
Relaxed 0.04 0.27 0.03 0.37 0.02 0.47
0.4 MCP 0.01 0.29 0.01 0.40 0.01 0.49
SIS-SCAD 0.01 0.38 0.01 0.47 0.00 0.55
RuleFit 0.11 0.25 0.09 0.40 0.06 0.52
RE-Lasso-100 0.49 0.13 0.46 0.25 0.43 0.43
RE-EN-100 0.61 0.12 0.60 0.23 0.57 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.28 0.18 0.23 0.31 0.17 0.47
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.23 0.22 0.28 0.33 0.28 0.49
Split-EN-10 0.29 0.21 0.38 0.31 0.36 0.47
Lasso 0.04 0.24 0.03 0.34 0.03 0.44
Elastic Net 0.06 0.23 0.06 0.34 0.05 0.46
Adaptive 0.03 0.25 0.03 0.34 0.02 0.45
Relaxed 0.03 0.27 0.03 0.35 0.02 0.46
0.3 MCP 0.01 0.24 0.01 0.36 0.01 0.45
SIS-SCAD 0.01 0.39 0.01 0.44 0.00 0.47
RuleFit 0.10 0.30 0.07 0.41 0.06 0.58
RE-Lasso-100 0.44 0.14 0.43 0.25 0.39 0.42
RE-EN-100 0.58 0.12 0.56 0.23 0.54 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.23 0.17 0.20 0.29 0.17 0.47
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.19 0.23 0.22 0.33 0.23 0.51
Split-EN-10 0.24 0.20 0.29 0.31 0.30 0.50
Lasso 0.03 0.23 0.02 0.29 0.02 0.43
Elastic Net 0.05 0.23 0.04 0.30 0.04 0.44
Adaptive 0.02 0.22 0.02 0.28 0.02 0.45
Relaxed 0.02 0.27 0.02 0.26 0.02 0.44
0.2 MCP 0.01 0.23 0.00 0.25 0.00 0.49
SIS-SCAD 0.01 0.29 0.00 0.36 0.00 0.57
RuleFit 0.07 0.28 0.05 0.43 0.04 0.61
RE-Lasso-100 0.44 0.14 0.39 0.24 0.34 0.42
RE-EN-100 0.56 0.13 0.51 0.23 0.49 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.17 0.17 0.15 0.28 0.13 0.46
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 68: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.33 0.26 0.37 0.36 0.29 0.50
Split-EN-10 0.42 0.24 0.46 0.35 0.40 0.50
Lasso 0.05 0.29 0.03 0.34 0.02 0.40
Elastic Net 0.09 0.31 0.07 0.36 0.05 0.44
Adaptive 0.04 0.28 0.03 0.33 0.02 0.39
Relaxed 0.04 0.36 0.03 0.36 0.02 0.40
0.4 MCP 0.01 0.37 0.01 0.40 0.01 0.42
SIS-SCAD 0.02 0.59 0.01 0.58 0.00 0.56
RuleFit 0.12 0.31 0.09 0.39 0.06 0.51
RE-Lasso-100 0.49 0.13 0.45 0.24 0.41 0.41
RE-EN-100 0.64 0.13 0.61 0.24 0.58 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.36 0.25 0.27 0.36 0.20 0.51
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.31 0.26 0.29 0.35 0.27 0.49
Split-EN-10 0.37 0.25 0.38 0.34 0.39 0.50
Lasso 0.04 0.29 0.03 0.33 0.02 0.39
Elastic Net 0.08 0.29 0.06 0.37 0.05 0.42
Adaptive 0.04 0.26 0.02 0.34 0.02 0.36
Relaxed 0.04 0.32 0.03 0.38 0.02 0.41
0.3 MCP 0.01 0.45 0.01 0.42 0.01 0.44
SIS-SCAD 0.01 0.63 0.01 0.54 0.00 0.51
RuleFit 0.10 0.34 0.07 0.41 0.05 0.51
RE-Lasso-100 0.47 0.14 0.42 0.24 0.37 0.40
RE-EN-100 0.60 0.13 0.58 0.23 0.55 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.31 0.24 0.22 0.35 0.18 0.50
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.26 0.27 0.27 0.36 0.21 0.47
Split-EN-10 0.34 0.24 0.35 0.34 0.33 0.50
Lasso 0.03 0.23 0.02 0.31 0.02 0.36
Elastic Net 0.06 0.26 0.05 0.35 0.04 0.40
Adaptive 0.02 0.19 0.02 0.31 0.01 0.30
Relaxed 0.03 0.22 0.02 0.31 0.01 0.38
0.2 MCP 0.01 0.35 0.00 0.39 0.00 0.39
SIS-SCAD 0.01 0.42 0.01 0.50 0.00 0.48
RuleFit 0.08 0.36 0.06 0.45 0.03 0.51
RE-Lasso-100 0.44 0.15 0.37 0.23 0.33 0.39
RE-EN-100 0.57 0.13 0.54 0.23 0.49 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.22 0.22 0.19 0.34 0.13 0.48
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 69: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.28 0.23 0.28 0.35 0.23 0.49
Split-EN-10 0.34 0.22 0.38 0.34 0.33 0.49
Lasso 0.04 0.23 0.03 0.29 0.02 0.39
Elastic Net 0.07 0.22 0.06 0.34 0.05 0.42
Adaptive 0.03 0.20 0.03 0.27 0.02 0.40
Relaxed 0.03 0.23 0.03 0.31 0.02 0.40
0.4 MCP 0.01 0.24 0.01 0.36 0.00 0.37
SIS-SCAD 0.01 0.39 0.01 0.40 0.00 0.51
RuleFit 0.10 0.28 0.07 0.40 0.05 0.54
RE-Lasso-100 0.42 0.14 0.40 0.24 0.35 0.40
RE-EN-100 0.58 0.13 0.57 0.24 0.54 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.25 0.20 0.20 0.32 0.14 0.47
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.27 0.23 0.29 0.33 0.24 0.51
Split-EN-10 0.36 0.21 0.37 0.32 0.36 0.51
Lasso 0.03 0.21 0.02 0.26 0.02 0.37
Elastic Net 0.07 0.23 0.05 0.28 0.05 0.40
Adaptive 0.03 0.20 0.02 0.24 0.02 0.36
Relaxed 0.03 0.24 0.02 0.29 0.01 0.40
0.3 MCP 0.01 0.36 0.01 0.37 0.00 0.44
SIS-SCAD 0.01 0.39 0.00 0.43 0.00 0.45
RuleFit 0.08 0.31 0.06 0.43 0.04 0.56
RE-Lasso-100 0.41 0.14 0.35 0.23 0.33 0.40
RE-EN-100 0.57 0.13 0.53 0.23 0.52 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.23 0.20 0.16 0.29 0.13 0.46
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.28 0.23 0.27 0.35 0.24 0.55
Split-EN-10 0.35 0.22 0.37 0.32 0.34 0.53
Lasso 0.02 0.21 0.02 0.25 0.02 0.41
Elastic Net 0.05 0.22 0.04 0.29 0.04 0.44
Adaptive 0.02 0.20 0.02 0.26 0.02 0.39
Relaxed 0.02 0.22 0.01 0.28 0.01 0.43
0.2 MCP 0.00 0.32 0.00 0.36 0.00 0.51
SIS-SCAD 0.01 0.30 0.00 0.34 0.00 0.46
RuleFit 0.05 0.26 0.04 0.38 0.03 0.59
RE-Lasso-100 0.40 0.15 0.30 0.22 0.29 0.40
RE-EN-100 0.56 0.13 0.48 0.22 0.48 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.17 0.18 0.13 0.28 0.12 0.46
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 70: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.5, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.45 0.20 0.48 0.30 0.45 0.48
Split-EN-10 0.41 0.22 0.54 0.28 0.51 0.47
Lasso 0.09 0.34 0.07 0.38 0.05 0.49
Elastic Net 0.12 0.31 0.10 0.38 0.08 0.50
Adaptive 0.09 0.32 0.07 0.38 0.05 0.48
Relaxed 0.08 0.41 0.07 0.39 0.05 0.50
0.4 MCP 0.03 0.36 0.02 0.44 0.01 0.50
SIS-SCAD 0.03 0.57 0.01 0.56 0.01 0.58
RuleFit 0.22 0.25 0.17 0.38 0.12 0.52
RE-Lasso-100 0.65 0.12 0.60 0.24 0.56 0.44
RE-EN-100 0.74 0.11 0.70 0.23 0.67 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.50 0.18 0.38 0.31 0.29 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.36 0.21 0.44 0.32 0.43 0.49
Split-EN-10 0.38 0.21 0.52 0.29 0.51 0.47
Lasso 0.07 0.29 0.06 0.39 0.05 0.50
Elastic Net 0.10 0.28 0.09 0.39 0.07 0.50
Adaptive 0.07 0.31 0.06 0.38 0.04 0.49
Relaxed 0.07 0.35 0.06 0.40 0.04 0.49
0.3 MCP 0.02 0.39 0.02 0.46 0.01 0.52
SIS-SCAD 0.03 0.54 0.01 0.55 0.01 0.57
RuleFit 0.18 0.24 0.14 0.39 0.10 0.52
RE-Lasso-100 0.58 0.12 0.56 0.25 0.53 0.43
RE-EN-100 0.68 0.11 0.67 0.23 0.64 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.41 0.18 0.34 0.32 0.26 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.34 0.22 0.39 0.33 0.38 0.51
Split-EN-10 0.41 0.21 0.46 0.30 0.45 0.49
Lasso 0.06 0.29 0.05 0.37 0.04 0.48
Elastic Net 0.09 0.28 0.07 0.37 0.06 0.48
Adaptive 0.06 0.30 0.04 0.37 0.04 0.47
Relaxed 0.06 0.31 0.05 0.38 0.04 0.48
0.2 MCP 0.01 0.37 0.01 0.50 0.01 0.51
SIS-SCAD 0.02 0.48 0.01 0.55 0.01 0.57
RuleFit 0.14 0.27 0.11 0.41 0.08 0.56
RE-Lasso-100 0.51 0.13 0.49 0.25 0.45 0.43
RE-EN-100 0.61 0.12 0.59 0.23 0.56 0.43
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.34 0.19 0.26 0.30 0.22 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 71: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.51 0.20 0.51 0.31 0.45 0.49
Split-EN-10 0.53 0.21 0.61 0.29 0.56 0.49
Lasso 0.08 0.31 0.06 0.38 0.04 0.44
Elastic Net 0.13 0.35 0.10 0.41 0.08 0.50
Adaptive 0.07 0.31 0.06 0.37 0.04 0.46
Relaxed 0.06 0.52 0.06 0.40 0.04 0.45
0.4 MCP 0.02 0.33 0.02 0.42 0.01 0.48
SIS-SCAD 0.03 0.71 0.02 0.70 0.01 0.69
RuleFit 0.22 0.28 0.17 0.41 0.12 0.53
RE-Lasso-100 0.63 0.12 0.58 0.23 0.53 0.41
RE-EN-100 0.74 0.11 0.70 0.22 0.67 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.54 0.22 0.45 0.40 0.33 0.55
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.45 0.22 0.50 0.33 0.42 0.50
Split-EN-10 0.52 0.21 0.59 0.30 0.53 0.49
Lasso 0.07 0.29 0.06 0.39 0.04 0.43
Elastic Net 0.12 0.31 0.11 0.43 0.07 0.49
Adaptive 0.07 0.29 0.05 0.37 0.04 0.44
Relaxed 0.06 0.43 0.05 0.40 0.04 0.44
0.3 MCP 0.02 0.38 0.02 0.48 0.01 0.49
SIS-SCAD 0.03 0.61 0.01 0.62 0.01 0.60
RuleFit 0.20 0.30 0.16 0.42 0.10 0.53
RE-Lasso-100 0.57 0.12 0.55 0.24 0.50 0.42
RE-EN-100 0.69 0.12 0.69 0.23 0.64 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.48 0.25 0.40 0.40 0.30 0.54
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.43 0.23 0.42 0.34 0.37 0.51
Split-EN-10 0.48 0.23 0.52 0.32 0.47 0.51
Lasso 0.06 0.28 0.04 0.35 0.03 0.41
Elastic Net 0.10 0.32 0.08 0.39 0.06 0.46
Adaptive 0.06 0.29 0.04 0.35 0.03 0.40
Relaxed 0.05 0.41 0.04 0.37 0.03 0.45
0.2 MCP 0.02 0.56 0.01 0.52 0.01 0.54
SIS-SCAD 0.02 0.61 0.01 0.60 0.01 0.63
RuleFit 0.15 0.31 0.12 0.42 0.08 0.55
RE-Lasso-100 0.51 0.13 0.48 0.24 0.43 0.41
RE-EN-100 0.65 0.13 0.62 0.24 0.57 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.40 0.26 0.32 0.39 0.24 0.52
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 72: Mean recall (RCL) and precision (PRC) for Scenario 3b, $\rho_{1}=$ 0.8, $\rho_{2}=$ 0.5, $n=$ 100, $p=$ 1000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.42 0.20 0.42 0.33 0.34 0.50
Split-EN-10 0.45 0.21 0.51 0.31 0.44 0.49
Lasso 0.06 0.26 0.05 0.33 0.03 0.40
Elastic Net 0.10 0.26 0.09 0.36 0.06 0.43
Adaptive 0.06 0.25 0.05 0.33 0.03 0.41
Relaxed 0.06 0.29 0.05 0.33 0.03 0.40
0.4 MCP 0.02 0.29 0.01 0.36 0.01 0.42
SIS-SCAD 0.02 0.50 0.01 0.48 0.01 0.49
RuleFit 0.18 0.26 0.13 0.37 0.08 0.50
RE-Lasso-100 0.54 0.13 0.52 0.24 0.46 0.41
RE-EN-100 0.67 0.12 0.65 0.23 0.60 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.41 0.20 0.31 0.33 0.22 0.48
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.42 0.20 0.41 0.33 0.34 0.51
Split-EN-10 0.51 0.20 0.50 0.32 0.45 0.50
Lasso 0.05 0.26 0.04 0.31 0.03 0.40
Elastic Net 0.09 0.27 0.08 0.34 0.06 0.43
Adaptive 0.05 0.26 0.04 0.32 0.03 0.40
Relaxed 0.05 0.31 0.04 0.30 0.03 0.41
0.3 MCP 0.01 0.34 0.01 0.34 0.01 0.51
SIS-SCAD 0.02 0.43 0.01 0.37 0.01 0.53
RuleFit 0.15 0.25 0.11 0.38 0.08 0.53
RE-Lasso-100 0.52 0.13 0.48 0.24 0.42 0.40
RE-EN-100 0.64 0.12 0.62 0.23 0.57 0.41
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.39 0.21 0.28 0.33 0.20 0.49
XGB 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbf{\pi_{1}}$ Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 0.44 0.22 0.41 0.35 0.35 0.55
Split-EN-10 0.52 0.20 0.50 0.33 0.44 0.53
Lasso 0.04 0.23 0.03 0.30 0.03 0.39
Elastic Net 0.08 0.26 0.07 0.33 0.06 0.44
Adaptive 0.05 0.24 0.03 0.29 0.03 0.39
Relaxed 0.04 0.26 0.03 0.31 0.02 0.40
0.2 MCP 0.01 0.49 0.00 0.39 0.00 0.45
SIS-SCAD 0.01 0.42 0.01 0.42 0.00 0.46
RuleFit 0.12 0.30 0.09 0.42 0.07 0.54
RE-Lasso-100 0.46 0.14 0.40 0.24 0.38 0.41
RE-EN-100 0.61 0.13 0.55 0.23 0.53 0.42
RGLM-100 0.00 0.00 0.00 0.00 0.00 0.00
RF-500 0.32 0.22 0.21 0.31 0.17 0.48
XGB 0.00 0.00 0.00 0.00 0.00 0.00
Table 73: ACC and TSL relative performances for GSE20347 and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 2.31 1.00 2.31 1.05 2.31 1.07 2.31 1.08
Split-EN-CV 2.31 1.03 2.31 1.00 2.31 1.00 2.31 1.00
Lasso 2.26 1.92 2.26 1.95 2.26 1.89 2.26 1.89
EN 2.31 1.12 2.31 1.13 2.31 1.15 2.31 1.15
Adaptive 2.24 3.92 2.22 4.64 2.21 4.98 2.21 5.10
Relaxed 2.25 5.36 2.24 5.27 2.26 5.38 2.25 4.61
MCP 2.21 3.99 2.21 3.94 2.20 3.97 2.21 3.86
SIS-SCAD 2.20 4.66 2.21 4.64 2.21 4.52 2.21 4.43
RuleFit 2.20 7.17 2.20 7.09 2.20 7.08 2.20 7.01
RE-Lasso-100 2.30 1.17 2.29 1.23 2.30 1.25 2.30 1.34
RE-EN-100 2.31 1.01 2.31 1.04 2.31 1.08 2.31 1.11
RGLM-100 2.29 2.00 2.30 1.92 2.29 2.07 2.29 2.06
RF-500 1.00 14.32 1.00 14.14 1.00 14.15 1.00 13.99
XGB 2.20 5.27 2.20 5.24 2.20 5.22 2.20 5.17
Table 74: ACC and TSL relative performances for GSE23400 (part one) and training proportion 0.35.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.05 1.12 1.06 1.08 1.06 1.04 1.06 1.05
Split-EN-CV 1.05 1.05 1.05 1.07 1.06 1.05 1.06 1.03
Lasso 1.04 1.20 1.05 1.20 1.05 1.21 1.05 1.18
EN 1.05 1.07 1.05 1.09 1.06 1.09 1.06 1.09
Adaptive 1.01 1.71 1.03 1.73 1.04 1.61 1.04 1.72
Relaxed 1.03 3.82 1.04 3.65 1.04 3.66 1.04 4.15
MCP 1.02 1.59 1.02 1.66 1.02 1.68 1.02 1.67
SIS-SCAD 1.01 1.80 1.01 1.95 1.01 2.00 1.01 2.00
RuleFit 1.00 3.28 1.00 3.59 1.00 3.62 1.00 3.64
RE-Lasso-100 1.05 1.01 1.06 1.01 1.06 1.02 1.06 1.01
RE-EN-100 1.05 1.00 1.06 1.00 1.06 1.00 1.06 1.00
RGLM-100 1.04 1.23 1.05 1.21 1.05 1.22 1.06 1.25
RF-500 1.04 1.30 1.05 1.41 1.05 1.46 1.05 1.49
XGB 1.00 2.18 1.00 2.37 1.00 2.43 1.00 2.45
Table 75: ACC and TSL relative performances for GSE23400 (part one) and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.03 1.11 1.03 1.08 1.03 1.03 1.03 1.03
Split-EN-CV 1.03 1.06 1.03 1.07 1.03 1.04 1.02 1.02
Lasso 1.02 1.13 1.02 1.15 1.02 1.16 1.02 1.16
EN 1.03 1.06 1.02 1.07 1.02 1.06 1.02 1.07
Adaptive 1.00 1.56 1.01 1.57 1.01 1.62 1.01 1.66
Relaxed 1.02 2.57 1.02 2.23 1.02 2.20 1.02 2.76
MCP 1.00 1.55 1.01 1.59 1.01 1.60 1.00 1.61
SIS-SCAD 1.01 1.72 1.01 1.78 1.01 1.80 1.01 1.82
RuleFit 1.00 2.13 1.00 2.36 1.00 2.37 1.00 2.34
RE-Lasso-100 1.03 1.00 1.03 1.00 1.03 1.02 1.03 1.02
RE-EN-100 1.03 1.03 1.03 1.01 1.03 1.00 1.03 1.00
RGLM-100 1.03 1.13 1.03 1.19 1.03 1.18 1.03 1.18
RF-500 1.03 1.33 1.02 1.44 1.02 1.44 1.02 1.45
XGB 1.01 1.78 1.00 1.91 1.00 1.96 1.00 1.98
Table 76: ACC and TSL relative performances for GSE23400 (part two) and training proportion 0.35.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.04 1.21 1.05 1.11 1.04 1.06 1.04 1.05
Split-EN-CV 1.04 1.14 1.04 1.10 1.05 1.08 1.04 1.05
Lasso 1.04 1.13 1.05 1.11 1.05 1.13 1.04 1.11
EN 1.05 1.09 1.04 1.08 1.04 1.07 1.04 1.07
Adaptive 1.04 1.30 1.05 1.21 1.05 1.27 1.04 1.25
Relaxed 1.03 3.27 1.04 2.79 1.04 3.14 1.04 3.13
MCP 1.04 1.32 1.04 1.31 1.04 1.33 1.03 1.32
SIS-SCAD 1.03 1.58 1.03 1.60 1.03 1.62 1.03 1.60
RuleFit 1.00 2.92 1.01 2.96 1.00 3.02 1.00 2.99
RE-Lasso-100 1.05 1.00 1.05 1.00 1.05 1.00 1.05 1.00
RE-EN-100 1.05 1.04 1.05 1.01 1.05 1.00 1.05 1.04
RGLM-100 1.04 1.03 1.05 1.09 1.04 1.14 1.04 1.14
RF-500 1.03 1.30 1.04 1.34 1.04 1.41 1.04 1.46
XGB 1.00 2.00 1.00 2.05 1.00 2.04 1.00 1.97
Table 77: ACC and TSL relative performances for GSE23400 (part two) and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.02 1.10 1.03 1.10 1.04 1.07 1.03 1.02
Split-EN-CV 1.03 1.07 1.03 1.07 1.03 1.06 1.03 1.03
Lasso 1.03 1.12 1.02 1.14 1.03 1.15 1.03 1.14
EN 1.03 1.08 1.03 1.04 1.03 1.08 1.03 1.05
Adaptive 1.04 1.23 1.03 1.23 1.04 1.28 1.03 1.24
Relaxed 1.03 2.83 1.03 2.98 1.03 3.41 1.02 2.82
MCP 1.04 1.21 1.03 1.27 1.04 1.28 1.03 1.28
SIS-SCAD 1.04 1.34 1.04 1.42 1.04 1.48 1.04 1.50
RuleFit 1.00 1.94 1.00 2.09 1.00 2.02 1.00 2.02
RE-Lasso-100 1.04 1.01 1.03 1.00 1.04 1.00 1.04 1.04
RE-EN-100 1.03 1.03 1.04 1.01 1.04 1.02 1.04 1.00
RGLM-100 1.02 1.00 1.03 1.06 1.03 1.08 1.03 1.10
RF-500 1.02 1.26 1.02 1.31 1.03 1.38 1.02 1.43
XGB 1.00 1.59 1.00 1.58 1.00 1.70 1.00 1.60
Table 78: ACC and TSL relative performances for GSE5364 (Esophageal) and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 2.12 1.06 2.13 1.20 2.19 1.24 2.15 1.32
Split-EN-CV 2.14 1.01 2.15 1.00 2.22 1.02 2.18 1.08
Lasso 2.00 1.70 2.01 1.84 2.06 1.96 2.03 2.10
EN 2.10 1.20 2.10 1.29 2.16 1.32 2.13 1.35
Adaptive 1.92 2.59 1.86 3.00 1.91 3.00 1.75 3.65
Relaxed 1.98 5.60 1.97 6.49 2.03 6.53 2.01 6.69
MCP 1.79 2.72 1.78 3.13 1.80 3.34 1.81 3.35
SIS-SCAD 1.81 2.68 1.80 3.05 1.85 3.20 1.83 3.25
RuleFit 1.82 6.36 1.82 7.05 1.87 7.39 1.85 7.54
RE-Lasso-100 2.11 1.16 2.12 1.17 2.17 1.18 2.16 1.23
RE-EN-100 2.14 1.00 2.15 1.02 2.21 1.00 2.20 1.00
RGLM-100 2.02 1.61 2.05 1.69 2.15 1.73 2.13 1.80
RF-500 1.00 4.55 1.00 5.04 1.00 5.27 1.00 5.38
XGB 1.84 2.85 1.84 3.17 1.89 3.32 1.87 3.38
Table 79: ACC and TSL relative performances for GSE25869 and training proportion 0.35.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.10 1.66 1.18 1.43 1.23 1.29 1.24 1.21
Split-EN-CV 1.10 1.51 1.19 1.34 1.23 1.21 1.25 1.15
Lasso 1.07 1.59 1.14 1.47 1.20 1.36 1.18 1.32
EN 1.09 1.50 1.17 1.36 1.22 1.28 1.22 1.23
Adaptive 1.00 1.29 1.00 1.36 1.00 1.41 1.00 1.45
Relaxed 1.07 4.05 1.14 3.40 1.19 3.37 1.17 3.48
MCP 1.04 1.39 1.06 1.46 1.09 1.45 1.09 1.46
SIS-SCAD 1.03 1.35 1.10 1.40 1.14 1.39 1.14 1.41
RuleFit 1.06 2.26 1.13 2.36 1.17 2.40 1.17 2.43
RE-Lasso-100 1.09 1.30 1.17 1.22 1.22 1.17 1.22 1.16
RE-EN-100 1.10 1.38 1.18 1.23 1.23 1.17 1.24 1.14
RGLM-100 1.08 1.00 1.17 1.00 1.20 1.00 1.21 1.00
RF-500 1.08 1.05 1.14 1.11 1.19 1.12 1.20 1.14
XGB 1.06 1.32 1.12 1.39 1.16 1.39 1.16 1.40
Table 80: ACC and TSL relative performances for GSE25869 and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.09 1.46 1.07 1.32 1.14 1.30 1.19 1.23
Split-EN-CV 1.08 1.34 1.07 1.30 1.14 1.23 1.20 1.23
Lasso 1.06 1.48 1.04 1.51 1.10 1.51 1.16 1.47
EN 1.08 1.38 1.06 1.33 1.12 1.33 1.17 1.33
Adaptive 1.00 1.27 1.00 1.35 1.00 1.51 1.00 1.59
Relaxed 1.08 3.39 1.03 3.59 1.09 3.70 1.14 3.31
MCP 1.04 1.35 1.02 1.42 1.08 1.51 1.10 1.49
SIS-SCAD 1.04 1.30 1.00 1.38 1.06 1.44 1.10 1.45
RuleFit 1.04 1.34 1.01 1.43 1.07 1.48 1.12 1.51
RE-Lasso-100 1.08 1.24 1.05 1.19 1.11 1.24 1.18 1.23
RE-EN-100 1.08 1.24 1.06 1.20 1.14 1.20 1.17 1.20
RGLM-100 1.09 1.02 1.06 1.00 1.13 1.00 1.17 1.00
RF-500 1.08 1.00 1.06 1.08 1.12 1.13 1.17 1.15
XGB 1.02 1.21 1.01 1.29 1.06 1.35 1.12 1.36
Table 81: ACC and TSL relative performances for GSE10245 and training proportion 0.35.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.14 1.12 1.15 1.05 1.15 1.05 1.14 1.07
Split-EN-CV 1.14 1.10 1.15 1.00 1.15 1.00 1.15 1.00
Lasso 1.12 1.43 1.12 1.47 1.12 1.50 1.11 1.46
EN 1.14 1.16 1.14 1.17 1.14 1.20 1.14 1.16
Adaptive 1.06 1.76 1.06 1.86 1.04 1.97 1.03 1.93
Relaxed 1.10 4.00 1.11 3.66 1.11 3.34 1.11 3.66
MCP 1.05 1.78 1.05 1.75 1.05 1.84 1.04 1.74
SIS-SCAD 1.05 1.78 1.05 1.74 1.05 1.79 1.05 1.71
RuleFit 1.00 4.38 1.00 4.39 1.00 4.42 1.01 4.22
RE-Lasso-100 1.14 1.10 1.15 1.09 1.14 1.23 1.13 1.18
RE-EN-100 1.15 1.00 1.15 1.00 1.15 1.01 1.14 1.05
RGLM-100 1.11 1.72 1.12 1.49 1.13 1.37 1.13 1.32
RF-500 1.11 1.48 1.10 1.51 1.10 1.54 1.10 1.50
XGB 1.00 1.99 1.00 1.99 1.00 1.99 1.00 1.90
Table 82: ACC and TSL relative performances for GSE10245 and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.14 1.13 1.15 1.04 1.15 1.06 1.15 1.05
Split-EN-CV 1.14 1.03 1.15 1.00 1.15 1.00 1.15 1.00
Lasso 1.12 1.36 1.11 1.43 1.12 1.44 1.11 1.50
EN 1.14 1.11 1.14 1.15 1.14 1.13 1.15 1.14
Adaptive 1.05 1.93 1.03 2.31 1.01 2.31 1.00 2.32
Relaxed 1.11 3.27 1.11 2.98 1.12 2.61 1.11 3.06
MCP 1.04 1.89 1.05 2.02 1.06 1.98 1.07 1.91
SIS-SCAD 1.02 2.02 1.03 2.22 1.05 2.16 1.05 2.06
RuleFit 1.00 4.32 1.00 4.75 1.01 4.73 1.01 4.55
RE-Lasso-100 1.14 1.09 1.14 1.15 1.13 1.29 1.12 1.26
RE-EN-100 1.14 1.00 1.14 1.07 1.13 1.11 1.14 1.12
RGLM-100 1.09 2.08 1.12 1.61 1.13 1.46 1.13 1.42
RF-500 1.10 1.57 1.11 1.67 1.10 1.68 1.12 1.63
XGB 1.00 2.32 1.00 2.54 1.00 2.56 1.01 2.47
Table 83: ACC and TSL relative performances for GSE5364 (Lung) and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.76 1.03 1.75 1.05 1.76 1.01 1.74 1.01
Split-EN-CV 1.76 1.00 1.76 1.00 1.75 1.00 1.77 1.00
Lasso 1.57 1.64 1.57 1.62 1.58 1.59 1.60 1.56
EN 1.70 1.21 1.68 1.19 1.69 1.16 1.70 1.12
Adaptive 1.52 1.75 1.48 1.83 1.49 1.81 1.44 1.80
Relaxed 1.54 4.99 1.54 4.33 1.54 4.25 1.54 5.00
MCP 1.47 1.92 1.45 1.95 1.43 1.94 1.41 1.96
SIS-SCAD 1.44 1.86 1.44 1.87 1.44 1.83 1.43 1.83
RuleFit 1.44 4.32 1.47 4.34 1.45 4.27 1.46 4.14
RE-Lasso-100 1.73 1.05 1.73 1.07 1.74 1.07 1.74 1.07
RE-EN-100 1.74 1.01 1.74 1.01 1.74 1.00 1.75 1.00
RGLM-100 1.69 1.34 1.74 1.33 1.75 1.29 1.75 1.31
RF-500 1.00 2.69 1.00 2.69 1.00 2.62 1.00 2.58
XGB 1.48 1.87 1.48 1.85 1.48 1.83 1.48 1.79
Table 84: ACC and TSL relative performances for GSE5364 (Thyroid) and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.09 1.23 1.12 1.12 1.11 1.10 1.10 1.08
Split-EN-CV 1.09 1.11 1.12 1.05 1.11 1.06 1.10 1.05
Lasso 1.04 1.38 1.05 1.38 1.06 1.43 1.05 1.34
EN 1.07 1.25 1.08 1.22 1.09 1.21 1.08 1.20
Adaptive 1.07 1.20 1.08 1.21 1.05 1.26 1.04 1.25
Relaxed 1.03 3.16 1.04 3.01 1.04 3.08 1.04 2.86
MCP 1.00 1.32 1.02 1.32 1.02 1.30 1.00 1.31
SIS-SCAD 1.00 1.33 1.00 1.34 1.00 1.29 1.00 1.29
RuleFit 1.03 1.99 1.04 1.94 1.05 1.91 1.04 1.90
RE-Lasso-100 1.09 1.09 1.11 1.05 1.11 1.07 1.11 1.08
RE-EN-100 1.10 1.07 1.12 1.03 1.12 1.00 1.11 1.00
RGLM-100 1.08 1.19 1.11 1.10 1.11 1.00 1.12 1.01
RF-500 1.10 1.00 1.12 1.00 1.13 1.01 1.12 1.05
XGB 1.02 1.19 1.03 1.19 1.04 1.15 1.03 1.16
Table 85: ACC and TSL relative performances for GSE21942 and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 2.23 1.31 2.22 1.09 2.21 1.13 2.20 1.05
Split-EN-CV 2.22 1.30 2.20 1.30 2.20 1.28 2.20 1.19
Lasso 2.23 1.26 2.21 1.25 2.20 1.30 2.19 1.29
EN 2.25 1.00 2.24 1.00 2.23 1.00 2.23 1.00
Adaptive 2.15 2.32 2.10 2.60 2.12 2.52 2.09 2.52
Relaxed 2.19 4.46 2.18 4.40 2.17 4.53 2.17 3.55
MCP 2.06 3.37 2.04 3.32 2.02 3.46 2.02 3.23
SIS-SCAD 2.03 3.61 2.02 3.52 2.02 3.50 2.01 3.29
RuleFit 2.01 8.05 2.00 7.84 1.99 7.80 1.98 7.34
RE-Lasso-100 2.24 1.08 2.21 1.10 2.21 1.16 2.20 1.15
RE-EN-100 2.22 1.22 2.19 1.25 2.19 1.33 2.19 1.35
RGLM-100 2.19 2.07 2.18 2.14 2.19 2.29 2.17 2.20
RF-500 1.00 8.23 1.00 8.03 1.00 7.99 1.00 7.51
XGB 2.01 4.27 2.00 4.18 1.99 4.17 1.98 3.91
Table 86: ACC and TSL relative performances for GSE14905 and training proportion 0.5.
$\mathbf{\boldsymbol{p}=100}$ $\mathbf{\boldsymbol{p}=250}$ $\mathbf{\boldsymbol{p}=500}$ $\mathbf{\boldsymbol{p}=1000}$
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.04 1.04 1.04 1.02 1.04 1.10 1.04 1.08
Split-EN-CV 1.05 1.01 1.05 1.01 1.04 1.02 1.04 1.05
Lasso 1.02 1.69 1.03 1.73 1.03 1.65 1.03 1.73
EN 1.05 1.08 1.04 1.15 1.04 1.17 1.04 1.20
Adaptive 1.02 2.02 1.02 2.29 1.02 2.34 1.01 2.43
Relaxed 1.02 4.87 1.03 3.69 1.03 4.19 1.02 4.21
MCP 1.02 2.84 1.02 2.86 1.02 2.79 1.02 2.78
SIS-SCAD 1.02 3.24 1.02 3.22 1.02 3.19 1.02 3.14
RuleFit 1.00 5.83 1.00 5.85 1.00 5.78 1.00 5.76
RE-Lasso-100 1.04 1.24 1.03 1.25 1.03 1.25 1.03 1.28
RE-EN-100 1.05 1.00 1.05 1.00 1.04 1.00 1.05 1.00
RGLM-100 1.03 1.66 1.03 1.42 1.03 1.40 1.03 1.41
RF-500 1.05 1.68 1.05 1.86 1.05 1.99 1.06 2.15
XGB 1.00 3.59 1.00 3.61 1.00 3.58 1.00 3.57